A Bibliography of Publications in Theoretical Computer Science: 2010–2014

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

22 March 2018 Version 1.92

Title word cross-reference

(1 + ϵ) [1193]. (1 + λ) [2326]. (1, λ) [2255]. (3 − 2/(Δ + 1)) [718]. (Δ + 1) [2344]. (n + k) [478]. (n, k) [1974]. (s, t) [985, 1157]. −β [908]. {0, 1} [1730]. 1 [1057, 83, 491, 1948, 2238, 1585, 2362]. 1 + ln 2 [1692]. 1/2 [540]. 1/3^{h} [2458]. 10 [896]. 2 [2361, 2396, 760, 1692, 2107, 1918, 2084, 1235, 2242, 879, 1889, 1872, 819, 1380, 1993, 335, 2362, 253]. 2.375 [1567]. 2/3 [818]. 2^{k} [205]. 2n [1143]. 2P_{3} [1033, 1406]. 2 × n [1839]. 3 [2087, 812, 1061, 1583, 934, 204, 471, 502, 2146, 1498, 1886, 810]. 4 [1818, 736, 1254, 204, 1380]. 4.31 [2247]. 5 [1560, 1249, 896, 2000, 2213]. 5n [494]. 6 [1072, 896, 1993]. 7 [498, 896, 2024]. 8

[1560, 896]. 9 [896]. 9k [1979]. _{hs}^{srt} [1243]. ; AC^{0} [743]. α [435]. b [131, 2246]. β ; [1977, 2218, 444]. c [793, 741]. C^{2} [761]. C^{∞} ; [292]. N P [1625]. PLS [1494]. T [111]. D ; [1114, 192, 126]. δ [634]. ℓ^{1} [2004, 467]. ℓ_{∞} ; [1587]. ϵ [1795, 634, 170, 622]. f [1340]. ; F_{q}((x^{−1})) [1977]. ; FFD(L) ≤ (11/9)OPT(L) + 6/9 [1907]. ; G(2^{m}, 4) [742]. H ; [2354, 992, 236, 243, 1252, 1063]. K ; [1509, 2241, 77, 1936, 2076, 1289, 1395, 711, ; 2040, 1858, 371, 1416, 1730, 2288, 734, 2238, ; 1085, 1341, 1791, 1035, 1953, 1219, 1342, 968, ; 367, 79, 1211, 739, 1951, 2229, 785, 478, 730, ; 1051, 901, 1407, 2398, 1212, 1703, 2169]. K_{2,2} ; [29]. L [1551, 988, 2076, 1929]. L(2, 1) ; [2353, 2251, 1860, 1675, 256, 551]. L_{2} [2358]. ; L_{p} [2150, 2031]. λ [1268, 445, 770, 343]. Λμ

[1165]. log(n) [1843]. M [850, 751, 1198]. Z^{2} [1865]. ML^{F} [991]. μ [2451, 621, 338, 624]. N [2324, 829, 1401, 144, 1219, 272, 478, 256, 730, 901, 1407, 1700, 2061, 1703, 251]. n − 3 [1974]. NC^{1} [988]. ≠ [1789]. n log n [265]. ν [1815]. O(1) [1240]. O(m(m + n log n) log(nC)) [1266]. O(n^{1+ϵ}) [696]. O(n^{2}) [1731, 555, 1240]. O(nm) [1312]. O ∗ (3.53^{3k}) [502]. ω [660, 900]. P [1685, 357, 356, 1083, 2324, 352, 1347, 1103, 144, 1058]. P_{2} [2047]. P_{3} [937]. P_{4} [1197]. P_{5} [1978]. ϕ [1553]. π [2278, 1242, 149]. q [2314, 2123]. QR [480]. R [1103, 407]. R^{3} [1138]. S [806, 80]. Σ_{2}^{0} [496]. SN [444]. ⋆ [1030]. T [1999, 80]. T_{S4} [340]. V [2136, 243, 1642]. w [1698]. x^{n} [641, 1073]. XA ⊆ BX [2104]. z [793]. Z^{d} [764].

-abelian [1342]. -ANT [491]. -approximation [1843, 1692, 1918, 2247, 1193, 1072, 1567, 818, 718]. -ary [1219, 272, 730, 901, 1407, 1703]. -axis [1249]. -balanced [1815]. -balls [2004]. -barrel [2218]. -block [2238, 2238]. -bounded [1051]. -calculus [1242, 621, 338, 2451, 2278, 445, 343, 149, 1165]. -Capture [2040]. -center [1235, 1058]. -Chinese [1953]. -CNF [739]. -color [760]. -colorable [204]. -coloring [498, 204, 2353, 2354, 2344]. -competitive [1553]. -complete [352]. -completions [1114]. -complexity [1494]. -Compressive [467]. -conjecture [1143]. -connected [1254, 2146, 335, 253]. -Connecting [2361]. -constraint [2084]. -context-free [968]. -convex [243, 1551]. -covers [1035]. -cube [1219, 1407, 2061]. -cubes [730, 901, 1700, 1703]. -cuts [985]. -cycles [1560, 1380, 1993, 2000, 2024]. -degenerate [2396]. -dimensional [2362]. -disconnectivity [1157]. -Diversity [1929]. -edge-connectivity [1692]. -edge-fault-tolerant [1974]. -expansion [1977]. -factorization [793]. -factors [1340]. -fixed [622]. -fixed-endpoint [77]. -free [1033, 1978, 1406]. -gaps [2076]. -gatherings [407]. -gram [2314, 2123]. -hardness [1625]. -heavy-hitters [2358]. -Hitting [126]. -in-a-row [734, 144]. -inconnected [367]. -index [1252, 1063]. -indicator [1795]. -indiscernibility [850]. -labeling [551, 1860]. -labelings [2251, 256, 1675]. -letter [879]. -matching [131]. -matchings [2246]. -means [1858, 1211]. -median [1057, 83]. -minor [236]. -minor-free [992]. -mismatches [2076]. -nomials [478]. -norm [2031]. -operator [1030]. -order [1642, 2136]. -orientations [80]. -Packing [2047]. -path [934]. -planarity [1948]. -player [251, 829, 1401]. -queens [2324]. -recognizable [806]. -reduction [770]. -regular [1416, 1730, 736]. -resource [1395]. -restricted [2398]. -sat [1872]. -set [1936, 711, 502]. -sided [2087]. -space [2362]. -sparse [1197]. -stack [1818]. -star [1974]. -subgraph [2169]. -surviving [1380, 1993]. -terms [624, 900]. -time [1731, 1266, 1312, 1999, 502]. -tournaments [741]. -transformation [908]. -tree [371]. -uniform [1585]. -valued [1289]. -variate [478]. -vertex [1730]. -wide [660]. -words [435, 292].

1-tape [1411]. 1/2 [1712].

2-symbol [1411]. 2008 [493]. 2010 [1351]. 218 [2292]. 2SAT [125].

3 [2055]. 3-SAT [2055]. 34-36 [2459]. 3462 [1073]. 380 [1712].

400 [1557]. 411 [2459]. 412 [1040, 1073]. 423 [1706]. 429 [2456]. 469 [1678]. 474 [1836].

524 [2457]. 552 [2461]. 557 [2460].

60th [1845].

= [1789].

A. [2280]. a/sync [1095, 1344]. abelian [697, 1342, 2110, 872, 1653, 1497]. Abelian-primitive [1653]. Above [1923, 1947]. absent [1296]. absolute [1907]. Absoluteness [1002]. Abstract [351, 2250, 1354, 867, 169, 679, 1451, 1836, 1044, 1173, 1476, 1111, 1989, 1540, 112]. Abstraction [344, 1130, 2189, 269, 533, 1422]. abstraction-refinement [1422]. abstractions [2050, 1127]. abstractly [1163]. accelerate [453]. Accelerated [2324, 2008, 626]. accepted [310]. Accepting [184, 1359]. Access [869]. accessible [50]. accretive [488]. accurate [483]. achievable [1135]. achieve [2384]. acoustic [1257]. ACP [1243]. across [1954]. activation [1964]. activations [213]. Active [212, 2117, 506, 1083, 2324, 68, 2116, 48]. Acyclic [267, 671, 1367, 800, 1847, 1290, 1945, 273, 2354, 1835, 1615, 1461]. acyclicity [1836, 1540]. Ad [2340, 2216, 448, 628, 1887]. ad-hoc [448]. adaptation [2009]. adaptations [2449]. Adapting [320]. Adaptive [241, 2004, 332, 704, 573, 1394, 1952, 784, 508, 1857, 895]. Adaptively [1584]. Adaptivity [1982]. add [169]. Addendum [1706, 1712]. Adding [2157, 1750, 444]. addition [708, 832, 2238]. additional [2399, 930]. additive [387, 987, 307, 2283]. additively [2161]. address [648, 1959]. adjacency [2315, 723]. adjacent [1698, 1984]. adjoint [31]. Administrative [869]. admissible [340]. Admissions [245]. Advances [755, 2327]. advantage [252]. adversary [784]. Advice [2097, 2353, 2359, 208, 578, 115]. Affine [757, 309, 969, 1634, 1500]. affine-invariant [969]. age [1115]. agencies [1307]. agent [866, 2236, 427, 1490]. agents [1586, 543, 1900, 122, 2060, 9, 1901, 18, 117, 1534, 2010]. aggregating [907]. aggregation [1773]. aggregations [1303]. aggressive [849]. aging [402]. agreement [1936, 711, 1751, 10, 2366]. AIMD [1427]. Akl [20]. ALC [1861]. alert [1321]. algebra [1682, 623, 861, 1060, 173, 1243, 486, 634, 2195, 1986]. Algebraic [1771, 1092, 147, 609, 391, 2408, 1696, 1579, 2309, 628, 2274, 1592, 956, 1114, 2158, 2018, 517, 2296]. algebras [776, 1125, 146, 32, 642, 898, 1806]. Algorithm [253, 2055, 406, 395, 1040, 1553, 1225, 1843, 1924, 764, 760, 987, 2448, 1739, 1303, 266, 1294, 1484, 2369, 27, 1692, 2009, 1754, 1731, 1972, 1485, 1642, 369, 1948, 1674, 2376, 640, 876, 1008, 939, 647, 1918, 627, 1266, 519, 878, 1736, 737, 2319, 181, 1734, 186, 265, 482, 803, 1378, 1193, 1860, 204, 219, 554, 1567, 1764, 1775, 2237, 375, 926, 739, 818, 669, 1240, 2412, 1197, 1312, 1268, 231, 424, 1823, 982, 1881, 2231, 2255, 119, 2165, 1596, 937, 718, 819, 229, 321, 502, 1463, 1054, 2385, 1498, 2445, 474, 1382, 335, 196, 2169, 1670, 1039, 2254, 1613]. Algorithmic [1331, 1741, 1315, 1071, 2303, 72, 1921, 594, 1308, 2168, 649, 330, 13, 550, 2394]. algorithmics [2138]. Algorithms [1398, 2245, 559, 1890, 2047, 2206, 2366, 6, 587, 935, 1503, 383, 1504, 794, 964, 197, 1133, 666, 1726, 1201, 1150, 1641, 2346, 2374, 784, 1891, 1217, 1248, 22, 17, 1548, 1639, 2347, 564, 697, 874, 1171, 1761, 1509, 2391, 82, 1152, 1305, 989, 320, 859, 1702, 2069, 401, 1638, 2441, 1230, 579, 2442, 126, 84, 1059, 1771, 2192, 2084, 242, 1875, 190, 315, 408, 716, 747, 1320, 1737, 1925, 2089, 2326, 19, 2103, 2298, 1173, 693, 2342, 79, 1063, 792, 1419, 1239, 1773, 1769, 1824, 943, 1098, 2217, 1440, 76, 10, 1323, 2164, 1880, 1206, 700, 927, 1892, 992, 1461, 2062, 2343]. algorithms [1018, 2132, 1928, 365, 2061, 2362, 1763, 1783, 569, 2340, 889]. Alice [1480]. aligned [453]. aligning [424]. alignment [1426, 186, 72, 1488, 703, 970, 1037, 1634]. alignment-free [1426]. alignments [2076]. all-pairs [1512]. All-Pairs-Shortest-Path-Distance [257]. All-Partial-Sums [474]. all-shortest-paths [121]. All-to-all [124]. AllDifferent [1522]. allocation [1876, 1961, 333, 2378]. allowing [1637]. Almost [1134, 1199, 1371, 2329, 49, 1615]. almost-confluent [49]. along [2369, 2258, 675]. Alpha [1145, 475]. alphabet [2046, 1281, 1379, 887, 1626]. alphabetic [1468]. alphabets [879, 1062, 896, 1997]. Alternating [2034, 175, 2333, 269, 1234, 2090]. alternation [2451]. Ambient [1324]. ambiguation [636]. Ambiguity [2183, 645]. ambiguous [235]. among [1661, 1205, 117]. Amortized [729, 1814]. Amount [268]. ample [2193]. amplification [1664]. analog [1169]. Analysis [163, 274, 1292, 959, 1698, 2326, 1054, 1365, 1176, 961, 380, 689, 1934, 2130, 165, 1739, 103, 1119, 2341, 867, 491, 2045, 626, 1008, 1308, 1836, 1540, 1071, 1044, 1513, 26, 173, 2256, 978, 533, 1311, 2300, 2194, 254, 19, 1611, 390, 2037, 132, 899, 16, 1623, 2254, 954, 1198, 869, 732, 162, 1803, 819, 1478, 125]. Analytic [2273, 725, 2109, 2296, 2094, 1690]. Analytical [1457]. Analyzing [400, 164, 683]. anarchy [458, 1876]. Anatomy [2277]. ancestor [1945]. Ancestors [1486]. anchor [1319, 2343]. anchor-free [2343]. anchors [1634]. Ancilla [1113]. Ancilla-driven [1113]. Andersson [2025]. angle [786]. angles [471]. anisotropic [1770]. annotated [424]. annotations [39]. annulus [1893]. anonymity [172]. anonymous [1586, 1228, 1868, 1429, 1816, 718, 187, 1424]. answer [272]. ant [744, 491, 491, 183]. antenna [1773]. antennas [1904]. Anti [1569]. Anti-magic [1569]. antiblockers [2065]. antidictionary [2091]. antimorphisms [2151]. ants [837]. any [762, 926, 1984]. aperiodic [87, 1025]. apex [1184]. Application [2269, 2348, 745, 1927, 921, 2320, 277, 2443, 2300, 1917, 2101, 767, 804, 127, 763]. Applications [2437, 583, 1435, 658, 1855, 2382, 493, 1921, 50, 2347, 815, 1890, 57, 1152, 320, 1059, 1071, 596, 474]. applied [1388, 1634]. Applying [737]. approach [1716, 1295, 1303, 1203, 1945, 354, 1146, 1372, 2094, 628, 2191, 1595, 2300, 2194, 1792, 254, 1141, 954, 852, 1106, 12, 372, 1511, 1422]. approaches [749, 2296, 1591]. Approval [1786]. Approximability [1976, 216, 986, 1202, 2315, 1929, 1090, 2456, 384, 1998]. Approximate [648, 1443, 1672, 1198, 2074, 798, 17, 1057, 1136, 290, 300, 1631, 652, 882, 1609, 1597, 2031, 1259]. approximated [1606]. approximately [1178]. Approximating [958, 238, 1053, 2290, 2080, 2441, 979, 727, 1196, 470, 193, 437, 971, 2371, 2389, 802, 1245, 1785, 2347, 1069, 1023, 1258, 371]. Approximation [1504, 666, 1150, 2374, 1761, 845, 531, 1702, 432, 1638, 1875, 945, 2089, 693, 1773, 1769, 1987, 1892, 1423, 2169, 1490, 2134, 694, 1582, 2346, 1843, 1919, 1924, 2391, 1774, 483, 1183, 1694, 1020, 1692, 1251, 2376, 1008, 1918, 315, 2247, 1193, 1072, 1949, 1567, 2237, 375, 818, 2217, 1323, 1206, 982, 992, 937, 718, 819, 2343, 918, 196, 389, 569]. approximations [696, 534, 1107, 366, 2213]. APX [1262]. APX-completeness [1262]. arbitrary [921, 820, 675, 995, 1946, 2345, 862, 424, 732, 1051]. arbitrary-neighborhood [862]. arborescences [2210]. Arboricity [1063]. Arc [1186, 2085, 36, 2178, 803, 424]. Arc- [2085]. arc-annotated [424]. Arc-disjoint [1186]. architectural [1298]. architecture [699, 1780]. arcs [1545]. area [123, 1408]. arguments [576]. Arithmetic [1267, 756, 2225]. arithmetical [827]. Arithmetics [431, 687]. arity [2260]. armed [511]. ARNN [1789]. Arnold [2332]. arrangement [875, 2299, 2153]. arrays [2166, 607, 931, 61, 897, 609, 1647]. Arrows [150, 1951]. art [161]. articulation [1255]. ary [1219, 272, 730, 901, 1407, 1703]. Aspects [1351, 2177, 953, 1315, 2186, 517, 1125, 749, 1467, 1457, 609, 1111]. assembly [180, 1330, 70, 55, 1790, 358, 359, 488]. assignment [858, 14, 830, 1494, 2061, 1992, 2059]. assignments [1730, 1580, 281]. associated [872]. associative [2119]. asymmetric [2268, 1711]. Asymptotic [2296, 1534, 804, 2094, 1313]. asymptotical [1746]. Asymptotics [1828]. Asynchronous [928, 1083, 1586, 663, 639, 118, 1901, 1300, 1942, 252, 1849, 822]. asynchrony [2060]. AT-free [933, 1620]. at-most-once [1753]. atlas [1849]. átomata [2199]. atomic [1978, 1528, 2175]. attack [2101]. Attractor [2419]. Attribute [1028, 1602, 2098, 2099, 743]. Attribute-based [1028, 2098, 2099]. Attributed [1044]. auction [2198]. Auctions [2216, 28]. augmentation [199, 224, 1692]. augmented [1910, 1699, 120]. authority [189]. authorization [2173]. Automata [1714, 1968, 1218, 951, 2333, 1606, 1529, 663, 1577, 1274, 420, 2328, 2400, 1480, 2178, 155, 152, 1290, 267, 671, 1294, 307, 1170, 378, 1870, 2309, 1335, 1288, 2161, 675, 680, 174, 1850, 1975, 2422, 454, 995, 1654, 655, 268, 273, 269, 2157, 110, 1234, 2090, 2192, 348, 656, 681, 2418, 1470, 325, 271, 353, 311, 42, 1278, 1092, 2174, 1094, 956, 913, 1835, 1281, 275, 682, 2159, 1009, 1849, 2421, 1282, 439, 1287, 546, 678, 1848, 1346, 1102, 392, 1347, 1348, 81, 1207, 1106, 2423, 1142, 1155, 3, 864, 813, 24, 1107, 887, 1626, 2014, 1794]. automata [1670, 1077, 2417, 86, 1291, 339, 2449]. Automata-theoretic [1968]. Automated [708, 2352]. Automatic [1722, 509, 2152, 499, 2105]. Automaticity [2331]. automaton [2172, 274, 1292, 265, 350]. automorphism [658]. automorphisms [2107]. Autonomous [2340]. availability [376, 449, 792]. Avalanche [1851]. avalanches [685]. Average [116, 989, 2123, 186, 125, 2068, 430, 1153, 750, 16, 1923]. Average-case [2123, 125, 1153]. averaging [970]. avoidability [337, 1342]. avoidance [1866]. Avoiding [667, 897, 609, 1498]. aware [2269, 1891, 1521, 2378, 1938]. awareness [1940]. axiom [33]. axiomatic [1146, 2368]. axiomatizability [615]. Axiomatization [633]. Axiomatizing [2180, 2156]. axis [1249, 1256, 2436]. axis-parallel [1256].

Backbone [892, 1673, 1875]. backbones [1673]. backtracking [142, 2200]. Backup [1235]. backward [76]. bacterial [165]. bad [1858, 1438]. Bakery [587]. balanced [1815, 1947, 1656, 1011, 1875, 1458, 1389, 1412, 814]. Balancing [612, 552, 1230, 450, 1622]. balloon [390]. balls [2004, 2096, 1881, 2435]. bandit [2006]. bandits [511, 2410]. Bandpass [377]. Bandwidth [1840, 933, 2144, 290, 1918, 682]. bar [267, 671]. bargaining [2370]. barrel [2218]. barrier [2347]. Base [2052, 2380, 431]. base-monotone [2380]. Base-object [2380]. Based [869, 708, 123, 1503, 180, 326, 1028, 1492, 1052, 154, 2172, 108, 1358, 1248, 1903, 2115, 1483, 815, 1708, 1387, 778, 1757, 2073, 845, 2167, 1963, 761, 1127, 1084, 532, 1836, 1540, 1071, 1561, 1087, 26, 1734, 2098, 1300, 1237, 190, 428, 2432, 1793, 597, 884, 1411, 2237, 2203, 33, 534, 926, 468, 914, 1101, 1897, 260, 2307, 599, 1701, 2226, 1155, 1349, 717, 1064, 2019, 975, 2099, 1324, 312, 2018, 282, 569, 1119]. Bases [1260, 2341, 2191, 877, 1014, 1413, 62, 2019]. basic [920, 1180]. Basing [104]. basis [1392, 167, 697]. batch [1702, 801, 2235, 549]. batching [92, 1758, 2041]. Baxter [2142]. Bayes [1527]. Bayesian [1302, 85, 2198]. BCCSP [2180]. BCI [1828]. BCK [1828]. BDI [866]. be [710, 1171, 1049, 66]. beams [65]. Bee [1308]. Behavior [1013, 2438, 354, 86]. behavioral [2270, 914, 974]. behaviors [182]. behaviour [152, 2159]. behavioural [1605, 2322]. Bell [1661]. below [834]. below-upper-bound [834]. bend [2363]. beneficial [1031]. benefits [402]. Benenson [864, 3]. Bergstra [614]. Besicovitch [676]. Best [568, 1205, 1382]. Best-order [568]. Better [1006, 409, 1961, 1838, 2233]. between [383, 1504, 420, 1870, 811, 2360, 988, 218, 2273, 2304, 1375, 171, 1342, 790, 702, 385, 1950, 1098, 1749, 723, 2145]. betweenness [950]. beyond [1581]. Bi [2217, 1761, 2119, 2420, 1709]. Bi-criteria [2217]. bi-ideals [2119]. bi-level [1761]. bi-threshold [2420]. Bialgebras [779]. bias [2009]. biased [144]. bicomplexes [441]. Bicriteria [449, 1999]. Bidder [1580]. bidirectional [134]. Bids [2216]. bifix [1275]. bifix- [1275]. big [1161]. big-step [1161]. bigraphs [1649]. bijective [2136, 640]. bimatrix [17, 1524, 1140]. Bin [25, 1744, 1201, 938, 2390, 1907, 240, 226, 2362]. bin-packing [1907]. Binary [2216, 15, 540, 906, 116, 1332, 1400, 1827, 1011, 796, 2142, 2458, 1310, 1279, 147, 609, 1389, 305, 1622, 162, 1364, 2385]. binary-minded [540]. binomial [761]. binomial-based [761]. bio [597, 534, 1007, 532]. bio-pathway [597, 534]. Bio-PEPA [1007, 532]. Bio-PEPAd [1007]. BioAmbients [1123]. biobjective [1582, 1880]. biochemical [154, 1131, 593, 533, 599]. biocomputing [943]. biological [1130, 1125, 1005, 153, 152, 562, 1119, 529, 1682, 595, 861, 1513, 439]. Biologically [1119, 69]. biology [1337, 1129, 155]. biomedical [851]. bipancyclicity [929]. bipartite [573, 2440, 1667, 1961, 1581, 1468, 1487, 1901, 463, 29, 2168, 1878, 2089, 2100, 2388, 98, 1038, 1706, 569]. Birth [299]. birthday [614, 1845]. bisection [69]. bisimilarity [2156]. bisimulation [957, 339, 1388, 343, 1365]. bisimulation-like [343]. Bisimulations [2161, 1242, 170, 629]. Bispecial [1238]. biswapped [1804]. bit [497, 2258, 1816]. bit-flip [2258]. Bitonic [542]. bits [648]. Black [1515, 1900, 1908, 2260, 102, 2143, 835, 2431]. black-and-white [2143]. Black-box [1515, 2260, 102]. blackjack [1982]. BlenX [40]. Block [563, 421, 839, 676, 2238, 1193, 2101, 1460, 1058, 2238]. Block-graph [563]. block-moves [1193]. blocked [1799]. Blockers [2065]. blockwise [56]. blowing [2361]. blue [560]. Board [105, 157, 159, 177, 404, 503, 512, 526, 2222, 2176, 2185, 2197, 2205, 51, 63, 74, 258, 403, 855, 865, 870, 890, 902, 905, 919, 923, 940, 944, 952, 492, 558, 570, 584, 591, 600, 613, 637, 672, 686, 691, 706, 719, 724, 754, 765, 773, 782, 816, 825, 842, 963, 1213, 1221, 1233, 1241, 1246, 1261, 1271, 1284, 1297, 1301, 1314, 1325, 1350, 1356, 1366, 1383, 1391, 1402, 1414, 1425, 1434, 1447, 1456, 1465, 1472, 972, 983, 993, 1003, 1010, 1019, 1026, 1032, 1041, 1045]. Board [1056, 1065, 1075, 1108, 1117, 1132, 1144, 1149, 1158, 1166, 1175, 1181, 1192, 1200, 1208, 1479, 1738, 1747, 1760, 1777, 1787, 1796, 1809, 1830, 1837, 1844, 1854, 1864, 1873, 1884, 1894, 1905, 1914, 1931, 1943, 1955, 1965, 1491, 1502, 1510, 1516, 1525, 1536, 1543, 1558, 1568, 1578, 1588, 1601, 1607, 1617, 1628, 1645, 1651, 1659, 1668, 1679, 1686, 1695, 1707, 1713, 1723, 1971, 2232, 2244, 2252, 2261, 2282, 2293, 2302, 2310, 2321, 2330, 2339, 2350, 2372, 2381, 2392, 2404, 2416, 2425, 2426, 2427, 1985, 1990, 2002]. Board [2013, 2027, 2038, 2049, 2054, 2066, 2067, 2082, 2093, 2102, 2111, 2120, 2129, 2137, 2147, 2154, 2162, 2171]. boards [1839]. Bob [1480]. body [2383]. body-hinge [2383]. Böhm [1165]. Bologna [1119]. Boltzmann [1815]. bombs [1176]. bond [457]. bondage [2168]. Boolean [107, 1806, 2053, 1274, 420, 789, 222, 1671, 1604, 595, 128, 456, 2419, 1611, 739, 743, 1848, 785, 1379, 1981, 1051, 1862, 24, 426, 1050, 1259, 203]. Boolean-width [789]. bootstraping [1852]. Borda [1786]. border [1392, 931]. Borel [2447]. both [794, 1552]. bottleneck [2036, 211, 1498]. bound [651, 222, 1947, 1907, 552, 1521, 1486, 834, 142, 743, 317, 10, 1412, 2408, 646, 1156]. Boundary [649, 1820]. Bounded [1781, 950, 873, 133, 38, 1134, 249, 874, 2163, 928, 1604, 2430, 1403, 2283, 995, 1304, 280, 2354, 217, 2432, 882, 2298, 2235, 248, 1346, 1051, 1283, 1805, 1886, 1000, 1259, 2362, 822]. Bounded-Cycle [822]. Bounded-degree [1781, 1259]. bounded-quantum-storage [2430]. Bounding [1483, 1021, 1303]. Bounds [606, 2319, 102, 700, 323, 494, 2113, 1201, 809, 2439, 2328, 1866, 2359, 735, 1926, 1657, 1900, 130, 293, 483, 409, 1961, 1635, 1676, 2039, 551, 1838, 577, 412, 2413, 1527, 1599, 791, 210, 2410, 1348, 2233, 1596, 1156, 1717]. box [1515, 2260, 102, 1859]. bracelets [1554]. Branch [1462, 696, 2059]. branch-decomposition [696]. Branching [2268, 957, 2269, 1355, 2190, 2193]. Brane [1126]. breadth [2283, 1599]. break [1521]. breakable [1834]. breaking [1457, 1898]. Breakpoint [1735]. bridge [224]. bridge-connectivity [224]. bridges [1255]. bridging [1103]. Brlek [1464, 1270, 826, 1556]. broadcast [1771, 1621, 1776, 787, 1521]. broadcasting [2016, 651, 115]. Broadcastings [379]. brokers [2099]. Brouwer [622]. Bruijn [1377]. Brushing [2399]. Brute [2071]. Brute-force [2071]. Brzozowski [1294]. BSTs [1622]. bubble [1024]. bubble-sort [1024]. Buchberger [737]. Büchi [2333, 429]. bucket [254]. Bucklin [1786]. budget [950, 2256]. buffer [784, 1693]. buffering [1775]. buffers [1034]. build [1296]. Building [1933, 1902, 1998]. Bulking [679, 680]. burnt [419]. Burrows [2136, 2069, 1377, 612]. busy [277, 2444]. Buy [2414]. Buyer [411]. BWT [1641]. bytecode [961, 169]. Byzantine [1935, 710, 246, 1751]. Byzantine-resilient [246].

CA [2424]. Cache [652]. Cache-oblivious [652]. CafeOBJ [1454]. calculating [1597]. Calculi [1119, 1129, 868, 113]. calculus [620, 1128, 259, 1242, 1124, 621, 338, 1455, 1452, 2451, 171, 95, 148, 1388, 991, 903, 1542, 2278, 445, 1165, 343, 520, 149, 1126, 857]. call [1161, 223]. call-by-need [1161]. Camellia [2101]. cameras [2393]. can [1171, 1049, 66]. Canadian [2127, 1669]. canalizing [1611]. Canonical [586, 41, 948, 167]. canonicity [586]. capacity [2360, 1029, 2345, 1755]. Capture [2040]. car [274, 1292]. Carathéodory [1913]. cardinalities [967]. cardinality [1744, 2096]. cards [1745]. carrier [1903]. carrier-based [1903]. Cartesian [2266, 1569]. cascades [1483]. Case [1599, 2438, 2411, 2095, 1276, 1161, 639, 126, 2123, 788, 1153, 943, 1749, 16, 1489, 125]. cases [266, 611, 1498]. cat [2332]. Catalan [1829]. catch [2358]. Categorial [1289, 1304]. Categorical [850, 32, 2272, 1967]. Categories [1683, 2266, 2051, 2273, 1110, 631]. Category [1963, 2275, 327]. Category-based [1963]. category-theoretic [327]. Caucal [1557, 5]. Causality [2265]. CCA [1701]. CCA-secure [1701]. CCS [106, 1147]. Celebrating [1845]. cell [530, 682, 1103]. cell-inspired [1103]. cells [2268, 2269]. Cellular [1870, 951, 275, 682, 1505, 420, 1480, 307, 1170, 378, 14, 274, 675, 680, 1850, 1975, 2422, 531, 348, 656, 681, 1470, 1292, 1094, 1849, 2421, 350, 678, 2423, 24, 223, 2417, 86]. Cenk [1073]. Centenary [2437]. center [1235, 1058]. central [1882, 11]. centrality [950]. centroid [1023]. Cerný [813]. Certain [1500, 1201, 1800]. certificate [216]. Certificateless [810, 416]. certification [2053, 728]. certifying [803, 1312]. CF [1596]. chain [1112, 1803, 1382]. chain-complete [1112]. Chains [720, 1328, 704, 1902, 1290, 1303, 311, 1766]. change [1689, 78]. changeable [1001]. changeover [2210]. changing [887, 1626]. channel [573, 1771, 2433]. channels [1147, 147, 2436]. chaotic [2329]. character [2201, 1954]. characterisations [1306]. Characteristic [1345, 220, 1204, 605, 479, 1265]. characteristics [1803]. characterizability [1658]. Characterization [2181, 1966, 329, 2281, 1612, 2048, 2158, 456, 2089]. Characterizations [969, 1009, 2119, 1496]. characterize [1685]. Characterizing [187, 1666]. characters [1332]. chasm [1267]. checkable [1922]. Checking [262, 1005, 153, 920, 1711, 169, 174, 1715, 533, 30, 1833, 2453, 2160, 868, 2308, 1422, 1478]. checkpoint [1309]. chemical [1792]. China [2437]. Chinese [1953]. chocolate [2435]. choice [2224, 2255]. Chomsky [1168]. choose [1480]. Chooser [1546, 294]. Choosing [73, 1205]. chord [99]. Chordal [1441, 281, 201, 284, 1240, 1224]. chordless [98]. chords [1560, 2000, 2024]. Chosen [1429, 2203]. Chosen-ciphertext [1429, 2203]. Christoffel [1740, 605, 388]. Chromatic [99, 1033, 1898]. church [1164, 1339, 2304, 49]. church-style [1164]. churn [1935, 1318, 1939]. churn-prone [1935]. ciliate [72]. ciliates [70]. cipher [2101]. ciphertext [1429, 2203]. ciphertexts [1028, 2203]. circadian [530]. circle [23]. circles [1865, 892, 470]. circuit [494, 222, 112]. circuits [1799, 128, 1430, 1267, 2063, 1862, 24, 314, 426]. circulants [742]. Circular [2388, 329, 803, 1238, 1488, 603, 2316]. circular-arc [803]. circulating [714]. clan [2246]. clan-popular [2246]. Class [240, 2170, 1551, 45, 490, 30, 227, 1570, 1148]. classes [1544, 1201, 1921, 1078, 1812, 807, 1676, 988, 2215, 1930, 2458, 509, 2143, 1775, 499, 1346, 1283]. Classical [1667, 1794, 146]. classification [587, 1531, 401, 846, 848, 743, 851]. Classifications [680]. classifier [1055]. classifiers [2113, 1364]. classifying [1624]. clause [1710]. clause-sets [1710]. clauses [128]. claw [932]. claw-free [932]. cleaning [2399, 1445]. clients [2214]. climbing [2289]. clique [1867, 1398, 1657, 812, 195, 943, 643, 2164, 1156]. clique-width [1657, 812, 2164]. cliquewidth [1078]. clock [530, 959, 824]. clonal [399]. closed [2266, 50, 378, 1195, 1756]. closer [2409]. Closest [2150, 2019]. closing [1975]. closure [1933, 752, 841, 608]. closures [2151]. Cloud [2289, 1755]. cluster [1071, 813]. Clustering [97, 2209, 2409, 2228, 1838, 847, 612, 1079]. clusters [1079]. CNF [1912, 739, 1615]. Co [562, 873, 2007, 1014, 1497]. co-comparability [873]. co-cycle [1014]. co-decomposition [1497]. Co-evolution [562]. co-training [2007]. Coalgebraic [780, 2305, 778, 172, 1115, 2278]. coalgebraically [781]. Coalgebras [777, 776]. coalitions [1507]. COCOON’2011 [1415]. code [1813, 2317, 1814]. Codes [136, 676, 1081, 134, 21, 1220, 2023, 1206]. coding [1533, 2018, 228]. codings [891]. coercions [991]. Cograph [1419]. cographs [1678, 1493, 2039, 418, 967]. coherence [521]. coin [2429, 1361]. collapse [1043]. collapses [2451]. collated [2296]. Collatz [2458, 807]. collecting [1553]. collection [2369]. collections [1641]. Collective [2283]. College [245]. Colloquium [1351]. Colony [183, 491]. Color [1908, 54, 760]. color-preserving [54]. colorable [204]. Colored [1633, 385, 1763]. Coloring [1067, 1226, 1655, 2353, 1225, 749, 873, 994, 238, 965, 892, 498, 1518, 2215, 129, 1428, 565, 2407, 1732, 2354, 204, 2143, 1920, 643, 1984, 2397, 2460, 1991, 2044, 2024, 2344]. colorings [1560, 2092, 1663, 1704, 747, 1459, 2000, 306]. colors [2285]. colour [2347]. coloured [243, 16]. Colouring [2042, 1673, 1745, 1816]. Column [1020, 377]. Colussi [2458]. Combination [1404, 873]. Combinatorial [323, 1810, 1676, 2206, 1819, 2382, 2095, 1515, 1307, 472, 1343, 566, 254, 2326, 1567, 1951, 1074, 2175]. combinatoric [1340]. Combinatorics [601, 2351, 110, 725, 2109, 655]. combinatory [770]. Combined [557, 1833, 1370, 1180, 1909, 2317]. Combining [70, 2005, 103]. Comment [2458]. commitment [102]. commitments [815]. commodities [410]. common [1504, 245, 1945, 191, 2230, 2043, 2241]. communicating [893, 356, 392]. Communication [348, 656, 951, 996, 1808, 1662, 1371, 712, 576, 1084, 681, 1942, 566, 1095, 1835, 275, 1665, 270, 2108, 2061]. community [436]. commutative [474]. Compact [2314, 1254, 272, 261, 257, 970]. comparability [873]. Comparative [2346]. Comparing [907, 146, 129, 44]. comparison [2139, 1080, 1426, 2193, 1115, 2133, 1634, 837, 328]. compartments [1124]. compass [2349]. Compatibility [1646, 1482, 2057, 2288, 2457, 1954]. compensation [2436]. competition [2216]. Competitive [1759, 808, 1693, 2122, 784, 1553, 2378, 1992]. competitiveness [1427, 1561]. competitiveness-based [1561]. Complementary [1507, 2028]. complementation [981]. Complete [2276, 2187, 904, 451, 1112, 690, 329, 1359, 1022, 65, 1768, 415, 1061, 1988, 687, 1388, 1647, 352, 2035, 518]. completely [1460]. completeness [1262, 1015, 1036, 80]. Completing [1074]. completion [1731, 449, 912, 657, 330, 1999, 1677, 1877, 168]. completions [1114]. Complex [188, 1354, 1416, 1857, 860, 1790, 1927, 354, 936, 783, 1423, 1259, 434]. complex-weighted [1423, 1259]. complexities [1135, 1515, 30, 1018]. Complexity [2387, 1392, 2134, 2312, 694, 2126, 37, 54, 396, 1399, 1729, 1023, 1669, 951, 353, 1194, 2103, 2056, 390, 1419, 1362, 2206, 1287, 643, 176, 488, 382, 548, 1867, 180, 2117, 494, 659, 1808, 602, 38, 872, 2021, 1505, 827, 949, 1934, 1188, 1436, 2353, 337, 2359, 749, 497, 1906, 2450, 2109, 965, 1657, 1275, 1506, 507, 987, 222, 1967, 561, 828, 1334, 1846, 2448, 1178, 559, 731, 2128, 925, 1662, 2182, 712, 1923, 1180, 34, 523, 989, 1741, 2215, 1850, 2260, 2228, 576, 1524, 1494, 296, 1410, 1137, 1909, 565, 1304, 910, 1085, 1341, 1487, 1791]. complexity [833, 348, 656, 681, 1852, 2418, 2210, 1486, 1139, 1953, 841, 1930, 59, 186, 1300, 1293, 688, 139, 2168, 292, 879, 447, 95, 336, 1418, 1704, 1280, 1013, 566, 1153, 1195, 638, 2143, 1182, 1281, 633, 1665, 968, 2159, 1786, 1994, 1282, 1572, 544, 891, 248, 1839, 685, 1420, 1816, 98, 740, 1017, 914, 2091, 1841, 1105, 722, 2001, 2408, 62, 732, 250, 646, 838, 1283, 1497, 1762, 1807, 1887, 2220, 1381, 1423, 1690, 203, 550, 2394, 78, 1806, 744, 2287]. Component [597, 1646, 2257]. Component-based [597]. components [1385, 880, 299]. componentwise [308]. composed [310]. composition [1602, 2401, 914, 256]. Compositional [663, 2322]. Compositions [21]. compound [1780, 854]. comprehensive [26, 1611, 1764]. Compressed [1632, 1394, 2314, 607, 1555, 1394]. compressing [1952, 1640]. Compression [2400, 1027, 2073, 2069, 84, 2072]. Compressive [467]. Comput [1040, 1557, 1678, 2457, 671, 2461, 2459, 2292, 951, 2456, 2460, 1073, 864, 1706, 1712]. Computability [1409, 1375, 1797, 442, 1016]. computable [495, 1188, 1534, 1409, 1501]. Computation [2437, 2245, 2253, 1047, 1855, 1097, 493, 1113, 1696, 2440, 589, 483, 2258, 1644, 1049, 351, 2046, 578, 626, 2230, 960, 2329, 2265, 75, 1155, 1269, 1966, 1767]. Computational [296, 2418, 362, 524, 561, 1940, 559, 34, 69, 1572, 1853, 914, 836, 40, 837, 3, 864, 2220, 1174, 749]. computationally [1359]. computations [112, 2081, 705, 1121, 853]. compute [540, 1371]. Computer [1836, 1337, 189, 263]. computer-verified [263]. Computing [180, 2170, 457, 2204, 1351, 355, 2177, 762, 1249, 1119, 529, 1417, 2351, 1307, 695, 481, 1374, 1719, 463, 309, 281, 1598, 1051, 383, 1133, 728, 1962, 1945, 2060, 369, 2135, 841, 1896, 934, 1792, 2159, 844, 1224, 1349, 1862, 426, 1767]. concatenated [1618]. concatenation [1105]. concave [1236]. concavity [1140]. concentric [1865]. concept [887, 1626]. concerned [449]. concerning [2402, 1798, 1287]. concretely [1163]. concurrency [619, 2279]. Concurrent [301, 617, 461, 344, 698, 1756, 2297]. condition [26, 1232, 544, 340]. Conditional [875, 1910, 2133, 2299, 1807, 259, 2401, 1520, 1429, 2300, 2203, 1453, 1313, 901]. conditionally [1424]. conditions [1550, 999, 1969, 1207, 2398]. Confident [2406]. confidentiality [690]. confidentiality-preserving [690]. configuration [883]. Confining [1498]. conflict [380, 1044, 2041]. conflict-free [380]. confluence [259, 1044, 2193, 1450, 68]. confluent [49]. congestion [1958]. congruence [957, 629]. congruent [1881]. congruential [1339]. conjecture [826, 1464, 1556, 1270, 1217, 604, 1143, 422, 611, 813, 2279, 1627, 465]. conjugate [2046]. conjunction [1698]. conjunctions [128]. Conjunctive [198, 443, 2360]. Connected [200, 831, 1276, 1251, 1254, 2020, 217, 1090, 2456, 1072, 1772, 567, 1916, 916, 889, 1460, 2398, 2146, 1058, 6, 335, 2169, 253, 366, 1191]. Connectedness [654]. Connecting [2361]. Connection [2360, 731, 1610, 640, 1375]. connections [373, 1344, 757]. Connectivity [1718, 1316, 224, 1729, 1692, 1520, 2451, 926, 739, 1749, 193, 885, 1437, 2377]. connectivity-based [926]. coNP [433, 1061]. coNP-complete [1061]. conquer [2296]. cons [166]. conscious [2278]. consecutive [2080, 2144]. Consensus [710, 749, 794, 116, 823, 2228, 1488, 235, 1079]. consequences [2283]. conservation [1732]. Conservative [849]. conserved [2241]. conserving [656]. Consistency [1343, 36, 1232, 30]. consistent [690, 2406, 2011, 535]. Constant [696, 1992, 2214, 1028, 1492, 1274, 1630, 1694, 1692, 2203, 1814]. Constant-competitive [1992]. Constant-factor [696]. constant-size [1028, 1492, 2203]. constant-space [1630]. Constants [67]. Constrained [1257, 1367, 1133, 1151, 371, 1631, 240, 2096, 219, 2037, 437, 694, 2325]. Constraint [720, 1522, 708, 659, 2050, 1203, 224, 1708, 1251, 1494, 2084, 703, 792, 1423, 141]. constraint-based [708]. constraints [1744, 530, 376, 1522, 1029, 1236, 449, 1195, 201, 945, 791, 664, 2325, 1156, 1382]. construct [1463]. Constructing [1135, 538, 1641, 2210]. Construction [439, 1492, 1757, 1875, 1320, 2383, 597, 1835, 2104, 2091, 2025, 416, 2223]. Constructions [2064, 974, 101]. Constructive [440]. constructors [1539]. contained [1608]. containing [718, 2024]. containment [935, 70, 713, 1091, 1190]. contains [394]. content [1554]. contention [700]. Context [1681, 2264, 1958, 429, 977, 1550, 1496, 968, 1824, 1983, 1997, 113, 888, 510]. context-free [429, 977, 1824, 1997, 888, 510]. contextual [1273, 214, 1277, 343, 2435]. contiguity [2039]. contiguous [400]. continuation [481]. continued [544]. Continuity [635]. Continuous [1444, 235, 596, 871, 283, 511, 532, 2277, 209, 1476, 1986, 635]. continuous-armed [511]. continuous-time [1476]. Contract [260]. Contract-based [260]. contracting [1562]. contraction [130, 4, 1712]. contractions [1614]. contracts [1307]. contributor [1053]. control [373, 2346, 1722, 590, 921, 530, 1163, 1786, 392, 1365, 223, 869]. Controlled [276, 1496, 738, 1574, 1268, 838, 1283]. controllers [1722]. Convergence [1358, 1183, 761, 1684, 246, 807, 2458, 1072, 1229]. convergent [1483]. Conversation [345]. converse [551]. conversion [662]. convertible [1564]. convex [123, 243, 1551, 2149, 2212, 139, 631, 2388, 610, 1892, 2146, 729]. convexities [1913]. Convexity [1522, 1478]. convolution [1379]. cool [629]. cooperating [287]. cooperation [1507]. cooperations [1250]. Cooperative [221, 140, 1372]. Coordinated [1029]. coordinates [1734, 1319]. coordination [1702]. Cop [386]. Copeland [1786]. copies [462]. Cops [302, 1446, 1572, 1616]. Coq [441]. core [411, 75, 1767]. cores [2342]. corner [1236]. Corona [1941]. Corrected [809]. correction [326, 2009, 2025]. Correctness [1756, 518, 627, 519, 585, 1298]. correlation [1671]. correlation-immune [1671]. correspondence [171]. Corrigendum [1464, 1040, 1678, 2457, 2461, 2459, 2292, 1836, 2456, 2460]. Cost [961, 2187, 373, 448, 1274, 2201, 47, 1619, 1692, 1742, 1911, 1251, 947, 1266, 2210, 1236, 1616, 1776, 1408]. cost-effective [373]. cost-sensitive [2187]. costs [727]. countable [547, 57]. counter [2050, 1335, 174, 2224]. counterexamples [2306]. counterfeit [1361]. counters [2012, 1198]. Counting [985, 988, 2336, 2086, 1811, 1385, 301, 234, 561, 1178, 1711, 2334, 2461, 232, 2337, 201, 2249, 1541, 1198, 176, 1423, 1259, 389]. counts [188]. coupled [530]. couplings [1821]. cover [77, 1934, 2374, 1924, 293, 1899, 1305, 577, 565, 1976, 934, 446, 18, 1872, 143, 889, 937, 718, 321, 1404, 1054, 6, 2169, 1191]. coverability [1474]. coverage [744, 2247, 1994]. Covering [29, 709, 1865, 938, 140, 2390, 845, 1871, 2096, 945, 897, 1881, 2132, 2127, 25]. covers [200, 1035, 1944, 742, 2131, 193, 730, 1703]. CPS [31]. CPS-translation [31]. crash [717]. crash-recovery [717]. Criteria [137, 2217]. criterion [737, 669]. critical [1044, 685, 2301, 1345]. cross [406, 1360]. cross-talk [1360]. crossing [2357]. crossings [786]. Crossover [1049, 2258, 1512, 1098]. crossovers [1048]. cryptanalytic [1803]. cryptographic [2434, 1688, 1802, 104, 101]. cryptography [2429, 1908, 2432, 1698, 2431, 2435, 810]. cryptosystem [1064]. cryptosystems [1584]. CSP [249, 106, 1832, 1925]. CSPs [1259]. CTL [598, 2308]. CTL* [621]. CTMDPs [1476]. CTRL [598]. cube [1517, 1219, 554, 1407, 2061]. cubes [1910, 1699, 753, 295, 538, 730, 901, 1463, 1700, 1703]. cubic [1570]. Curry [2304]. curve [2284, 1595, 1698, 2386]. curves [1696, 966, 675, 1594, 761, 1727]. customers [83]. cut [809, 880, 1190, 119, 1462, 1196]. Cut-and-Join [1196]. Cutcake [829]. cuts [985, 1648, 316, 2303, 922]. cutting [698, 1892]. Cycle [1678, 1493, 1150, 1926, 2290, 530, 803, 1014, 99, 386, 813, 2420, 644, 918, 1061, 822]. cycles [1358, 2353, 1815, 735, 1560, 296, 736, 1012, 26, 753, 2424, 1459, 1984, 1380, 1993, 2000, 2024, 1700, 389]. cyclic [1602, 1287, 678, 1268, 805]. cyclical [2375]. cylinder [2231, 1767]. cylindrical [1592].

D [1437, 1889, 471]. D0L [833, 1143, 1238]. dangerous [1519]. Dantzig [467]. Data [2409, 572, 2214, 619, 1904, 701, 53, 50, 2369, 1137, 1252, 557, 1771, 1071, 788, 847, 1689, 956, 2012, 19, 1773, 2122, 1623, 75, 722, 1408, 797, 1805, 510, 78, 1455]. data-analysis [1623]. databases [690]. dates [1432, 732]. Davis [1710]. DBDH [2240]. Dcpo [168, 586]. Dcpo-completion [168]. Deadline [280, 2368]. Deadlines [2220]. deal [945]. deallocation [343]. decay [237, 2459]. Decidability [2318, 1015, 722, 1975, 2192, 324]. decidable [2452, 917]. Deciding [1547, 1708, 499, 546, 1951, 337, 1524, 1362]. decision [67, 1721, 103, 2364, 1968, 2191, 96, 210]. decoding [1426, 134, 2023]. decompilation [1160]. decomposable [276, 76]. Decomposing [1489, 239, 1514]. decomposition [123, 1331, 1912, 858, 2048, 696, 1592, 2086, 1497]. decompositions [220, 746, 1220, 2184]. decompression [2069]. decompressor [394]. decreasing [986, 1907]. dedicated [614]. defect [396]. definability [2181, 827]. Definable [46]. defined [311, 1346, 1120, 4, 1712]. definitions [58, 769, 1477]. deformations [871]. degeneracy [1524, 248]. degenerate [2396]. Degree [1468, 1566, 1000, 2170, 1781, 2312, 1560, 695, 1614, 26, 2354, 1583, 2298, 669, 437, 1991, 2044, 1498, 1886, 1423, 1259]. degree- [1498]. Degree-bounded [1000, 2354]. degree-two [1423]. degrees [496, 1863]. Dejean [896, 611]. Delaunay [1378, 653]. Delete [1040, 395]. deletion [357, 421, 1930, 2089, 1363, 783]. Delimited [1163]. delivery [2348, 873, 1702, 801, 1029, 375, 1187]. demand [39, 1062, 231, 954]. deniable [2443]. Dense [2224, 200, 1736, 971, 1478, 389]. Dense-choice [2224]. dense-time [1478]. Density [1613, 1814]. deontic [1355]. dependence [2341, 921]. dependencies [1693, 1969, 1121, 545]. dependent [2272, 1527, 1154, 1187]. depending [1898]. deployment [427]. depth [1442, 1604, 1267, 542, 1576, 314]. derandomization [1153, 500]. Derek [381]. Derivation [626, 1688, 2182]. derivational [1182]. derivative [761]. deriving [1161]. descents [192]. description [2172, 1861, 914, 975]. description-based [914]. Descriptional [1282, 1602, 2109]. descriptive [255]. Design [530, 2045, 1411, 1625, 666, 1752, 1831, 1263, 2036, 368, 727, 18]. designing [1737, 473]. destructible [488]. detailed [1198]. detect [242]. Detecting [2074, 153, 1620, 764, 1206]. detection [1392, 1070, 1044, 2078]. detector [711, 1942]. deteriorating [376, 285, 2237, 1239, 1432, 1677]. deterioration [1769, 279]. deteriorative [1358]. Determinacy [443, 1720]. determinants [1333, 2331]. determination [748]. Determining [1033, 2242, 731]. determinism [1243]. Deterministic [373, 996, 726, 1868, 1663, 179, 1737, 223, 2172, 1274, 651, 987, 564, 1068, 928, 1759, 268, 271, 265, 133, 439, 1941, 210, 1348, 1670]. Determinization [1453, 1606]. device [1018]. devices [1321]. DFA [1265]. diagnosability [2401, 2300, 2133, 2299, 2125, 1313]. diagnosis [326, 1716, 2133]. diagnostic [1804]. diagram [2141, 12, 1763]. diagrammatic [1450]. diagrams [1750, 30, 229]. Diameter [2227, 986, 1673, 1484, 1962, 294, 239]. diameters [660]. dichotomic [1823]. dichotomy [1259, 548]. dictionary [2073, 1555]. dictionary-based [2073]. diffeomorphisms [23]. difference [2331, 2183]. different [1290, 2418, 400]. differentiable [1218]. differential [520, 2175]. digit [2284, 379]. Digital [469, 1865, 1269]. digitization [1467]. digits [751]. Digraph [746, 1215, 1545, 1571]. digraphs [1186, 1139, 2354, 1441, 1157, 389]. Dilworth [2457, 2057, 2288]. dimension [762, 1963, 360, 677, 188, 654]. Dimensional [1061, 1856, 1840, 772, 246, 226, 110, 1470, 1797, 1378, 2174, 554, 678, 1489, 379, 469, 2031, 2362]. dimensionality [658]. dimensions [911, 1627]. dimer [2298]. Diophantine [1276, 976]. dipath [1847]. direct [167, 585]. Directed [1847, 1367, 1330, 800, 1436, 1652, 2448, 1484, 1945, 273, 834, 702, 385, 835, 193, 437, 885, 1461]. Directional [675, 1773]. directions [2422, 353, 332]. disc [1865, 1780]. disc-ring [1780]. disconnectivity [1157]. discounted [2413]. discounting [2011]. Discovery [121, 260]. Discrete [1846, 683, 756, 1820, 2450, 755, 2448, 2332, 490, 465, 1107, 471, 635, 2386, 1365, 1500]. discretisation [532]. discretized [761]. Disjoint [742, 2387, 1186, 2085, 735, 2391, 103, 1869, 1139, 753, 1944, 2131, 999, 730, 1460, 1461, 1703, 253, 1951]. Disjunction [688, 1473, 198]. disk [498, 1072, 1000, 366, 2213]. disks [2247]. disoriented [1868]. dispersal [216]. dissemination [2015]. dissociation [934]. Distance [1269, 383, 1504, 794, 1225, 302, 1439, 1399, 1021, 300, 2378, 2123, 144, 702, 385, 703, 1098, 1672, 2078, 469, 2145, 2386, 257, 228]. distance-2 [1225]. distance-aware [2378]. distance-restricted [144]. Distances [759, 955, 2454]. distant [2370]. Distinct [334, 1027, 758]. distinguishability [1529]. distinguishing [1808]. distortion [647, 463]. Distributed [689, 707, 2192, 1224, 2344, 587, 1935, 413, 82, 189, 287, 1896, 41, 1378, 1298, 1816, 705, 836, 1107, 1199]. distribution [2434, 2429, 2332, 631, 1527, 1566, 2433, 1591, 664, 2436]. distribution-dependent [1527]. Distributional [132]. distributions [2323, 1850, 1468, 960]. distributive [1806]. distributivity [1384]. divergence [1353, 360]. Diversity [1929]. divide [2296]. divide-and-conquer [2296]. divisionless [68]. DNA [1328, 1080, 1790, 530, 368, 1737]. document [1633, 2229, 2449]. Domain [146, 2009, 394, 936, 2277, 868]. domain-free [868]. domains [2266]. Dominating [932, 2445, 1398, 1676, 1251, 834, 1072, 567, 324, 1916, 971, 2371, 374, 1928, 1886, 366, 2213]. domination [1399, 1262, 1565, 550, 2394, 1418]. domino [322]. double [1484, 2198, 1819, 2105, 1670, 1196]. double-loop [1484]. doubled [1980]. Doubly [2240, 2087, 219]. doubly-logarithmic [2087]. doughnuts [1811]. Down [2447]. dP [1120]. DPLL [1022]. DQBF [2053]. drawbacks [402]. Drawing [786, 390, 2363, 12, 2146]. drawings [1236]. Drawn [734]. drift [1172]. drifting [209]. driven [1113, 705, 2368]. drunk [1446]. drunkenness [1616]. DSA [1018]. dual [1517, 1340, 175, 422, 937, 2019, 1381]. dual-cube-like [1517]. duality [249, 605, 887]. Duchêne [422]. duo [2121]. duo-preservation [2121]. duplication [192, 1566, 2126]. DVcube [1780]. Dyck [1825]. Dykstra [219]. Dynamic [2087, 639, 479, 190, 1939, 883, 2091, 1444, 572, 1553, 1843, 1922, 1483, 82, 1972, 859, 1252, 580, 1718, 2250, 1063, 2367, 1240, 1244, 1197, 270, 2420]. dynamical [683, 1203, 2320, 677, 1500]. Dynamically [535]. dynamics [683, 156, 1846, 675, 1, 534, 1849, 164, 582].

E-complete [2035]. each [1336]. Earley [1268]. EAs [2326]. Easy [2157]. Economic [1755]. Edge [929, 1984, 2000, 238, 1416, 1730, 1910, 1692, 1742, 28, 1974, 129, 1470, 1139, 2020, 753, 747, 201, 2357, 1552, 1772, 193, 783, 971, 1733, 901, 2398, 1928, 1886, 459, 2394, 2445]. Edge-bipancyclicity [929]. edge-connected [2020]. edge-connectivity [193]. edge-disjoint [1139, 753]. edge-fault [1910]. edges [640, 1562, 29, 2363, 897, 1038, 1706, 1700]. edit [383, 1504, 300, 1672, 2078]. editing [1305, 248, 1419]. Editorial [105, 157, 159, 177, 404, 503, 512, 526, 2222, 2176, 2185, 2197, 2205, 2340, 953, 1315, 1047, 51, 63, 74, 258, 403, 855, 865, 870, 890, 902, 905, 919, 923, 940, 944, 952, 492, 558, 570, 584, 591, 600, 613, 637, 672, 686, 691, 706, 719, 724, 754, 765, 773, 782, 816, 825, 842, 963, 1213, 1221, 1233, 1241, 1246, 1261, 1271, 1284, 1297, 1301, 1314, 1325, 1350, 1356, 1366, 1383, 1391, 1402, 1414, 1425, 1434, 1447, 1456, 1465, 1472, 972, 983, 993, 1003, 1010, 1019]. Editorial [1026, 1032, 1041, 1045, 1056, 1065, 1075, 1108, 1117, 1132, 1144, 1149, 1158, 1166, 1175, 1181, 1192, 1200, 1208, 1479, 1738, 1747, 1760, 1777, 1787, 1796, 1809, 1830, 1837, 1844, 1854, 1864, 1873, 1884, 1894, 1905, 1914, 1931, 1943, 1955, 1965, 1491, 1502, 1510, 1516, 1525, 1536, 1543, 1558, 1568, 1578, 1588, 1601, 1607, 1617, 1628, 1645, 1651, 1659, 1668, 1679, 1686, 1695, 1707, 1713, 1723, 1971, 2232, 2244, 2252, 2261, 2282, 2293, 2302, 2310, 2321, 2330, 2339, 2350, 2372, 2381, 2392, 2404, 2416, 2425, 2426]. Editorial [2427, 1985, 1990, 2002, 2013, 2027, 2038, 2049, 2054, 2066, 2067, 2082, 2093, 2102, 2111, 2120, 2129, 2137, 2147, 2154, 2162, 2171]. Editors’ [1526, 2003, 2405, 2373, 206, 114]. Effective [441, 911, 373, 1112, 1396, 1512, 1592, 2104]. Effectively [1069]. effectiveness [1753]. efficiency [2202, 102, 10, 729]. Efficient [1813, 794, 1273, 1903, 1637, 1136, 483, 1644, 1711, 830, 1263, 472, 217, 2249, 2342, 460, 79, 2081, 2308, 1478, 127, 1212, 510, 2213, 395, 1040, 1225, 1902, 234, 82, 1228, 401, 640, 2157, 2191, 1564, 803, 2337, 1321, 884, 1623, 1224, 2311, 1422, 1424, 2223, 1044]. efficiently [762]. election [820, 1868, 1786, 1231]. Electric [698]. element [18]. Elementary [858, 1772, 1849, 2257]. Elements [2274, 616, 929, 742, 1411, 305, 553]. elephants [837]. eligibility [791]. elimination [809, 1592, 1173, 922]. elliptic [2284]. embed [554]. embedded [2088, 1991, 1649]. Embedding [555, 753, 295, 2335, 901, 1840, 647, 1236, 1552, 341, 2231, 918, 1888]. embeddings [54, 1759, 463]. emptiness [2090, 42]. Encapsulating [1355]. encoded [1133, 1329, 1136, 334]. encoding [1671, 1946, 851]. encodings [1048]. encryption [1028, 2240, 27, 1429, 2443, 2203, 1701, 2226, 416]. end [874]. endgame [278]. endpoint [77]. Energy [1902, 1386, 1321, 2311, 1862, 48, 314, 1508, 693, 1776, 787, 426, 453]. Energy-efficient [1902, 2311]. enforcing [1086]. Enhanced [1871, 660]. enough [349]. enriched [2051]. entangled [1666]. Entanglement [1436, 1662]. Entities [2340]. entity [2307]. entrant [912]. Entropy [311, 1469, 1017, 233]. enumerable [539]. enumerating [736, 931, 1367]. Enumeration [1481, 1818, 1792, 284, 496, 1676, 1485, 195, 127, 1212]. Enumerative [1671]. environment [1646]. episturmian [288, 793, 11]. Equal [123, 1891, 1547, 1697, 624, 2235, 305, 1462, 1382]. Equal-area [123]. equal-length [1462, 1382]. equalities [1145]. equality [2325]. Equation [107]. Equational [661, 946]. Equations [603, 327, 1276, 194, 976, 981, 2291]. equidistributed [763]. equilateral [2379].

Equilibria [417, 798, 543, 17, 715, 1621, 1792]. Equilibrium [1559, 85, 1612, 2384, 1625, 2037]. equilogical [2271]. Equitable [1459, 306]. equivalence [958, 2117, 67, 1708, 1145, 532, 1091, 137, 1365]. equivalence-based [1708]. Equivalences [861, 2270, 2453, 2322]. equivalent [860, 770]. erasing [645]. Eric [1845]. Erratum [1557, 671, 951, 233, 864]. Error [91, 1639, 1599, 926, 897, 740, 1206, 1229]. error-detecting [1206]. error-locating [897]. Error-repair [91]. error-tolerant [1639]. errors [483, 882]. escaping [1451]. especially [1142]. essentially [2071, 205]. establish [1313]. Estimating [743, 1595]. estimation [595, 761, 597, 763]. Euclidean [89, 470, 1766]. Euler [1334, 12]. Eulerian [1195, 1287, 1017]. Eulerianity [885]. evaluation [39, 1161, 2157, 588]. evasion [2040, 303]. Evasive [2291]. even [2032, 1656, 1206]. event [2268, 531, 325, 1101, 2263, 149, 1365]. event-based [1101]. event-recording [325]. Events [2264, 1409]. every [394, 2158]. eviction [1959]. evident [104]. evolution [562, 1084, 1470, 1575]. Evolutionary [85, 859, 2253, 847, 1047, 1362, 1171, 1049, 485, 2326, 2255]. Evolvability [1433]. Evolving [733, 1468]. Exact [383, 798, 1696, 467, 230, 2258, 1022, 290, 2225, 1206, 927, 2195, 876, 84, 647, 2, 2292, 2191, 2084, 171, 1860, 1925, 739, 1591, 2445]. example [587, 1934]. excessive [1800]. exchange [124, 2370, 2443, 670, 75, 2099, 2041]. excursion [1337]. execution [2324, 515]. executions [1756]. exemplar [2315]. Exhaustive [1817, 2175, 1822]. Existence [255, 1172]. Existential [2304]. existentially [222]. expanding [744]. expansion [1977, 544]. expansions [2296, 877]. expansive [170]. expansiveness [677]. expansivity [1975]. expectation [2258]. Expectations [2390]. expected [2087, 1746]. experimental [354]. experiments [262]. expert [208]. experts [209]. Explicit [1827]. Exploit [2202]. exploration [511, 508, 1779, 118, 1495, 1093, 1440]. Exploring [1482, 1519]. explosion [1986]. exponent [906, 2301]. exponential [1725, 1919, 1593, 2441, 1918, 142, 658]. exposure [530]. expressions [1538, 267, 671, 1711, 840, 232, 598]. Expressive [785, 1537, 1169, 981]. Expressiveness [2192, 501, 109, 2450]. Extended [55, 1252, 997, 128, 1268]. Extending [2076, 543, 1114, 1969, 2297]. extension [2387, 2221, 1007, 605, 1917, 598]. extensional [1043]. Extensions [2295, 1369, 2452, 1022, 1474, 586, 627, 687, 49]. external [214, 1277]. externalities [1559, 2365, 2202]. Extraconnectivity [1219]. Extracting [2030, 1413]. extraction [332]. extractor [1397]. extractors [1006, 101]. Extremal [1207, 266, 1345]. extremely [1321].

F [1160]. faces [506]. facets [167].

Facilitating [1752]. facility

[140, 2234, 1774, 1567, 1523]. factor [1177, 1275, 655, 514, 696]. factor-free [1275]. factorially [814]. Factorization [1548, 57, 1345, 793, 482]. factorizations [1740]. Factors [220, 313, 1204, 655, 1340, 1238]. Failure [1353, 1112, 2400, 711, 1942]. Failure-divergence [1353]. failures [924, 823, 716, 235]. fair [962, 821, 2099]. fairness [485, 598]. families [994, 1304, 969, 2243, 1950, 758, 193, 549]. family [2166, 968]. fan [1862]. fan-in [1862]. Farey [434]. farthest [1763]. fashion [2221]. Fast [1133, 1639, 1922, 298, 2140, 1770, 1656, 315, 115, 1860, 1980, 669, 459, 1883, 1053, 2369, 94, 1736, 1872].

Fast-mixed [1883]. Faster [935, 1555, 2164, 992]. fat [2393]. Father [725]. Fault [713, 1552, 367, 1804, 1407, 1024, 2153, 1700, 708, 1716, 2234, 1750, 1484, 1910, 1520, 1974, 1752, 640, 979, 889, 718, 6]. fault-containing [718]. Fault-free [1700]. fault-tolerance [708]. Fault-tolerant [1552, 367, 1484, 640, 979, 889, 6]. faults [1517, 122, 1552, 901]. faulty [369, 296, 640, 929, 1944, 742, 553, 1700]. feasibility [2128, 75]. feasible [1249]. Featherweight [1987]. Feedback [1878, 2388, 1421, 2396, 2165, 1733, 569]. Feedforward [854]. Ferruh [1073]. few [1562, 2363]. few-bend [2363]. Fewest [906]. Fibonacci [2292, 1179, 2, 2106, 1622, 1142]. field [1444, 486, 475]. fields [423, 1489]. FIFO [1775]. Fighting [1151]. fill [2135, 289]. fill-in [2135, 289]. filling [795, 1039]. filtration [1609, 1245]. final [2172]. Find [395, 1040]. Finding [1215, 1631, 191, 736, 1255, 1488, 1038, 1706, 2312, 800, 1067, 130, 2391, 2352, 878, 1430, 98]. fine [988]. Finite [828, 2095, 415, 1390, 2187, 541, 2113, 1531, 1393, 1577, 2400, 1721, 1295, 1336, 1603, 338, 1288, 1654, 268, 1234, 2090, 833, 41, 353, 133, 2089, 1281, 1009, 439, 1489, 455, 1794, 1670, 1155]. finite-memory [2090]. finite-state [2400]. finitely [511]. finitely-armed [511]. firefighter [438]. fires [1151]. firing [2286, 2424]. First [904, 621, 878, 2250, 1907, 129]. First-Fit [129]. First-order [904, 621]. Fisher [582]. Fit [1907, 199, 129]. Fitting [2323]. Fixed [2379, 27, 1927, 2148, 269, 77, 1112, 1042, 1079, 965, 130, 2369, 932, 626, 769, 695, 181, 1236, 608, 2256, 1554, 750, 1814, 622]. fixed-density [1814]. Fixed-orientation [2379]. Fixed-parameter [1927]. Fixed-point [2148, 1112, 626]. fixed-points [769]. Fixing [630]. fixpoint [261]. Flajolet [725]. flip [2258]. flips [430]. flocking [166, 717]. Flood [1839]. floorplans [2142]. flow [573, 387, 1508, 821, 860, 912, 698, 2194, 150, 982, 2311, 2062, 2196, 1477, 1650]. flows [2085, 218, 2250]. Fluid [960, 173, 2195]. folded [1552, 2023]. folding [1762]. Following [2386]. foraging [165]. forbidden [2042, 311, 2089, 897, 1058]. Forbidding [1086, 287]. force [2071]. Forest [1634, 1545, 965, 2366, 1700, 694]. forests [57]. Foreword [2267, 1435, 1629, 2405, 2373, 206, 1526]. form [1289, 1698, 71]. formable [187]. Formal [1714, 341, 1354, 593, 1542, 954, 767, 2368]. formalisms [920]. Formalization [1298, 588]. formalized [441]. format [956]. formation [798, 85, 757]. formations [950, 1082]. formats [616, 1384, 629]. forms [2148, 479, 948]. formula [360, 1156, 975]. formulae [2053]. formulas [222, 1827, 1615, 205]. Formulations [1880]. forwarding [1754]. Foundational [1125, 1467]. Foundations [593, 1047, 1052]. founded [769]. Four [1658, 349, 620, 1797, 1267, 350]. four-valued [620]. Fourier [1433]. FPLs [1821]. FPRAS [389]. FPT [1657]. FPTAS [2062]. FPTASs [1499]. fractal [966, 2106]. fractals [2107, 1570]. fraction [544]. fragment [904, 2452]. frames [338, 796, 2432]. framework [1354, 326, 1933, 1750, 1790, 1682, 594, 173, 771, 768, 1173, 468, 1542, 2173, 852, 2307]. Fréchet [2386]. Free [490, 380, 394, 146, 809, 1033, 1275, 361, 1618, 932, 1972, 1426, 1339, 2107, 1909, 429, 977, 2304, 933, 1620, 1088, 1293, 1978, 1036, 968, 2125, 1824, 1406, 868, 1312, 1997, 634, 62, 250, 2116, 992, 922, 2343, 887, 888, 416, 1700, 510, 236, 1338, 1839]. free-cut [809]. Free-Flood-It [1839]. freeness [90, 2402]. frequencies [911, 228]. frequency [858, 14, 1961]. frequent [237, 2459, 217]. Front [228]. Frontiers [2138]. FRR [392]. FTSS [639]. full [1135, 1388]. Fully [1240, 1197, 389, 1877]. function [494, 1546, 806, 1662, 807, 1486, 2212, 1050, 2458]. functional [2440, 777, 263]. Functions [1471, 15, 2355, 1416, 1730, 1671, 477, 1172, 1299, 839, 2406, 1988, 1684, 1013, 1153, 1611, 1950, 1571, 743, 1189, 1534, 469, 1269, 1051, 1862, 763, 426, 489, 1690, 203]. functors [2273]. Further [1739, 1610]. fusion [2025]. Fuzzy [846, 1469, 848, 851, 420, 847, 844, 2119, 1365]. Fuzzy-rough [848].

Gaddum [1514]. Gaddum-type [1514]. gain [2345]. gains [387]. Game [189, 1311, 950, 1716, 2221, 2095, 381, 1774, 829, 294, 2455, 1524, 1308, 910, 1445, 2048, 144, 456, 2237, 1572, 1951, 386, 2035, 88, 699]. Game-theoretic [1311, 1716, 2048, 456]. Games [1625, 1728, 1531, 182, 1655, 1546, 458, 85, 47, 1958, 17, 417, 1667, 1720, 2448, 140, 1386, 734, 1183, 1729, 251, 1401, 745, 411, 247, 218, 1968, 894, 1523, 2036, 1612, 1561, 1715, 2034, 1718, 1025, 1621, 1140, 422, 1030, 18, 2037, 209, 2447, 1476, 303, 917, 164, 1717, 1714]. GandALF [1714, 2446]. Gao [482]. gap [2050, 73, 811]. gap-order [2050]. gapped [2294]. gaps [2076, 1216, 630, 1634]. gateway [1257]. Gathering [1901, 2393, 1586, 1368, 1904, 413, 1070, 252]. gatherings [407]. gcd [728, 1597]. gene [180, 1330, 70, 55, 2315]. General [2011, 380, 1053, 1580, 1427, 552, 1009, 12, 1313, 763, 1804, 1677]. generalization [1911, 2106, 211]. generalizations [2283, 1704]. Generalized [1826, 676, 2072, 1861, 2377, 2332, 124, 1204, 1745, 356, 908, 1869, 606, 2286, 884, 1224, 1981, 1460, 1705, 282, 335, 1338, 331, 480]. generalizes [1305]. generate [229]. generated [329, 1566, 1591, 1074, 634, 1471, 1283, 1626]. Generating [1554, 729, 2385]. Generation [1812, 1810, 1813, 1820, 690, 276, 1817, 271, 1094, 1824, 1822, 1574, 1823, 2175, 887]. generators [2355, 425, 101]. Generic [974, 1364, 113, 2184, 416]. Genetic [1359, 2253, 2254, 181, 2259, 1098, 598, 1050]. genomes [2315, 1705]. genotype [1173]. geocasting [448]. geoleader [1868]. geometric [2323, 2247, 473, 1141, 187, 1048]. Geometrical [1276, 351]. Geometry [514, 755, 521, 522, 836]. gets [1267]. Ginsburg [1798]. given [425, 1598, 936]. global [458, 967]. Glynn [2267]. GMPLS [1223]. Gödel [111, 904]. Goles’ [1853, 1845]. gominoes [1817]. Gomory [1014]. Good [699, 1438, 2216]. GoS [2033]. gossip [2438]. gossiping [117, 1521]. GPGCD [1597]. grade [1691]. gradient [2289]. gram [2314, 2123]. Grammar [867, 287, 392]. grammars [1602, 1289, 1277, 241, 977, 1304, 1184, 456, 863, 1824, 738, 1574, 1983, 1268, 198, 785, 1379, 1981, 838, 1283]. granular [846, 1469]. granularity [804]. granulation [1681]. granules [846, 853]. Graph [1921, 2130, 1729, 583, 1435, 863, 1867, 180, 1113, 1544, 1545, 986, 22, 54, 1080, 1067, 1439, 508, 563, 731, 1068, 2334, 2461, 1676, 1305, 2215, 320, 1519, 1405, 1913, 1253, 94, 1663, 1487, 1044, 1614, 1486, 1930, 1780, 880, 2251, 1184, 649, 1096, 2100, 468, 658, 284, 1440, 76, 2287, 799, 319, 916, 488, 1106, 12, 646, 500, 1038, 1706, 438, 1024, 2219, 1000, 1783]. graph-based [180, 468]. graph-theoretic [1080]. graphical [683, 458, 417, 47]. Graphs [1648, 812, 1367, 1957, 77, 1330, 2361, 1186, 2021, 1226, 2026, 772, 800, 1847, 1436, 985, 1227, 2088, 1652, 5, 1557, 457, 1151, 873, 994, 2396, 1678, 1493, 55, 1333, 1812, 1033, 965, 507, 892, 789, 1438, 1416, 1730, 1482, 2057, 2288, 2457, 1673, 2448, 200, 662, 1399, 1560, 2092, 1940, 1262, 1610, 83, 875, 2360, 1961, 1962, 932, 498, 1228, 1581, 2042, 1945, 2215, 786, 1631, 432, 2283, 551, 1973, 1974, 1948, 291, 1638, 1946, 191, 894, 120, 695, 2248, 273, 1976, 736, 1086, 933, 1012, 1620, 1060, 696, 1901, 1732, 1254, 463, 195, 29]. graphs [281, 1235, 1236, 217, 2168, 929, 803, 95, 1378, 336, 1418, 1878, 1860, 1072, 1190, 1583, 934, 204, 201, 2131, 304, 2143, 2383, 980, 649, 835, 1979, 2089, 227, 1565, 2103, 1784, 2298, 2342, 1569, 2299, 2125, 1772, 2388, 1406, 99, 1980, 2367, 117, 2081, 248, 289, 1185, 98, 1240, 2287, 1389, 1459, 535, 1984, 915, 1197, 1312, 2144, 1224, 885, 319, 684, 2064, 80, 971, 256, 1675, 1258, 2291, 1514, 992, 1461, 2371, 2377, 1380, 2153, 2398, 1993, 2146, 1991, 2044, 1054, 1649, 2000, 644, 1498, 1886, 2024, 459, 1883, 1058, 1157, 236, 335, 802, 434, 1888, 550, 253, 306, 366, 2213]. graphs [11]. Gray [1813, 1814]. greatest [2104]. Greedy [1888, 2073, 2230, 2146, 335]. Grey [1859]. Grey-box [1859]. grid [744, 696, 2081, 787, 470, 802, 548]. grids [1226, 1368, 1518, 1656, 295, 2335, 1623, 469]. Grobman [1375]. Gröbner [1392, 62]. grooming [666, 947]. group [798, 1857, 1927, 2107, 979, 88, 887, 1191]. group-Steiner [979]. groupoids [2273]. Groups [1778, 1220, 194, 697, 833, 500, 1626]. Growth [250, 806, 299, 1898]. guarantees [2390]. Guard [894]. Guarding [466, 386, 2035]. guards [2145]. guessing [2187, 2359]. Guest [2405, 2373, 206, 1526, 2003]. Guide [2109]. guided [269].

Hadamard [1397]. hairpin [1082, 657, 330]. Hajós [95]. Half [1821]. Half-turn [1821]. Hamilton [736]. Hamiltonian [1061, 2026, 1520, 26, 803, 753, 901, 553, 1700, 389]. Hamiltonicity [291, 915]. Hamming [1856, 1051]. Hankel [2331]. Hanoi [1784]. haplotyping [1137]. Harary [1772]. hard [1053, 2032, 318, 1656, 2083, 2059, 7, 1211]. hardcore [2355]. harder [301, 66, 881]. Hardness [920, 2149, 741, 897, 800, 1473, 1610, 2036, 1090, 2456, 1625]. Hartman [1375]. hash [1950]. hashing [1549].

Hausdorff [702, 385]. having [2057, 2288, 2457, 1336]. head [353, 1278, 233, 1282]. heavy [2358]. Hedge [2449]. height [1274]. held [1119]. helps [47]. hemirings [1654]. Henkin [2053]. heptagrid [350]. hereditary [1960]. Herman [2319]. Hermitian [2023]. heteroscedastic [212]. heuristics [2346, 402, 2256, 693, 1174]. hexagonal [470]. hexagons [1821]. Heyting [1806]. HHD [1312]. HHD-free [1312]. hidden [2384]. hiding [617]. hierarchical [830, 87, 450, 227, 954, 436, 2226, 1460, 372, 2132]. hierarchically [1842]. hierarchies [1148, 42]. hierarchy [1289, 1168, 5, 1557, 1661, 1234, 2451, 2447, 1120]. high [1084, 2414]. high-level [1084]. higher [1242, 1299, 110, 1988, 1378, 2174, 2101, 653, 1413, 113]. higher-dimensional [110, 1378]. higher-order [1299, 1988, 2101, 113]. highest [430]. Highlights [1452]. Highly [461]. highway [1410]. hinge [2383, 471]. Histogram [2115]. Histogram-based [2115]. Hitchhiker [2109]. hitters [2358]. hitting [2376, 834, 143, 126]. HKZ [2019]. HKZ-bases [2019]. Hoc [2340, 448, 628, 1887]. HOL [1299, 588, 341]. hole [1066, 1900, 835]. holes [2349]. homogeneous [2296]. homology [441]. homomorphism [1571, 500]. homomorphisms [1374]. homonyms [1751]. Honey [1308]. Honey-Bee [1308]. honor [2267]. hop [1321, 2203]. Hopcroft [266, 989]. horizon [1534, 1966]. Horn [128]. Hu [1014]. hull [1544]. Hybrid [156, 155, 1722, 772, 823, 37, 1716, 1942, 1243, 1609, 1244, 705, 1107, 917, 2062, 152, 1642]. hybridization [1080]. hyenas [2040]. Hyper [1437, 1594, 265, 1514, 1437]. hyper-D-width [1437]. hyper-minimizing [265]. Hyper-T-width [1437]. hyper-Wiener [1514]. hyperbolic [87, 350]. Hyperbolicity [181, 60]. hypercodes [2318]. hypercube [1520, 1780, 2131, 2133, 2231, 553]. hypercube-like [1520, 1780, 2131, 553]. hypercubes [1869, 296, 1944, 2227, 1552, 999, 1458, 2335, 2397, 2460]. hyperelliptic [1698]. hyperfinite [514]. hypergraph [1223, 7, 918]. hypergraphs [1442, 1503, 1548, 1899, 98, 299, 1437]. hypermutations [400]. hyperplane [321]. hyperplane-cover [321]. hyperproperties [2196]. Hypertree [1442]. Hypertree-depth [1442]. Hypervolume [1052, 1053]. Hypervolume-based [1052]. hypothesis [1376].

Iasi [1119]. ICTAC [1351, 2177]. ID [1492]. ID-based [1492]. IDD [599]. IDD-based [599]. ideal [1506]. ideals [1813, 2119]. idempotent [2161]. identical [665, 1370, 1462, 1999, 2062]. identification [1716, 1400, 2430, 1265, 1804, 2018]. Identity [815, 2018, 2167, 428, 884, 2203, 2063, 2226, 282, 1535]. Identity-based [815, 2167, 428, 2203, 2226]. iff [1492]. ignorance [1302, 47]. II [1813, 615, 680, 517, 1650]. II. [2287]. Illustration [485]. im [1749]. image [378, 475, 468, 852, 474]. images [2115, 1908, 473]. immune [1671, 400]. Immunity [888]. impact [448]. Impartial [1655]. impatient [2083]. imperative [175, 1988, 585]. imperfect [1715, 1116]. implement [1942]. implementation [331, 1390, 1592, 263, 2431, 2061]. implementing [1935]. implicational [167]. implications [1052, 1480]. Implicit [2440, 34, 523]. Implicitization [1594]. importance [2259]. imposing [1523]. impossibility [1936, 1664]. imprecise [702]. improve [712]. Improved [986, 651, 293, 564, 1152, 244, 579, 2442, 412, 408, 990, 1320, 791, 1916, 1323, 2023, 1596, 1763, 39, 1949, 375, 2412, 71, 2165, 646, 196, 389, 1039]. improvement [2448]. Improvements [2019, 436]. Improving [1397]. in-place [135]. Inapproximability [194, 1587, 668, 2149, 216]. Incidence [2397, 2460, 2239]. inclusion [1330, 1711, 1969]. incompatible [2243, 549]. incompleteness [362]. inconnected [367]. inconsistency [1736]. Inconstancy [541]. Increasing [1614, 1553]. Incremental [1732, 215, 2130, 409, 1961]. incurred [926]. indefinite [2435]. independence [1583, 1148, 915]. Independent [1978, 1031, 1421, 1503, 1194, 1979, 538, 2064, 2165, 2371, 1463, 1498]. Independent-valued [1031]. indeterminates [1110]. index [2140, 1252, 652, 1310, 1063, 99, 1800, 1514]. indexes [1632]. Indexing [2403, 1640]. indicator [1795]. indicators [1358]. indices [1394, 1507]. indiscernibility [850]. individuals [540, 486]. Induced [1012, 2312, 1215, 1960, 965, 1581, 2107, 1620, 1570, 1406]. Inductive [495, 2383]. Inefficiency [2037]. inequalities [1174]. inequality [2104, 1002, 1514]. inevitable [1431]. infeasibility [533]. infected [286]. inference [495, 2407]. Inferring [1647, 78]. infinitary [1452]. Infinite [602, 1618, 93, 541, 202, 906, 1602, 872, 949, 2058, 642, 396, 1720, 1169, 194, 1179, 57, 2364, 1608, 1390, 311, 2159, 758, 359, 1120, 1966]. infinite-horizon [1966]. infinite-state [1720]. Influence [1137, 2202, 2313, 374]. Influence-and-Exploit [2202]. Information [617, 202, 853, 2015, 1477, 1650, 97, 665, 1370, 562, 1059, 1715, 2194, 150, 147, 1681, 211, 1349, 2017, 2196]. inheritance [1044]. initial [827, 949, 687, 81, 2005]. initialization [1647]. injective [1704]. inner [1662]. input [1169, 2192]. input/output [2192]. Inscribed [465]. insecurity [2187]. insertion [357, 421, 1496, 1363]. insertion-deletion [1363]. insights [2126]. Inspired [1119, 400, 69, 1103, 852]. instance [1858]. instances [1385, 760, 1090, 2456]. instantaneous [440]. integer [497, 1334, 945, 231, 1877]. integers [1721, 811]. integral [1587, 263, 1500]. integrals [1590]. Integrated [1131, 326]. integration [797]. intelligence [161]. inter [1161, 682]. inter-cell [682]. inter-deriving [1161]. interaction [163, 356, 109, 521, 2389, 514]. interactions [164]. Interactive [1349, 1584, 853, 2223]. interconnection [227, 1313]. Interface [619, 2454, 2309, 1831, 1879]. interfaces [169]. interference [1316, 1368, 927]. interleaved [2023]. interleaving [1711, 232]. intermediate [1164]. intermittent [716]. Internal [1160]. International [1351]. Internet [1311]. Interplay [1749]. Interpolation [1539, 477]. interpretation [867, 2272, 169, 1340, 1836, 1173, 1111, 1540]. interpretations [590, 1475]. intersecting [193]. intersection [764, 1295, 1336, 298, 291, 1085, 1341, 232, 915, 2064, 1627]. intersections [1727]. Interval [2450, 659, 77, 1890, 1731, 1913, 244, 1253, 1235, 336, 1418, 304, 1949, 2225, 670, 98]. Intervals [1011, 2055, 333, 2241, 48]. intractable [1563]. Intrinsic [951, 348]. Introducing [1841]. introduction [2303, 2003, 779]. intruder [2187, 103, 894]. Intuitionistic [520]. Invariant [247, 422, 1030, 969, 1995]. Invariants [107, 307, 860]. Inverse [1111, 425, 1061, 1061]. Inverse-limit [1111]. Inversion [228]. inversions [1637]. inverting [1641]. investments [1530]. invisible [1232, 1616, 1701]. involution [2318]. involving [1253]. irinotecan [530]. IRNs [1131]. irredundancy [1418]. Irredundant [2077]. Irreversible [662]. ISAAC [2245]. Isabelle [588, 341]. Isabelle/HOL [588, 341]. isolated [309, 1599]. isolation [1804]. isomorphism [910, 499]. Isomorphisms [2152, 1783]. IsRewritten [737]. Issue [583, 1435, 2206, 614, 1351, 2446, 1047]. issues [37, 54, 1975]. item [462]. items [2348, 1553, 237, 2459]. iterated [1289, 608, 657, 330, 1142, 1054]. Iterative [214, 84, 2012, 775, 1405, 61, 1597]. iterativity [2052]. Itineraries [23].

Jacobi [1820, 1727]. Jan [614]. Java [169, 1987]. Jena [1119]. Jiang [186]. job [665, 315, 408, 2243, 375, 549, 1805, 141]. jobs [1891, 2236, 376, 285, 2237, 2037, 1239, 2235, 1432, 48, 1462, 1805, 1999, 2062, 453, 1677, 1382, 1877]. join [310, 1196]. jump [1949]. Jumping [1451]. junction [1292]. just [2236]. just-in-time [2236].

Kadanoff [1851]. Keep [894]. Kernel [735, 1926, 1736, 1778, 2357, 1979, 567, 1916, 1535]. Kernelization [2415]. Kernels [2047, 834, 1743, 2008]. key [2434, 2429, 1688, 1322, 830, 2432, 2443, 2433, 1064, 2436, 416]. keyword [1429]. KGC [810]. kinds [1290]. kinetic [1005]. kings [59]. Kleene [946, 454]. knapsack [2097, 1761, 316, 2212, 881, 1889, 1982, 88, 142]. knowledge [2010, 2307]. knowledge-based [2307]. knowledge-seeking [2010]. known [665]. Kolmogorov [827, 566, 1420, 1807]. Kriegspiel [278].

L [331]. L-machines [331]. label [1223, 67]. label-equivalence [67]. labeled [93, 1060]. labeling [1937, 1741, 551, 1860]. labelings [2251, 256, 1675]. labelling [110, 1569]. labels [508, 2015]. Lachish [494]. ladder [127]. lambda [259, 1828, 1452, 1541, 868, 520]. lambda-calculus [259, 520]. lamplighter [1626]. landscape [858]. landscapes [1050, 2257]. Language [1082, 981, 1283, 1726, 2130, 1850, 841, 585, 2104, 49, 352, 1346, 1164]. language-theoretic [352]. Languages [953, 2186, 2172, 1538, 214, 1289, 547, 2201, 1687, 53, 67, 329, 1275, 1506, 1276, 2182, 1782, 1277, 1474, 1339, 241, 1909, 429, 1550, 1085, 1341, 1791, 1060, 1293, 311, 1689, 1279, 1280, 2317, 1658, 768, 1238, 898, 1100, 1672, 1823, 1120, 1995, 1105, 1997, 767, 250, 797, 888, 510, 78]. large [667, 2323, 1171, 695, 960, 1566, 239, 2015]. largest [696]. Last [611, 2412]. last-step [2412]. late [1462]. Latency [2163, 2108, 1250, 579, 1773]. Latency-bounded [2163]. Latency-optimal [2108]. lateness [1432, 1805]. lattice [762, 1401, 1138, 1011, 1376, 2159]. lattice-structured [1376]. lattices [946, 1112, 2439, 2095, 995, 1698, 1806]. law [1976, 1258]. layer [2357, 1733]. layouts [1223]. Lazy [625, 25]. LCS [1133]. leader [820, 1231]. Leaf [876, 230]. learnability [1304, 1534, 1528]. learners [505, 1658]. Learning [2113, 1529, 325, 1376, 1624, 1533, 743, 209, 211, 213, 15, 212, 182, 214, 1532, 2007, 506, 2406, 215, 509, 1689, 2012, 545, 1055, 510]. least [1053, 752, 1608, 219, 1991]. least-squares [219]. leaves [1212]. Left [287]. Left-forbidding [287]. lemma [1066, 622, 1420]. lemmas [113]. Lempel [2140]. length [1133, 1329, 1216, 1891, 1547, 1136, 1426, 1869, 2442, 1496, 1801, 999, 2243, 334, 2235, 813, 763, 1462, 1382]. lengths [2439, 2059]. Less [2403, 1462]. Let [1445]. letter [879, 1379]. letters [896]. Letting [1480]. level [1761, 1752, 1084, 1257, 1460, 654, 810]. levels [38, 2033]. Lexicographically [752]. LFSR [1237]. library [368, 767]. lifetime [1323, 1994]. Lightweight [1641]. like [2166, 1517, 1520, 1780, 2131, 343, 553]. likelihood [19]. Limit [2420, 86, 1376, 1849, 1111]. Limitations [358]. limited [1532, 2356, 1600, 316, 682, 2235, 1983]. limits [1896]. line [1352, 466, 1093, 882, 1880, 223]. Linear [380, 1586, 38, 1799, 697, 2046, 525, 748, 1461, 569, 2355, 1722, 37, 2450, 764, 965, 378, 1276, 1250, 2121, 14, 57, 621, 976, 1927, 1642, 174, 369, 1172, 2296, 1948, 1043, 401, 2190, 191, 308, 2004, 519, 2123, 812, 1255, 956, 1184, 2357, 2152, 377, 969, 2237, 1769, 567, 1916, 7, 1099, 289, 279, 2173, 2091, 35, 982, 2231, 2001, 2291, 8, 149, 2306, 1700, 1677, 2338, 515, 518]. linear- [969]. linear-size [2355]. linear-time [37, 174, 401, 2123, 8, 2306]. linear-time-branching-time [2190]. linearity [2039]. Linearization [1735, 1375]. Linearly [1517, 1188]. link [1176, 924, 823, 2442, 716, 231, 1998]. link-length [2442]. linkable [1492]. linked [1492, 2059, 1455]. links [120, 2384]. lion [2040]. Lipschitz [490]. list [749, 1068, 1732, 2251, 1941, 306]. Listing [50, 201]. lists [1134, 2005]. lived [116]. liveness [2196]. LL [785]. load [1619, 1694, 9, 552, 1230, 1875, 450]. load-balanced [1875]. loading [231]. Local [458, 498, 1850, 2189, 2071, 747, 1420, 1426, 1044, 1901, 1320, 446, 1054, 2280]. Localities [1095]. locality [1663, 1666]. Localization [1324, 1372, 1319, 2342, 2343]. locally [1902, 1922, 1603, 753, 554, 969, 538]. Locating [1439, 1204, 122, 2200, 897]. location [140, 2234, 2380, 1567, 220, 1523]. Locator [2166]. locus [123]. locus-based [123]. log [1590]. log-sine [1590]. logarithmic [2087, 2412, 1501]. logarithmic-space [1501]. Logic [1714, 38, 689, 1799, 620, 1042, 2305, 530, 1299, 1768, 1043, 1988, 442, 973, 525, 2276, 1411, 1861, 35, 596, 340, 770, 1155, 1862, 501, 975, 515, 518, 2364, 514, 632]. Logical [587, 331, 771, 535]. logics [904, 772, 37, 2450, 1603, 778, 1604, 174, 995, 1539, 1715, 688, 1833, 780, 1115, 46, 2173, 176]. long [116, 1353, 296, 120, 166, 1591, 253]. long-lived [116]. long-range [120, 166]. longest [1177, 802]. look [2409, 2334, 2461]. lookahead [2243]. loop [1484, 1972, 975]. loop-free [1972]. looping [1128, 857]. loopless [2385]. Loops [1291]. Loosely [1231]. Loosely-stabilizing [1231]. loss [192]. losses [387]. lotteries [127]. Lovász [1420]. low [1856, 1840, 2414]. low-dimensional [1856]. Lower [2439, 1412, 1348, 1201, 2328, 245, 2359, 651, 1967, 1635, 1947, 552, 1521, 412, 142, 317, 10, 2233, 646, 1156]. lowest [602, 1945]. Lowness [1403]. LP [1156]. LRM [1394]. LRM-Trees [1394]. LSH [2128]. LTL [1717]. LTS [632]. ludics [524]. Lumpability [1127, 2195]. lunches [490]. Lyndon [1814]. Lyndons [1642].

M [2280]. M2M [1199]. machine [2208, 2236, 1702, 1911, 376, 449, 285, 1154, 375, 1187, 1239, 1769, 2235, 92, 1573, 279, 1805, 1404, 2132, 141, 112]. machines [2117, 331, 964, 1919, 361, 225, 665, 1370, 893, 1249, 1335, 1694, 414, 351, 801, 262, 41, 450, 96, 133, 990, 1802, 1, 233, 1411, 792, 1691, 1853, 678, 2444, 8, 549, 1462, 1999, 2132, 1382, 1490, 2224]. magic [1569]. Magnus [381]. Mahler [1590]. majority [2204, 1417]. makes [1522]. makespan [1919, 449, 792, 92, 2311]. management [784, 280, 1693, 700, 1055]. Many [2131, 1703, 1517, 618, 1603, 1944, 252, 367, 1767]. many-core [1767]. Many-to-many [2131, 1703, 1944]. many-to-one [367]. many-valued [1603]. map [2332, 1469, 1338]. Mapping [2349, 2121]. mappings [130]. maps [472, 229, 1811]. margin [2008]. Marked [1143, 1389]. market [2198, 582]. marketing [2365]. markets [2216, 1755]. Markov [475, 720, 1290, 1303, 2191, 960, 311, 76, 1591, 210, 2410, 1476, 1966, 2008, 1245]. Markovian [958, 1007, 173, 2195, 1986]. marriage [138]. Masking [65]. matchable [1038, 1706]. matchgates [56]. Matching [2221, 2047, 1458, 2395, 1210, 648, 2379, 2079, 1216, 1652, 131, 1630, 1637, 564, 875, 1699, 2401, 1635, 298, 1340, 2294, 652, 1555, 1643, 2075, 882, 1879, 1036, 1609, 818, 886, 819, 1627, 2031, 2022, 1061]. matchings [1134, 2440, 1333, 1178, 2391, 1842, 1581, 2065, 695, 581, 462, 2056, 2217, 2246]. mathematical [1846, 372]. matrices [257, 480]. Matrix [1363, 2239, 1593, 537, 1524, 90, 2345, 2402, 219, 1981, 137, 2415, 2014]. matroid [2341]. matroidal [1511]. max [407, 238, 821, 1428, 1190, 693, 2412, 2219, 2008, 2441, 125, 1582, 2259, 1994]. MAX-2SAT [125]. max-cut [1190]. max-edge-coloring [238]. Max-lifetime [1994]. max-margin [2008]. max-min [821]. maxima [878]. maxima-finding [878]. Maximal [15, 778, 1381, 1367, 602, 906, 1202, 1948, 2248, 195, 668, 999, 21, 136]. maximal-exponent [906]. maximality [2317]. maximally [487, 1038, 1706]. maximally-matchable [1038, 1706]. Maximin [1786]. maximization [2198, 2202]. Maximizing [1619, 9, 1694]. Maximum [573, 1947, 1879, 19, 1406, 1503, 2312, 2440, 1960, 387, 1560, 2121, 2080, 1581, 1494, 1583, 1920, 79, 818, 943, 1432, 1323, 2366, 2145, 2389, 819, 1805, 1999, 1991, 2044, 1498, 196, 876, 1978]. MaxSat [1923]. may [710]. MDPs [2413]. mean [2290, 486]. means [618, 1858, 1211]. Measurable [1126, 1069]. Measure [475, 1188, 515]. measurements [1373, 2289]. measures [202, 1590, 2139, 2448, 139, 1528, 596, 1437, 203]. measuring [852]. MeCBIC [1119]. Mechanism [18]. mechanisms [1774, 2182, 1523, 1996]. media [1770]. median [451, 1057, 83]. medians [574, 409]. Mediated [556]. Mediating [2333]. meet [2101]. meet-in-the-middle [2101]. meets [699]. Membership [1963, 2242, 1672]. Membrane [1119, 1121]. membranes [2117, 1083, 2324, 68, 2116, 1685]. Memory [2016, 1532, 116, 699, 2356, 1521, 118, 2090, 1896, 215, 1756]. memory-aware [1521]. Memory-limited [1532]. memoryful [174]. memoryless [1775]. mercurial [815]. merging [135]. mesh [2133, 1199, 2108, 2061]. message [711, 1754, 1749, 1941, 297]. message-passing [711, 1941]. metabolism [530]. method [2359, 467, 1912, 728, 1093, 1917, 1609, 851, 1622, 343, 1597, 1803, 2041]. methodology [2352]. methods [1027, 1546, 1263, 481, 1071, 2048, 566, 2408]. Metric [589, 1856, 2113, 1395, 327, 1057, 176, 312, 1888, 4, 1712]. metric-space [327]. Metrics [633]. Metropolis [143]. Meyniel [1067]. Mezei [32]. microarray [1071]. middle [2101]. midpoint [757]. migration [166]. Min [453, 407, 821, 1139, 1872, 2412, 693, 1887]. Min-energy [453]. min-max [2412]. min-min [1139]. Min-power [1887]. Mind [1689, 78]. minded [540]. Minimal [1676, 610, 2301, 1408, 2306, 357, 507, 1507, 1296, 1731, 356, 1790, 308, 877, 542, 439, 76, 669, 192]. minimality [1912, 1207, 1106, 2048]. minimally [2383, 1710]. Minimax [2411, 1250, 1031]. Minimization [487, 2145, 1223, 1919, 1294, 579, 2357, 1670]. minimize [792, 1432, 927, 1462, 1999, 1677]. minimized [265]. Minimizing [277, 912, 385, 92, 1805, 1877, 2333, 794, 265]. Minimum [1236, 1776, 1893, 787, 1856, 1781, 985, 1248, 1079, 1924, 2347, 2290, 1587, 1619, 811, 1694, 1692, 1742, 432, 1251, 9, 2442, 647, 695, 2135, 1266, 2230, 1614, 2210, 463, 1090, 2456, 216, 1793, 1072, 934, 1621, 2286, 69, 1773, 926, 289, 669, 183, 193, 1881, 2233, 2371, 1000, 2169, 366, 2213, 2126, 1735]. minimum-colour [2347]. minimum-cost [1692]. minimum-radius [1781]. minimum-weighted [366]. Minimum-width [1893]. Mining [2384, 50, 1485, 217]. MinMax [2241]. MinMax-profiles [2241]. minor [935, 696, 332, 992, 236]. minority [684]. minors [1442, 1620]. minterm [839]. minterm-transitive [839]. mismatch [300]. mismatches [1329, 2076]. missing [1137]. mixed [494, 2021, 2088, 1638, 581, 1565, 1883, 550]. MLF [1164]. MLL [519]. Mobile [2340, 1586, 448, 1505, 1902, 1903, 893, 122, 2060, 628, 166, 2393, 2337, 926, 117, 717, 1229, 187]. mobility [448]. modal [1042, 621, 1087, 780, 2276]. Modalities [1115, 176]. Model [174, 1715, 2059, 2453, 323, 1444, 1316, 1005, 153, 784, 690, 192, 1387, 274, 860, 1789, 475, 2430, 530, 320, 1520, 1043, 552, 237, 2459, 1008, 2148, 1071, 1292, 1300, 771, 533, 2300, 1802, 1833, 1566, 2133, 460, 2299, 69, 2160, 297, 1851, 783, 822, 568, 599, 2308, 717, 1231, 1422, 1804, 1478, 2257, 1649, 453, 2344, 1424, 1767, 43]. Model-checking [1715, 2453, 2308, 1478]. model-generation [690]. Modeling [683, 1244, 1121, 1328, 1125, 849, 954, 853]. Modelling [1119, 1084, 1354, 1682, 1360]. Models [2437, 1855, 493, 2181, 1130, 2114, 689, 1160, 749, 1080, 1170, 559, 523, 529, 595, 2364, 960, 1896, 533, 597, 687, 1575, 1755, 1103, 434, 328]. Moderately [1919, 2441]. modern [749, 341]. modified [467]. Modular [1360, 1107, 2086, 1675, 1451]. Modularity [1449, 2095, 1450]. Modulated [2070]. modules [2052, 1131]. modulo [641, 751, 1672, 260, 1073]. Moessner [2281]. molecule [457]. moments [1051]. Monad [346]. monadic [263]. monads [775, 631]. Monge [1037]. monitorable [2179]. Monitoring [1318, 1444, 1531]. monodromy [481]. Monoid [2014, 346, 62]. Monoid-matrix [2014]. Monoidal [1110]. monoids [1799, 60, 136, 1338]. monomer [2298]. monomer-dimer [2298]. Monomials [2063, 1765]. monopolar [2103]. monopolies [1483, 883]. monotone [182, 329, 2380, 656, 1153]. monotonic [704, 1474]. monotonicity [746]. monotony [1848]. Moore [1294, 989]. Morphic [1496]. morphing [1957]. morphism [2402, 1142]. morphisms [966, 1264, 2174, 1585, 645, 1074, 1142]. Morse [2320, 2331]. mosaic [2142]. Moss [2305]. most [497, 369, 812, 1753, 1984, 2024]. mostly [618]. motif [1413, 2294]. motifs [1631, 439, 2077]. motions [385]. Motzkin [16]. Move [228]. Move-to-Front [228]. moves [1193]. moving [233]. MPNL [1721]. MR4UM [1750]. Muller [2034, 2447, 977]. Multi [2207, 2347, 1087, 1808, 461, 243, 772, 1358, 994, 1646, 267, 671, 1682, 995, 427, 1372, 66, 485, 1771, 353, 1237, 1278, 1879, 554, 2133, 2203, 1099, 436]. multi-agent [427]. multi-channel [1771]. multi-coloring [994]. multi-coloured [243]. multi-component [1646]. multi-dimensional [772, 554]. multi-head [353, 1278]. multi-hop [2203]. multi-interface [1879]. multi-mesh [2133]. Multi-modal [1087]. multi-objective [1358, 66, 485]. multi-parameter [1099]. multi-party [1808]. Multi-Path [2347]. multi-robot [1372]. multi-scale [1682, 436]. multi-secret [1237]. multi-tape [1278]. Multi-target [2207]. multi-tilde-bar [267, 671]. multi-valued [995]. multi-word [461]. multiagent [2040]. multicast [47, 371, 1776, 1887]. multicast/broadcast [1776]. multicasting [1199]. multicolorings [2144]. multicore [699]. multicover [190]. multicut [2021]. multicuts [800]. multicyclic [299]. Multidimensional [1975, 1303, 1863, 1343, 2335, 510]. multidirectional [1904]. multigrid [763]. multihead [1335, 42]. multilinearity [2063]. multiobjective [1052]. multiparty [576]. multiple [167, 2204, 1857, 28, 2045, 577, 1523, 2151, 336, 1418, 1129, 2424, 664, 970, 2366, 332, 2297, 48, 2389, 1018, 2385, 370, 510]. multiple-interval [336, 1418]. multiple-strip [370]. Multiplication [641, 1073, 497, 537, 2284, 1981]. multiplicative [90]. multiplicities [1598]. multiplicity [1070, 945]. multiprocess [1757]. multiprocessor [2311]. multiscale [1125, 1467]. multiset [487, 190, 2184]. multisingular [1244]. multivalued [545]. multivariate [1765, 477]. Multiway [135]. Murat [1073]. Musings [521]. mutation [2258, 2259]. mutual [1140]. Myhill [391].

Narrowing [10]. Nash

[543, 17, 715, 1621, 2037, 1625]. Natural [1575, 2130, 1721, 393, 844, 2184]. naturally [1805]. Nature [852]. Nature-inspired [852]. Near [1317, 2413, 800, 1775, 1571, 852]. Near-optimal [1317, 2413, 800, 1775]. near-unanimity [1571]. nearest [728, 848, 1598, 936]. Nearly [2039, 1753]. Nearly- [2039]. Necessary [1550]. necklaces [1814]. need [1161]. neg-weighted [83]. Negative [401, 2365, 431]. Neighborhood [1154, 862, 546, 1898]. neighborhoods [1921, 2257]. neighbour [848]. neighbourhood [470]. Nelson [973]. Nerode [391]. nested [1611, 424, 1756, 2241, 312]. net [519, 1087, 1100, 1268]. net-based [1087]. nets [380, 1273, 1537, 2178, 1474, 1084, 1719, 171, 1344, 1244, 722, 520]. Network [2221, 1521, 1625, 373, 798, 666, 683, 950, 121, 1317, 234, 122, 369, 594, 1427, 1263, 2036, 727, 2342, 76, 2278, 1964, 677, 1408, 188, 332, 286, 2061]. network-conscious [2278]. Networks [2340, 1359, 1315, 2348, 1781, 2214, 1517, 213, 996, 573, 448, 2085, 1505, 1904, 1933, 246, 387, 1903, 651, 2178, 1169, 2323, 1484, 1774, 2401, 575, 124, 14, 2163, 821, 2341, 1972, 726, 595, 1520, 1251, 1963, 1759, 640, 579, 2313, 277, 1495, 1468, 846, 2370, 628, 1131, 2418, 1718, 1875, 959, 593, 824, 1318, 2020, 660, 1219, 2337, 716, 2300, 1321, 1879, 2419, 227, 367, 1257, 1566, 1776, 2133, 1773, 1994, 2384, 598, 2160, 1575, 787, 1848, 437, 1323, 379, 783, 1800, 119, 599, 1313, 854, 1460, 927, 1804, 718, 2015, 374, 1407, 1024, 1887, 2022, 1199, 2108, 553, 223, 434]. networks [2008, 2041, 1362]. neural [1169, 846, 1793, 71, 2118, 854, 332]. neutrality [1050]. Newman [627]. next [1152, 557, 199, 129]. Next-Fit [129]. next-smaller-values [557]. NFAs [840, 2183]. niche [85]. Nim [2166, 1973]. Nim-like [2166]. Nimbers [1431]. NL [518]. NL-complete [518]. NLC [1078, 1184]. NLC-width [1078]. no [99]. Node [1460, 924, 727, 2089, 1552, 999, 1992]. Node-disjoint [1460, 999]. nodes [369]. noise [212, 1316, 743]. noisy [2411, 531, 761, 1008, 1147]. nomials [478]. nominal [1299]. Non [140, 1857, 2417, 1172, 1138, 2363, 1824, 68, 2435, 170, 1778, 1936, 2172, 986, 1637, 1007, 1532, 276, 1850, 2422, 376, 2313, 832, 2238, 1663, 412, 1664, 2193, 716, 1238, 1698, 446, 1584, 1017, 1848, 1534, 1528, 610, 1666, 2119, 2223]. Non-adaptive [1857]. non-adjacent [1698]. non-approximability [986]. non-associative [2119]. non-atomic [1528]. non-availability [376]. Non-confluence [68]. Non-contextual [2435]. non-convex [610]. Non-cooperative [140]. non-deterministic [2172]. Non-existence [1172]. Non-expansive [170]. non-interactive [1584, 2223]. Non-lattice-periodic [1138]. non-learnability [1534]. non-locality [1663, 1666]. non-Markovian [1007]. non-monotony [1848]. non-oblivious [446]. non-overlapping [1637]. Non-planar [2363]. non-probabilistic [2193]. non-progressive [2313]. non-pushy [1238]. Non-redundant [1824]. non-repetitive [1017]. non-signaling [1664]. non-silent [716]. non-standard [832, 2238]. non-topological [1936]. non-U-shaped [1532]. Non-uniform [2417, 1778, 276, 1850, 2422]. non-utilitarian [412]. noncomputability [393]. nonconstructivity [268]. nondeterminism [2270, 175]. Nondeterministic [1294, 1293, 871, 990, 585, 1665]. Nonemptiness [2285]. nonexistence [255]. noninterference [328]. nonlinear [529, 1375, 1950]. nonlocal [1728]. nonmasking [708]. nonpowerful [1260]. nonsense [1099]. nonseparating [1979]. nonsignalling [1667]. nonsplitting [496]. Nonterminal [838, 1283]. Nontriviality [1725]. nonuniform [2419, 2423]. Nordhaus [1514]. norm [1587, 2031]. Normal [1577, 479, 71]. normalisation [445]. normalization [991, 35, 516]. notation [636]. Note [2073, 670, 1316, 1270, 1034, 2090, 482, 292, 428, 750, 898, 1073, 1104, 2110, 889, 445, 1795, 2338]. notion [543]. Notions [60, 53, 44, 2223]. novel [1924, 1780, 554, 588]. NP [2187, 904, 451, 1053, 1262, 1022, 1789, 65, 433, 2059, 1036, 7, 1211, 80]. NP-complete [2187, 904, 451, 1022, 65]. NP-Completeness [80, 1262, 1036]. NP-hard [1053, 2059, 7, 1211]. NTRU [2439]. Number [908, 2187, 2214, 1544, 540, 949, 1079, 1033, 288, 2057, 2288, 2457, 2411, 1610, 2315, 2068, 430, 1913, 1608, 2, 2292, 2248, 656, 796, 880, 1793, 1583, 934, 896, 1949, 2225, 750, 431, 1983, 1196, 653, 915, 1898, 1462, 1500, 389]. number-conserving [656]. numberings [495]. numbers [872, 1577, 1721, 1204, 605, 310, 544, 2184, 1501]. numeration [832, 2238]. numeric [1696, 1592].

OBDD [497, 1248]. OBDD-based [1248].

OBDDs [1016]. Obituary [1285]. object [961, 2380]. object-oriented [961]. objective [1358, 66, 485]. objects [1834, 344]. oblivious [1779, 118, 1901, 895, 652, 252, 446, 187]. obnoxious [1774]. observation [1229]. observationally [1988]. Observations [716]. observers [355, 362]. observing [355]. obstruction [1067]. obtained [308, 1088]. Obtaining [1562]. OCaml [1163]. occasion [614]. occurrence [1719]. occurrences [1142]. odd [1012, 992]. odd- [992]. ODEs [534]. offline [1034]. offspring [2255]. old [1440]. omnidirectional [2393]. On-demand [39]. On-line [882, 223]. once [1753, 1984]. One [1808, 1335, 1983, 730, 2031, 394, 1066, 174, 1496, 1470, 1658, 367, 1009, 1599, 1098, 678, 1379, 813, 8, 24]. one-cluster [813]. one-counter [174]. One-dimensional [2031, 1470, 678]. one-letter [1379]. one-point [1098]. One-reversal [1335]. One-round [1808]. one-shot [1658]. One-sided [1983]. one-tape [8]. One-to-one [730]. one-way [1009, 24]. ones [1872]. Online [225, 2390, 414, 578, 909, 801, 2209, 316, 2212, 450, 2243, 2235, 1440, 210, 2444, 372, 549, 795, 370, 1677, 2362, 1214, 2353, 784, 1891, 2097, 2139, 665, 1370, 2411, 874, 1635, 2341, 129, 1838, 226, 2378, 2004, 1746, 408, 693, 1764, 1775, 791, 792, 1239, 1691, 2033, 1379, 1018, 2132, 365, 1382, 1039]. only [2356]. onto [1888]. open [1532, 1069, 1702, 1798, 1783]. opening [2223]. operating [2256]. operational [779, 148]. operations [1274, 2440, 1180, 1909, 1232, 1114, 1281]. operator [165, 608, 1030, 2280]. operators [267, 671, 1512, 442, 598, 98, 2175, 635]. optical [666, 277, 2133, 2061, 65]. Optimal [2214, 1368, 1904, 1959, 246, 1721, 130, 874, 1484, 1842, 1779, 1521, 2191, 792, 1239, 1476, 2436, 2132, 365, 2061, 1717, 1176, 1027, 798, 2079, 800, 1317, 574, 1720, 987, 555, 1688, 820, 1580, 1307, 2004, 2071, 2142, 788, 1753, 2413, 1488, 1775, 2335, 1258, 48, 229, 2062, 474, 1199, 2108, 1229]. Optimality [242, 2073, 1505]. Optimally [213, 1772]. optimistic [2099]. Optimization [537, 2206, 2382, 1052, 1358, 165, 467, 22, 163, 530, 1253, 66, 485, 336, 2326, 596, 2444, 1996, 411, 183, 2325]. optimizer [491, 162]. Optimizing [947, 1954, 478, 1758]. Oracle [717, 1424]. Oracle-based [717]. oracles [112, 1564, 1701]. orbit [752]. orchestrations [2269]. orchestrators [260]. Order [543, 2075, 1813, 775, 904, 2050, 1242, 307, 435, 621, 1642, 2136, 1299, 1988, 2101, 653, 297, 35, 568, 113, 871]. Order-preserving [2075, 307, 297]. Ordered [1226, 2452, 925, 1937, 2119, 1212]. ordering [1451, 1099]. orderings [57]. orders [871, 457, 2450, 1949, 2152]. organized [1131]. organizing [1469]. orientable [1649]. orientation [2379, 1545]. orientations [2248, 2100, 80]. oriented [961, 1831]. orienting [1638]. Origins [353]. orthogonal [1797, 2363]. oscillations [154]. oscillators [153]. oscillatory [152]. other [1852]. outerplanar [2361, 438]. outliers [1785]. output [2192]. overlap [1330, 70]. overlap-inclusion [1330]. overlapping [1637]. overlaps [1217]. overlay [1933, 1090, 2456]. overloops [1291]. overview [780]. ownership [28]. Ozbudak [1073].

P [2117, 2114, 1789, 310, 1793, 71, 68, 2118, 2116]. P/T [310]. PAC [505, 2413, 1527, 1528]. PAC-Bayes [1527]. PAC-learners [505]. packed [2079]. packet [1693, 1764, 1775]. Packing [2348, 2047, 1744, 1201, 564, 1907, 240, 226, 1012, 1797, 408, 2250, 930, 502, 370, 2362, 1889, 25]. PADS [954]. Paging [1505]. Paired [1944, 1399]. paired-domination [1399]. pairing [416]. pairing-free [416]. Pairs [257, 1829, 2452, 1512, 1044, 2017].

Pairwise [1250, 1482, 2057, 2288, 2457, 186]. Palindrome [1643]. Palindromes [2316, 288, 1136, 1608, 603, 49]. Palindromic [2151, 891, 1995, 396, 452]. palindromization [1338]. PALS [1298]. pancake [430]. pancakes [419]. pancyclicity [1910, 1974]. paper [1270, 1073, 2431]. Paperfolding [966]. papers [1119]. parabolas [763]. paraconsistent [973]. paradigm [1319]. paradigms [1071]. Parallel [2069, 552, 832, 480, 285, 392, 180, 487, 964, 547, 320, 1230, 801, 277, 2148, 2238, 418, 408, 1256, 2326, 792, 1769, 2235, 92, 705, 2444, 332, 549, 270, 1404, 1690, 2061, 1767, 1877, 1490]. parallel-batch [549]. Parallel-machine [285]. Parameter [595, 932, 1927, 597, 1099, 596]. Parameterisation [1989]. parameterised [1924, 107]. parameterizable [1485]. parameterizations [834]. Parameterized [1329, 1538, 1177, 1509, 2391, 1923, 1305, 126, 565, 1953, 1930, 1418, 1920, 1786, 1994, 2160, 1421, 303, 1887, 1886, 935, 1504, 1657, 1947, 2228, 1252, 2441, 2442, 1035, 1732, 931, 96, 336, 1419, 248, 2164, 2366, 2165, 992, 1733, 321, 502, 1928]. Parameterizing [880]. Parametric [271, 2037, 1717]. parametrized [2052]. parentheses [312]. parity [2448, 1386, 2034, 205, 1463, 203]. Parry [872]. parses [1574]. parsimony [2336]. Parsing [1379, 1981, 1027, 867, 2073, 91]. part [1813, 1477, 1650]. Partial [788, 474, 1806, 1531, 337, 667, 1135, 1217, 1066, 1653, 2058, 2124, 97, 665, 845, 2406, 585, 1566, 2257, 548]. partial-monitoring [1531]. Partially [82, 907, 925, 1374]. participants [1802]. particle [163, 162]. parties [2297]. Partition [1730, 576, 2374, 1334, 224, 2230, 1418]. Partitioning [1545, 2026, 1922, 371, 1642, 1253, 1656, 7, 2164, 916, 2219]. partitions [1822]. partizan [251, 1951]. parts [1514]. party [1808]. passage [960]. passage-time [960]. passing [711, 1147, 1941, 297, 1700]. Passively [893]. past [1187]. past-sequence-dependent [1187]. Path [2347, 28, 257, 77, 1926, 387, 764, 778, 1972, 1451, 1512, 2442, 463, 1944, 934, 1195, 742, 2131, 1552, 1920, 2081, 188, 730, 2017, 1703, 1882]. path-based [778]. Pathological [1031]. paths [2353, 121, 735, 1906, 82, 1869, 296, 790, 999, 98, 1460, 1461, 901, 729, 236, 802, 253]. pathway [597, 534]. pathways [1328, 1360]. pathwidth [2361, 249]. patrolling [2145]. pattern [602, 1652, 1866, 1635, 1643, 1094, 609, 1298, 851, 1997, 2031, 1381, 2061, 78]. patterned [1790]. Patterns [763, 2074, 676, 1179, 298, 1790, 191, 255, 2294, 1591, 187, 1260]. Paxos [1758]. payments [2211]. payoff [1524]. PCA [332]. pearl [1162]. pebbles [2157]. Peek [36]. peer [1938]. peers [1205]. penalties [1567]. penalty [1508]. Pennies [2221]. PEPA [1007, 532]. PEPAd [1007]. per-character [2201]. per-flow [821]. perception [1349]. percolation [1852]. perfect [1332, 1678, 1493, 1333, 278, 1563, 1930, 2143, 1036, 1950, 284, 1627]. perfectness [1749]. Performance [1008, 2256, 705, 2139, 788, 1897, 1938]. Period [2332]. Periodic [1389, 2150, 2074, 1228, 2045, 1138, 1470, 322, 792]. Periodicity [1217]. periods [2030]. permutation [444, 1549, 463, 650, 1925, 289, 982]. permutations [602, 1394, 1952, 1826, 192, 1021, 1618, 1088, 2142, 1189, 2241]. permuted [2218]. Perron [1820]. persistent [1332]. personalized [124]. perspective [2053, 973]. perturbation [1008]. perturbed [195, 2371]. Petri [380, 1273, 1537, 2178, 1474, 1344, 1100, 1244, 1268, 722]. Pfaffian [2100]. phenomenon [1963]. pheromone [164]. Philippe [725]. Phutball [318]. phylogenetic [1954]. phylogeny [1332, 1639, 1296, 1563]. physical [331, 1316, 2344]. physics [1170, 2265]. pi [171, 1388]. pi-calculus [171]. Picker [1546, 294]. pickup [873]. pile [1851]. Piles [43, 1170]. pipelining [1758]. Pisot [283, 877]. Pitts [2280]. Pivots [1333]. pixel [468]. place [135]. placement [2214, 1257, 1408]. plain [394]. plan [866]. planar [1957, 1696, 985, 2087, 892, 1673, 1560, 2092, 498, 695, 2248, 1012, 696, 1139, 95, 204, 980, 1979, 2363, 567, 1916, 1211, 1459, 1984, 1380, 1993, 2044, 729, 2000, 644, 2024]. planarity [2088, 1948, 1562]. Planarization [1733]. plane [466, 555, 87, 1254, 1595, 1236, 1822, 350, 758, 319, 80, 2146, 1783]. planefilling [966]. planes [756, 1820]. planning [525, 1755]. Plateaus [66]. platforms [1084]. play [1445]. played [116, 1973]. player [251, 829, 1401, 2032]. Playing [182, 278]. plural [1648]. plus [1316]. PMC [2300, 2299, 1804]. PNL [1299]. Pnueli [176]. point [2379, 1112, 1042, 712, 1373, 2289, 626, 2148, 269, 181, 702, 385, 757, 1098, 1853, 516, 2019, 2031, 622]. point-to-point [712]. points [130, 27, 1600, 769, 1255, 608, 750]. polar [2103]. polarised [171]. policies [1959, 2006, 2099]. Policy [869]. Polish [636, 1445]. polycubes [1138]. polygon [123, 2349]. polygonal [2145]. polygons [89, 2387, 701]. polyhedron [1023, 1892]. Polynomial [1473, 1251, 1265, 1096, 419, 2343, 1783, 77, 985, 1033, 987, 237, 2459, 479, 308, 309, 482, 1737, 1658, 1599, 936, 544, 1576, 1489, 231, 1051, 1805, 918, 389, 1670]. Polynomial-time [1251, 419, 2343, 1737, 231]. polynomials [641, 1765, 1593, 1827, 976, 1598, 1073, 1597]. Polyominoes [862, 243, 1551, 1818, 758]. polytope [406]. polytopes [1627]. pooling [368]. Popular [581, 462, 2246]. population [709, 1934, 1171, 2337, 556, 2255, 1231]. populations [1098, 486]. pos [83]. pos/neg [83]. pos/neg-weighted [83]. poset [1813]. Posets [192, 547, 168]. position [729]. positional [926]. positioning [726, 926]. Positive [677, 1559, 1857, 2202, 1689, 2012, 1489, 374, 510, 78]. positive-dimensional [1489]. possible [1110, 967, 1382]. Post [436]. Post-processing [436]. Postman [1953]. potential [1183, 1621]. Power [1896, 1316, 2346, 1537, 1169, 1507, 1940, 1834, 1976, 280, 446, 10, 193, 785, 981, 40, 1258, 250, 3, 864, 1887]. power-free [250]. power-law [1976, 1258]. powers [906, 2026, 230, 1697, 2030, 670]. PPTL [2368]. practical [1052, 2364, 75]. practice [1053]. PRE [2223]. precedence [1195, 1382]. Precise [2194]. precision [1600]. preclusion [875, 1699, 1458, 886, 2022]. precomputation [2284]. predicate [585]. predicative [517]. predicted [1594]. Predicting [508]. prediction [208, 848]. predictions [530]. predistribution [1322]. Preemptive [279, 1746]. Preface [1109, 571, 536, 1448, 1076, 766, 614, 1159, 1286, 342, 464, 152, 2262, 52, 513, 1748, 1932, 528, 1956, 817, 2155, 364, 692, 1415, 2446, 1724, 984, 1272, 774, 405, 674, 1895, 476, 1589, 1222, 1874, 504, 264, 145, 1660, 2428, 1209, 1915, 64, 151, 158, 160, 178, 185, 397, 484, 527, 592, 673, 843, 856, 941, 1004, 1046, 1118, 1167, 1326, 1327, 1357, 1466, 1680, 1788, 2112, 398, 347, 363, 1247, 100, 1885, 114]. preference [1134, 1684]. prefix [394, 1275, 361, 1021, 1636, 134, 62]. prefix- [1275]. prefix-free [394, 361]. prefix-search [1636]. prefixes [1088, 16]. preorder [920]. preperfect [49]. preprocessing [1587, 851]. Prequential [207]. prescribed [2312, 1545]. presentations [586]. presented [1119]. preservation [2121, 445]. preserves [444]. preserving [690, 54, 2178, 307, 2075, 297, 392]. presheaf [2273]. pressure [73]. previous [1177, 557]. previous- [557]. Price [2060, 2216, 458, 1876]. Pricing [410, 1755, 1559]. primal [937]. primal-dual [937]. primality [2184]. Primary [1999, 998]. Primary-secondary [1999]. prime [813]. primes [2291]. primitive [1653, 45, 1260]. Principles [1454]. Priority [22, 615, 197]. priors [1527]. prismoid [768]. privacy [1938, 1664]. privately [1371]. privileged [1841]. Probabilistic [1005, 172, 534, 781, 2220, 957, 617, 1529, 589, 475, 2188, 1779, 2189, 2192, 1513, 2193, 1089, 956, 585, 1802, 1833, 2421, 2023, 2043, 170, 455, 2306, 339]. probabilistic-nondeterministic [585]. probabilistically [716]. Probabilities [631, 2270, 531]. probability [13, 1807, 207, 1409]. probable [1174]. Problem [2220, 2395, 1856, 77, 451, 448, 694, 283, 2374, 560, 2375, 2353, 1395, 245, 138, 1134, 2211, 2126, 2097, 2359, 749, 1248, 413, 873, 131, 467, 238, 1480, 760, 371, 1697, 1399, 224, 1761, 1587, 2315, 2121, 83, 858, 2236, 1400, 745, 1745, 1731, 1152, 1947, 1305, 2215, 1512, 1929, 2376, 1410, 244, 1428, 237, 2459, 2442, 876, 1636, 939, 1669, 1976, 90, 2090, 1266, 2084, 1035, 1375, 1746, 1139, 1309, 2212, 2402, 1797, 803, 42, 1798, 2247, 216, 322, 1978, 1753, 1195, 2143, 2286, 2059, 597, 2259, 1949, 1036, 1565, 1982, 377, 1567, 79, 69, 499, 324, 1916, 1211]. problem [739, 818, 943, 2424, 2324, 900, 183, 231, 352, 1998, 1880, 1528, 2233, 889, 732, 2366, 2165, 2116, 2017, 382, 2389, 937, 1762, 321, 438, 502, 125, 2219, 2062, 1928, 365, 1058, 6, 2169, 1882, 1735, 1953, 2127, 889]. Problems [583, 1342, 1503, 794, 2207, 1726, 2021, 1150, 986, 1926, 1960, 22, 1785, 1398, 387, 1657, 1922, 561, 2149, 1370, 1532, 1203, 2347, 73, 564, 2285, 2380, 1741, 1123, 1518, 320, 1515, 65, 1522, 565, 1852, 2071, 834, 2168, 190, 96, 1091, 1798, 447, 1361, 336, 912, 881, 1194, 2250, 384, 979, 741, 649, 727, 2089, 219, 25, 1691, 1571, 899, 1287, 643, 2287, 393, 193, 1323, 1964, 2164, 836, 465, 2423, 119, 2431, 1258, 1064, 1461, 1887, 1018, 1423, 1883, 569, 2134, 1435, 2325]. procedure [690, 67, 2364]. procedures [1721, 103, 1605, 1154]. Process [1119, 1306, 1125, 823, 745, 1682, 861, 173, 956, 1243, 1129, 2453, 486, 634, 2195, 1986, 1097]. process-algebraic [956]. Processes [1122, 2270, 710, 2188, 531, 1405, 1084, 2191, 210, 362, 170, 1966, 1245]. processing [315, 480, 2033, 436]. processors [280, 693, 1359, 1362]. product [2239, 1352, 1662, 551, 1569, 1675]. production [1029]. Productivity [58, 625]. products [772, 1548, 2365, 2028]. profiles [2241]. profiling [2314]. profinite [841]. profits [2134]. Program [2455, 108, 921, 1988]. Programming [953, 2186, 1922, 777, 2121, 859, 699, 175, 190, 2259, 372, 1050, 569]. programs [961, 866, 326, 156, 1757, 1782, 2189, 1752, 1605, 945, 270, 975]. progressive [2313]. projected [1280]. Projection [2364, 1768, 501]. projections [1468, 2048]. projective [1603]. Prolog [1162]. prone [1935]. Proof [1556, 2305, 1603, 973, 1936, 1006, 1579, 561, 1774, 1604, 444, 1768, 1523, 1454, 519, 171, 2276, 1420, 1542, 49, 1074, 343, 2368, 518, 1709]. proof-nets [171]. proof-theory [1604]. proofs [1912]. propagating [1574]. proper [77, 1741]. Properties [1014, 380, 1545, 121, 1708, 2360, 1082, 2179, 1487, 721, 1605, 650, 2317, 72, 649, 850, 1819, 1346, 1037, 2291, 343, 646, 654, 553]. property [1847, 1473, 307, 378, 1975, 433, 688, 1280, 1015, 35]. Proportional [582]. propositional [1295, 1768, 128, 501, 2364]. pros [166]. protects [2458]. protein [703, 2059]. proteins [1685, 2218]. Protocol [2187, 1935, 689, 103, 1745, 959, 1232, 1231]. protocols [2438, 709, 1934, 1940, 1387, 1371, 2337, 1802, 556, 104, 2436]. Prouhet [1697]. provability [361]. Provable [282, 1064]. provably [830, 1049]. provably-secure [830]. prove [2359]. provenance [1455]. Proving [721, 1513, 2408]. provision [1691]. Proxy [1701, 1429, 2203]. Proxy-invisible [1701]. Pseudo [805, 1051, 101]. pseudo-Boolean [1051]. Pseudo-cyclic [805]. pseudo-random [101]. pseudopalindromic [608]. pseudorandom [2355]. pseudorandomness [888]. Pseudospectra [1593]. PSPACE [318, 1015, 1685]. PSPACE-completeness [1015]. PSPACE-hard [318]. PTAS [1079, 2096, 2043]. Public [2429, 2432, 1064, 416]. Public-key [2432]. Pure [511, 1625]. purposes [2434]. Pursuing [94]. pursuit [2040, 303]. pursuit-evasion [303]. pushdown [1274, 913, 1282, 1155, 2308, 797]. pushy [1238]. Putnam [1710]. pyramid [660].

QKD [2430]. QoS [2269]. QSAT [179]. quadratic [872, 1835, 317]. qualitative [154, 1130, 1203]. Quality [436]. quantification [1681]. quantified [222, 2452, 2303, 2001]. quantifier [1592, 922]. quantifier-free [922]. quantifiers [2053]. Quantifying [116]. Quantitative [953, 2186, 2196, 2190, 2194]. Quantitatively [962]. Quantum [2429, 1662, 34, 112, 1361, 1665, 2433, 2434, 1113, 1006, 2328, 1667, 2115, 2430, 1663, 1089, 2432, 473, 1009, 942, 740, 1666, 2431, 1155, 2435, 2436, 1794]. quantum-proof [1006]. quartet [1639]. quartile [851]. Quasi [1027, 590, 976, 913, 871, 2088, 1975]. Quasi-distinct [1027]. quasi-expansivity [1975]. Quasi-interpretations [590]. quasi-orders [871]. Quasi-polynomials [976]. Quasi-rocking [913]. quasi-upward [2088]. quasiperiodic [2301]. quasiperiodicities [242]. queens [2324]. queries [443, 121, 2204, 2087, 1439, 1417, 1633, 2242, 272, 2403, 942, 1623, 545]. Query [154, 443, 507, 987, 1770, 1240, 740].

Query-based [154]. query-time [1240]. questions [1304]. queue [1553]. queuing [301]. Quick [2230, 607, 132]. Quickselect [899]. Quiescence [117]. quorum [1413]. quotas [245]. Quotient [1506, 1048].

Rabin [1579, 2279]. Radii [2209]. radio [1368, 1317, 651, 115, 2015]. radius [794, 1781, 1692, 1051]. radix [435]. Rainbow [799, 731, 1610, 2056, 2233, 2377]. Ramsey [2042]. Ramsey-type [2042]. Random [2239, 305, 1210, 2055, 2439, 1828, 1022, 1899, 1842, 276, 291, 577, 1550, 481, 271, 1564, 1824, 1983, 2277, 1576, 915, 1823, 299, 2064, 1701, 1054, 1409, 1424, 101]. Randomised [2016, 2376, 2256, 1816]. randomization [2097]. Randomized [964, 197, 784, 2438, 406, 1079, 244, 552, 1230, 402, 883, 389, 1174]. randomness [2016, 1396, 1306, 1403, 44, 360, 207]. range [704, 574, 2087, 1152, 120, 1633, 166, 899]. rank [2107]. Ranking [1561, 1937, 1364]. rare [1821]. rate [980, 286, 438, 1380, 1993]. rates [1005, 250]. ratio [1746, 1949, 819]. Rational [1336, 477, 93, 841, 81]. rationality [543]. Rauzy [2107, 1570]. raw [1319]. ray [2207]. Raz [494]. Raz-type [494]. re [1429, 912, 2203, 1701]. re-encryption [1429, 2203, 1701]. re-entrant [912]. Reachability [1123, 1836, 1540, 1720, 155, 1474]. Reachability-based [1836, 1540]. reachable [529]. Reaction [1102, 1337, 1089, 1346, 2322, 1471]. reactions [440]. read [2356, 1232, 2063]. read-only [2356]. read-restricted [2063]. Real [2188, 61, 2431, 1630, 1730, 2323, 1962, 913, 1833, 2225, 1970, 544, 1298, 1501]. Real-reward [2188]. Real-time [61, 1630, 913, 1833, 1970, 1298]. real-world [1962]. Realizability [523, 2280, 517]. Realization [2423, 489]. rearrangement [1328, 874, 414]. Reasoning [866, 632]. reassignment [225]. reception [1776].

Receptor [1613]. Recognisability [642]. Recognition [273, 1726, 1687, 418, 1980, 1185, 2287, 1197, 1381, 548]. recognizable [806, 423, 2158]. Recognizing [2028, 2103, 1312]. recoloring [2149]. reconfigurability [1194]. Reconfiguration [2251, 745, 447, 1408]. Reconstructing [243, 701, 1227, 507]. Reconstruction [304, 760, 594, 593]. recording [325]. Recovering [120]. recovery [1202, 668, 717]. Rectangle [930, 795, 1039]. rectangular [1893, 787]. rectilinear [555, 2442]. Recurrence [2058, 455]. recurrences [2296]. recurrent [996, 1169, 2105]. recurring [308]. recursion [1967, 522]. Recursive [769, 327, 2406, 742, 254]. Recursively [4, 1712, 539]. red [560]. red/blue [560]. redistribution [1322]. reducibilities [1725]. Reducing [1876, 2260, 1605, 999]. reduction [2333, 658, 535, 770, 804]. reductions [1396, 1305, 1710]. Redundancy [877]. redundant [1824, 1048]. Reed [2023]. Refactoring [1148]. reference [2432]. references [343]. refined [2445]. Refinement [1650, 1352, 1353, 2189, 269, 533, 1422]. reflection [757]. reflexive [67, 1374]. regenerator [947]. region [386, 2145]. Regions [1344, 2380]. register [1942]. registers [1939]. regression [2009, 2004, 1535]. Regret [2410, 2412, 210]. Regular [2407, 997, 1997, 2312, 1538, 67, 329, 1275, 1416, 1730, 1276, 1711, 1909, 1085, 1341, 1791, 840, 736, 232, 415, 1901, 272, 1279, 1280, 1658, 1939, 1406, 598, 738, 1574, 1672, 1823, 392, 1105]. regular-controlled [738, 1574]. regularities [2068]. regularity [53]. regularization [1595]. regulated [863]. regulatory [1131, 598, 535]. reinforcement [2168]. rejection [1508, 909, 285, 141]. related [1226, 1919, 1694, 414, 1304, 2094, 1798, 336, 330, 899, 643, 609, 465, 1883, 1500].

Relating [2270, 585]. relation [1191, 1719]. relational [1043, 847]. Relations [723, 1652, 1336, 1190, 767]. relationship [420, 10]. relationships [1950]. relative [2140, 911]. relativises [1789]. Relativized [1081]. release [1150, 1432, 732]. relevant [16]. reliable [369, 1749]. relocation [732]. Remarks [2402, 1446]. remedian [1739]. Remembering [118]. removable [316, 2212]. renaming [297]. rendezvous [928]. Renewal [978, 805]. Reoptimization [1960, 2208]. repair [530, 91]. Repairing [1243, 2201, 122]. repeats [2078]. repetends [2458]. repetition [606]. repetitions [906, 2315, 998]. repetitive [1640, 1017]. representability [1547]. representable [1060]. Representation [1687, 2095, 1204, 1288, 908, 580, 1087, 1254, 2142, 272, 1148, 319]. representations [1602, 257, 1189, 352]. representative [1945]. Representing [771]. reproduction [165]. reputation [1101]. required [540, 1869]. rerouting [1906]. research [2422]. resemblance [852]. reset [1287]. resilience [2347]. resilient [246]. resizable [2348]. resolution [2307]. resolved [907]. resolving [422]. Resource [333, 1055, 659, 1395, 199, 954]. resources [590, 1667, 768]. response [582]. restart [2235]. restless [2410]. restrained [1262]. restrict [2099]. Restricted [895, 645, 996, 994, 561, 893, 1520, 640, 2451, 144, 1878, 30, 2131, 1772, 2063, 1282, 2217, 198, 2398, 78]. Restriction [2051, 2141, 2322]. restrictions [2399, 1137, 315, 930]. result [324]. results [744, 986, 199, 1057, 194, 1610, 1587, 14, 1581, 244, 2056, 1187, 897, 330, 1362, 1440, 62]. resumable [376]. retract [106]. retransmission [1897]. retransmission-based [1897]. retrieval [1136, 1059, 1633, 1771, 2229]. Return [638]. returns [2110]. Reutenauer [1464, 1270, 826, 1556]. reveals [1719]. Revenue [2198, 2202]. Reversal [1279, 1335, 1341, 1670]. reversals [1282]. reversibility [378]. Reversible [1740, 1405, 307, 61, 1411, 678, 489]. revision [866]. revisit [2135]. revisited [2395, 1214, 39, 111, 2224, 1967, 1644, 240, 228, 1375, 1449, 1087, 322, 2259, 1100, 1347, 1001, 1335]. Revisiting [1735, 2272, 1230]. Revocable [2226, 1492]. revocable-iff-linked [1492]. Revolutionaries [1438]. reward [2188]. Rewriting [1328, 797, 443, 487, 261, 1330, 259, 1452, 721, 1449, 415, 1182, 863, 1453]. rhombus [759]. rich [2295]. richness [1995, 452]. ridge [1535]. Riemann [481]. right [786, 1983]. rigid [385, 2383, 23]. rigidity [1074]. Rigo [422]. Rigorous [1606]. ring [1492, 1250, 2167, 1779, 427, 1780, 428, 252, 1816, 231, 1424]. rings [1900, 62]. Riordan [609]. risk [373]. RNA [186, 1634, 2389]. road [711]. robber [1439, 94, 386, 2200]. robber-region [386]. robbers [1446, 1616, 1572, 302]. Robot [2289, 246, 1372, 1373, 166, 837, 1229]. robots [744, 1227, 1902, 1779, 118, 2393, 252, 717, 187]. Robust [2234, 1253, 1470, 1970, 189, 384]. Robustness [2006, 924, 596]. rocking [913]. Role [869, 116, 281]. Role-Based [869]. roommates [2211, 939, 1134]. Root [1430, 1598]. rooted [390, 229, 312]. roots [811, 309, 544]. Rosser [1339, 49]. rotation [1948, 226, 2110]. Rotations [131, 891, 23, 471]. Rough [844, 1683, 846, 1469, 848, 849, 852, 1511]. round [1808]. rounds [1951]. router [2122]. routing [85, 1903, 371, 1250, 1484, 1972, 1251, 1963, 640, 2036, 1734, 808, 698, 367, 1776, 1224, 335, 1888]. routings [2041]. row [734, 144]. RSA [27, 2098, 750]. Rudin [752]. Rule [1384, 616, 326, 1328, 274, 1850, 1127]. rule-based [326, 1127]. Rules [340, 356, 444, 1983, 1268, 2118]. Rumor [575]. run [1133, 1329, 1136, 1801, 334]. run-length [1133, 1329, 1136, 334]. Running [1674, 1353]. runs [604, 2030]. Runtime [491, 162, 2179, 2254]. Russian [1745].

Safe [1967, 632, 1072]. Safety

[1722, 707, 2447, 2196]. Sahlqvist [1042]. Salesman [2134, 2220]. same [2239]. sample [2113, 2009]. samplers [1815]. samples [2411]. Sampling [1290, 1210, 985, 1828, 2411, 1938, 201, 2336]. Sand [1170, 1851, 43]. sandpile [2148, 685]. sat [1872, 2055, 1022, 2441, 1162, 341, 2116]. satisfaction [1203, 1522, 2084, 1423]. Satisfiability [1615, 904, 1295, 1861, 1420, 1806]. satisfy [2083]. saturated [2181]. Sauer [2408]. scalability [943, 1055]. scalable [1945, 189]. scalar [2284]. scale [2323, 1682, 436]. scanner [701]. scattered [547, 1900, 57, 2152]. scenarios [1084, 1087, 1573, 1196]. scheduler [820, 2368]. schedulers [2191, 1476, 1477, 1650]. Schedules [2365, 48]. Scheduling [1508, 376, 1432, 1625, 1214, 964, 659, 1150, 2346, 2375, 962, 1891, 2208, 225, 665, 1370, 874, 2236, 1702, 414, 1911, 909, 801, 277, 1029, 280, 450, 1797, 315, 449, 408, 912, 1917, 285, 1764, 2237, 2243, 1154, 791, 2037, 375, 792, 1187, 1239, 1769, 2235, 1691, 2033, 2122, 1573, 1389, 705, 279, 954, 982, 2311, 372, 549, 1999, 1404, 453, 1677, 1382, 141, 1490]. schemas [733]. schemata [91]. scheme [1251, 2135, 428, 884, 918, 282, 389]. schemes [1028, 1908, 830, 2167, 2418, 102, 1564, 945, 1224, 424, 992, 416, 1001, 1490]. Schnyder [1734]. Schröder [1826]. Schützenberger [454]. Sci [1040, 1557, 1678, 2457, 671, 2461, 2459, 2292, 951, 2456, 2460, 1073, 864, 1706, 1712]. Science [1836, 1337]. SCJ [2336]. score [2081]. scores [1454]. Scott [2266, 1043]. screening [368]. SDP [1503]. SDP-based [1503]. seals [104]. search [704, 2139, 1900, 1443, 1429, 1636, 472, 2071, 402, 2256, 835, 446, 1154, 1880, 305, 1622, 770, 2019, 1054, 365, 1174, 1917]. Searchable [788]. Searching [583, 1435, 1256, 2207, 1329, 2070, 925, 831, 1742, 2123, 746, 459, 1883]. Second [2216, 775, 2334, 2461, 35, 543]. second-order [775]. Second-Price [2216]. secondary [186, 998, 2218, 1999]. secret [1908, 1237, 1001]. Secure [1492, 2430, 824, 830, 726, 1429, 967, 150, 2203, 1584, 1701, 282]. Security [707, 689, 2223, 20]. seed [1006, 1644]. seeds [2045, 1470]. seeking [1373, 2010]. segment [949]. segmentation [2115, 475, 468, 474]. segments [827, 466, 687, 2005, 1763]. Select [132]. Selection [2356, 323, 572, 1785, 1057, 73, 2163, 1020, 1277, 401, 244, 1071, 79, 1786, 399, 76, 1795]. selective [2215]. Self [820, 1937, 1070, 359, 1324, 709, 1933, 1225, 1940, 1068, 1790, 2140, 358, 1469, 824, 2337, 1378, 716, 1072, 818, 117, 488, 119, 718, 819, 1626, 1500]. self-affine [1500]. Self-assembly [359, 1790, 358, 488]. self-awareness [1940]. self-destructible [488]. self-index [2140]. Self-Localization [1324]. self-organizing [1469]. self-similar [1626]. self-stabilization [1378]. Self-stabilizing [820, 1937, 1070, 709, 1933, 1225, 1068, 824, 2337, 716, 1072, 818, 117, 119, 718, 819]. Selfish [1744, 938, 964, 543, 1250, 1876, 189, 9, 1263, 1311, 18]. selfish-computer [189]. selfishness [1619, 1876]. sell [2414]. Semantic [1987, 532, 515, 328]. Semantics [1604, 2180, 440, 261, 108, 1128, 1967, 1353, 867, 1161, 2455, 1709, 585, 779, 780, 148, 588, 486, 2278, 2322, 975]. Semi [1214, 665, 1691, 2033, 1370, 1203, 976, 1779, 450, 2132, 1888, 1603]. semi-dynamical [1203]. semi-linear [976]. semi-metric [1888]. Semi-online [1214, 665, 1691, 2033, 1370, 450, 2132].

Semi-projective [1603].

semi-synchronous [1779]. Semigroup [81]. semigroups [425, 90, 415, 2402, 2119, 474]. semilinear [1798]. semiring [2161, 1015]. semirings [2158]. Sensing [467]. sensitive [2187]. sensitivity [839, 1487, 311, 646]. Sensor [2340, 2341, 1315, 1444, 1781, 996, 1387, 726, 1251, 1875, 959, 824, 2337, 716, 1257, 1776, 1773, 1994, 927, 1521]. sentences [1574]. separable [2266]. separated [1842]. Separation [714, 2181, 1418]. sequence [2387, 2080, 1827, 1485, 1426, 308, 2331, 133, 72, 1154, 33, 1187, 970, 2105, 1037]. sequence-based [33]. sequence-dependent [1154]. sequences [541, 202, 827, 949, 1128, 562, 313, 751, 2140, 2331, 1237, 638, 1640, 968, 284, 470, 2229, 1591, 424, 1471, 2385, 857]. sequent [1542]. sequential [620, 1793]. sequentialization [519]. Series [547, 423, 2158, 330, 970, 365]. Series-parallel [547]. servers [406, 860]. service [1150, 2269, 1831, 1691, 914, 1755]. service-oriented [1831]. services [260]. sessions [2187]. Set [871, 1191, 2287, 1421, 323, 1367, 1503, 1936, 1358, 50, 711, 2396, 466, 2452, 2163, 298, 932, 1251, 2376, 87, 1035, 834, 181, 190, 315, 1072, 1194, 1979, 446, 18, 567, 324, 1916, 1849, 10, 970, 852, 889, 767, 971, 2165, 1511, 2371, 1733, 502, 2031, 1498, 1928, 1886, 6, 569, 517, 126, 1978, 2445]. set-based [1358]. setoids [2275]. sets [2379, 337, 1547, 50, 539, 1483, 806, 1857, 1069, 1676, 845, 1790, 976, 1683, 1299, 997, 310, 1265, 2193, 2284, 1798, 272, 880, 967, 1878, 702, 385, 1710, 2388, 1576, 1489, 1412, 844, 81, 814, 2064, 654, 374, 1498, 1500, 366, 2213]. setting [2279]. setups [1154]. seven [1991]. Several [1370, 1954]. SFTs [1863]. shallow [341]. Shape [1210, 2378, 1851]. shaped [1532, 83, 1236]. Shapiro [752]. shared [1896]. Sharing [2017, 47, 1729, 1237, 113, 1001]. Sharp [291]. Sheng [1285]. shifts [1393, 1870]. shop [1702, 912, 982, 2062]. shoreline [1256]. Short [1564, 1006, 1590, 1902, 444, 1193, 1737, 1459, 1984]. short-seed [1006]. shortcuts [1766]. shorter [49]. Shortest [387, 257, 790, 2439, 121, 1906, 82, 1972, 1512, 2043, 2017, 236]. shortest-path [1972]. shot [1658]. shrinking [433]. shuffle [1331, 124, 2029, 2041]. shuffle-exchange [124, 2041]. Shuffled [1687]. sided [2087, 1983]. sieve [2281]. sign [748, 1260]. signal [1316, 860, 351]. signaling [1664]. signalling [1360]. signals [562, 854, 1324]. signature [1492, 2167, 1564, 428, 2297, 1424]. signatures [56, 472, 2098, 460]. signcryption [884, 282]. signed [2055]. signer [428]. significant [497]. silent [716]. similar [2140, 2123, 2287, 1626]. Similarity [1782]. Simple [1330, 1630, 1600, 76, 101, 214, 395, 1040, 1227, 55, 355, 1940, 354, 181, 1734, 2142, 204, 2383, 2237, 2063, 1489, 279, 352, 260, 1498, 1677, 355, 2254]. Simplicity [361, 1505]. Simplification [860]. Simulating [678, 862]. Simulation [2114, 1652, 955, 1124, 24, 2180, 2454, 990, 1129]. simulations [172]. simultaneity [1121]. simultaneous [1008, 2242]. simultaneously [2385]. sine [1590]. Single [1582, 1573, 1018, 236, 141, 1150, 2236, 1618, 2032, 376, 1138, 1321, 1154, 375, 1187, 1239, 92, 279, 1805]. single-hop [1321]. Single-machine [141, 2236, 1187]. singular [1598, 1599]. singularity [1595]. SINR [1316, 2346]. sixtieth [614]. Size [2328, 138, 371, 426, 1028, 1492, 2355, 425, 1274, 1171, 2203, 1348, 2255, 1156]. Size-constrained [371]. size-cost [1274]. sizes [665, 1483]. skeleton [392]. skew [551]. skip [1068, 1941]. Skolem [283]. slicing [108]. sliding [2358, 676]. slim [2393]. slow [274]. slow-to-stop [274]. small [1673, 267, 671, 811, 1161, 537, 1963, 144, 2336, 478, 1997, 1627, 644, 351]. small-step [1161]. small-world [1963].

Smaller [1288, 557, 840]. Smallest [205]. Smart [505]. smooth [879]. SMT [1162]. snap [712, 1754]. snap-stabilizing [712, 1754]. snapshot [161]. snowflake [2106]. Sobriety [2271]. Social [1958, 2370, 213, 683, 575, 2163, 1963, 2313, 2384, 374]. sofic [1870]. software [1354, 1385, 1352]. Solomon [2023]. solution [77, 327, 1697, 1587, 2104, 2116]. Solutions [2211, 618, 1927, 65, 1470, 179, 1599, 2336, 2338]. Solvability [448]. solvable [377]. solve [1480, 1266, 2324]. solved [1532]. solver [341]. Solving [2121, 1753, 69, 1872, 2447, 1203, 1708, 2191, 1162, 2250, 748]. somatic [400]. Some [1057, 1446, 836, 62, 1544, 710, 561, 2452, 2028, 2182, 444, 1518, 2215, 627, 242, 336, 2143, 727, 609, 98, 465, 2291]. sort [1024, 1917]. sortable [419]. sorted [2452]. sorters [542]. Sorting [572, 1705, 1394, 1952, 1034, 430, 1193, 1062]. SOS [616]. sound [690]. source [369, 1373, 1591, 236]. sources [1367]. Space [2117, 1635, 2337, 1623, 1586, 1223, 327, 1473, 1630, 764, 1122, 893, 712, 820, 65, 309, 1093, 2403, 2229, 1881, 1986, 1409, 1501, 2362]. Space-efficient [2337, 1623]. space/time [2229]. spaces [2150, 2113, 406, 1395, 2271, 676, 1057, 368, 1376, 850, 1888, 4, 1712]. spanner [1320]. spanners [432, 2283]. Spanning [876, 1781, 1186, 560, 1248, 1842, 1869, 2442, 273, 384, 538, 183, 2086, 1463, 1000]. Sparse [477, 1927, 1591, 2092, 1020, 432, 1524, 309, 275, 1197, 885, 2291, 1054, 2338]. Spatial [857, 1203, 2240, 2114]. Spatially [1131]. Special [1351, 655, 583, 2206, 892, 45, 2215, 2446, 1090, 2456, 1047, 1435]. Specialization [85]. specification [2157, 33]. Specifications [917, 1355, 623, 1146, 974, 596, 1966, 1989]. Specifying [1832, 1542]. spectrum [2190]. speed [280, 1689]. speed-up [1689].

Speeding [2078]. Speedup [393]. speedups [2326]. Sperner [622]. sphere [1008, 1892]. spheres [1051]. spies [1438]. spike [1793]. Spiking [71, 2118, 1793]. Spin [1416]. Splicing [1168, 67, 329, 184]. Split [1867]. Splitting [496, 945, 799, 231]. spread [2313]. spreading [575]. spy [1438, 1438]. spy-bad [1438]. Spy-good [1438]. squad [2286, 2424]. Square [1088, 465, 811, 1618, 1341, 1791, 2249, 2363, 1819, 2362]. Square-free [1088, 1618]. square-orthogonal [2363]. Squares [2124, 667, 1066, 2, 2292, 2286, 1889, 219, 334]. SSS [707]. Stability [165, 924, 2409, 138, 43, 1071, 2419]. stability-based [1071]. Stabilization [821, 1378, 707]. stabilizing [708, 709, 116, 1933, 1225, 1843, 1068, 712, 1754, 713, 820, 1937, 1070, 715, 824, 2337, 716, 1072, 818, 117, 1941, 119, 1231, 718, 819]. stable [1134, 2211, 131, 1178, 2065, 939, 1437]. stack [1818]. stacks [419]. stage [2062]. stamp [1762]. standard [832, 2238, 793, 687, 460]. Star [1339, 1545, 241, 1974, 1085, 1791, 1293, 1236, 929, 501]. Star-free [1339, 1293]. star-shaped [1236]. starting [883]. State [1180, 1909, 1085, 1341, 1791, 1281, 1105, 1794, 1529, 2400, 161, 1720, 828, 1295, 1750, 1387, 821, 1288, 41, 1293, 133, 1093, 1280, 1116, 1591, 1155, 1986]. state-based [1387]. state-space [1986]. stateless [42]. states [1113, 529, 175, 350, 349, 1666, 1794]. Static [744]. Stationary [13]. statistical [1071, 533]. statistics [1252, 2094]. status [965]. steganography [1859]. Steiner [1191, 1856, 1843, 979, 970]. step [1161, 2250, 2412]. Stepped [2107]. stigmergetic [164]. STM [1232]. Stochastic [1129, 684, 683, 1959, 1128, 156, 1720, 1008, 1746, 533, 1982, 2006]. stochastics [1126]. Stone [1115]. stop [274]. storage [1935, 2430, 1408]. storeless [1161]. stories [1367]. strata [2411]. strategic [950, 1612]. strategically [1140]. strategically-zero-sum [1140]. Strategies [473, 487, 1902, 1720, 1443, 2202, 1561, 1025, 402]. Strategy [1774, 39, 2448, 745, 2378, 1523, 1609]. Strategy-proof [1774, 1523]. stratified [2411]. stream [58, 262, 320]. Streaming [1726, 1881, 237, 2459, 568, 733]. streams [1169, 2334, 2461, 635]. Strict [1866, 1661, 108, 1289]. strictly [2146]. String [1216, 648, 794, 2070, 1641, 1188, 2079, 2359, 1785, 1630, 1637, 1136, 2121, 1743, 1871, 2230, 2071, 652, 1062, 882, 1182, 863, 1609, 1413]. string-matching [1637]. string-regulated [863]. strings [1133, 1329, 2314, 55, 539, 1021, 401, 978, 1488, 334, 1576]. strip [1202, 408, 668, 370]. Strong [1396, 1699, 991, 297, 35, 516, 886, 2022, 2270, 1070, 1255, 445]. Stronger [2223, 1156]. strongly [50, 2450, 1276, 87]. structural [2360, 2048, 1605, 1280, 779, 2183]. structure [5, 1557, 1481, 1401, 2030, 988, 2320, 557, 186, 1978, 1184, 2275, 1835, 703, 1851, 1634, 2420, 1548]. structured [1921, 623, 191, 1376, 2122, 1989]. Structures [1810, 2268, 1829, 1721, 1846, 589, 1890, 925, 921, 276, 1770, 1252, 769, 997, 2094, 2451, 788, 1092, 19, 499, 2263, 359, 436, 270, 2218, 149, 518]. structuring [1146]. stuck [648]. study [1529, 2346, 772, 1820, 1480, 1161, 1764, 1050]. Sturm [605]. Sturmian [93, 638, 1141, 1345, 1841, 1104, 1074]. style [2304, 1164]. sub [1186, 468, 2412, 2213]. sub- [2213]. sub-logarithmic [2412]. sub-pixel [468]. subclass [760]. subclasses [1550, 1978]. subdivision [1215, 728]. subgraph [1960, 1657, 1947, 1252, 217, 741, 2089, 1036, 2233, 2169, 1783]. subgraphs [2312, 498, 2042, 1406, 916]. subject [449]. subjective [1681]. subjective-information [1681]. Sublinear [2298, 2440, 697]. Sublinear-time [2298]. submodular [1567]. subpermutations [650]. subregular [1277]. subset [1922, 1020, 605, 1364, 1795]. substitutability [510]. substitutions [2239, 1343]. substring [2123, 2072]. substrings [2249]. substructural [688]. substructure [1258]. subsums [2080]. subtraction [1025, 422, 1030]. subtree [83]. subtree-shaped [83]. subtrees [1504]. Subword [833, 1142, 1135, 1486, 1017, 1002]. subwords [2320]. Succinct [580, 1946, 788, 1189, 2146]. Succinctness [232, 37, 1794]. such [887, 1626]. Sufficient [2398, 340]. suffix [1369, 1275, 1909, 655, 607]. suffix- [1275]. suffix-free [1909]. Sum [751, 2209, 794, 994, 1140, 79, 2219]. sum-max [2219]. Sum-Radii [2209]. Summarizing [970]. sums [1808, 811, 79, 658, 788, 474]. sup [1475]. sup-interpretations [1475]. super [1843, 939, 2020, 2218]. super-secondary [2218]. super-stabilizing [1843]. super-stable [939]. supergraph [741]. superimposition [388]. Supermartingales [208]. superstring [2043]. Supervised [2007]. supplier [411]. supply [954]. support [1594, 2158]. suppressions [213]. surely [2329]. surface [1257]. surface-level [1257]. surfaces [2258, 2107, 1594, 481, 1649]. surfers [2083]. surjective [1374]. survey [2434, 359, 1475, 781]. Survivable [1964]. surviving [980, 286, 438, 1380, 1993]. SVD [1020]. SVR [849]. swap [2395, 300]. swaps [132]. Swarm [161, 163, 162]. swarms [166]. sweep [1093]. sweep-line [1093]. Switching [1530, 1722, 1078, 1311]. Sylvester [480]. symbol [1411]. Symbolic [103, 1696, 1248, 2455, 1592, 2194, 588, 1388, 677]. symbolic-numeric [1696, 1592]. Symmetric [2284, 1728, 1821, 56, 2148, 219, 1862, 426, 2183]. symmetries [417, 252, 1995]. Symmetry [1898]. synapses [2118]. sync [1095, 1344]. synchronisation [153]. synchronization [461, 1317, 2156, 959, 824, 2286, 2424]. synchronized [313, 1278]. synchronizer [1897]. synchronizing [1778]. Synchronous [823, 835, 1900, 1751, 1779, 10, 297, 767, 119]. synchrony [1749]. Syntactic [1275]. Synthesis [533, 1475, 663, 1722, 2341, 1970, 489]. Synthesizing [1790]. system [111, 2114, 1160, 1385, 1551, 530, 1768, 1948, 1709, 33, 1121, 1055, 1381, 2368]. systematic [2352]. SystemC [1752]. Systems [2340, 707, 2186, 1354, 871, 487, 357, 2117, 261, 957, 617, 1935, 154, 1130, 1125, 1005, 683, 1722, 1168, 2305, 50, 329, 711, 2050, 1716, 1416, 152, 355, 1119, 1337, 356, 1927, 713, 354, 1682, 2320, 1751, 1868, 189, 639, 1127, 310, 1083, 832, 2238, 1496, 2304, 287, 1513, 715, 415, 309, 1390, 41, 1089, 1143, 1232, 1942, 533, 1793, 1243, 1939, 1095, 1182, 2225, 148, 633, 2276, 1970, 2298, 1599, 1786, 926, 544, 2324, 431, 903, 1298, 1542, 184, 678, 1244, 1453, 1101, 49, 71, 2322, 392, 1120, 1103, 1756, 1363, 677, 68, 488, 2001, 1471]. systems [805, 853, 1349, 781, 2308, 2118, 2116, 1685, 1478, 2420, 1365, 1500, 2338, 1477, 1650, 155, 107]. Szilard [2182].

T [310, 1437]. table [1306]. tableaux [1825]. tables [455]. Tackling [1986]. tag [354, 678]. tail [1174]. Takesue [307]. Takesue-type [307]. Taking [252]. talk [1360]. Tamari [1011]. TAMC [493]. tamper [104]. tamper-evident [104]. tandem [2077, 2078]. tape [1278, 1411, 8]. tardiness [1154]. Target [1994, 323, 2207, 2163]. targets [1367]. task [1876]. tasks [964]. Taylor [20]. TCP [1427, 218]. TCTL [2178].

TCTL-preserving [2178]. TDoA [2343]. teaching [2408]. team [1077]. technique [2327, 1313]. techniques [1124, 537, 1258]. Teleportation [1116]. Telling [1367]. Temperature [1891, 358]. Temporal [2364, 1709, 689, 2450, 778, 530, 1768, 525, 1833, 176, 596, 340, 501]. temporary [215]. tension [1266]. term [261, 1449, 1453]. terminal [81]. termination [1451, 625, 1513, 1832]. terms [1828, 624, 1541, 900]. ternary [1618, 1700, 2061]. terrains [928]. test [500]. testable [969]. Testing [650, 885, 1016, 1157, 1354, 958, 2088, 1765, 1857, 1927, 2188, 1948, 1300, 133, 2063]. text [2073, 1632]. texts [46]. textures [1683]. their [1602, 1538, 2182, 605, 1360, 241, 1304, 1468, 880, 1658, 687, 1574, 2119, 1766]. theme [20]. Theor [951]. theorem [1042, 32, 1334, 2095, 454, 2352, 1375, 1165, 2279, 1259, 2166, 1579, 1345]. theorems [1112, 738, 1661, 391]. Theoret [1040, 1557, 1678, 2457, 671, 2461, 2459, 2292, 2456, 2460, 1073, 864, 1706, 1712]. theoretic [327, 1716, 1080, 1968, 2048, 1311, 2275, 456, 352, 1106, 12]. Theoretical [1351, 2177, 1836, 195, 1047, 1082, 771, 2307, 1052]. Theories [687, 619, 103, 2274]. Theory [493, 2437, 2199, 583, 1435, 2206, 1855, 8, 946, 1130, 161, 1352, 2452, 777, 2095, 2309, 1603, 2009, 2272, 1604, 921, 679, 699, 1831, 771, 2274, 978, 973, 1917, 2237, 147, 1155, 1511, 2018, 1806, 517, 146]. ‘thermodynamic’ [1372]. theta [452]. theta-palindromic [452]. thin [756]. Three [14, 2422, 1627, 1066, 2285, 670, 377, 379, 469, 1064, 1514]. three-dimensional [379, 469]. Three-matching [1627]. three-squares [1066]. Threshold [1001, 1483, 2167, 606, 2418, 463, 1430, 1584, 314, 2420]. threshold-based [1483]. thresholding [474]. thresholds [1022, 291]. throughput [693]. Thue [2320, 2331, 49]. Tiara [1068]. tie [1457]. Tight [1934, 199, 1657, 1900, 1907, 577, 732, 2039, 1612]. Tighter [1527]. Tightness [494]. tilde [267, 671]. tile [1790, 758]. tiles [2285, 966, 87, 1819]. tiling [1551]. Tilings [1570, 1821, 759, 1138, 862, 379]. Time [2116, 380, 1725, 1586, 38, 1537, 1904, 985, 710, 37, 1630, 764, 2087, 1033, 987, 1355, 555, 1508, 697, 2236, 1731, 333, 174, 369, 65, 1251, 1948, 1674, 401, 577, 2190, 277, 1495, 1918, 1266, 696, 2123, 2193, 960, 808, 1797, 449, 1255, 912, 1737, 2250, 1243, 913, 638, 2286, 1833, 61, 419, 1970, 2011, 2298, 142, 377, 2033, 289, 1298, 1816, 1240, 2229, 1312, 231, 970, 982, 2231, 1476, 1814, 2444, 2311, 1051, 8, 1107, 1461, 635, 1805, 1999, 502, 1478, 2343, 2306, 365, 918, 1677, 1877, 1174, 515]. time-abstract [1476]. time-convexity [1478]. time-determinism [1243]. Time-free [2116]. time-varying [1495]. Timed [2443, 339, 2178, 2189, 454, 1968, 903]. timed-arc [2178]. times [1150, 1934, 1899, 801, 1187, 143]. TMTO [1803]. Toggling [442]. tokens [1519, 714]. tolerance [708, 1750, 1752, 2325, 1407, 1024, 2153, 1229]. tolerant [1639, 2234, 1484, 1974, 640, 979, 1552, 367, 889, 6]. tool [525]. toolpaths [1249]. top [1509, 2229]. top- [2229, 1509]. topic [1090, 2456]. Topological [1957, 1936, 1378, 1, 1111, 1649, 146]. topologies [1754, 227]. Topology [2179, 1137, 676]. topos [1112]. Topswops [317]. tori [295]. torus [2300, 554, 379, 2231, 2022]. tossing [2429]. Total [1560, 2092, 1991, 2044, 2024, 277, 1801, 449, 912, 1999, 1677, 1877, 2394]. tour [873, 607]. Touring [2387]. tournament [382]. tournaments [741, 569]. tours [1842]. trace [1336]. traceable [2026, 460]. Traced [681]. traces [1959, 778]. Tracing [1455]. track [2369]. Tractability [1202, 1927, 1929]. tractable [932, 1522]. Tradeoffs [745, 20, 1938, 2229, 426]. trading [65]. traffic [274, 947, 1292]. training [2007, 401, 2008]. trajectories [421]. trajectory [108, 588]. trajectory-based [108]. transaction [1757, 1752]. transaction-based [1757]. transactional [699, 1756]. transactions [1353, 1756]. transcendence [1977]. Transducers [134, 1602, 1091]. transduction [860]. transductions [46]. Transfer [1101, 895]. transform [1433, 612, 1269, 2136]. transformability [1096]. transformation [1778, 2130, 908, 308, 1044]. transformations [331, 661, 1084, 1377, 473, 670, 2449]. transformer [709, 585]. transformers [346]. transience [455]. transition [1390, 633]. transitions [2400, 311, 210]. transitive [1933, 338, 839, 340]. Translating [128]. translation [621, 31, 758]. translations [2178, 1709, 385]. translocations [1705]. transmembrane [2218]. transmission [1749]. transposition [451, 1021]. transversal [644]. transversals [1678, 1493]. trapdoor [815]. trapezoidal [1481]. traps [2034]. traveler [1669]. Traveling [2220, 65, 382, 2134]. Traveller [2127]. traversal [1228]. Treapsort [1674]. Tree [1393, 1579, 2048, 2029, 383, 1856, 1602, 1150, 560, 1248, 661, 371, 1870, 83, 1754, 1827, 369, 2283, 1137, 626, 997, 273, 423, 271, 186, 217, 384, 1658, 2249, 1954, 183, 305, 1348, 1105, 2200, 838, 1283, 2025, 2385, 1000, 1992, 1673, 1191, 876, 118, 1291]. tree-breadth [2283]. Tree-shifts [1393]. Tree-Walking [1291]. tree-width [217]. trees [383, 1504, 1781, 1586, 1602, 907, 1843, 642, 1369, 249, 925, 1400, 1692, 1842, 1937, 831, 1443, 2156, 580, 1656, 1059, 93, 1011, 1374, 254, 2159, 390, 1954, 538, 2086, 239, 1622, 1165, 2366, 723, 1463, 312, 1499, 1212, 1394]. Treewidth [289, 1078, 2135, 2367, 565]. trends [154]. triangle [2379, 1033, 2125, 1185]. triangle-free [1033, 2125]. triangles [2092, 2334, 2461]. triangular [229]. triangulation [76, 669]. triangulations [653]. triconnected [319]. trie [2094]. tries [978]. trimming [1499]. triple [1417]. triplet [1736]. trivalent [1812, 229]. trivial [949]. tropical [1015]. truncated [2323]. Trust [1938, 2263, 1101, 810]. Trust-aware [1938]. truth [1306]. truth-table [1306]. Truthful [1996, 964, 1694]. truthfulness [412]. TSP [89, 221, 1582, 196]. Turing [2117, 2437, 361, 351, 1863, 990, 1802, 1, 233, 1411, 678, 362, 8]. turn [1821]. twins [1015]. Twisted [1727, 1113, 753, 295, 554, 538, 553]. Two [2236, 506, 2167, 1869, 1373, 226, 999, 758, 465, 88, 89, 1799, 406, 2057, 2457, 2452, 225, 665, 1370, 1560, 2391, 1179, 298, 1180, 1702, 414, 753, 1091, 2131, 2357, 1281, 1775, 1691, 2433, 2033, 1282, 1984, 916, 1103, 1881, 2105, 1413, 1016, 1675, 2145, 1461, 2062, 2000, 2132, 2024, 1423, 416, 1382, 1490, 1794, 253, 1733]. Two-agent [2236, 1490]. Two-dimensional [226]. two-disjoint [2131]. Two-group [88]. two-layer [2357, 1733]. two-machine [1702]. two-patterns [298]. Two-robot [1373]. two-sorted [2452]. two-stage [2062]. two-valued [2105]. two-variable [1799]. two-way [1179, 1091, 1281, 2433, 1282, 1794]. typability [627, 868]. Type [868, 494, 1393, 1385, 777, 307, 1711, 2272, 2042, 169, 2304, 1595, 771, 2100, 1701, 1514, 2014]. type-based [1701]. type-checking [1711]. type-free [2304]. type-theoretical [771]. Typed [1128, 149, 171]. Types [1987, 871, 1289, 345, 2317].

U [1532]. UDGs [747]. UML [1750, 30]. Unambiguous [1585, 1264]. unanimity [1571]. unary [1281, 499]. unate [426]. unbiased [2260]. unbordered [2046].

Unbounded [740, 1062, 549, 78, 1097]. Unbounded-error [740]. uncapacitated [2234]. uncertain [1005]. uncertainty [1253, 525]. Uncountable [509]. undecidability [86, 1015]. undecidable [2452, 1002]. undeniable [1564]. underlying [1545, 2342]. underpinnings [195]. understand [182]. underwater [1257]. undirected [1962, 1228, 1621]. unfolding [1273, 2327]. Unguardedness [618]. unhelpful [1171]. uni [246]. uni-dimensional [246]. Uniform [427, 446, 1778, 1316, 1395, 225, 2258, 1899, 2417, 276, 1850, 2422, 833, 450, 973, 1691, 1585, 915, 1823, 1229]. unifying [921, 2241]. Union [1040, 103, 1085, 1341, 1791, 395]. Union-Find-Delete [1040]. unions [78]. Unique [1549, 1563, 1847, 1612, 2247, 99, 205]. unison [639]. unit [616, 167, 498, 1838, 2096, 2247, 1072, 2243, 2033, 1000, 366, 2213]. unit-length [2243]. unital [2436]. unitary [368]. units [1877]. univariate [1598, 1597]. Universal [539, 1094, 2329, 2010, 361, 1757, 1172, 351, 1950, 350, 1853, 546, 1534, 1807, 1806]. Universality [951, 1793, 348]. universe [1099]. unknown [508, 1519, 2015]. unlabeled [312]. unordered [383, 1504, 2046, 722]. unranked [1348]. unsatisfiable [1710]. unsharp [1155]. Untangled [704]. untyped [1541, 343]. unusual [516]. unweighted [446]. update [1240]. update-time [1240]. updates [2415]. Updating [965, 2418, 2424]. upper [809, 222, 293, 483, 551, 1486, 834, 743, 2233]. uptimals [948]. Upward [2088, 2248]. use [1317, 893, 1174]. useful [1049]. uselessness [942]. users [1505, 1311]. Using [2434, 1296, 169, 1649, 683, 467, 1290, 267, 671, 1708, 860, 1642, 2136, 2455, 2349, 1519, 908, 1594, 1252, 2294, 1044, 472, 29, 2151, 533, 1802, 2342, 69, 2433, 669, 1389, 1244, 209, 2086, 1379, 2116, 1364, 332, 372, 471, 1996, 1365]. usually [1099]. utilitarian [412]. utilities [1580].

V [514]. Validating [2449]. validation [599]. validity [1071]. valuations [596]. Value [1147, 1152, 1304, 2435]. Value-passing [1147]. valued [1289, 620, 1603, 995, 1031, 2105]. values [557, 532]. variability [2227]. Variable [1426, 1799, 1216, 462, 693]. variables [1340, 2277, 205, 1409]. variance [2094]. variant [2136]. variants [1669, 400, 7]. variate [478]. Variations [20]. Varieties [423, 1600]. various [1812, 1013, 1576]. varying [1495]. Vector [1077, 2439]. vectors [2150]. vehicle [1150]. verifiability [428]. Verifiable [1237]. Verification [2050, 1714, 628, 154, 169, 2179, 2455, 341, 1244, 1996, 2306]. Verified [1599, 263]. verify [598]. Verifying [931, 270]. Versatile [1743, 1232]. versions [465]. versus [906, 1505, 138, 565, 2238, 677, 2435]. Vertex [1973, 1421, 2374, 2396, 1924, 1922, 1730, 200, 293, 731, 1610, 1305, 1023, 565, 2407, 1976, 1374, 1060, 1930, 1878, 934, 2388, 1872, 2164, 2165, 937, 718, 1404, 1054]. vertex-connection [731, 1610]. vertex-labeled [1060]. vertex-surjective [1374]. vertices [1067, 312, 1058]. via [2113, 2181, 2204, 2139, 361, 1439, 508, 1204, 1020, 2341, 625, 481, 586, 172, 311, 991, 1433, 2447, 2449, 1156, 1888, 101]. Video [664, 2348]. Vidyasagar [1528]. view [1853, 516, 436, 2241, 2184]. views [443]. Virtual [1319, 1232, 1759, 1875]. visibility [1227, 1254]. visibly [797]. vision [1901]. visual [1908, 2141, 852]. voltage [693]. volume [878]. Voronoi [2141, 1763]. voter [1786]. vs [2016, 2390, 2191, 10]. Vulnerability [2020].

W [320]. W-Stream [320]. w.r.t [1725]. walk [1195]. Walking [1291]. walks [1590, 1899, 577, 16, 143]. Wang [186, 2285]. wave [1851]. wavelength [2061]. wavelet [1059]. way [590, 1179, 1091, 1281, 1009, 2433, 1282, 24, 1794]. ways [758]. WDM [745]. Weak [2174, 339, 1571, 2180, 1725, 1974, 1636, 721, 1797, 1321, 629]. weak-pancyclicity [1974]. weakest [711, 1942]. Weakly [1264, 1799, 713, 1665]. Web [2083, 914, 260]. Weight [1978, 1546, 1960, 432, 877, 424, 819, 2169]. Weighted [995, 1654, 2047, 2412, 1606, 985, 507, 1508, 564, 83, 1400, 2161, 831, 1443, 454, 2156, 126, 1340, 2157, 385, 633, 2276, 1824, 46, 2160, 231, 437, 1142, 1499, 1423, 1259, 918, 1877, 366]. weighted-degree [437]. weights [1553]. well [494, 769]. well-founded [769]. well-mixed [494]. Weyl [676]. Wheeler [2136, 2069, 1377, 612]. where [1455]. which [1480, 2158, 1984]. white [1908, 2143]. who [1455]. wide [73, 660]. wide-gap [73]. wider [1267]. width [15, 2113, 573, 1078, 1657, 789, 563, 762, 812, 217, 1698, 1893, 2164, 1016, 1437, 1437]. width- [1698]. width-two [1016]. Wiener [1514]. wildcards [2403, 1632]. Window [218, 1029]. Window-games [218]. windows [2358]. Winner [1523, 1951]. Winner-imposing [1523]. Winning [1025, 1507]. Winskel [2267]. Wireless [2340, 1315, 2345, 2348, 1781, 1904, 413, 1317, 726, 1251, 579, 1771, 1875, 959, 1318, 148, 367, 1257, 1776, 1773, 1994, 926, 903, 787, 1323, 927, 1887, 1199, 2108]. withdrawal [909]. within [1427]. without [2361, 1473, 965, 892, 1560, 821, 118, 2284, 1564, 1533, 1459, 1268, 1701, 1380, 1993, 2000, 4, 1712]. Wolfe [467]. word [461, 2201, 2068, 1179, 1204, 2030, 752, 1618, 1608, 1801, 1310, 546, 900, 2106]. words [906, 872, 1820, 1331, 337, 667, 1135, 1217, 1066, 1653, 1547, 2058, 2124, 396, 1481, 601, 1296, 220, 435, 1179, 57, 752, 1618, 45, 605, 1218, 2, 2292, 793, 833, 796, 1088, 1377, 292, 879, 1342, 1737, 896, 998, 1287, 603, 1141, 609, 1345, 1389, 1823, 388, 1841, 1104, 610, 2110, 612, 814, 1074, 2301, 137, 1814, 2316, 2295, 11]. Working [2458]. Workshop [1119]. Workspace [738, 2356]. world [2323, 1962, 1963, 1232]. worlds [1110]. worst [2438, 788, 1784, 199]. worst-case [788]. wrapped [1124]. Wright [32].

X [262]. X-machines [262]. XCSF [1055]. XML [1711, 733, 2449, 797]. XOR [1022]. XOR-SAT [1022].

Yu [1285].

ZebraNet [1934]. Zeckendorf [1204]. Zeno [630]. zero [616, 1140, 936]. Zhang [186]. Zielonka [1835]. zippers [862]. Ziv [2140].

References

Kurka:1997:TDT

[1] Petr Kůrka. On topological dynamics of Turing machines. Theoretical Computer Science, 174(1–2): 203–216, March 15, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                              //www.elsevier.com/cgi-bin/cas/
                                              tree/store/tcs/cas_sub/browse/browse.
                                              cgi?year=1997&volume=174&issue=

1-2&aid=2228. See erratum [233].

Fraenkel:1999:ENS

[2] Aviezri S. Fraenkel and Jamie Simpson. The exact number of squares in Fibonacci words. Theoretical Computer Science, 218(1):95–106, April 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

                                              tronic). URL http://www.elsevier.
                                              com/cas/tree/store/tcs/sub/1999/

218/1/3061.pdf. See corrigendum [2292].

Soloveichik:2005:CPB

[3] David Soloveichik and Erik Winfree. The computational power of Benenson automata. Theoretical Computer Science, 344(2–3):279–297, November 17, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [864].

vanBreugel:2007:RDM

[4] Franck van Breugel, Claudio Hermida, Michael Makkai, and James Worrell. Recursively defined metric spaces without contraction. Theoretical Computer Science, 380(1–2):143–163, June 21, 2007. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See addendum [1712].

Blumensath:2008:SGC

[5] Achim Blumensath. On the structure of graphs in the Caucal hierarchy. Theoretical Computer Science, 400(1–3): 19–45, June 9, 2008. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [1557].

Zhang:2009:ACS

[6] Zhao Zhang, Xiaofeng Gao, and Weili Wu. Algorithms for connected set cover problem and fault-tolerant connected set cover problem. Theoretical Computer Science, 410(8–10):812– 817, March 1, 2009. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See note [889].

Lyaudet:2010:NHL

[7] Laurent Lyaudet. NP-hard and linear variants of hypergraph partitioning. Theoretical Computer Science, 411(1):

10–21, January 1, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tadaki:2010:TOT

[8] Kohtaro Tadaki, Tomoyuki Yamakami, and Jack C. H. Lin. Theory of onetape linear-time Turing machines. Theoretical Computer Science, 411(1):22– 43, January 1, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Epstein:2010:MML

[9] Leah Epstein and Rob van Stee. Maximizing the minimum load for selfish agents. Theoretical Computer Science, 411(1):44–57, January 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mostefaoui:2010:NPV

[10] Achour Mostéfaoui, Michel Raynal, and Corentin Travers. Narrowing power vs efficiency in synchronous set agreement: Relationship, algorithms and lower bound. Theoretical Computer Science, 411(1):58–69, January 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deLuca:2010:GCE

[11] Aldo de Luca and Luca Q. Zamboni. On graphs of central episturmian words. Theoretical Computer Science, 411(1):70–90, January 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stapleton:2010:GTA

[12] Gem Stapleton, John Howse, and Peter Rodgers. A graph theoretic approach to general Euler diagram drawing. Theoretical Computer Science, 411(1):91– 112, January 1, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Muller:2010:SAP

[13] Markus Müller. Stationary algorithmic probability. Theoretical Computer Science, 411(1):113–130, January 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chrobak:2010:TRF

[14] Marek Chrobak and Jiří Sgall. Three results on frequency assignment in linear cellular networks. Theoretical Computer Science, 411(1):131–137, January 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anthony:2010:MWL

[15] Martin Anthony and Joel Ratsaby. Maximal width learning of binary functions. Theoretical Computer Science, 411(1):138–147, January 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Merlini:2010:RPC

[16] D. Merlini and R. Sprugnoli. The relevant prefixes of coloured Motzkin walks: an average case analysis. Theoretical Computer Science, 411(1):148– 163, January 1, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bosse:2010:NAA

[17] Hartwig Bosse, Jaroslaw Byrka, and Evangelos Markakis. New algorithms for approximate Nash equilibria in bimatrix games. Theoretical Computer Science, 411(1):164–173, January 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2010:MDS

[18] Xiang-Yang Li, Zheng Sun, WeiZhao Wang, Xiaowen Chu, ShaoJie Tang, and Ping Xu. Mechanism design for set cover games with selfish element agents. Theoretical Computer Science, 411(1):174–187, January 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Laube:2010:MLA

[19] Ulrich Laube and Markus E. Nebel. Maximum likelihood analysis of algorithms and data structures. Theoretical Computer Science, 411(1):188– 212, January 1, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DArco:2010:VTA

[20] Paolo D’Arco, Alfredo De Santis, Anna Lisa Ferrara, and Barbara Masucci. Variations on a theme by Akl and Taylor: Security and tradeoffs. Theoretical Computer Science, 411(1):213– 227, January 1, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2010:CMC

[21] Yun Liu. Compositions of maximal codes. Theoretical Computer Science, 411(1):228–238, January 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Borodin:2010:PAG

[22] Allan Borodin, Joan Boyar, Kim S. Larsen, and Nazanin Mirmohammadi.

Priority algorithms for graph optimization problems. Theoretical Computer Science, 411(1):239–258, January 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Richeson:2010:IRR

[23] David Richeson, Paul Winkler, and Jim Wiseman. Itineraries of rigid rotations and diffeomorphisms of the circle. Theoretical Computer Science, 411(1):259– 265, January 1, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Terrier:2010:SOW

[24] Véronique Terrier. Simulation of oneway cellular automata by Boolean circuits. Theoretical Computer Science, 411(1):266–276, January 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lin:2010:LCP

[25] Mingen Lin, Yang Yang, and Jinhui Xu. On Lazy Bin Covering and Packing problems. Theoretical Computer Science, 411(1):277–284, January 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hasan:2010:CAD

[26] Md. Kamrul Hasan, Mohammad Kaykobad, Young-Koo Lee, and Sungyoung Lee. A comprehensive analysis of degree based condition for Hamiltonian cycles. Theoretical Computer Science, 411(1):285–287, January 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chmielowiec:2010:FPR

[27] Andrzej Chmielowiec. Fixed points of the RSA encryption algorithm. Theoretical Computer Science, 411(1):288– 292, January 1, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Du:2010:PAM

[28] Ye Du, Rahul Sami, and Yaoyun Shi. Path auctions with multiple edge ownership. Theoretical Computer Science, 411(1):293–300, January 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hochbaum:2010:CEB

[29] Dorit S. Hochbaum and Asaf Levin. Covering the edges of bipartite graphs using K_{2,2} graphs. Theoretical Computer Science, 411(1):v–x, January 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kaneiwa:2010:CCC

[30] Ken Kaneiwa and Ken Satoh. On the complexities of consistency checking for restricted UML class diagrams. Theoretical Computer Science, 411(2):301– 323, January 2, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fujita:2010:CTA

[31] Ken etsu Fujita. CPS-translation as adjoint. Theoretical Computer Science, 411(2):324–340, January 2, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bloom:2010:MWT

[32] S. L. Bloom and Z. Ésik. A Mezei– Wright theorem for categorical algebras. Theoretical Computer Science, 411(2):341–359, January 2, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lin:2010:ASS

[33] Lan Lin, Stacy J. Prowell, and Jesse H. Poore. An axiom system for sequencebased specification. Theoretical Computer Science, 411(2):360–376, January 2, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DalLago:2010:QIC

[34] Ugo Dal Lago, Andrea Masini, and Margherita Zorzi. Quantum implicit computational complexity. Theoretical Computer Science, 411(2):377– 409, January 2, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pagani:2010:SNP

[35] Michele Pagani and Lorenzo Tortora de Falco. Strong normalization property for second order linear logic. Theoretical Computer Science, 411(2):410– 444, January 2, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bodirsky:2010:PAC

[36] Manuel Bodirsky and Hubie Chen. Peek arc consistency. Theoretical Computer Science, 411(2):445–453, January 2, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bozzelli:2010:CSI

[37] Laura Bozzelli and Ruggero Lanotte. Complexity and succinctness issues for linear-time hybrid logics. Theoretical Computer Science, 411(2):454– 469, January 2, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baillot:2010:LLL

[38] Patrick Baillot and Damiano Mazza. Linear logic by levels and bounded time complexity. Theoretical Computer Science, 411(2):470–503, January 2, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alpuente:2010:DSA

[39] M. Alpuente, S. Escobar, B. Gramlich, and S. Lucas. On-demand strategy annotations revisited: an improved on-demand evaluation strategy. Theoretical Computer Science, 411(2):504– 541, January 2, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Romanel:2010:CPB

[40] Alessandro Romanel and Corrado Priami. On the computational power of BlenX. Theoretical Computer Science, 411(2):542–565, January 2, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hierons:2010:CFS

[41] Robert M. Hierons. Canonical finite state machines for distributed systems. Theoretical Computer Science, 411(2): 566–580, January 2, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2010:SMA

[42] Oscar H. Ibarra, Juhani Karhumäki, and Alexander Okhotin. On stateless multihead automata: Hierarchies and the emptiness problem. Theoretical Computer Science, 411(3):581– 593, January 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duong:2010:SSP

[43] Phan Thi Ha Duong and Tran Thi Thu Huong. On the stability of Sand Piles Model. Theoretical Computer Science, 411(3):594–601, January 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kastermans:2010:CNR

[44] Bart Kastermans and Steffen Lempp. Comparing notions of randomness. Theoretical Computer Science, 411(3): 602–616, January 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Czeizler:2010:SCP

[45] Elena Czeizler, Lila Kari, and Shinnosuke Seki. On a special class of primitive words. Theoretical Computer Science, 411(3):617–630, January 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mathissen:2010:DTW

[46] Christian Mathissen. Definable transductions and weighted logics for texts. Theoretical Computer Science, 411(3): 631–659, January 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bilo:2010:WIH

[47] Vittorio Bilò, Angelo Fanelli, Michele Flammini, and Luca Moscardelli. When ignorance helps: Graphical multicast cost sharing games. Theoretical Computer Science, 411(3):660– 671, January 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tian:2010:EOS

[48] Wanyong Tian, Minming Li, and Enhong Chen. Energy optimal schedules for jobs with multiple active intervals. Theoretical Computer Science, 411(3): 672–676, January 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

ODunlaing:2010:SPP

[49] Colm ÓDúnlaing and Natalie Schluter. A shorter proof that palindromes are not a Church–Rosser language, with extensions to almost-confluent and preperfect Thue systems. Theoretical Computer Science, 411(3):677– 690, January 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boley:2010:LCS

[50] Mario Boley, Tamás Horváth, Axel Poigné, and Stefan Wrobel. Listing closed sets of strongly accessible set systems with applications to data mining. Theoretical Computer Science, 411 (3):691–700, January 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2010:EBa

[51] Anonymous. Editorial Board. Theoretical Computer Science, 411(3):i– vi, January 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Csuhaj-Varju:2010:P

[52] Erzsébet Csuhaj-Varjú and Zoltán Ésik. Preface. Theoretical Computer Science, 411(4–5):701, January 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bjorklund:2010:NRD

[53] Henrik Björklund and Thomas Schwentick. On notions of regularity for data languages. Theoretical Computer Science, 411(4–5):702–715, January 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brevier:2010:CIC

[54] Gaëlle Brevier, Romeo Rizzi, and Stéphane Vialette. Complexity issues in color-preserving graph embeddings. Theoretical Computer Science, 411(4– 5):716–729, January 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brijder:2010:ESG

[55] Robert Brijder, Miika Langille, and Ion Petre. Extended strings and graphs for simple gene assembly. Theoretical Computer Science, 411(4–5):730– 738, January 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cai:2010:BSS

[56] Jin-Yi Cai and Pinyan Lu. On blockwise symmetric signatures for matchgates. Theoretical Computer Science, 411(4–5):739–750, January 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Colcombet:2010:FFI

[57] Thomas Colcombet. Factorization forests for infinite words and applications to countable scattered linear orderings. Theoretical Computer Science, 411(4–5):751–764, January 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Endrullis:2010:PSD

[58] Jörg Endrullis, Clemens Grabmayer, Dimitri Hendriks, Ariya Isihara, and Jan Willem Klop. Productivity of stream definitions. Theoretical Computer Science, 411(4–5):765–782, January 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hemaspaandra:2010:CK

[59] Edith Hemaspaandra, Lane A. Hemaspaandra, Till Tantau, and Osamu Watanabe. On the complexity of kings. Theoretical Computer Science, 411(4– 5):783–798, January 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hoffmann:2010:NHM

[60] Michael Hoffmann and Richard M. Thomas. Notions of hyperbolicity in monoids. Theoretical Computer Science, 411(4–5):799–811, January 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kutrib:2010:RTR

[61] Martin Kutrib and Andreas Malcher. Real-time reversible iterative arrays. Theoretical Computer Science, 411(4– 5):812–822, January 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sattler-Klein:2010:SCR

[62] Andrea Sattler-Klein. Some complexity results for prefix Gröbner bases in free monoid rings. Theoretical Computer Science, 411(4–5):823–833, January 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2010:EBb

[63] Anonymous. Editorial Board. Theoretical Computer Science, 411(4–5):i– vi, January 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2010:Pa

[64] G. Rozenberg. Preface. Theoretical Computer Science, 411(6):835–836, February 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dolev:2010:MTB

[65] Shlomi Dolev and Hen Fitoussi. Masking traveling beams: Optical solutions for NP-complete problems, trading space for time. Theoretical Computer Science, 411(6):837–853, February 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Friedrich:2010:PCH

[66] Tobias Friedrich, Nils Hebbinghaus, and Frank Neumann. Plateaus can be harder in multi-objective optimization. Theoretical Computer Science, 411(6): 854–864, February 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonizzoni:2010:CLE

[67] Paola Bonizzoni. Constants and labelequivalence: a decision procedure for reflexive regular splicing languages. Theoretical Computer Science, 411(6): 865–877, February 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Porreca:2010:NCD

[68] Antonio E. Porreca, Giancarlo Mauri, and Claudio Zandron. Non-confluence in divisionless P systems with active membranes. Theoretical Computer Science, 411(6):878–887, February 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2010:SMB

[69] Xingchang Liu, Xiaofan Yang, Shenglin Li, and Yong Ding. Solving the minimum bisection problem using a biologically inspired computational model. Theoretical Computer Science, 411(6): 888–896, February 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brijder:2010:COC

[70] Robert Brijder and Hendrik Jan Hoogeboom. Combining overlap and containment for gene assembly in ciliates. Theoretical Computer Science, 411(6):897–905, February 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pan:2010:SNP

[71] Linqiang Pan and Gheorghe Păun. Spiking neural P systems: an improved normal form. Theoretical Computer Science, 411(6):906–918, February 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Keil:2010:APC

[72] J. Mark Keil, Jing Liu, and Ian Mc-Quillan. Algorithmic properties of ciliate sequence alignment. Theoretical Computer Science, 411(6):919– 925, February 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2010:CSP

[73] Tianshi Chen, Jun He, Guoliang Chen, and Xin Yao. Choosing selection pressure for wide-gap problems. Theoretical Computer Science, 411(6):926– 934, February 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2010:EBc

[74] Anonymous. Editorial Board. Theoretical Computer Science, 411(6):i– vi, February 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pichler:2010:TPF

[75] Reinhard Pichler and Vadim Savenkov. Towards practical feasibility of core computation in data exchange. Theoretical Computer Science, 411(7–9): 935–957, February 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mezzini:2010:SAM

[76] Mauro Mezzini and Marina Moscarini. Simple algorithms for minimal triangulation of a graph and backward selection of a decomposable Markov network. Theoretical Computer Science, 411(7–9):958–966, February 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Asdre:2010:PSF

[77] Katerina Asdre and Stavros D. Nikolopoulos. A polynomial solution to the k-fixed-endpoint path cover problem on proper interval graphs. Theoretical Computer Science, 411(7–9):967– 975, February 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deBrecht:2010:MCC

[78] Matthew de Brecht and Akihiro Yamamoto. Mind change complexity of inferring unbounded unions of restricted pattern languages from positive data. Theoretical Computer Science, 411(7–9):976–985, February 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lin:2010:EAS

[79] Tien-Ching Lin and D. T. Lee. Efficient algorithms for the sum selection problem and k maximum sums problem. Theoretical Computer Science, 411(7–9):986–994, February 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sadasivam:2010:NCO

[80] Sadish Sadasivam and Huaming Zhang. NP-Completeness of s t-orientations for plane graphs. Theoretical Computer Science, 411(7–9):995–1003, February 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Render:2010:SAR

[81] Elaine Render and Mark Kambites. Semigroup automata with rational initial and terminal sets. Theoretical Computer Science, 411(7–9):1004– 1012, February 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cicerone:2010:PDE

[82] Serafino Cicerone, Gianlorenzo D’Angelo, Gabriele Di Stefano, and Daniele Frigioni. Partially dynamic efficient algorithms for distributed shortest paths. Theoretical Computer Science, 411 (7–9):1013–1037, February 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cheng:2010:PNW

[83] Yukun Cheng, Liying Kang, and Changhong Lu. The pos/neg-weighted 1-median problem on tree graphs with subtree-shaped customers. Theoretical Computer Science, 411(7–9):1038– 1044, February 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fomin:2010:ICE

[84] Fedor V. Fomin, Serge Gaspers, Dieter Kratsch, Mathieu Liedloff, and Saket Saurabh. Iterative compression and exact algorithms. Theoretical Computer Science, 411(7–9):1045– 1053, February 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berenbrink:2010:EEB

[85] Petra Berenbrink and Oliver Schulte. Evolutionary equilibrium in Bayesian routing games: Specialization and niche formation. Theoretical Computer Science, 411(7–9):1054–1074, February 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DiLena:2010:ULB

[86] P. Di Lena and L. Margara. On the undecidability of the limit behavior of Cellular Automata. Theoretical Computer Science, 411(7–9):1075– 1084, February 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goodman-Strauss:2010:HSA

[87] C. Goodman-Strauss. A hierarchical strongly aperiodic set of tiles in the hyperbolic plane. Theoretical Computer Science, 411(7–9):1085–1093, February 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2010:TGK

[88] Zhenbo Wang, Wenxun Xing, and Shu-Cherng Fang. Two-group knapsack game. Theoretical Computer Science, 411(7–9):1094–1103, February 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abrahamson:2010:ETT

[89] Jeff Abrahamson and Ali Shokoufandeh. Euclidean TSP on two polygons. Theoretical Computer Science, 411(7–9):1104–1114, February 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gawrychowski:2010:PFM

[90] Pawel Gawrychowski, Marin Gutan, and Andrzej Kisielewicz. On the problem of freeness of multiplicative matrix semigroups. Theoretical Computer Science, 411(7–9):1115–1120, February 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gomez-Rodriguez:2010:ERP

[91] Carlos Gómez-Rodríguez, Miguel A. Alonso, and Manuel Vilares. Errorrepair parsing schemata. Theoretical Computer Science, 411(7–9):1121– 1139, February 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lu:2010:MMS

[92] Shenpeng Lu, Haodi Feng, and Xiuqian Li. Minimizing the makespan on a single parallel batching machine. Theoretical Computer Science, 411(7–9):1140– 1145, February 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gast:2010:ILT

[93] Nicolas Gast and Bruno Gaujal. Infinite labeled trees: From rational to Sturmian trees. Theoretical Computer Science, 411(7–9):1146–1166, February 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fomin:2010:PFR

[94] Fedor V. Fomin, Petr A. Golovach, Jan Kratochvíl, Nicolas Nisse, and Karol Suchan. Pursuing a fast robber on a graph. Theoretical Computer Science, 411(7–9):1167–1181, February 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Iwama:2010:CHC

[95] Kazuo Iwama, Kazuhisa Seto, and Suguru Tamaki. The complexity of the Hajós calculus for planar graphs. Theoretical Computer Science, 411(7–9): 1182–1191, February 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ibarra:2010:DPP

[96] Oscar H. Ibarra, Igor Potapov, and Hsu-Chun Yen. On decision problems for parameterized machines. Theoretical Computer Science, 411(7–9):1192– 1201, February 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bodlaender:2010:CPI

[97] Hans L. Bodlaender, Michael R. Fellows, Pinar Heggernes, Federico Mancini, Charis Papadopoulos, and Frances Rosamond. Clustering with partial information. Theoretical Computer Science, 411(7–9):1202–1211, February 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mezzini:2010:CFC

[98] Mauro Mezzini. On the complexity of finding chordless paths in bipartite graphs and some interval operators in graphs and hypergraphs. Theoretical Computer Science, 411(7–9):1212– 1220, February 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Machado:2010:CIG

[99] R. C. S. Machado, C. M. H. de Figueiredo, and K. Vusković. Chromatic index of graphs with no cycle with a unique chord. Theoretical Computer Science, 411(7–9):1221– 1234, February 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yung:2010:P

[100] Moti Yung. Preface. Theoretical Computer Science, 411(10):1235, March 4, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zimand:2010:SEC

[101] Marius Zimand. Simple extractors via constructions of cryptographic pseudo-random generators. Theoretical Computer Science, 411(10):1236– 1250, March 4, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Horvitz:2010:BEB

[102] Omer Horvitz and Jonathan Katz. Bounds on the efficiency of blackbox commitment schemes. Theoretical Computer Science, 411(10):1251– 1260, March 4, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chevalier:2010:SPA

[103] Yannick Chevalier and Michael Rusinowitch. Symbolic protocol analysis in the union of disjoint intruder theories: Combining decision procedures. Theoretical Computer Science, 411(10):

1261–1282, March 4, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Moran:2010:BCP

[104] Tal Moran and Moni Naor. Basing cryptographic protocols on tamperevident seals. Theoretical Computer Science, 411(10):1283–1310, March 4, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2010:EBd

[105] Anonymous. Editorial board. Theoretical Computer Science, 411(10):i– vi, March 4, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

He:2010:CRC

[106] Jifeng He and Tony Hoare. CSP is a retract of CCS. Theoretical Computer Science, 411(11–13):1311–1337, March 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Orzan:2010:IPB

[107] Simona Orzan and Tim A. C. Willemse. Invariants for Parameterised Boolean Equation Systems. Theoretical Computer Science, 411(11–13):1338– 1371, March 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barraclough:2010:TBS

[108] Richard W. Barraclough, David Binkley, Sebastian Danicic, Mark Harman, Robert M. Hierons, Ákos Kiss, Mike Laurence, and Lahcen Ouarbya. A trajectory-based strict semantics for program slicing. Theoretical Computer Science, 411(11–13):1372–1386, March 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fu:2010:EI

[109] Yuxi Fu and Hao Lu. On the expressiveness of interaction. Theoretical Computer Science, 411(11–13):1387– 1451, March 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gaucher:2010:CLH

[110] Philippe Gaucher. Combinatorics of labelling in higher-dimensional automata. Theoretical Computer Science, 411(11–13):1452–1483, March 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alves:2010:GSR

[111] Sandra Alves, Maribel Fernández, Mário Florido, and Ian Mackie. Gödel’s system T revisited. Theoretical Computer Science, 411(11–13):1484–1500, March 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hines:2010:QCO

[112] Peter Hines. Quantum circuit oracles for Abstract Machine computations. Theoretical Computer Science, 411(11–13):1501–1520, March 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schmidt-Schauss:2010:GCL

[113] Manfred Schmidt-Schauß and David Sabel. On generic context lemmas for higher-order calculi with sharing. Theoretical Computer Science, 411(11–13): 1521–1541, March 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shvartsman:2010:EP

[114] Alexander Shvartsman and Pascal Felber. Editors’ preface. Theoretical Computer Science, 411(14–15):1543, March 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ilcinkas:2010:FRB

[115] David Ilcinkas, Dariusz R. Kowalski, and Andrzej Pelc. Fast radio broadcasting with advice. Theoretical Computer Science, 411(14–15):1544– 1557, March 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Becker:2010:ALL

[116] Florent Becker, Sergio Rajsbaum, Ivan Rapaport, and Eric Rémila. Average long-lived binary consensus: Quantifying the stabilizing role played by memory. Theoretical Computer Science, 411(14–15):1558–1566, March 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Masuzawa:2010:QSS

[117] Toshimitsu Masuzawa and Sébastien Tixeuil. Quiescence of self-stabilizing gossiping among mobile agents in graphs. Theoretical Computer Science, 411(14–15):1567–1582, March 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flocchini:2010:RMT

[118] Paola Flocchini, David Ilcinkas, Andrzej Pelc, and Nicola Santoro. Remembering without memory: Tree exploration by asynchronous oblivious robots. Theoretical Computer Science, 411(14–15):1583–1598, March 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sauerwald:2010:SSA

[119] Thomas Sauerwald and Dirk Sudholt. A self-stabilizing algorithm for cut problems in synchronous networks. Theoretical Computer Science, 411(14–15):1599–1612, March 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fraigniaud:2010:RLR

[120] Pierre Fraigniaud, Emmanuelle Lebhar, and Zvi Lotker. Recovering the long-range links in augmented graphs. Theoretical Computer Science, 411(14–15):1613–1625, March 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bilo:2010:DNP

[121] Davide Bilò, Thomas Erlebach, Matúš Mihalák, and Peter Widmayer. Discovery of network properties with allshortest-paths queries. Theoretical Computer Science, 411(14–15):1626– 1637, March 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cooper:2010:LRF

[122] Colin Cooper, Ralf Klasing, and Tomasz Radzik. Locating and repairing faults in a network with mobile agents. Theoretical Computer Science, 411(14–15):1638–1647, March 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Adjiashvili:2010:EAL

[123] D. Adjiashvili and D. Peleg. Equal-area locus-based convex polygon decomposition. Theoretical Computer Science, 411(14–15):1648–1667, March 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chou:2010:AAP

[124] Well Y. Chou and Chiuyuan Chen. Allto-all personalized exchange in generalized shuffle-exchange networks. Theoretical Computer Science, 411(16–18): 1669–1684, March 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Watanabe:2010:ACA

[125] Osamu Watanabe and Masaki Yamamoto. Average-case analysis for the MAX-2SAT problem. Theoretical Computer Science, 411(16–18):1685– 1697, March 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fernau:2010:PAH

[126] Henning Fernau. Parameterized algorithms for d-Hitting Set: The weighted case. Theoretical Computer Science, 411(16–18):1698–1713, March 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yamanaka:2010:EEA

[127] Katsuhisa Yamanaka, Shin ichi Nakano, Yasuko Matsui, Ryuhei Uehara, and Kento Nakada. Efficient enumeration of all ladder lotteries and its application. Theoretical Computer Science, 411(16–18):1714–1722, March 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gaintzarain:2010:TPE

[128] Jose Gaintzarain, Montserrat Hermo, Paqui Lucio, and Marisa Navarro. Translating propositional extended conjunctions of Horn clauses into Boolean circuits. Theoretical Computer Science, 411(16–18):1723–1733, March 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ehmsen:2010:CFF

[129] Martin R. Ehmsen, Lene M. Favrholdt, Jens S. Kohrt, and Rodica Mihai. Comparing First-Fit and Next-Fit for online edge coloring. Theoretical Computer Science, 411(16–18):1734–1741, March 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chang:2010:OBF

[130] Ching-Lueh Chang and Yuh-Dauh Lyuu. Optimal bounds on finding fixed points of contraction mappings. Theoretical Computer Science, 411(16–18): 1742–1749, March 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Borbelova:2010:RSM

[131] Viera Borbel’ová and Katarína Cechlárová. Rotations in the stable b-matching problem. Theoretical Computer Science, 411(16–18):1750–1762, March 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mahmoud:2010:DAS

[132] Hosam M. Mahmoud. Distributional analysis of swaps in Quick Select. Theoretical Computer Science, 411(16–18):

1763–1769, March 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ipate:2010:BST

[133] Florentin Ipate. Bounded sequence testing from deterministic finite state machines. Theoretical Computer Science, 411(16–18):1770–1784, March 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Giambruno:2010:TBD

[134] Laura Giambruno and Sabrina Mantaci. Transducers for the bidirectional decoding of prefix codes. Theoretical Computer Science, 411(16–18):1785– 1792, March 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Geffert:2010:MPM

[135] Viliam Geffert and Jozef Gajdoš. Multiway in-place merging. Theoretical Computer Science, 411(16–18):1793– 1808, March 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shin:2010:CMM

[136] Sujin Shin. Codes and maximal monoids. Theoretical Computer Science, 411(16–18):1809–1817, March 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salomaa:2010:CME

[137] Arto Salomaa. Criteria for the matrix equivalence of words. Theoretical Computer Science, 411(16–18):1818– 1827, March 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Biro:2010:SVS

[138] Péter Biró, David F. Manlove, and Shubham Mittal. Size versus stability in the marriage problem. Theoretical Computer Science, 411(16–18): 1828–1841, March 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hrubes:2010:CCM

[139] P. Hrubeš, S. Jukna, A. Kulikov, and P. Pudlák. On convex complexity measures. Theoretical Computer Science, 411(16–18):1842–1854, March 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cardinal:2010:NCF

[140] Jean Cardinal and Martin Hoefer. Noncooperative facility location and covering games. Theoretical Computer Science, 411(16–18):1855–1876, March 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2010:SMS

[141] Liqi Zhang, Lingfa Lu, and Jinjiang Yuan. Single-machine scheduling under the job rejection constraint. Theoretical Computer Science, 411(16–18): 1877–1882, March 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2010:ETL

[142] Xin Li and Tian Liu. On exponential time lower bound of Knapsack under backtracking. Theoretical Computer Science, 411(16–18):1883– 1888, March 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nonaka:2010:HCT

[143] Yoshiaki Nonaka, Hirotaka Ono, Kunihiko Sadakane, and Masafumi Yamashita. The hitting and cover times of Metropolis walks. Theoretical Computer Science, 411(16–18):1889– 1894, March 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hong:2010:BDR

[144] Cheng W. Hong. The biased, distancerestricted n-in-a-row game for small p. Theoretical Computer Science, 411 (16–18):1895–1897, March 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mislove:2010:P

[145] Michael Mislove. Preface. Theoretical Computer Science, 411(19):1899, April 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Battenfeld:2010:CFA

[146] Ingo Battenfeld. Comparing free algebras in Topological and Classical Domain Theory. Theoretical Computer Science, 411(19):1900–1917, April 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Martin:2010:AIT

[147] Keye Martin, Ira S. Moskowitz, and Gerard Allwein. Algebraic information theory for binary channels. Theoretical Computer Science, 411(19):1918– 1927, April 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lanese:2010:OSC

[148] Ivan Lanese and Davide Sangiorgi. An operational semantics for a calculus for wireless systems. Theoretical Computer Science, 411(19):1928–1948, April 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Varacca:2010:TES

[149] Daniele Varacca and Nobuko Yoshida. Typed event structures and the linear π-calculus. Theoretical Computer Science, 411(19):1949–1973, April 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2010:ASI

[150] Peng Li and Steve Zdancewic. Arrows for secure information flow. Theoretical Computer Science, 411(19):1974– 1994, April 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2010:Pb

[151] G. Rozenberg. Preface. Theoretical Computer Science, 411(20):1995– 1996, April 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cannata:2010:PHA

[152] Nicola Cannata, Emanuela Merelli, and Irek Ulidowski. Preface: Hybrid automata and oscillatory behaviour in biological systems. Theoretical Computer Science, 411(20):1997–1998, April 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bartocci:2010:DSB

[153] Ezio Bartocci, Flavio Corradini, Emanuela Merelli, and Luca Tesei. Detecting synchronisation of biological oscillators by model checking. Theoretical Computer Science, 411(20):1999– 2018, April 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ballarini:2010:QBV

[154] Paolo Ballarini and Maria Luisa Guerriero. Query-based verification of qualitative trends and oscillations in biochemical systems. Theoretical Computer Science, 411(20):2019– 2036, April 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Campagna:2010:HAR

[155] Dario Campagna and Carla Piazza. Hybrid automata, reachability, and Systems Biology. Theoretical Computer Science, 411(20):2037–2051, April 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bortolussi:2010:HDS

[156] Luca Bortolussi and Alberto Policriti. Hybrid dynamics of stochastic programs. Theoretical Computer Science, 411(20):2052–2077, April 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2010:EBe

[157] Anonymous. Editorial board. Theoretical Computer Science, 411(20):i– vi, March 4, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2010:Pc

[158] G. Rozenberg. Preface. Theoretical Computer Science, 411(20):1995– 1996, April 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2010:EBf

[159] Anonymous. Editorial board. Theoretical Computer Science, 411(21): i–vi, May 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2010:Pd

[160] G. Rozenberg. Preface. Theoretical Computer Science, 411(21):2079– 2080, May 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonabeau:2010:SIT

[161] Eric Bonabeau, David Corne, Joshua Knowles, and Riccardo Poli. Swarm intelligence theory: a snapshot of the state of the art. Theoretical Computer Science, 411(21):2081–2083, May 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sudholt:2010:RAB

[162] Dirk Sudholt and Carsten Witt. Runtime analysis of a binary particle swarm optimizer. Theoretical Computer Science, 411(21):2084–2100, May 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2010:API

[163] Ying ping Chen and Pei Jiang. Analysis of particle interaction in particle swarm optimization. Theoretical Computer Science, 411(21):2101–2115, May 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vrancx:2010:ADS

[164] Peter Vrancx, Katja Verbeeck, and Ann Nowé. Analyzing the dynamics of stigmergetic interactions through pheromone games. Theoretical Computer Science, 411(21):2116–2126, May 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Biswas:2010:SAR

[165] Arijit Biswas, Swagatam Das, Ajith Abraham, and Sambarta Dasgupta. Stability analysis of the reproduction operator in bacterial foraging optimization. Theoretical Computer Science, 411(21):2127–2139, May 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gokce:2010:PCF

[166] Fatih Gökçe and Erol Şahin. The pros and cons of flocking in the long-range “migration” of mobile robot swarms. Theoretical Computer Science, 411(21): 2140–2154, May 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bertet:2010:MFC

[167] K. Bertet and B. Monjardet. The multiple facets of the canonical direct unit implicational basis. Theoretical Computer Science, 411(22–24): 2155–2166, May 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhao:2010:DCP

[168] Dongsheng Zhao and Taihe Fan. Dcpocompletion of posets. Theoretical Computer Science, 411(22–24):2167– 2173, May 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeFrancesco:2010:UAI

[169] Nicoletta De Francesco, Giuseppe Lettieri, and Luca Martini. Using abstract interpretation to add type checking for interfaces in Java bytecode verification. Theoretical Computer Science, 411(22–24):2174–2201, May 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tini:2010:NEB

[170] Simone Tini. Non-expansive ϵ bisimulations for probabilistic processes. Theoretical Computer Science, 411(22–24): 2202–2222, May 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Honda:2010:ECB

[171] Kohei Honda and Olivier Laurent. An exact correspondence between a typed pi-calculus and polarised proof-nets. Theoretical Computer Science, 411(22– 24):2223–2238, May 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hasuo:2010:PAC

[172] Ichiro Hasuo, Yoshinobu Kawabe, and Hideki Sakurada. Probabilistic anonymity via coalgebraic simulations. Theoretical Computer Science, 411(22– 24):2239–2259, May 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hayden:2010:FAF

[173] Richard A. Hayden and Jeremy T. Bradley. A fluid analysis framework for a Markovian process algebra. Theoretical Computer Science, 411(22–24): 2260–2297, May 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Demri:2010:MCM

[174] Stéphane Demri, Ranko Lazić, and Arnaud Sangnier. Model checking memoryful linear-time logics over onecounter automata. Theoretical Computer Science, 411(22–24):2298–2316, May 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hesselink:2010:ASD

[175] Wim H. Hesselink. Alternating states for dual nondeterminism in imperative programming. Theoretical Computer Science, 411(22–24):2317–2330, May 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rabinovich:2010:CMT

[176] Alexander Rabinovich. Complexity of metric temporal logics with counting and the Pnueli modalities. Theoretical Computer Science, 411(22–24): 2331–2342, May 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2010:EBg

[177] Anonymous. Editorial board. Theoretical Computer Science, 411(25):i– vi, May 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2010:Pe

[178] G. Rozenberg. Preface. Theoretical Computer Science, 411(25):2343– 2344, May 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ishdorj:2010:DSQ

[179] Tseren-Onolt Ishdorj, Alberto Leporati, Linqiang Pan, Xiangxiang Zeng, and Xingyi Zhang. Deterministic solutions to QSAT. Theoretical Computer Science, 411(25):2345–2358, May 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alhazov:2010:CGB

[180] Artiom Alhazov, Chang Li, and Ion Petre. Computing the graph-based parallel complexity of gene assembly. Theoretical Computer Science, 411(25): 2359–2367, May 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hayes:2010:HFP

[181] Christina Hayes and Tomáš Gedeon. Hyperbolicity of the fixed point set for the simple genetic algorithm. Theoretical Computer Science, 411(25):2368– 2383, May 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Apolloni:2010:PMG

[182] Bruno Apolloni, Simone Bassis, Sabrina Gaito, Dario Malchiodi, and Italo Zoppis. Playing monotone games to understand learning behaviors. Theoretical Computer Science, 411(25):2384– 2405, May 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Neumann:2010:ACO

[183] Frank Neumann and Carsten Witt. Ant Colony Optimization and the minimum spanning tree problem. Theoretical Computer Science, 411(25):2406– 2413, May 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mitrana:2010:ASS

[184] Victor Mitrana, Ion Petre, and Vladimir Rogojin. Accepting splicing systems. Theoretical Computer Science, 411(25):2414–2422, May 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2010:Pf

[185] G. Rozenberg. Preface. Theoretical Computer Science, 411(25):i–vi, May 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Herrbach:2010:ACJ

[186] Claire Herrbach, Alain Denise, and Serge Dulucq. Average complexity of the Jiang–Wang–Zhang pairwise tree alignment algorithm and of a RNA secondary structure alignment algorithm. Theoretical Computer Science, 411(26– 28):2423–2432, June 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yamashita:2010:CGP

[187] Masafumi Yamashita and Ichiro Suzuki. Characterizing geometric patterns formable by oblivious anonymous mobile robots. Theoretical Computer Science, 411(26–28):2433–2453, June 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shanker:2010:CND

[188] O. Shanker. Complex network dimension and path counts. Theoretical Computer Science, 411(26–28):2454– 2458, June 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dolev:2010:GAR

[189] Shlomi Dolev, Elad M. Schiller, Paul G. Spirakis, and Philippas Tsigas. Game authority for robust and scalable distributed selfish-computer systems. Theoretical Computer Science, 411(26– 28):2459–2466, June 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hua:2010:DPB

[190] Qiang-Sheng Hua, Yuexuan Wang, Dongxiao Yu, and Francis C. M. Lau. Dynamic programming based algorithms for set multicover and multiset multicover problems. Theoretical Computer Science, 411(26–28):2467– 2474, June 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fertin:2010:FCS

[191] Guillaume Fertin, Danny Hermelin, Romeo Rizzi, and Stéphane Vialette. Finding common structured patterns in linear graphs. Theoretical Computer Science, 411(26–28):2475–2486, June 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bouvel:2010:PPD

[192] Mathilde Bouvel and Elisa Pergola. Posets and permutations in the duplication–loss model: Minimal permutations with d descents. Theoretical Computer Science, 411(26–28):2487– 2501, June 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nutov:2010:AMP

[193] Zeev Nutov. Approximating minimum power covers of intersecting families and directed edge-connectivity problems. Theoretical Computer Science, 411(26–28):2502–2512, June 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2010:IRE

[194] WenBin Chen, Dengpan Yin, and Zhengzhang Chen. Inapproximability results for equations over infinite groups. Theoretical Computer Science, 411(26–28):2513–2519, June 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hendrix:2010:TUM

[195] William Hendrix, Matthew C. Schmidt, Paul Breimyer, and Nagiza F. Samatova. Theoretical underpinnings for maximal clique enumeration on perturbed graphs. Theoretical Computer Science, 411(26–28):2520–2536, June 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2010:IAA

[196] Tongquan Zhang, Ying Yin, and Jianping Li. An improved approximation algorithm for the maximum TSP. Theoretical Computer Science, 411(26–28): 2537–2541, June 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Angelopoulos:2010:RPA

[197] Spyros Angelopoulos and Allan Borodin. Randomized priority algorithms. Theoretical Computer Science, 411(26–28): 2542–2558, June 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Okhotin:2010:CGR

[198] Alexander Okhotin and Christian Reitwießner. Conjunctive grammars with restricted disjunction. Theoretical Computer Science, 411(26–28):2559– 2571, June 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boyar:2010:TRN

[199] Joan Boyar, Leah Epstein, and Asaf Levin. Tight results for Next Fit and Worst Fit with resource augmentation. Theoretical Computer Science, 411(26–28):2572–2580, June 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cardinal:2010:CVC

[200] Jean Cardinal and Eythan Levy. Connected vertex covers in dense graphs. Theoretical Computer Science, 411(26– 28):2581–2590, June 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kijima:2010:LSC

[201] Shuji Kijima, Masashi Kiyomi, Yoshio Okamoto, and Takeaki Uno. On listing, sampling, and counting the chordal graphs with edge constraints. Theoretical Computer Science, 411(26–28): 2591–2601, June 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Antunes:2010:IMI

[202] Luís Antunes and André Souto. Information measures for infinite sequences. Theoretical Computer Science, 411(26– 28):2602–2611, June 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2010:PCM

[203] Zhiqiang Zhang and Yaoyun Shi. On the parity complexity measures of Boolean functions. Theoretical Computer Science, 411(26–28):2612–2618, June 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kawarabayashi:2010:SAC

[204] Ken ichi Kawarabayashi and Kenta Ozeki. A simple algorithm for 4coloring 3-colorable planar graphs. Theoretical Computer Science, 411(26– 28):2619–2622, June 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tarui:2010:SFP

[205] Jun Tarui. Smallest formulas for the parity of 2^{k} variables are essentially unique. Theoretical Computer Science, 411(26–28):2623–2627, June 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gyorfi:2010:GEF

[206] László Györfi, György Turán, and Thomas Zeugmann. Guest Editors’ foreword. Theoretical Computer Science, 411(29–30):2629–2631, June 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vovk:2010:PRP

[207] Vladimir Vovk and Alexander Shen. Prequential randomness and probability. Theoretical Computer Science, 411(29–30):2632–2646, June 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chernov:2010:SPE

[208] Alexey Chernov, Yuri Kalnishkan, Fedor Zhdanov, and Vladimir Vovk. Supermartingales in prediction with expert advice. Theoretical Computer Science, 411(29–30):2647–2669, June 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mukherjee:2010:LCE

[209] Indraneel Mukherjee and Robert E. Schapire. Learning with continuous experts using drifting games. Theoretical Computer Science, 411(29–30): 2670–2683, June 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ortner:2010:ORB

[210] Ronald Ortner. Online regret bounds for Markov decision processes with deterministic transitions. Theoretical Computer Science, 411(29–30):2684– 2695, June 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shamir:2010:LGI

[211] Ohad Shamir, Sivan Sabato, and Naftali Tishby. Learning and generalization with the information bottleneck. Theoretical Computer Science, 411(29–30):2696–2711, June 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Antos:2010:ALH

[212] András Antos, Varun Grover, and Csaba Szepesvári. Active learning in heteroscedastic noise. Theoretical Computer Science, 411(29–30):2712– 2728, June 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Angluin:2010:OLS

[213] Dana Angluin, James Aspnes, and Lev Reyzin. Optimally learning social networks with activations and suppressions. Theoretical Computer Science, 411(29–30):2729–2740, June 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Becerra-Bonache:2010:ILS

[214] Leonor Becerra-Bonache, John Case, Sanjay Jain, and Frank Stephan. Iterative learning of simple external contextual languages. Theoretical Computer Science, 411(29–30):2741–2756, June 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jain:2010:ILT

[215] Sanjay Jain, Steffen Lange, Samuel E. Moelius III, and Sandra Zilles. Incremental learning with temporary memory. Theoretical Computer Science, 411(29–30):2757–2772, June 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Izumi:2010:AIM

[216] Tomoko Izumi, Taisuke Izumi, Hirotaka Ono, and Koichi Wada. Approximability and inapproximability of the minimum certificate dispersal problem. Theoretical Computer Science, 411(31–33):2773–2783, June 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Horvath:2010:EFC

[217] Tamás Horváth and Jan Ramon. Efficient frequent connected subgraph mining in graphs of bounded tree-width. Theoretical Computer Science, 411(31– 33):2784–2797, June 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Efraimidis:2010:WGB

[218] Pavlos S. Efraimidis, Lazaros Tsavlidis, and George B. Mertzios. Windowgames between TCP flows. Theoretical Computer Science, 411(31–33):2798– 2817, June 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2010:DAC

[219] Jiao fen Li, Xi yan Hu, and Lei Zhang. Dykstra’s algorithm for constrained least-squares doubly symmetric matrix problems. Theoretical Computer Science, 411(31–33):2818–2826, June 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chuan:2010:FCW

[220] Wai-Fong Chuan and Hui-Ling Ho. Factors of characteristic words: Location and decompositions. Theoretical Computer Science, 411(31–33):2827– 2846, June 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Armon:2010:CT

[221] Amitai Armon, Adi Avidor, and Oded Schwartz. Cooperative TSP. Theoretical Computer Science, 411(31–33): 2847–2863, June 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Buning:2010:UBC

[222] H. Kleine Büning and A. Remshagen. An upper bound for the circuit complexity of existentially quantified Boolean formulas. Theoretical Computer Science, 411(31–33):2864– 2870, June 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ye:2010:DLC

[223] Deshi Ye, Xin Han, and Guochuan Zhang. Deterministic on-line call control in cellular networks. Theoretical Computer Science, 411(31–33):2871– 2877, June 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2010:BCA

[224] Yen-Chiu Chen, Hsin-Wen Wei, Pei-Chi Huang, Wei-Kuan Shih, and Tsan sheng Hsu. The bridge-connectivity augmentation problem with a partition constraint. Theoretical Computer Science, 411(31–33):2878–2889, June 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cao:2010:OSR

[225] Qian Cao and Zhaohui Liu. Online scheduling with reassignment on two uniform machines. Theoretical Computer Science, 411(31–33):2890– 2898, June 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Epstein:2010:TDO

[226] Leah Epstein. Two-dimensional online bin packing with rotation. Theoretical Computer Science, 411(31–33): 2899–2911, June 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lai:2010:CHG

[227] Pao-Lien Lai, Hong-Chun Hsu, Chang-Hsiung Tsai, and Iain A. Stewart. A class of hierarchical graphs as topologies for interconnection networks. Theoretical Computer Science, 411(31–33): 2912–2924, June 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gagie:2010:MFD

[228] Travis Gagie and Giovanni Manzini. Move-to-front, Distance Coding, and Inversion Frequencies revisited. Theoretical Computer Science, 411(31–33): 2925–2944, June 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Vidal:2010:OAG

[229] Samuel Alexandre Vidal. An optimal algorithm to generate rooted trivalent diagrams and rooted triangular maps. Theoretical Computer Science, 411(31–33):2945–2967, June 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brandstadt:2010:ELP

[230] Andreas Brandstädt, Van Bang Le, and Dieter Rautenbach. Exact leaf powers. Theoretical Computer Science, 411(31–33):2968–2977, June 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nong:2010:PTA

[231] Q. Q. Nong, T. C. E. Cheng, and C. T. Ng. A polynomial-time algorithm for the weighted link ring loading problem with integer demand splitting. Theoretical Computer Science, 411(31–33):2978–2986, June 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gelade:2010:SRE

[232] Wouter Gelade. Succinctness of regular expressions with interleaving, intersection and counting. Theoretical Computer Science, 411(31–33):2987– 2998, June 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kurka:2010:EET

[233] Petr Kůrka. Erratum to: Entropy of Turing machines with moving head. Theoretical Computer Science, 411(31– 33):2999–3000, June 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [1].

Busch:2010:ECN

[234] Costas Busch and Marios Mavronicolas. An efficient counting network. Theoretical Computer Science, 411(34–36): 3001–3030, July 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mizrahi:2010:CCA

[235] Tal Mizrahi and Yoram Moses. Continuous consensus with ambiguous failures. Theoretical Computer Science, 411(34–36):3031–3041, July 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yuster:2010:SSS

[236] Raphael Yuster. Single source shortest paths in H-minor free graphs. Theoretical Computer Science, 411(34–36): 3042–3047, July 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Feigenblat:2010:FIP

[237] Guy Feigenblat, Ofra Itzhaki, and Ely Porat. The frequent items problem, under polynomial decay, in the streaming model. Theoretical Computer Science, 411(34–36):3048–3054, July 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See corrigendum [2459].

Bourgeois:2010:AME

[238] N. Bourgeois, G. Lucarelli, I. Milis, and V. Th. Paschos. Approximating the max-edge-coloring problem. Theoretical Computer Science, 411(34–36): 3055–3067, July 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ravaux:2010:DTL

[239] Romain Ravaux. Decomposing trees with large diameter. Theoretical Computer Science, 411(34–36):3068–3072, July 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Epstein:2010:CCP

[240] Leah Epstein, Csanád Imreh, and Asaf Levin. Class constrained bin packing revisited. Theoretical Computer Science, 411(34–36):3073–3089, July 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Drewes:2010:ASG

[241] Frank Drewes, Berthold Hoffmann, Dirk Janssens, and Mark Minas. Adaptive star grammars and their languages. Theoretical Computer Science, 411(34– 36):3090–3109, July 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Groult:2010:OSA

[242] R. Groult and G. Richomme. Optimality of some algorithms to detect quasiperiodicities. Theoretical Computer Science, 411(34–36):3110– 3122, July 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bains:2010:RCM

[243] Adam Bains and Therese Biedl. Reconstructing h v-convex multi-coloured polyominoes. Theoretical Computer Science, 411(34–36):3123–3128, July 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Epstein:2010:IRR

[244] Leah Epstein and Asaf Levin. Improved randomized results for the interval selection problem. Theoretical Computer Science, 411(34–36):3129– 3135, July 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Biro:2010:CAP

[245] Péter Biró, Tamás Fleiner, Robert W. Irving, and David F. Manlove. The College Admissions problem with lower and common quotas. Theoretical Computer Science, 411(34–36):3136–3153, July 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bouzid:2010:OBR

[246] Zohir Bouzid, Maria Gradinariu Potop-Butucaru, and Sébastien Tixeuil. Optimal Byzantine-resilient convergence in uni-dimensional robot networks. Theoretical Computer Science, 411(34–36): 3154–3168, July 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duchene:2010:IG

[247] Eric Duchêne and Michel Rigo. Invariant games. Theoretical Computer Science, 411(34–36):3169–3180, July 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mathieson:2010:PCE

[248] Luke Mathieson. The parameterized complexity of editing graphs for bounded degeneracy. Theoretical Computer Science, 411(34–36):3181–3187, July 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carvalho:2010:CDT

[249] Catarina Carvalho, Víctor Dalmau, and Andrei Krokhin. CSP duality and trees of bounded pathwidth. Theoretical Computer Science, 411(34–36):

3188–3208, July 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shur:2010:GRC

[250] Arseny M. Shur. Growth rates of complexity of power-free languages. Theoretical Computer Science, 411(34–36): 3209–3223, July 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cincotti:2010:PPG

[251] Alessandro Cincotti. N-player partizan games. Theoretical Computer Science, 411(34–36):3224–3234, July 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Klasing:2010:TAS

[252] Ralf Klasing, Adrian Kosowski, and Alfredo Navarra. Taking advantage of symmetries: Gathering of many asynchronous oblivious robots on a ring. Theoretical Computer Science, 411(34– 36):3235–3246, July 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhou:2010:ATD

[253] Shan Zhou, Hao Li, and Guanghui Wang. Algorithm for two disjoint long paths in 2-connected graphs. Theoretical Computer Science, 411(34–36): 3247–3254, July 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kuba:2010:CAA

[254] Markus Kuba and Alois Panholzer. A combinatorial approach to the analysis of bucket recursive trees. Theoretical Computer Science, 411(34–36):

3255–3273, July 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Freydenberger:2010:END

[255] Dominik D. Freydenberger and Daniel Reidenbach. Existence and nonexistence of descriptive patterns. Theoretical Computer Science, 411(34–36): 3274–3286, July 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shao:2010:LCG

[256] Zhendong Shao and Roberto Solis-Oba. L(2, 1)-labelings on the composition of n graphs. Theoretical Computer Science, 411(34–36):3287–3292, July 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ferragina:2010:CRA

[257] Paolo Ferragina, Igor Nitto, and Rossano Venturini. On compact representations of All-Pairs-Shortest-Path-Distance matrices. Theoretical Computer Science, 411(34–36):3293–3300, July 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2010:EBh

[258] Anonymous. Editorial Board. Theoretical Computer Science, 411(34–36): i–vi, July 17, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blanqui:2010:CLC

[259] Frédéric Blanqui, Claude Kirchner, and Colin Riba. On the confluence of lambda-calculus with conditional rewriting. Theoretical Computer Science, 411(37):3301–3327, August 7, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Padovani:2010:CBD

[260] Luca Padovani. Contract-based discovery of Web services modulo simple orchestrators. Theoretical Computer Science, 411(37):3328–3347, August 7, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alpuente:2010:CFS

[261] M. Alpuente, M. Comini, S. Escobar, M. Falaschi, and J. Iborra. A compact fixpoint semantics for term rewriting systems. Theoretical Computer Science, 411(37):3348–3371, August 7, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hierons:2010:CES

[262] Robert M. Hierons. Checking experiments for stream X-machines. Theoretical Computer Science, 411(37):3372– 3385, August 7, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

OConnor:2010:CVM

[263] Russell O’Connor and Bas Spitters. A computer-verified monadic functional implementation of the integral. Theoretical Computer Science, 411(37): 3386–3402, August 7, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maneth:2010:P

[264] Sebastian Maneth. Preface. Theoretical Computer Science, 411(38–39): 3403, August 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Holzer:2010:AHM

[265] Markus Holzer and Andreas Maletti. An n log n algorithm for hyperminimizing a (minimized) deterministic automaton. Theoretical Computer Science, 411(38–39):3404–3413, August 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Castiglione:2010:ECH

[266] G. Castiglione, A. Restivo, and M. Sciortino. On extremal cases of Hopcroft’s algorithm. Theoretical Computer Science, 411(38–39):3414– 3422, August 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Caron:2010:AAS

[267] Pascal Caron, Jean-Marc Champarnaud, and Ludovic Mignot. Acyclic automata and small expressions using multi-tilde-bar operators. Theoretical Computer Science, 411(38–39):3423– 3435, August 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [671].

Freivalds:2010:AND

[268] Rusins Freivalds. Amount of nonconstructivity in deterministic finite automata. Theoretical Computer Science, 411(38–39):3436–3443, August 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ganty:2010:FPG

[269] Pierre Ganty, Nicolas Maquet, and Jean-François Raskin. Fixed point guided abstraction refinement for alternating automata. Theoretical Computer Science, 411(38–39):3444–3459, August 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Touili:2010:VPP

[270] Tayssir Touili and Mohamed Faouzi Atig. Verifying parallel programs with dynamic communication structures. Theoretical Computer Science, 411(38–39):3460–3468, August 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heam:2010:PRG

[271] Pierre-Cyrille Héam, Cyril Nicaud, and Sylvain Schmitz. Parametric random generation of deterministic tree automata. Theoretical Computer Science, 411(38–39):3469–3480, August 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Inaba:2010:CRA

[272] Kazuhiro Inaba and Haruo Hosoya. Compact representation for answer sets of n-ary regular queries. Theoretical Computer Science, 411(38–39):3481– 3492, August 28, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fujiyoshi:2010:RDA

[273] Akio Fujiyoshi. Recognition of directed acyclic graphs by spanning tree automata. Theoretical Computer Science, 411(38–39):3493–3506, August 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Clarridge:2010:ACA

[274] Adam Clarridge and Kai Salomaa. Analysis of a cellular automaton model for car traffic with a slow-to-stop rule. Theoretical Computer Science, 411 (38–39):3507–3515, August 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kutrib:2010:CAS

[275] Martin Kutrib and Andreas Malcher. Cellular automata with sparse communication. Theoretical Computer Science, 411(38–39):3516–3526, August 28, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Denise:2010:CNU

[276] A. Denise, Y. Ponty, and M. Termier. Controlled non-uniform random generation of decomposable structures. Theoretical Computer Science, 411(40–42):3527–3552, September 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Flammini:2010:MTB

[277] Michele Flammini, Gianpiero Monaco, Luca Moscardelli, Hadas Shachnai, Mordechai Shalom, Tami Tamir, and Shmuel Zaks. Minimizing total busy time in parallel scheduling with application to optical networks. Theoretical Computer Science, 411(40–42):3553– 3562, September 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ciancarini:2010:PPK

[278] Paolo Ciancarini and Gian Piero Favini. Playing the perfect Kriegspiel endgame. Theoretical Computer Science, 411(40–42):3563–3577, September 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ng:2010:PSS

[279] C. T. Ng, Shisheng Li, T. C. E. Cheng, and Jinjiang Yuan. Preemptive scheduling with simple linear deterioration on a single machine. Theoretical Computer Science, 411(40–42):3578– 3586, September 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Han:2010:DSP

[280] Xin Han, Tak-Wah Lam, Lap-Kei Lee, Isaac K. K. To, and Prudence W. H. Wong. Deadline scheduling and power management for speed bounded processors. Theoretical Computer Science, 411(40–42):3587–3600, September 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hof:2010:CRA

[281] Pim van ’t Hof, Daniël Paulusma, and Johan M. M. van Rooij. Computing role assignments of chordal graphs. Theoretical Computer Science, 411(40– 42):3601–3613, September 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yu:2010:PSI

[282] Gang Yu, Xiaoxiao Ma, Yong Shen, and Wenbao Han. Provable secure identity based generalized signcryption scheme. Theoretical Computer Science, 411(40–42):3614–3624, September 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bell:2010:CSP

[283] Paul C. Bell, Jean-Charles Delvenne, Raphaël M. Jungers, and Vincent D. Blondel. The continuous Skolem–Pisot problem. Theoretical Computer Science, 411(40–42):3625–3634, September 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Matsui:2010:EPS

[284] Yasuko Matsui, Ryuhei Uehara, and Takeaki Uno. Enumeration of the perfect sequences of a chordal graph. Theoretical Computer Science, 411(40– 42):3635–3641, September 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2010:PMS

[285] Shisheng Li and Jinjiang Yuan. Parallel-machine scheduling with deteriorating jobs and rejection. Theoretical Computer Science, 411(40–42):3642– 3650, September 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2010:SRI

[286] Weifan Wang, Stephen Finbow, and Ping Wang. The surviving rate of an infected network. Theoretical Computer Science, 411(40–42):3651–3660, September 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goldefus:2010:LFC

[287] Filip Goldefus, Tomás Masopust, and Alexander Meduna. Left-forbidding cooperating distributed grammar systems. Theoretical Computer Science, 411(40–42):3661–3667, September 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bucci:2010:NEP

[288] Michelangelo Bucci, Aldo de Luca, and Alessandro De Luca. On the number of episturmian palindromes. Theoretical Computer Science, 411(40–42):3668– 3684, September 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Meister:2010:TMF

[289] Daniel Meister. Treewidth and minimum fill-in on permutation graphs in linear time. Theoretical Computer Science, 411(40–42):3685–3700, September 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cygan:2010:EAB

[290] Marek Cygan and Marcin Pilipczuk. Exact and approximate bandwidth. Theoretical Computer Science, 411(40– 42):3701–3713, September 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Efthymiou:2010:STH

[291] Charilaos Efthymiou and Paul G. Spirakis. Sharp thresholds for Hamiltonicity in random intersection graphs.

Theoretical Computer Science, 411(40– 42):3714–3730, September 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huang:2010:NCW

[292] Yun Bao Huang and William D. Weakley. A note on the complexity of C^{∞}- words. Theoretical Computer Science, 411(40–42):3731–3735, September 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2010:IUB

[293] Jianer Chen, Iyad A. Kanj, and Ge Xia. Improved upper bounds for vertex cover. Theoretical Computer Science, 411(40–42):3736–3756, September 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Csernenszky:2010:PCD

[294] András Csernenszky. The Picker– Chooser diameter game. Theoretical Computer Science, 411(40–42):3757– 3762, September 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lai:2010:ETG

[295] Pao-Lien Lai and Chang-Hsiung Tsai. Embedding of tori and grids into twisted cubes. Theoretical Computer Science, 411(40–42):3763–3773, September 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dvorak:2010:CCL

[296] Tomás Dvorák and Václav Koubek. Computational complexity of long paths and cycles in faulty hypercubes.

Theoretical Computer Science, 411(40– 42):3774–3786, September 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Okun:2010:SOP

[297] Michael Okun. Strong order-preserving renaming in the synchronous message passing model. Theoretical Computer Science, 411(40–42):3787–3794, September 6, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cohen:2010:FSI

[298] Hagai Cohen and Ely Porat. Fast set intersection and two-patterns matching. Theoretical Computer Science, 411 (40–42):3795–3800, September 6, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ravelomanana:2010:BGM

[299] Vlady Ravelomanana. Birth and growth of multicyclic components in random hypergraphs. Theoretical Computer Science, 411(43):3801–3813, October 9, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dombb:2010:ASM

[300] Yair Dombb, Ohad Lipsky, Benny Porat, Ely Porat, and Asaf Tsur. The approximate swap and mismatch edit distance. Theoretical Computer Science, 411(43):3814–3822, October 9, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Busch:2010:CCH

[301] Costas Busch and Srikanta Tirthapura. Concurrent counting is harder than queuing. Theoretical Computer Science, 411(43):3823–3833, October 9, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonato:2010:CRD

[302] Anthony Bonato, Ehsan Chiniforooshan, and Pawel Pralat. Cops and Robbers from a distance. Theoretical Computer Science, 411(43):3834–3844, October 9, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Scott:2010:PPE

[303] Allan Scott and Ulrike Stege. Parameterized pursuit-evasion games. Theoretical Computer Science, 411(43):3845– 3858, October 9, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kiyomi:2010:RIG

[304] Masashi Kiyomi, Toshiki Saitoh, and Ryuhei Uehara. Reconstruction of interval graphs. Theoretical Computer Science, 411(43):3859–3866, October 9, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pasanen:2010:RBS

[305] Tomi A. Pasanen. Random binary search tree with equal elements. Theoretical Computer Science, 411(43): 3867–3872, October 9, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhu:2010:EEL

[306] Junlei Zhu and Yuehua Bu. Equitable and equitable list colorings of graphs. Theoretical Computer Science, 411(43):3873–3876, October 9, 2010.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Caterina:2010:OPP

[307] Gianluca Caterina and Bruce Boghosian. An order-preserving property of additive invariants for Takesue-type reversible cellular automata. Theoretical Computer Science, 411(43):3877– 3881, October 9, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gao:2010:MPS

[308] Zhi-Han Gao and Fang-Wei Fu. The minimal polynomial of a sequence obtained from the componentwise linear transformation of a linear recurring sequence. Theoretical Computer Science, 411(44–46):3883–3893, October 25, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Herrero:2010:CIR

[309] María Isabel Herrero, Gabriela Jeronimo, and Juan Sabia. Computing isolated roots of sparse polynomial systems in affine space. Theoretical Computer Science, 411(44–46):3894–3904, October 25, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frisco:2010:SNA

[310] Pierluigi Frisco and Oscar H. Ibarra. On sets of numbers accepted by P/T systems composed by join. Theoretical Computer Science, 411(44–46):3905– 3916, October 25, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huss:2010:ESL

[311] Wilfried Huss, Ecaterina Sava, and Wolfgang Woess. Entropy sensitivity of languages defined by infinite automata, via Markov chains with forbidden transitions. Theoretical Computer Science, 411(44–46):3917–3922, October 25, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wu:2010:MRT

[312] Chan-Shuo Wu and Guan-Shieng Huang. A metric for rooted trees with unlabeled vertices based on nested parentheses. Theoretical Computer Science, 411(44–46):3923–3931, October 25, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carpi:2010:FSS

[313] Arturo Carpi and Valerio D’Alonzo. On factors of synchronized sequences. Theoretical Computer Science, 411(44–46): 3932–3937, October 25, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Uchizawa:2010:EDT

[314] Kei Uchizawa, Takao Nishizeki, and Eiji Takimoto. Energy and depth of threshold circuits. Theoretical Computer Science, 411(44–46):3938–3946, October 25, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huo:2010:FAA

[315] Yumei Huo and Joseph Y.-T. Leung. Fast approximation algorithms for job scheduling with processing set restrictions. Theoretical Computer Science, 411(44–46):3947–3955, October 25, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Han:2010:ORK

[316] Xin Han and Kazuhisa Makino. Online removable knapsack with limited cuts. Theoretical Computer Science, 411 (44–46):3956–3964, October 25, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Morales:2010:QLB

[317] Linda Morales and Hal Sudborough. A quadratic lower bound for Topswops. Theoretical Computer Science, 411 (44–46):3965–3970, October 25, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dereniowski:2010:PPH

[318] Dariusz Dereniowski. Phutball is PSPACE-hard. Theoretical Computer Science, 411(44–46):3971–3978, October 25, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ota:2010:PGR

[319] Shunsuke Ota, Ehab Morsy, and Hiroshi Nagamochi. A plane graph representation of triconnected graphs. Theoretical Computer Science, 411(44–46): 3979–3993, October 25, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Demetrescu:2010:APA

[320] Camil Demetrescu, Bruno Escoffier, Gabriel Moruz, and Andrea Ribichini.

Adapting parallel algorithms to the W-Stream model, with applications to graph problems. Theoretical Computer Science, 411(44–46):3994–4004, October 25, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2010:PAH

[321] Jianxin Wang, Wenjun Li, and Jianer Chen. A parameterized algorithm for the hyperplane-cover problem. Theoretical Computer Science, 411(44–46): 4005–4009, October 25, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jeandel:2010:PDP

[322] Emmanuel Jeandel. The periodic domino problem revisited. Theoretical Computer Science, 411(44–46):4010– 4016, October 25, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ackerman:2010:CMB

[323] Eyal Ackerman, Oren Ben-Zwi, and Guy Wolfovitz. Combinatorial model and bounds for target set selection. Theoretical Computer Science, 411 (44–46):4017–4022, October 25, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lozin:2010:DRD

[324] Vadim V. Lozin. A decidability result for the dominating set problem. Theoretical Computer Science, 411(44–46): 4023–4027, October 25, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grinchtein:2010:LER

[325] Olga Grinchtein, Bengt Jonsson, and Martin Leucker. Learning of eventrecording automata. Theoretical Computer Science, 411(47):4029–4054, October 29, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alpuente:2010:IFD

[326] M. Alpuente, D. Ballis, F. Correa, and M. Falaschi. An integrated framework for the diagnosis and correction of rule-based programs. Theoretical Computer Science, 411(47):4055– 4101, October 29, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Birkedal:2010:CTS

[327] Lars Birkedal, Kristian Støvring, and Jacob Thamsborg. The categorytheoretic solution of recursive metricspace equations. Theoretical Computer Science, 411(47):4102–4122, October 29, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanderMeyden:2010:CSM

[328] Ron van der Meyden and Chenyi Zhang. A comparison of semantic models for noninterference. Theoretical Computer Science, 411(47):4123– 4147, October 29, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonizzoni:2010:CRC

[329] Paola Bonizzoni, Clelia De Felice, and Rosalba Zizza. A characterization of (regular) circular languages generated by monotone complete splicing systems. Theoretical Computer Science, 411(48):4149–4161, November 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manea:2010:SAR

[330] Florin Manea. A series of algorithmic results related to the iterated hairpin completion. Theoretical Computer Science, 411(48):4162–4178, November 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anderson:2010:PIL

[331] Neal G. Anderson. On the physical implementation of logical transformations: Generalized L-machines. Theoretical Computer Science, 411(48): 4179–4199, November 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tan:2010:AMM

[332] Kok Kiong Tan, Jian Cheng Lv, Zhang Yi, and Sunan Huang. Adaptive multiple minor directions extraction in parallel using a PCA neural network. Theoretical Computer Science, 411(48): 4200–4215, November 1, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Darmann:2010:RAT

[333] Andreas Darmann, Ulrich Pferschy, and Joachim Schauer. Resource allocation with time intervals. Theoretical Computer Science, 411(49):4217– 4234, November 5, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2010:DSR

[334] J. J. Liu. Distinct squares in run-length encoded strings. Theoretical Computer Science, 411(49):4235–4241, November 5, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2010:GGR

[335] Huaming Zhang and Xin He. A generalized greedy routing algorithm for 2-connected graphs. Theoretical Computer Science, 411(49):4242–4252, November 5, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jiang:2010:PCS

[336] Minghui Jiang. On the parameterized complexity of some optimization problems related to multiple-interval graphs. Theoretical Computer Science, 411(49):4253–4262, November 5, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blakeley:2010:CDA

[337] Brandon Blakeley, F. Blanchet-Sadri, Josh Gunter, and Narad Rampersad. On the complexity of deciding avoidability of sets of partial words. Theoretical Computer Science, 411(49):4263– 4271, November 5, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DAgostino:2010:CTF

[338] Giovanna D’Agostino and Giacomo Lenzi. On the μ-calculus over transitive and finite transitive frames. Theoretical Computer Science, 411(50):4273– 4290, November 12, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lanotte:2010:WBP

[339] Ruggero Lanotte, Andrea Maggiolo-Schettini, and Angelo Troina. Weak bisimulation for Probabilistic Timed Automata. Theoretical Computer Science, 411(50):4291–4322, November 12, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rybakov:2010:RAT

[340] Vladimir Rybakov. Rules admissible in transitive temporal logic T_{S4}, sufficient condition. Theoretical Computer Science, 411(50):4323–4332, November 12, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Maric:2010:FVM

[341] Filip Marić. Formal verification of a modern SAT solver by shallow embedding into Isabelle/HOL. Theoretical Computer Science, 411(50):4333– 4356, November 12, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Boyland:2010:P

[342] John Tang Boyland and Giuseppe Castagna. Preface. Theoretical Computer Science, 411(51–52):4357, December 4, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sumii:2010:BLP

[343] Eijiro Sumii. A bisimulation-like proof method for contextual properties in untyped λ-calculus with references and deallocation. Theoretical Computer Science, 411(51–52):4358– 4378, December 4, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Filipovic:2010:ACO

[344] Ivana Filipović, Peter O’Hearn, Noam Rinetzky, and Hongseok Yang. Abstraction for concurrent objects. Theoretical Computer Science, 411(51–52): 4379–4398, December 4, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Caires:2010:CT

[345] Luís Caires and Hugo Torres Vieira. Conversation types. Theoretical Computer Science, 411(51–52):4399–4440, December 4, 2010. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jaskelioff:2010:MTM

[346] Mauro Jaskelioff and Eugenio Moggi. Monad transformers as monoid transformers. Theoretical Computer Science, 411(51–52):4441–4466, December 4, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Woods:2011:Pa

[347] Damien Woods, Turlough Neary, and Anthony K. Seda. Preface. Theoretical Computer Science, 412(1–2):1, January 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goles:2011:CCI

[348] E. Goles, P.-E. Meunier, I. Rapaport, and G. Theyssier. Communication complexity and intrinsic universality in cellular automata. Theoretical Computer Science, 412(1–2):2–21, January 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See erratum [951].

Ollinger:2011:FSE

[349] N. Ollinger and G. Richard. Four states are enough! Theoretical Computer Science, 412(1–2):22–32, January 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Margenstern:2011:UCA

[350] Maurice Margenstern. A universal cellular automaton on the heptagrid of the hyperbolic plane with four states. Theoretical Computer Science, 412(1–2): 33–56, January 2, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Durand-Lose:2011:AGC

[351] Jérôme Durand-Lose. Abstract geometrical computation 4: Small Turing universal signal machines. Theoretical Computer Science, 412(1–2):57– 67, January 2, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Okhotin:2011:SCP

[352] Alexander Okhotin. A simple P - complete problem and its languagetheoretic representations. Theoretical Computer Science, 412(1–2):68– 82, January 2, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Holzer:2011:CMH

[353] Markus Holzer, Martin Kutrib, and Andreas Malcher. Complexity of multihead finite automata: Origins and directions. Theoretical Computer Science, 412(1–2):83–96, January 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeMol:2011:CBS

[354] Liesbeth De Mol. On the complex behavior of simple tag systems — an experimental approach. Theoretical Computer Science, 412(1–2):97– 112, January 2, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cavaliere:2011:COS

[355] Matteo Cavaliere and Peter Leupold. Computing by observing: Simple systems and simple observers. Theoretical Computer Science, 412(1–2):113– 123, January 2, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Csuhaj-Varju:2011:GCS

[356] Erzsébet Csuhaj-Varjú and Sergey Verlan. On generalized communicating P systems with minimal interaction rules. Theoretical Computer Science, 412(1– 2):124–135, January 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alhazov:2011:SMI

[357] Artiom Alhazov, Alexander Krassovitskiy, Yurii Rogozhin, and Sergey Verlan. P systems with minimal insertion and deletion. Theoretical Computer Science, 412(1–2):136–144, January 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Doty:2011:LSA

[358] David Doty, Matthew J. Patitz, and Scott M. Summers. Limitations of selfassembly at temperature 1. Theoretical Computer Science, 412(1–2):145– 158, January 2, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Patitz:2011:SAI

[359] Matthew J. Patitz and Scott M. Summers. Self-assembly of infinite structures: a survey. Theoretical Computer Science, 412(1–2):159–165, January 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lutz:2011:DFR

[360] Jack H. Lutz. A divergence formula for randomness and dimension. Theoretical Computer Science, 412(1–2):166– 177, January 2, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calude:2011:SPU

[361] Cristian S. Calude. Simplicity via provability for universal prefix-free Turing machines. Theoretical Computer Science, 412(1–2):178–182, January 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sutner:2011:CPO

[362] Klaus Sutner. Computational processes, observers and Turing incompleteness. Theoretical Computer Science, 412(1–2):183–190, January 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Woods:2011:Pb

[363] Damien Woods, Turlough Neary, and Anthony K. Seda. Preface. Theoretical Computer Science, 412(1–2):v– x, January 2, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Du:2011:Pa

[364] Ding-Zhu Du, Xiaodong Hu, and Panos M. Pardalos. Preface. Theoretical Computer Science, 412(3):191, January 21, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Xu:2011:OAO

[365] Yinfeng Xu, Wenming Zhang, and Feifeng Zheng. Optimal algorithms for the online time series search problem. Theoretical Computer Science, 412(3): 192–197, January 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zou:2011:NAM

[366] Feng Zou, Yuexuan Wang, Xiao-Hua Xu, Xianyue Li, Hongwei Du, Pengjun Wan, and Weili Wu. New approximations for minimum-weighted dominating sets and minimum-weighted connected dominating sets on unit disk graphs. Theoretical Computer Science, 412(3):198–208, January 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2011:FTR

[367] Deying Li, Yuexuan Wang, Qinghua Zhu, and Huiqiang Yang. Faulttolerant routing: k-inconnected manyto-one routing in wireless networks. Theoretical Computer Science, 412(3): 209–216, January 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gao:2011:DLS

[368] Suogang Gao, Zengti Li, Jiangchen Yu, Xiaofeng Gao, and Weili Wu. DNA library screening, pooling design and unitary spaces. Theoretical Computer Science, 412(3):217–224, January 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ding:2011:LTA

[369] Wei Ding and Guoliang Xue. A linear time algorithm for computing a most reliable source on a tree network with faulty nodes. Theoretical Computer Science, 412(3):225–232, January 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ye:2011:OMS

[370] Deshi Ye, Xin Han, and Guochuan Zhang. Online multiple-strip packing. Theoretical Computer Science, 412(3): 233–239, January 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cai:2011:SCT

[371] Zhipeng Cai, Randy Goebel, and Guohui Lin. Size-constrained tree partitioning: Approximating the multicast k-tree routing problem. Theoretical Computer Science, 412(3):240– 245, January 21, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tan:2011:OHS

[372] Zhiyi Tan and An Zhang. Online hierarchical scheduling: an approach using mathematical programming. Theoretical Computer Science, 412(3):246– 256, January 21, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alvarez-Miranda:2011:DRC

[373] Eduardo Álvarez-Miranda, Xujin Chen, Jie Hu, Xiaodong Hu, and Alfredo Candia-Véjar. Deterministic risk control for cost-effective network connections. Theoretical Computer Science, 412(3):257–264, January 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2011:PID

[374] Feng Wang, Hongwei Du, Erika Camacho, Kuai Xu, Wonjun Lee, Yan Shi, and Shan Shan. On positive influence dominating sets in social networks. Theoretical Computer Science, 412(3): 265–269, January 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2011:IAA

[375] Peihai Liu and Xiwen Lu. An improved approximation algorithm for single machine scheduling with job delivery. Theoretical Computer Science, 412(3):270– 274, January 21, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fan:2011:SRD

[376] Baoqiang Fan, Shisheng Li, Li Zhou, and Liqi Zhang. Scheduling resumable deteriorating jobs on a single machine with non-availability constraints. Theoretical Computer Science, 412(4– 5):275–280, February 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2011:TCB

[377] Zhong Li and Guohui Lin. The three column Bandpass problem is solvable in linear time. Theoretical Computer Science, 412(4–5):281–299, February 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ceccherini-Silberstein:2011:RCI

[378] Tullio Ceccherini-Silberstein and Michel Coornaert. On the reversibility and the closed image property of linear cellular automata. Theoretical Computer Science, 412(4–5):300–306, February 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Okazaki:2011:BDT

[379] Ryotaro Okazaki, Hirotaka Ono, Taizo Sadahiro, and Masafumi Yamashita. Broadcastings and digit tilings on three-dimensional torus networks. Theoretical Computer Science, 412(4–5): 307–319, February 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alimonti:2011:LTA

[380] Paola Alimonti, Esteban Feuerstein, Luigi Laura, and Umberto Nanni. Linear time analysis of properties of conflict-free and general Petri nets. Theoretical Computer Science, 412(4– 5):320–338, February 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2011:MMD

[381] Li-Jui Chen, Jinn-Jy Lin, Min-Zheng Shieh, and Shi-Chun Tsai. More on the Magnus–Derek game. Theoretical Computer Science, 412(4–5):339– 344, February 4, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Thielen:2011:CTT

[382] Clemens Thielen and Stephan Westphal. Complexity of the traveling tournament problem. Theoretical Computer Science, 412(4–5):345–351, February 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Akutsu:2011:EAC

[383] Tatsuya Akutsu, Daiji Fukagawa, Atsuhiro Takasu, and Takeyuki Tamura. Exact algorithms for computing the tree edit distance between unordered trees. Theoretical Computer Science, 412(4–5):352–364, February 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kasperski:2011:ARS

[384] Adam Kasperski and Paweł Zieliński. On the approximability of robust spanning tree problems. Theoretical Computer Science, 412(4–5):365– 374, February 4, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Knauer:2011:MWD

[385] C. Knauer, K. Kriegel, and F. Stehn. Minimizing the weighted directed Hausdorff distance between colored point sets under translations and rigid motions. Theoretical Computer Science, 412(4–5):375–382, February 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nagamochi:2011:CRG

[386] Hiroshi Nagamochi. Cop–robber guarding game with cycle robberregion. Theoretical Computer Science, 412(4–5):383–390, February 4, 2011.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brandenburg:2011:SPM

[387] Franz J. Brandenburg and Mao cheng Cai. Shortest path and maximum flow problems in networks with additive losses and gains. Theoretical Computer Science, 412(4–5):391–401, February 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Paquin:2011:SCW

[388] Geneviève Paquin and Christophe Reutenauer. On the superimposition of Christoffel words. Theoretical Computer Science, 412(4–5):402– 418, February 4, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2011:IFP

[389] Jinshan Zhang and Fengshan Bai. An improved fully polynomial randomized approximation scheme (FPRAS) for counting the number of Hamiltonian cycles in dense digraphs. Theoretical Computer Science, 412(4–5):419– 429, February 4, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lin:2011:CAB

[390] Chun-Cheng Lin, Hsu-Chun Yen, Sheung-Hung Poon, and Jia-Hao Fan. Complexity analysis of balloon drawing for rooted trees. Theoretical Computer Science, 412(4–5):430–447, February 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nichols:2011:AMN

[391] Warren D. Nichols and Robert G. Underwood. Algebraic Myhill–Nerode Theorems. Theoretical Computer Science, 412(4–5):448–457, February 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pardubska:2011:PCG

[392] Dana Pardubská, Martin Plátek, and Friedrich Otto. Parallel communicating grammar systems with regular control and skeleton preserving FRR automata. Theoretical Computer Science, 412(4–5):458–477, February 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Monroe:2011:SNP

[393] Hunter Monroe. Speedup for natural problems and noncomputability. Theoretical Computer Science, 412(4–5): 478–481, February 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Andreev:2011:EDP

[394] Mikhail Andreev, Ilya Razenshteyn, and Alexander Shen. Not every domain of a plain decompressor contains the domain of a prefix-free one. Theoretical Computer Science, 412(4–5):482– 486, February 4, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ben-Amram:2011:SEU

[395] Amir Ben-Amram and Simon Yoffe. A simple and efficient Union–Find–Delete algorithm. Theoretical Computer Science, 412(4–5):487–492, February 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See corrigendum [1040].

Brlek:2011:CPD

[396] S. Brlek and C. Reutenauer. Complexity and palindromic defect of infinite words. Theoretical Computer Science, 412(4–5):493–497, February 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2011:Pa

[397] G. Rozenberg. Preface. Theoretical Computer Science, 412(6):499, February 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Timmis:2011:P

[398] Jon Timmis, Paul Andrews, and Andrew Hone. Preface. Theoretical Computer Science, 412(6):500–501, February 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

McEwan:2011:CS

[399] Chris McEwan and Emma Hart. On clonal selection. Theoretical Computer Science, 412(6):502–516, February 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jansen:2011:ADV

[400] Thomas Jansen and Christine Zarges. Analyzing different variants of immune inspired somatic contiguous hypermutations. Theoretical Computer Science, 412(6):517–533, February 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Elberfeld:2011:NSA

[401] Michael Elberfeld and Johannes Textor. Negative selection algorithms on strings with efficient training and linear-time classification. Theoretical Computer Science, 412(6):534–542, February 16, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jansen:2011:BDA

[402] Thomas Jansen and Christine Zarges. On benefits and drawbacks of aging strategies for randomized search heuristics. Theoretical Computer Science, 412(6):543–559, February 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBa

[403] Anonymous. Editorial Board. Theoretical Computer Science, 412(6):i– vi, February 16, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBb

[404] Anonymous. Editorial board. Theoretical Computer Science, 412(7):i– vi, February 25, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kaklamanis:2011:P

[405] Christos Kaklamanis and Martin Skutella. Preface. Theoretical Computer Science, 412(7):561–562, February 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bein:2011:RAT

[406] Wolfgang Bein, Kazuo Iwama, Jun Kawahara, Lawrence L. Larmore, and James A. Oravec. A randomized algorithm for two servers in cross polytope spaces. Theoretical Computer Science, 412(7):563–572, February 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Armon:2011:MMG

[407] Amitai Armon. On min–max rgatherings. Theoretical Computer Science, 412(7):573–582, February 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hurink:2011:IOA

[408] J. L. Hurink and J. J. Paulus. Improved online algorithms for parallel job scheduling and strip packing. Theoretical Computer Science, 412(7):583– 593, February 25, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chrobak:2011:BBI

[409] Marek Chrobak and Mathilde Hurand. Better bounds for incremental medians. Theoretical Computer Science, 412(7): 594–601, February 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krauthgamer:2011:PC

[410] Robert Krauthgamer, Aranyak Mehta, and Atri Rudra. Pricing commodities. Theoretical Computer Science, 412(7): 602–613, February 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dimitrov:2011:BSG

[411] Nedialko B. Dimitrov and C. Greg Plaxton. Buyer–supplier games: Optimization over the core. Theoretical Computer Science, 412(7):614–625, February 25, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gamzu:2011:ILB

[412] Iftah Gamzu. Improved lower bounds for non-utilitarian truthfulness. Theoretical Computer Science, 412(7):626– 632, February 25, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonifaci:2011:DWG

[413] Vincenzo Bonifaci, Peter Korteweg, Alberto Marchetti-Spaccamela, and Leen Stougie. The distributed wireless gathering problem. Theoretical Computer Science, 412(8–10):633–641, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dosa:2011:OSR

[414] György Dósa, Yuxin Wang, Xin Han, and He Guo. Online scheduling with rearrangement on two related machines. Theoretical Computer Science, 412(8– 10):642–653, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gray:2011:FCR

[415] R. Gray and A. Malheiro. Finite complete rewriting systems for regular semigroups. Theoretical Computer Science, 412(8–10):654–661, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yang:2011:CPK

[416] Guomin Yang and Chik How Tan. Certificateless public key encryption: a new generic construction and two pairing-free schemes. Theoretical Computer Science, 412(8–10):662–674, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brandt:2011:EGG

[417] Felix Brandt, Felix Fischer, and Markus Holzer. Equilibria of graphical games with symmetries. Theoretical Computer Science, 412(8–10): 675–685, March 4, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hsieh:2011:PRC

[418] Sun-Yuan Hsieh. On parallel recognition of cographs. Theoretical Computer Science, 412(8–10):686–694, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Labarre:2011:PTS

[419] Anthony Labarre and Josef Cibulka. Polynomial-time sortable stacks of burnt pancakes. Theoretical Computer Science, 412(8–10):695–702, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Betel:2011:RBF

[420] Heather Betel and Paola Flocchini. On the relationship between fuzzy and Boolean cellular automata. Theoretical Computer Science, 412(8–10):703– 713, March 4, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cui:2011:BID

[421] Bo Cui, Lila Kari, and Shinnosuke Seki. Block insertion and deletion on trajectories. Theoretical Computer Science, 412(8–10):714–728, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Larsson:2011:IDS

[422] Urban Larsson, Peter Hegarty, and Aviezri S. Fraenkel. Invariant and dual subtraction games resolving the Duchêne–Rigo conjecture. Theoretical Computer Science, 412(8–10):729– 735, March 4, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fulop:2011:VRT

[423] Zoltán Fülöp and Magnus Steinby. Varieties of recognizable tree series over fields. Theoretical Computer Science, 412(8–10):736–752, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ouangraoua:2011:NAA

[424] Aïda Ouangraoua, Valentin Guignon, Sylvie Hamel, and Cedric Chauve. A new algorithm for aligning nested arcannotated sequences under arbitrary weight schemes. Theoretical Computer Science, 412(8–10):753–764, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beaudry:2011:SIS

[425] Martin Beaudry and Markus Holzer. On the size of inverse semigroups given by generators. Theoretical Computer Science, 412(8–10):765–772, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Uchizawa:2011:SET

[426] Kei Uchizawa, Eiji Takimoto, and Takao Nishizeki. Size–energy tradeoffs for unate circuits computing symmetric Boolean functions. Theoretical Computer Science, 412(8–10):773– 782, March 4, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Elor:2011:UMA

[427] Yotam Elor and Alfred M. Bruckstein. Uniform multi-agent deployment on a ring. Theoretical Computer Science, 412(8–10):783–795, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hwang:2011:NIB

[428] Jung Yeon Hwang. A note on an identity-based ring signature scheme with signer verifiability. Theoretical Computer Science, 412(8–10):796– 804, March 4, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Esik:2011:BCF

[429] Zoltán Ésik and Szabolcs Iván. Büchi context-free languages. Theoretical Computer Science, 412(8–10):805–821, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cibulka:2011:AHN

[430] Josef Cibulka. On average and highest number of flips in pancake sorting. Theoretical Computer Science, 412(8–10): 822–834, March 4, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Masakova:2011:ANS

[431] Z. Masáková, E. Pelantová, and T. Vávra. Arithmetics in number systems with a negative base. Theoretical Computer Science, 412(8–10):835– 845, March 4, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dragan:2011:AMW

[432] Feodor F. Dragan, Fedor V. Fomin, and Petr A. Golovach. Approximation of minimum weight spanners for sparse graphs. Theoretical Computer Science, 412(8–10):846–852, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Glasser:2011:SPN

[433] Christian Glaßer, Christian Reitwießner, and Victor Selivanov. The shrinking property for NP and coNP. Theoretical Computer Science, 412(8– 10):853–864, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2011:FGM

[434] Zhongzhi Zhang and Francesc Comellas. Farey graphs as models for complex networks. Theoretical Computer Science, 412(8–10):865–875, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chuan:2011:WRO

[435] Wai-Fong Chuan, Chun-Yu Chen, Hui-Ling Ho, and Fang-Yi Liao. αwords and the radix order. Theoretical Computer Science, 412(8–10):876– 891, March 4, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pons:2011:PPH

[436] Pascal Pons and Matthieu Latapy. Post-processing hierarchical community structures: Quality improvements and multi-scale view. Theoretical Computer Science, 412(8–10):892– 900, March 4, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nutov:2011:ADW

[437] Zeev Nutov. Approximating directed weighted-degree constrained networks. Theoretical Computer Science, 412(8– 10):901–912, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2011:SRO

[438] Weifan Wang, Xubin Yue, and Xuding Zhu. The surviving rate of an outerplanar graph for the firefighter problem. Theoretical Computer Science, 412(8– 10):913–921, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marschall:2011:CMD

[439] Tobias Marschall. Construction of minimal deterministic finite automata from biological motifs. Theoretical Computer Science, 412(8–10):922–930, March 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aguado:2011:CSI

[440] Joaquín Aguado and Michael Mendler. Constructive semantics for instantaneous reactions. Theoretical Computer Science, 412(11):931–961, March 11, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dominguez:2011:EHB

[441] César Domínguez and Julio Rubio. Effective homology of bicomplexes, formalized in Coq. Theoretical Computer Science, 412(11):962–970, March 11, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Japaridze:2011:TOC

[442] Giorgi Japaridze. Toggling operators in computability logic. Theoretical Computer Science, 412(11):971– 1004, March 11, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Afrati:2011:DQR

[443] Foto N. Afrati. Determinacy and query rewriting for conjunctive queries and views. Theoretical Computer Science, 412(11):1005–1021, March 11, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

David:2011:SPA

[444] René David. A short proof that adding some permutation rules to β preserves SN. Theoretical Computer Science, 412(11):1022–1026, March 11, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Santo:2011:NPS

[445] José Espírito Santo. A note on preservation of strong normalisation in the λ-calculus. Theoretical Computer Science, 412(11):1027–1032, March 11, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Levin:2011:UUS

[446] Asaf Levin and Uri Yovel. Uniform unweighted set cover: The power of non-oblivious local search. Theoretical Computer Science, 412(12–14):1033– 1053, March 18, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ito:2011:CRP

[447] Takehiro Ito, Erik D. Demaine, Nicholas J. A. Harvey, Christos H. Papadimitriou, Martha Sideri, Ryuhei Uehara, and Yushi Uno. On the complexity of reconfiguration problems. Theoretical Computer Science, 412(12–14):1054–1065, March 18, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baldoni:2011:IMG

[448] R. Baldoni, A. Fernández Anta, K. Ioannidou, and A. Milani. The impact of mobility on the geocasting problem in mobile ad-hoc networks: Solvability and cost. Theoretical Computer Science, 412(12–14):1066– 1080, March 18, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Huo:2011:BSC

[449] Yumei Huo and Hairong Zhao. Bicriteria scheduling concerned with makespan and total completion time subject to machine availability constraints. Theoretical Computer Science, 412(12–14):1081–1091, March 18, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hou:2011:OSO

[450] Li ying Hou and Liying Kang. Online and semi-online hierarchical scheduling for load balancing on uniform machines. Theoretical Computer Science, 412(12–14):1092–1098, March 18, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bader:2011:TMP

[451] Martin Bader. The transposition median problem is NP-complete. Theoretical Computer Science, 412(12–14): 1099–1110, March 18, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Starosta:2011:TPR

[452] Štěpán Starosta. On theta-palindromic richness. Theoretical Computer Science, 412(12–14):1111–1121, March 18, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wu:2011:MES

[453] Weiwei Wu, Minming Li, and Enhong Chen. Min-energy scheduling for aligned jobs in accelerate model. Theoretical Computer Science, 412(12–14): 1122–1139, March 18, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Droste:2011:KST

[454] Manfred Droste and Karin Quaas. A Kleene–Schützenberger theorem for weighted timed automata. Theoretical Computer Science, 412(12–14):1140– 1153, March 18, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tracol:2011:RTF

[455] Mathieu Tracol. Recurrence and transience for finite probabilistic tables. Theoretical Computer Science, 412(12–14):1154–1168, March 18, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kountouriotis:2011:GTC

[456] Vassilis Kountouriotis, Christos Nomikos, and Panos Rondogiannis. A gametheoretic characterization of Boolean grammars. Theoretical Computer Science, 412(12–14):1169–1183, March 18, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bocker:2011:CBO

[457] S. Böcker, Q. B. A. Bui, and A. Truss. Computing bond orders in molecule graphs. Theoretical Computer Science, 412(12–14):1184–1195, March 18, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ben-Zwi:2011:LGP

[458] Oren Ben-Zwi and Amir Ronen. Local and global price of anarchy of graphical games. Theoretical Computer Science, 412(12–14):1196–1207, March 18, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yang:2011:FES

[459] Boting Yang. Fast edge searching and fast searching on graphs. Theoretical Computer Science, 412(12–14):1208– 1219, March 18, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Libert:2011:ETS

[460] Benoît Libert and Moti Yung. Efficient traceable signatures in the standard model. Theoretical Computer Science, 412(12–14):1220–1242, March 18, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Attiya:2011:HCM

[461] Hagit Attiya and Eshcar Hillel. Highly concurrent multi-word synchronization. Theoretical Computer Science, 412(12–14):1243–1262, March 18, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kavitha:2011:PMV

[462] Telikepalli Kavitha and Meghana Nasre. Popular matchings with variable item copies. Theoretical Computer Science, 412(12–14):1263–1274, March 18, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heggernes:2011:CMD

[463] Pinar Heggernes, Daniel Meister, and Andrzej Proskurowski. Computing minimum distortion embeddings into a path for bipartite permutation graphs and threshold graphs. Theoretical Computer Science, 412(12–14):1275– 1297, March 18, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brimkov:2011:P

[464] Valentin E. Brimkov, Reneta P. Barneva, and Petra Wiederhold. Preface. Theoretical Computer Science, 412(15):1299–1300, March 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sagols:2011:TDV

[465] Feliú Sagols and Raúl Marín. Two discrete versions of the Inscribed Square Conjecture and some related problems. Theoretical Computer Science, 412(15): 1301–1312, March 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brimkov:2011:GSL

[466] Valentin E. Brimkov, Andrew Leach, Michael Mastroianni, and Jimmy Wu. Guarding a set of line segments in the plane. Theoretical Computer Science, 412(15):1313–1324, March 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Borghi:2011:EOC

[467] Alexandre Borghi, Jérôme Darbon, and Sylvain Peyronnet. Exact optimization for the ℓ^{1}-Compressive Sensing problem using a modified Dantzig–Wolfe method. Theoretical Computer Science, 412(15):1325–1337, March 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Malmberg:2011:GBF

[468] F. Malmberg, J. Lindblad, N. Sladoje, and I. Nyström. A graph-based framework for sub-pixel image segmentation. Theoretical Computer Science, 412(15): 1338–1349, March 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Strand:2011:DDF

[469] Robin Strand, Benedek Nagy, and Gunilla Borgefors. Digital distance functions on three-dimensional grids. Theoretical Computer Science, 412(15):

1350–1363, March 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Nagyand:2011:AEC

[470] Benedek Nagyand and Robin Strand. Approximating Euclidean circles by neighbourhood sequences in a hexagonal grid. Theoretical Computer Science, 412(15):1364–1377, March 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Thibault:2011:DRU

[471] Yohan Thibault, Akihiro Sugimoto, and Yukiko Kenmochi. 3D discrete rotations using hinge angles. Theoretical Computer Science, 412(15):1378– 1391, March 25, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gosselin:2011:ESC

[472] Stéphane Gosselin, Guillaume Damiand, and Christine Solnon. Efficient search of combinatorial maps using signatures. Theoretical Computer Science, 412(15): 1392–1405, March 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Le:2011:SDG

[473] Phuc Q. Le, Abdullahi M. Iliyasu, Fangyan Dong, and Kaoru Hirota. Strategies for designing geometric transformations on quantum images. Theoretical Computer Science, 412(15): 1406–1418, March 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Xie:2011:OAP

[474] Xie Xie, Jiu-Lun Fan, and Yin Zhu. The optimal All-Partial-Sums algorithm in commutative semigroups and its applications for image thresholding segmentation. Theoretical Computer Science, 412(15):1419–1433, March 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dalmau:2011:AMM

[475] Oscar Dalmau and Mariano Rivera. Alpha Markov Measure Field model for probabilistic image segmentation. Theoretical Computer Science, 412(15): 1434–1441, March 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kotsireas:2011:P

[476] Ilias S. Kotsireas, Bernard Mourrain, and Victor Y. Pan. Preface. Theoretical Computer Science, 412(16):1443– 1444, April 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cuyt:2011:SIM

[477] Annie Cuyt and Wen shin Lee. Sparse interpolation of multivariate rational functions. Theoretical Computer Science, 412(16):1445–1456, April 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pebay:2011:OVN

[478] Philippe Pébay, J. Maurice Rojas, and David C. Thompson. Optimizing nvariate (n + k)-nomials for small k. Theoretical Computer Science, 412(16): 1457–1469, April 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frandsen:2011:DNF

[479] Gudmund Skovbjerg Frandsen and Piotr Sankowski. Dynamic normal forms and dynamic characteristic polynomial. Theoretical Computer Science, 412(16): 1470–1483, April 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kourniotis:2011:PPG

[480] M. Kourniotis, M. Mitrouli, and D. Triantafyllou. Parallel QR processing of Generalized Sylvester matrices. Theoretical Computer Science, 412(16): 1484–1491, April 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galligo:2011:CMC

[481] André Galligo and Adrien Poteaux. Computing monodromy via continuation methods on random Riemann surfaces. Theoretical Computer Science, 412(16):1492–1507, April 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hoppen:2011:NGA

[482] Carlos Hoppen, Virginia M. Rodrigues, and Vilmar Trevisan. A note on Gao’s algorithm for polynomial factorization. Theoretical Computer Science, 412(16): 1508–1522, April 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chevillard:2011:EAC

[483] S. Chevillard, J. Harrison, M. Joldeş, and Ch. Lauter. Efficient and accurate computation of upper bounds of approximation errors. Theoretical Computer Science, 412(16):1523– 1543, April 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2011:Pb

[484] G. Rozenberg. Preface. Theoretical Computer Science, 412(17):1545, April 8, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Friedrich:2011:IFE

[485] Tobias Friedrich, Christian Horoba, and Frank Neumann. Illustration of fairness in evolutionary multi-objective optimization. Theoretical Computer Science, 412(17):1546–1556, April 8, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

McCaig:2011:IPM

[486] Chris McCaig, Rachel Norman, and Carron Shankland. From individuals to populations: a mean field semantics for process algebra. Theoretical Computer Science, 412(17):1557–1580, April 8, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Alhazov:2011:MSM

[487] Artiom Alhazov and Sergey Verlan. Minimization strategies for maximally parallel multiset rewriting systems. Theoretical Computer Science, 412(17): 1581–1591, April 8, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Reif:2011:CGS

[488] John H. Reif, Sudheer Sahu, and Peng Yin. Complexity of graph selfassembly in accretive systems and self-destructible systems. Theoretical Computer Science, 412(17):1592– 1605, April 8, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yang:2011:RSR

[489] Guowu Yang, Fei Xie, William N. N. Hung, Xiaoyu Song, and Marek A. Perkowski. Realization and synthesis of reversible functions. Theoretical Computer Science, 412(17):1606– 1613, April 8, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jiang:2011:FLD

[490] Pei Jiang and Ying ping Chen. Free lunches on the discrete Lipschitz class. Theoretical Computer Science, 412(17): 1614–1628, April 8, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Doerr:2011:RAA

[491] Benjamin Doerr, Frank Neumann, Dirk Sudholt, and Carsten Witt. Runtime analysis of the 1-ANT ant colony optimizer. Theoretical Computer Science, 412(17):1629–1644, April 8, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBc

[492] Anonymous. Editorial Board. Theoretical Computer Science, 412(17):iii– viii, April 8, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Agrawal:2011:TAM

[493] Manindra Agrawal and Angsheng Li. Theory and applications of models of computation (TAMC 2008). Theoretical Computer Science, 412(18):

1645, April 15, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amano:2011:WMF

[494] Kazuyuki Amano and Jun Tarui. A well-mixed function with circuit complexity 5n: Tightness of the Lachish–Raz-type bounds. Theoretical Computer Science, 412(18):1646– 1651, April 15, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ambos-Spies:2011:IIC

[495] Klaus Ambos-Spies, Serikzhan Badaev, and Sergey Goncharov. Inductive inference and computable numberings. Theoretical Computer Science, 412(18): 1652–1668, April 15, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arslanov:2011:SNE

[496] M. M. Arslanov, S. B. Cooper, I. Sh. Kalimullin, and M. I. Soskova. Splitting and nonsplitting in the Σ_{2}^{0} enumeration degrees. Theoretical Computer Science, 412(18):1669–1685, April 15, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bollig:2011:OCM

[497] Beate Bollig. On the OBDD complexity of the most significant bit of integer multiplication. Theoretical Computer Science, 412(18):1686–1695, April 15, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Czyzowicz:2011:LCP

[498] J. Czyzowicz, S. Dobrev, H. González-Aguilar, R. Kralovic, E. Kranakis, J. Opatrny, L. Stacho, and J. Urrutia. Local 7-coloring for planar subgraphs of unit disk graphs. Theoretical Computer Science, 412(18):1696– 1704, April 15, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2011:DIP

[499] Jiamou Liu and Mia Minnes. Deciding the isomorphism problem in classes of unary automatic structures. Theoretical Computer Science, 412(18):1705– 1717, April 15, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tang:2011:DGT

[500] Linqing Tang. On the derandomization of the graph test for homomorphism over groups. Theoretical Computer Science, 412(18):1718–1728, April 15, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tian:2011:EPP

[501] Cong Tian and Zhenhua Duan. Expressiveness of propositional projection temporal logic with star. Theoretical Computer Science, 412(18):1729– 1744, April 15, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2011:TPA

[502] Jianxin Wang, Qilong Feng, and Jianer Chen. An O ∗ (3.53^{3k})-time parameterized algorithm for the 3-set packing problem. Theoretical Computer Science, 412(18):1745–1753, April 15, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBd

[503] Anonymous. Editorial board. Theoretical Computer Science, 412(18):iii– viii, April 15, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lugosi:2011:P

[504] Gábor Lugosi and Sandra Zilles. Preface. Theoretical Computer Science, 412 (19):1755, April 22, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Darnstadt:2011:SPL

[505] Malte Darnstädt and Hans Ulrich Simon. Smart PAC-learners. Theoretical Computer Science, 412(19):1756– 1766, April 22, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dasgupta:2011:TFA

[506] Sanjoy Dasgupta. Two faces of active learning. Theoretical Computer Science, 412(19):1767–1781, April 22, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bshouty:2011:RWG

[507] Nader H. Bshouty and Hanna Mazzawi. Reconstructing weighted graphs with minimal query complexity. Theoretical Computer Science, 412(19):1782– 1790, April 22, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cesa-Bianchi:2011:PLU

[508] Nicolò Cesa-Bianchi, Claudio Gentile, and Fabio Vitale. Predicting the labels of an unknown graph via adaptive exploration. Theoretical Computer Science, 412(19):1791–1804, April 22, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jain:2011:UAC

[509] Sanjay Jain, Qinglong Luo, Pavel Semukhin, and Frank Stephan. Uncountable automatic classes and learning. Theoretical Computer Science, 412(19): 1805–1820, April 22, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yoshinaka:2011:ELM

[510] Ryo Yoshinaka. Efficient learning of multiple context-free languages with multidimensional substitutability from positive data. Theoretical Computer Science, 412(19):1821–1831, April 22, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bubeck:2011:PEF

[511] Sébastien Bubeck, Rémi Munos, and Gilles Stoltz. Pure exploration in finitely-armed and continuous-armed bandits. Theoretical Computer Science, 412(19):1832–1852, April 22, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBe

[512] Anonymous. Editorial board. Theoretical Computer Science, 412(19):iii– viii, April 22, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Curien:2011:P

[513] Pierre-Louis Curien. Preface. Theoretical Computer Science, 412(20):1853– 1859, April 29, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Girard:2011:GIV

[514] Jean-Yves Girard. Geometry of Interaction V: Logic in the hyperfinite factor. Theoretical Computer Science, 412(20):1860–1883, April 29, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deCarvalho:2011:SME

[515] D. de Carvalho, M. Pagani, and L. Tortora de Falco. A semantic measure of the execution time in linear logic. Theoretical Computer Science, 412(20): 1884–1902, April 29, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Paolini:2011:SNU

[516] Luca Paolini, Elaine Pimentel, and Simona Ronchi Della Rocca. Strong normalization from an unusual point of view. Theoretical Computer Science, 412(20):1903–1915, April 29, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vandenBerg:2011:APA

[517] Benno van den Berg and Ieke Moerdijk. Aspects of predicative algebraic set theory, II: Realizability. Theoretical Computer Science, 412(20):1916– 1940, April 29, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deNaurois:2011:CLL

[518] Paulin Jacobé de Naurois and Virgile Mogbil. Correctness of linear logic proof structures is NL-complete. Theoretical Computer Science, 412(20):

1941–1957, April 29, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Guerrini:2011:LAM

[519] Stefano Guerrini. A linear algorithm for MLL proof net correctness and sequentialization. Theoretical Computer Science, 412(20):1958–1978, April 29, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tranquilli:2011:IDN

[520] Paolo Tranquilli. Intuitionistic differential nets and lambda-calculus. Theoretical Computer Science, 412(20):1979– 1997, April 29, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goubault-Larrecq:2011:MAG

[521] Jean Goubault-Larrecq. Musings around the geometry of interaction, and coherence. Theoretical Computer Science, 412(20):1998–2014, April 29, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Haghverdi:2011:TGR

[522] Esfandiar Haghverdi. Towards a geometry of recursion. Theoretical Computer Science, 412(20):2015–2028, April 29, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DalLago:2011:RMI

[523] Ugo Dal Lago and Martin Hofmann. Realizability models and implicit complexity. Theoretical Computer Science, 412(20):2029–2047, April 29, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Terui:2011:CL

[524] Kazushige Terui. Computational ludics. Theoretical Computer Science, 412(20):2048–2071, April 29, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kanovich:2011:LLT

[525] Max Kanovich and Jacqueline Vauzeilles. Linear logic as a tool for planning under temporal uncertainty. Theoretical Computer Science, 412(20): 2072–2092, April 29, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBf

[526] Anonymous. Editorial board. Theoretical Computer Science, 412(20):iii– viii, April 29, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2011:Pc

[527] G. Rozenberg. Preface. Theoretical Computer Science, 412(21):2093, May 6, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Degano:2011:P

[528] Pierpaolo Degano. Preface. Theoretical Computer Science, 412(21):2094, May 6, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dang:2011:CRS

[529] Thao Dang, Colas Le Guernic, and Oded Maler. Computing reachable states for nonlinear biological models. Theoretical Computer Science, 412(21): 2095–2107, May 6, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeMaria:2011:DOP

[530] Elisabetta De Maria, François Fages, Aurélien Rizk, and Sylvain Soliman. Design, optimization and predictions of a coupled model of the cell cycle, circadian clock, DNA repair system, irinotecan metabolism and exposure control under temporal logic constraints. Theoretical Computer Science, 412(21):2108–2127, May 6, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Didier:2011:AEP

[531] Frédéric Didier, Thomas A. Henzinger, Maria Mateescu, and Verena Wolf. Approximation of event probabilities in noisy cellular processes. Theoretical Computer Science, 412(21):2128– 2141, May 6, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Galpin:2011:SEB

[532] Vashti Galpin and Jane Hillston. A semantic equivalence for Bio-PEPA based on discretisation of continuous values. Theoretical Computer Science, 412(21):2142–2161, May 6, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jha:2011:SIA

[533] Sumit Kumar Jha and Christopher James Langmead. Synthesis and infeasibility analysis for stochastic models of biochemical systems using statistical model checking and abstraction refinement. Theoretical Computer Science, 412(21):2162–2187, May 6, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2011:PAO

[534] Bing Liu, David Hsu, and P. S. Thiagarajan. Probabilistic approximations of ODEs based bio-pathway dynamics. Theoretical Computer Science, 412(21): 2188–2206, May 6, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Naldi:2011:DCR

[535] Aurélien Naldi, Elisabeth Remy, Denis Thieffry, and Claudine Chaouiya. Dynamically consistent reduction of logical regulatory graphs. Theoretical Computer Science, 412(21):2207– 2218, May 6, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:P

[536] Anonymous. Preface. Theoretical Computer Science, 412(21):iii–viii, May 6, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Drevet:2011:OTS

[537] Charles-Éric Drevet, Md. Nazrul Islam, and Éric Schost. Optimization techniques for small matrix multiplication. Theoretical Computer Science, 412(22): 2219–2236, May 13, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2011:CIS

[538] Yi-Jiun Liu, James K. Lan, Well Y. Chou, and Chiuyuan Chen. Constructing independent spanning trees for locally twisted cubes. Theoretical Computer Science, 412(22):2237– 2252, May 13, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calude:2011:URE

[539] Cristian S. Calude, André Nies, Ludwig Staiger, and Frank Stephan. Universal recursively enumerable sets of strings. Theoretical Computer Science, 412(22): 2253–2261, May 13, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aupy:2011:NBM

[540] Guillaume Aupy and Olivier Bournez. On the number of binary-minded individuals required to compute \sqrt{1/2}. Theoretical Computer Science, 412(22): 2262–2267, May 13, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Allouche:2011:IFI

[541] Jean-Paul Allouche and Laurence Maillard-Teyssier. Inconstancy of finite and infinite sequences. Theoretical Computer Science, 412(22):2268– 2281, May 13, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Levi:2011:BSM

[542] Tamir Levi and Ami Litman. Bitonic sorters of minimal depth. Theoretical Computer Science, 412(22):2282– 2295, May 13, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bilo:2011:ENR

[543] Vittorio Bilò and Michele Flammini. Extending the notion of rationality of selfish agents: Second Order Nash equilibria. Theoretical Computer Science, 412(22):2296–2311, May 13, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mantzaflaris:2011:CFE

[544] Angelos Mantzaflaris, Bernard Mourrain, and Elias Tsigaridas. On continued fraction expansion of real roots of polynomial systems, complexity and condition numbers. Theoretical Computer Science, 412(22):2312–2330, May 13, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Puente:2011:LMD

[545] Víctor Lavín Puente. On learning multivalued dependencies with queries. Theoretical Computer Science, 412(22): 2331–2339, May 13, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mitankin:2011:DWN

[546] Petar Mitankin, Stoyan Mihov, and Klaus U. Schulz. Deciding word neighborhood with universal neighborhood automata. Theoretical Computer Science, 412(22):2340–2355, May 13, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bedon:2011:SPL

[547] Nicolas Bedon and Chloé Rispal. Series-parallel languages on scattered and countable posets. Theoretical Computer Science, 412(22):2356–2369, May 13, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

deSa:2011:CDP

[548] Vinícius G. P. de Sá, Guilherme D. da Fonseca, Raphael C. S. Machado, and Celina M. H. de Figueiredo. Complexity dichotomy on partial grid recognition. Theoretical Computer Science, 412(22):2370–2379, May 13, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tian:2011:OSU

[549] Ji Tian, T. C. E. Cheng, C. T. Ng, and Jinjiang Yuan. Online scheduling on unbounded parallel-batch machines with incompatible job families. Theoretical Computer Science, 412(22): 2380–2386, May 13, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhao:2011:ACM

[550] Yancai Zhao, Liying Kang, and Moo Young Sohn. The algorithmic complexity of mixed domination in graphs. Theoretical Computer Science, 412(22):2387–2392, May 13, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Duan:2011:NUB

[551] Ziming Duan, Pingli Lv, Lianying Miao, Zhengke Miao, and Cuiqi Wang. New upper bounds on the L(2, 1)- labeling of the skew and converse skew product graphs. Theoretical Computer Science, 412(22):2393–2397, May 13, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Even:2011:PRL

[552] Guy Even and Moti Medina. Parallel randomized load balancing: a lower bound for a more general model. Theoretical Computer Science, 412(22): 2398–2408, May 13, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yang:2011:HPT

[553] Xiaofan Yang, Qiang Dong, Erjie Yang, and Jianqiu Cao. Hamiltonian properties of twisted hypercube-like networks with more faulty elements. Theoretical Computer Science, 412(22):2409– 2417, May 13, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2011:NAE

[554] Tseng-Kui Li, Chia-Jui Lai, and Chang-Hsiung Tsai. A novel algorithm to embed a multi-dimensional torus into a locally twisted cube. Theoretical Computer Science, 412(22):2418– 2424, May 13, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Catusse:2011:ERP

[555] Nicolas Catusse, Victor Chepoi, and Yann Vaxès. Embedding into the rectilinear plane in optimal O(n^{2}) time. Theoretical Computer Science, 412(22): 2425–2433, May 13, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Michail:2011:MPP

[556] Othon Michail, Ioannis Chatzigiannakis, and Paul G. Spirakis. Mediated population protocols. Theoretical Computer Science, 412(22):2434– 2450, May 13, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fischer:2011:CDS

[557] Johannes Fischer. Combined data structure for previous- and nextsmaller-values. Theoretical Computer Science, 412(22):2451–2456, May 13, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBg

[558] Anonymous. Editorial Board. Theoretical Computer Science, 412(22):iii– viii, May 13, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2011:ACC

[559] Jianer Chen and S. Barry Cooper. Algorithms, complexity and computational models. Theoretical Computer Science, 412(23):2457–2458, May 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bereg:2011:RBS

[560] Sergey Bereg, Minghui Jiang, Boting Yang, and Binhai Zhu. On the red/ blue spanning tree problem. Theoretical Computer Science, 412(23):2459– 2467, May 20, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cai:2011:CPC

[561] Jin-Yi Cai, Pinyan Lu, and Mingji Xia. A computational proof of complexity of some restricted counting problems. Theoretical Computer Science, 412(23):

2468–2485, May 20, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carbone:2011:CEI

[562] A. Carbone and L. Dib. Co-evolution and information signals in biological sequences. Theoretical Computer Science, 412(23):2486–2495, May 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chang:2011:BGW

[563] Maw-Shang Chang, Ling-Ju Hung, Ton Kloks, and Sheng-Lung Peng. Blockgraph width. Theoretical Computer Science, 412(23):2496–2502, May 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2011:IDA

[564] Jianer Chen, Qilong Feng, Yang Liu, Songjian Lu, and Jianxin Wang. Improved deterministic algorithms for weighted matching and packing problems. Theoretical Computer Science, 412(23):2503–2512, May 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fiala:2011:PCC

[565] Jirí Fiala, Petr A. Golovach, and Jan Kratochvíl. Parameterized complexity of coloring problems: Treewidth versus vertex cover. Theoretical Computer Science, 412(23):2513–2523, May 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kaplan:2011:KCC

[566] Marc Kaplan and Sophie Laplante. Kolmogorov complexity and combinatorial methods in communication complexity. Theoretical Computer Science, 412(23):2524–2535, May 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lokshtanov:2011:LKP

[567] Daniel Lokshtanov, Matthias Mnich, and Saket Saurabh. A linear kernel for a planar connected dominating set. Theoretical Computer Science, 412(23): 2536–2543, May 20, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sarma:2011:BOS

[568] Atish Das Sarma, Richard J. Lipton, and Danupon Nanongkai. Bestorder streaming model. Theoretical Computer Science, 412(23):2544– 2555, May 20, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanZuylen:2011:LPB

[569] Anke van Zuylen. Linear programming based approximation algorithms for feedback set problems in bipartite tournaments. Theoretical Computer Science, 412(23):2556–2561, May 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBh

[570] Anonymous. Editorial Board. Theoretical Computer Science, 412(23):iii– viii, May 20, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Albers:2011:P

[571] Susanne Albers and Alberto Marchetti-Spaccamela. Preface. Theoretical Computer Science, 412(24):2563, May 27, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anagnostopoulos:2011:SSD

[572] Aris Anagnostopoulos, Ravi Kumar, Mohammad Mahdian, and Eli Upfal. Sorting and selection on dynamic data. Theoretical Computer Science, 412(24): 2564–2576, May 27, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Azar:2011:MBF

[573] Yossi Azar, Aleksander Madry, Thomas Moscibroda, Debmalya Panigrahi, and Aravind Srinivasan. Maximum bipartite flow in networks with adaptive channel width. Theoretical Computer Science, 412(24):2577–2587, May 27, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brodal:2011:TOR

[574] Gerth Stølting Brodal, Beat Gfeller, Allan Grønlund Jørgensen, and Peter Sanders. Towards optimal range medians. Theoretical Computer Science, 412(24):2588–2601, May 27, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chierichetti:2011:RSS

[575] Flavio Chierichetti, Silvio Lattanzi, and Alessandro Panconesi. Rumor spreading in social networks. Theoretical Computer Science, 412(24):2602– 2610, May 27, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Draisma:2011:PAM

[576] Jan Draisma, Eyal Kushilevitz, and Enav Weinreb. Partition arguments in multiparty communication complexity. Theoretical Computer Science, 412(24): 2611–2622, May 27, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Elsasser:2011:TBC

[577] Robert Elsässer and Thomas Sauerwald. Tight bounds for the cover time of multiple random walks. Theoretical Computer Science, 412(24):2623– 2641, May 27, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Emek:2011:OCA

[578] Yuval Emek, Pierre Fraigniaud, Amos Korman, and Adi Rosén. Online computation with advice. Theoretical Computer Science, 412(24):2642– 2656, May 27, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fanghanel:2011:IAL

[579] Alexander Fanghänel, Thomas Kesselheim, and Berthold Vöcking. Improved algorithms for latency minimization in wireless networks. Theoretical Computer Science, 412(24):2657– 2667, May 27, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Farzan:2011:SRD

[580] Arash Farzan and J. Ian Munro. Succinct representation of dynamic trees. Theoretical Computer Science, 412(24): 2668–2678, May 27, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kavitha:2011:PMM

[581] Telikepalli Kavitha, Julián Mestre, and Meghana Nasre. Popular mixed matchings. Theoretical Computer Science, 412(24):2679–2690, May 27, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2011:PRD

[582] Li Zhang. Proportional response dynamics in the Fisher market. Theoretical Computer Science, 412(24):2691– 2698, May 27, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fomin:2011:SIT

[583] Fedor V. Fomin, Pierre Fraigniaud, Stephan Kreutzer, and Dimitrios M. Thilikos. Special issue on “Theory and Applications of Graph Searching Problems”. Theoretical Computer Science, 412(24):2699, May 27, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBi

[584] Anonymous. Editorial Board. Theoretical Computer Science, 412(24):iii– viii, May 27, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Keimel:2011:RDP

[585] K. Keimel, A. Rosenbusch, and T. Streicher. Relating direct and predicate transformer partial correctness semantics for an imperative probabilisticnondeterministic language. Theoretical Computer Science, 412(25):2701– 2713, June 3, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gehrke:2011:CEC

[586] Mai Gehrke and Jacob Vosmaer. Canonical extensions and canonicity via dcpo presentations. Theoretical Computer Science, 412(25):2714–2723, June 3, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abraham:2011:LCD

[587] Uri Abraham. Logical classification of distributed algorithms (Bakery algorithms as an example). Theoretical Computer Science, 412(25):2724– 2745, June 3, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2011:NFS

[588] Yongjian Li, William N. N. Hung, and Xiaoyu Song. A novel formalization of symbolic trajectory evaluation semantics in Isabelle/HOL. Theoretical Computer Science, 412(25):2746–2765, June 3, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calvert:2011:MSP

[589] Wesley Calvert. Metric structures and probabilistic computation. Theoretical Computer Science, 412(25):2766– 2775, June 3, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonfante:2011:QIW

[590] G. Bonfante, J.-Y. Marion, and J.-Y. Moyen. Quasi-interpretations a way to control resources. Theoretical Computer Science, 412(25):2776–2796, June 3, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBj

[591] Anonymous. Editorial Board. Theoretical Computer Science, 412(25):iii– viii, June 3, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2011:Pd

[592] G. Rozenberg. Preface. Theoretical Computer Science, 412(26):2797, June 10, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heiner:2011:FFR

[593] Monika Heiner and Adelinde Uhrmacher. Foundations of formal reconstruction of biochemical networks. Theoretical Computer Science, 412(26):2798– 2799, June 10, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Durzinsky:2011:AFN

[594] Markus Durzinsky, Annegret Wagler, and Robert Weismantel. An algorithmic framework for network reconstruction. Theoretical Computer Science, 412(26):2800–2815, June 10, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dimitrova:2011:PEB

[595] Elena Dimitrova, Luis David García-Puente, Franziska Hinkelmann, Abdul S. Jarrah, Reinhard Laubenbacher, Brandilyn Stigler, Michael Stillman, and Paola Vera-Licona. Parameter estimation for Boolean models of biological networks. Theoretical Computer Science, 412(26):2816–2826, June 10, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rizk:2011:CVT

[596] Aurélien Rizk, Grégory Batt, François Fages, and Sylvain Soliman. Continuous valuations of temporal logic specifications with applications to parameter optimization and robustness measures. Theoretical Computer Science, 412(26): 2827–2839, June 10, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Koh:2011:CBC

[597] Geoffrey Koh, David Hsu, and P. S. Thiagarajan. Component-based construction of bio-pathway models: The parameter estimation problem. Theoretical Computer Science, 412(26): 2840–2853, June 10, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mateescu:2011:CEC

[598] Radu Mateescu, Pedro T. Monteiro, Estelle Dumas, and Hidde de Jong. CTRL: Extension of CTL with regular expressions and fairness operators to verify genetic regulatory networks. Theoretical Computer Science, 412(26): 2854–2883, June 10, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schwarick:2011:IBM

[599] Martin Schwarick and Alexej Tovchigrechko. IDD-based model validation of biochemical networks. Theoretical Computer Science, 412(26):2884– 2908, June 10, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBk

[600] Anonymous. Editorial Board. Theoretical Computer Science, 412(26):iii– viii, June 10, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Carpi:2011:CW

[601] Arturo Carpi and Clelia De Felice. Combinatorics on words. Theoretical Computer Science, 412(27):2909– 2910, June 16, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Avgustinovich:2011:IPL

[602] S. V. Avgustinovich, A. Frid, T. Kamae, and P. Salimov. Infinite permutations of lowest maximal pattern complexity. Theoretical Computer Science, 412(27):2911–2921, June 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Masse:2011:EPC

[603] A. Blondin Massé, S. Brlek, A. Garon, and S. Labbé. Equations on palindromes and circular words. Theoretical Computer Science, 412(27):2922– 2930, June 16, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Crochemore:2011:RC

[604] Maxime Crochemore, Lucian Ilie, and Liviu Tinta. The “runs” conjecture. Theoretical Computer Science, 412(27): 2931–2941, June 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dominguez:2011:ECD

[605] Manuel Domínguez, Thomas Noll, and Marco Castrillón. An extension of Christoffel duality to a subset of Sturm numbers and their characteristic words. Theoretical Computer Science, 412(27): 2942–2954, June 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fiorenzi:2011:BGR

[606] Francesca Fiorenzi, Pascal Ochem, and Elise Vaslet. Bounds for the generalized repetition threshold. Theoretical Computer Science, 412(27):2955– 2963, June 16, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Grossi:2011:QTS

[607] Roberto Grossi. A quick tour on suffix arrays and compressed suffix arrays. Theoretical Computer Science, 412(27): 2964–2973, June 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jamet:2011:FPI

[608] D. Jamet, G. Paquin, G. Richomme, and L. Vuillon. On the fixed points of the iterated pseudopalindromic closure operator. Theoretical Computer Science, 412(27):2974–2987, June 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Merlini:2011:AAS

[609] D. Merlini and R. Sprugnoli. Algebraic aspects of some Riordan arrays related to binary words avoiding a pattern. Theoretical Computer Science, 412(27):2988–3001, June 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Provencal:2011:MNC

[610] Xavier Provençal. Minimal non-convex words. Theoretical Computer Science, 412(27):3002–3009, June 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rao:2011:LCD

[611] Michaël Rao. Last cases of Dejean’s conjecture. Theoretical Computer Science, 412(27):3010–3018, June 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Restivo:2011:BCW

[612] Antonio Restivo and Giovanna Rosone. Balancing and clustering of words in the Burrows–Wheeler transform. Theoretical Computer Science, 412(27): 3019–3032, June 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBl

[613] Anonymous. Editorial Board. Theoretical Computer Science, 412(27):iii– viii, June 16, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bethke:2011:PID

[614] Inge Bethke, Alban Ponse, and Piet H. Rodenburg. Preface: This issue is dedicated to Jan Bergstra on the occasion of his sixtieth birthday. Theoretical Computer Science, 412(28):3033– 3034, June 20, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aceto:2011:API

[615] Luca Aceto, Taolue Chen, Anna Ingolfsdottir, Bas Luttik, and Jaco van de Pol. On the axiomatizability of priority II. Theoretical Computer Science, 412(28):3035–3044, June 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Aceto:2011:SRF

[616] Luca Aceto, Matteo Cimini, Anna Ingolfsdottir, MohammadReza Mousavi, and Michel A. Reniers. SOS rule formats for zero and unit elements. Theoretical Computer Science, 412(28): 3045–3071, June 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Andres:2011:IHP

[617] Miguel E. Andrés, Catuscia Palamidessi, Peter van Rossum, and Ana Sokolova. Information hiding in probabilistic concurrent systems. Theoretical Computer Science, 412(28):3072– 3089, June 20, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baeten:2011:UMM

[618] Jos C. M. Baeten and Bas Luttik. Unguardedness mostly means many solutions. Theoretical Computer Science, 412(28):3090–3100, June 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bauer:2011:ITC

[619] Sebastian S. Bauer, Rolf Hennicker, and Martin Wirsing. Interface theories for concurrency and data. Theoretical Computer Science, 412(28):3101– 3121, June 20, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bergstra:2011:CFV

[620] Jan A. Bergstra and Jaco van de Pol. A calculus for four-valued sequential logic. Theoretical Computer Science, 412(28):3122–3128, June 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cranen:2011:LTC

[621] Sjoerd Cranen, Jan Friso Groote, and Michel Reniers. A linear translation from CTL* to the first-order modal μ-calculus. Theoretical Computer Science, 412(28):3129–3139, June 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanDalen:2011:BFP

[622] Dirk van Dalen. Brouwer’s ϵ-fixed point and Sperner’s lemma. Theoretical Computer Science, 412(28):3140– 3144, June 20, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Diaconescu:2011:ASS

[623] Razvan Diaconescu and Ionut Tutu. On the algebra of structured specifications. Theoretical Computer Science, 412(28): 3145–3174, June 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Endrullis:2011:ET

[624] Jörg Endrullis, Clemens Grabmayer, Jan Willem Klop, and Vincent van Oostrom. On equal μ-terms. Theoretical Computer Science, 412(28):3175– 3202, June 20, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Endrullis:2011:LPT

[625] Jörg Endrullis and Dimitri Hendriks. Lazy productivity via termination. Theoretical Computer Science, 412(28): 3203–3225, June 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Esparza:2011:DTA

[626] Javier Esparza, Stefan Kiefer, and Michael Luttenberger. Derivation tree analysis for accelerated fixed-point computation. Theoretical Computer Science, 412(28):3226–3241, June 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Geuvers:2011:CNT

[627] Herman Geuvers and Robbert Krebbers. The correctness of Newman’s typability algorithm and some of its extensions. Theoretical Computer Science, 412(28):3242–3261, June 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ghassemi:2011:VMA

[628] Fatemeh Ghassemi, Wan Fokkink, and Ali Movaghar. Verification of mobile ad hoc networks: an algebraic approach. Theoretical Computer Science, 412(28): 3262–3282, June 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

vanGlabbeek:2011:CCF

[629] R. J. van Glabbeek. On cool congruence formats for weak bisimulations. Theoretical Computer Science, 412(28):

3283–3302, June 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hofner:2011:FZG

[630] Peter Höfner and Bernhard Möller. Fixing Zeno gaps. Theoretical Computer Science, 412(28):3303–3322, June 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jacobs:2011:PDM

[631] Bart Jacobs. Probabilities, distribution monads, and convex categories. Theoretical Computer Science, 412(28): 3323–3336, June 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Luttgen:2011:SRL

[632] Gerald Lüttgen and Walter Vogler. Safe reasoning with Logic LTS. Theoretical Computer Science, 412(28): 3337–3357, June 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Larsen:2011:MWT

[633] Kim G. Larsen, Uli Fahrenberg, and Claus Thrane. Metrics for weighted transition systems: Axiomatization and complexity. Theoretical Computer Science, 412(28):3358–3369, June 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rodenburg:2011:FPA

[634] Piet H. Rodenburg, Jan Willem Klop, Karst Koymans, and Jos Vrancken. The free process algebra generated by δ and ϵ. Theoretical Computer Science, 412(28):3370–3377, June 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tucker:2011:COC

[635] J. V. Tucker and J. I. Zucker. Continuity of operators on continuous and discrete time streams. Theoretical Computer Science, 412(28):3378– 3403, June 20, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Visser:2011:APN

[636] Albert Visser. On the ambiguation of Polish notation. Theoretical Computer Science, 412(28):3404–3411, June 20, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBm

[637] Anonymous. Editorial Board. Theoretical Computer Science, 412(28):iii– viii, June 20, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kim:2011:RTC

[638] Dong Han Kim. Return time complexity of Sturmian sequences. Theoretical Computer Science, 412(29):3413– 3417, July 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dubois:2011:DFA

[639] Swan Dubois, Maria Potop-Butucaru, and Sébastien Tixeuil. Dynamic FTSS in asynchronous systems: The case of unison. Theoretical Computer Science, 412(29):3418–3439, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fan:2011:EFT

[640] Jianxi Fan, Xiaohua Jia, Baolei Cheng, and Jia Yu. An efficient fault-tolerant routing algorithm in bijective connection networks with restricted faulty edges. Theoretical Computer Science, 412(29):3440–3450, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cenk:2011:MPM

[641] Murat Cenk and Ferruh Özbudak. Multiplication of polynomials modulo x^{n}. Theoretical Computer Science, 412 (29):3451–3462, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See note [1073].

Blumensath:2011:RAI

[642] Achim Blumensath. Recognisability for algebras of infinite trees. Theoretical Computer Science, 412(29):3463– 3486, July 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Marx:2011:CCC

[643] Dániel Marx. Complexity of clique coloring and related problems. Theoretical Computer Science, 412(29):3487– 3500, July 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Xia:2011:SCT

[644] Ge Xia and Yong Zhang. On the small cycle transversal of planar graphs. Theoretical Computer Science, 412(29): 3501–3509, July 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Reidenbach:2011:RAE

[645] Daniel Reidenbach and Johannes C. Schneider. Restricted ambiguity of erasing morphisms. Theoretical Computer Science, 412(29):3510–3523, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sun:2011:ILB

[646] Xiaoming Sun. An improved lower bound on the sensitivity complexity of graph properties. Theoretical Computer Science, 412(29):3524–3529, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fomin:2011:EAM

[647] Fedor V. Fomin, Daniel Lokshtanov, and Saket Saurabh. An exact algorithm for minimum distortion embedding. Theoretical Computer Science, 412(29):3530–3536, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amir:2011:ASM

[648] Amihood Amir, Estrella Eisenberg, Orgad Keller, Avivit Levy, and Ely Porat. Approximate string matching with stuck address bits. Theoretical Computer Science, 412(29):3537–3544, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Korpelainen:2011:BPG

[649] Nicholas Korpelainen, Vadim V. Lozin, Dmitriy S. Malyshev, and Alexander Tiskin. Boundary properties of graphs for algorithmic graph problems. Theoretical Computer Science, 412(29): 3545–3554, July 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hoppen:2011:TPP

[650] Carlos Hoppen, Yoshiharu Kohayakawa, Carlos Gustavo Moreira, and Rudini Menezes Sampaio. Testing permutation properties through subpermutations. Theoretical Computer Science, 412(29):3555–3567, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brito:2011:ILB

[651] Carlos Fisch Brito and Shailesh Vaya. Improved lower bound for deterministic broadcasting in radio networks. Theoretical Computer Science, 412(29): 3568–3578, July 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hon:2011:COI

[652] Wing-Kai Hon, Tak-Wah Lam, Rahul Shah, Siu-Lung Tam, and Jeffrey Scott Vitter. Cache-oblivious index for approximate string matching. Theoretical Computer Science, 412(29):3579– 3588, July 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mitsche:2011:NHO

[653] Dieter Mitsche, Maria Saumell, and Rodrigo I. Silveira. On the number of higher order Delaunay triangulations. Theoretical Computer Science, 412(29): 3589–3597, July 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Turetsky:2011:CPD

[654] Daniel Turetsky. Connectedness properties of dimension level sets. Theoretical Computer Science, 412(29):3598– 3603, July 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fici:2011:SFC

[655] Gabriele Fici. Special factors and the combinatorics of suffix and factor automata. Theoretical Computer Science, 412(29):3604–3615, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goles:2011:CCN

[656] E. Goles, A. Moreira, and I. Rapaport. Communication complexity in numberconserving and monotone cellular automata. Theoretical Computer Science, 412(29):3616–3628, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kopecki:2011:IHC

[657] Steffen Kopecki. On iterated hairpin completion. Theoretical Computer Science, 412(29):3629–3638, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manjunath:2011:ADR

[658] Madhusudan Manjunath and Vikram Sharma. Applications of dimensionality reduction and exponential sums to graph automorphism. Theoretical Computer Science, 412(29):3639– 3649, July 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Angelelli:2011:CIS

[659] Enrico Angelelli and Carlo Filippi. On the complexity of interval scheduling with a resource constraint. Theoretical Computer Science, 412(29):3650– 3657, July 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hsieh:2011:WDE

[660] Hsien-Jone Hsieh and Dyi-Rong Duh. ω-wide diameters of enhanced pyramid networks. Theoretical Computer Science, 412(29):3658–3675, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bozapalidis:2011:ETT

[661] Symeon Bozapalidis, Zoltán Fülöp, and George Rahonis. Equational tree transformations. Theoretical Computer Science, 412(29):3676–3692, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Centeno:2011:ICG

[662] Carmen C. Centeno, Mitre C. Dourado, Lucia Draque Penso, Dieter Rautenbach, and Jayme L. Szwarcfiter. Irreversible conversion of graphs. Theoretical Computer Science, 412(29):3693– 3700, July 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Baudru:2011:CSA

[663] Nicolas Baudru. Compositional synthesis of asynchronous automata. Theoretical Computer Science, 412(29):3701– 3716, July 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Patt-Shamir:2011:VDU

[664] Boaz Patt-Shamir and Dror Rawitz. Video distribution under multiple constraints. Theoretical Computer Science, 412(29):3717–3730, July 1, 2011.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cao:2011:SOS

[665] Qian Cao, Zhaohui Liu, and T. C. E. Cheng. Semi-online scheduling with known partial information about job sizes on two identical machines. Theoretical Computer Science, 412(29): 3731–3737, July 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Antonakopoulos:2011:AAG

[666] Spyridon Antonakopoulos and Lisa Zhang. Approximation algorithms for grooming in optical network design. Theoretical Computer Science, 412(29): 3738–3751, July 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Blanchet-Sadri:2011:ALS

[667] F. Blanchet-Sadri, Ilkyoo Choi, and Robert Mercas. Avoiding large squares in partial words. Theoretical Computer Science, 412(29):3752–3758, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jiang:2011:IMS

[668] Minghui Jiang. Inapproximability of maximal strip recovery. Theoretical Computer Science, 412(29):3759– 3774, July 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mezzini:2011:FMT

[669] Mauro Mezzini. Fast minimal triangulation algorithm using minimum degree criterion. Theoretical Computer Science, 412(29):3775–3787, July 1, 2011.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lenz:2011:NPT

[670] Daniel Lenz, Zuzana Masáková, and Edita Pelantová. Note on powers in three interval exchange transformations. Theoretical Computer Science, 412(29):3788–3794, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Caron:2011:EAA

[671] Pascal Caron, Jean-Marc Champarnaud, and Ludovic Mignot. Erratum to “Acyclic automata and small expressions using multi-tilde-bar operators” [Theoret. Comput. Sci. 411 (38–39) (2010) 3423–3435]. Theoretical Computer Science, 412(29):3795–3796, July 1, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See [267].

Anonymous:2011:EBn

[672] Anonymous. Editorial Board. Theoretical Computer Science, 412(29):iii– viii, July 1, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2011:Pe

[673] G. Rozenberg. Preface. Theoretical Computer Science, 412(30):3797, July 8, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kari:2011:P

[674] Jarkko Kari. Preface. Theoretical Computer Science, 412(30):3798–3799, July 8, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Delacourt:2011:DDA

[675] M. Delacourt, V. Poupet, M. Sablik, and G. Theyssier. Directional dynamics along arbitrary curves in cellular automata. Theoretical Computer Science, 412(30):3800–3821, July 8, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Capobianco:2011:GBW

[676] Silvio Capobianco. Generalized Besicovitch and Weyl spaces: Topology, patterns, and sliding block codes. Theoretical Computer Science, 412(30): 3822–3837, July 8, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pivato:2011:PEV

[677] Marcus Pivato. Positive expansiveness versus network dimension in symbolic dynamical systems. Theoretical Computer Science, 412(30):3838–3855, July 8, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Morita:2011:SRT

[678] Kenichi Morita. Simulating reversible Turing machines and cyclic tag systems by one-dimensional reversible cellular automata. Theoretical Computer Science, 412(30):3856–3865, July 8, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Delorme:2011:BAT

[679] M. Delorme, J. Mazoyer, N. Ollinger, and G. Theyssier. Bulking I: an abstract theory of bulking. Theoretical Computer Science, 412(30):3866– 3880, July 8, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Delorme:2011:BIC

[680] M. Delorme, J. Mazoyer, N. Ollinger, and G. Theyssier. Bulking II: Classifications of cellular automata. Theoretical Computer Science, 412(30):3881– 3905, July 8, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goles:2011:TCC

[681] Eric Goles, Pierre Guillon, and Ivan Rapaport. Traced communication complexity of cellular automata. Theoretical Computer Science, 412(30):3906– 3916, July 8, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kutrib:2011:CAL

[682] Martin Kutrib and Andreas Malcher. Cellular automata with limited intercell bandwidth. Theoretical Computer Science, 412(30):3917–3931, July 8, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Barrett:2011:MAS

[683] Chris Barrett, Harry B. Hunt III, Madhav V. Marathe, S. S. Ravi, Daniel J. Rosenkrantz, and Richard E. Stearns. Modeling and analyzing social network dynamics using stochastic discrete graphical dynamical systems. Theoretical Computer Science, 412(30): 3932–3946, July 8, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rouquier:2011:SMG

[684] Jean-Baptiste Rouquier, Damien Regnault, and Éric Thierry. Stochastic minority on graphs. Theoretical Computer Science, 412(30):3947–3963, July 8, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Mejia:2011:CSC

[685] Carolina Mejia and J. Andres Montoya. On the complexity of sandpile critical avalanches. Theoretical Computer Science, 412(30):3964–3974, July 8, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBo

[686] Anonymous. Editorial Board. Theoretical Computer Science, 412(30):iii– viii, July 8, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Krynicki:2011:TIS

[687] Michal Krynicki, Jerzy Tomasik, and Konrad Zdanowski. Theories of initial segments of standard models of arithmetics and their complete extensions. Theoretical Computer Science, 412(31): 3975–3991, July 15, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Horcik:2011:DPC

[688] Rostislav Horcík and Kazushige Terui. Disjunction property and complexity of substructural logics. Theoretical Computer Science, 412(31):3992– 4006, July 15, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Basin:2011:DTL

[689] David Basin, Carlos Caleiro, Jaime Ramos, and Luca Viganò. Distributed temporal logic for the analysis of security protocol models. Theoretical Computer Science, 412(31):4007– 4043, July 15, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Biskup:2011:SCM

[690] Joachim Biskup and Lena Wiese. A sound and complete model-generation procedure for consistent and confidentialitypreserving databases. Theoretical Computer Science, 412(31):4044–4072, July 15, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBp

[691] Anonymous. Editorial Board. Theoretical Computer Science, 412(31):iii– viii, July 15, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Du:2011:Pb

[692] Ding-Zhu Du, Yingfei Dong, and Zhao Zhang. Preface. Theoretical Computer Science, 412(32):4073, July 22, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2011:AAV

[693] Minming Li. Approximation algorithms for variable voltage processors: Min energy, max throughput and online heuristics. Theoretical Computer Science, 412(32):4074–4080, July 22, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bazgan:2011:CAC

[694] Cristina Bazgan, Basile Couëtoux, and Zsolt Tuza. Complexity and approximation of the Constrained Forest problem. Theoretical Computer Science, 412(32):4081–4091, July 22, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Franke:2011:CLM

[695] Robert Franke, Ignaz Rutter, and Dorothea Wagner. Computing large matchings in planar graphs with fixed minimum degree. Theoretical Computer Science, 412(32):4092–4099, July 22, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gu:2011:CFA

[696] Qian-Ping Gu and Hisao Tamaki. Constant-factor approximations of branch-decomposition and largest grid minor of planar graphs in O(n^{1+ϵ}) time. Theoretical Computer Science, 412(32):4100–4109, July 22, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2011:LST

[697] Li Chen and Bin Fu. Linear and sublinear time algorithms for the basis of abelian groups. Theoretical Computer Science, 412(32):4110–4122, July 22, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kelner:2011:ERC

[698] Jonathan Kelner and Petar Maymounkov. Electric routing and concurrent flow cutting. Theoretical Computer Science, 412(32):4123–4135, July 22, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Eidenbenz:2011:GPT

[699] Raphael Eidenbenz and Roger Wattenhofer. Good programming in transactional memory: Game theory meets multicore architecture. Theoretical Computer Science, 412(32):4136–4150, July 22, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Schneider:2011:BCM

[700] Johannes Schneider and Roger Wattenhofer. Bounds on contention management algorithms. Theoretical Computer Science, 412(32):4151–4160, July 22, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Biedl:2011:RPS

[701] Therese Biedl, Stephane Durocher, and Jack Snoeyink. Reconstructing polygons from scanner data. Theoretical Computer Science, 412(32):4161– 4172, July 22, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Knauer:2011:DHD

[702] Christian Knauer, Maarten Löffler, Marc Scherfenberg, and Thomas Wolle. The directed Hausdorff distance between imprecise point sets. Theoretical Computer Science, 412(32):4173– 4186, July 22, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Li:2011:PSA

[703] Shuai Cheng Li and Yen Kaow Ng. On protein structure alignment under distance constraint. Theoretical Computer Science, 412(32):4187–4199, July 22, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Arroyuelo:2011:UMC

[704] Diego Arroyuelo, Francisco Claude, Reza Dorrigiv, Stephane Durocher, Meng He, Alejandro López-Ortiz, J. Ian Munro, Patrick K. Nicholson, Alejandro Salinger, and Matthew Skala. Untangled monotonic chains and adaptive range search. Theoretical Computer Science, 412(32):4200– 4211, July 22, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Narang:2011:PDD

[705] Ankur Narang and Rudrapatna K. Shyamasundar. Performance driven distributed scheduling of parallel hybrid computations. Theoretical Computer Science, 412(32):4212–4225, July 22, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBq

[706] Anonymous. Editorial Board. Theoretical Computer Science, 412(32):iii– viii, July 22, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Datta:2011:SSS

[707] Ajoy K. Datta, Franck Petit, and Rachid Guerraoui. Stabilization, Safety, and Security of Distributed Systems (SSS 2009). Theoretical Computer Science, 412(33):4227, July 29, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abujarad:2011:ACB

[708] F. Abujarad and S. S. Kulkarni. Automated constraint-based addition of nonmasking and stabilizing faulttolerance. Theoretical Computer Science, 412(33):4228–4246, July 29, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beauquier:2011:SST

[709] Joffroy Beauquier, Janna Burman, and Shay Kutten. A self-stabilizing transformer for population protocols with covering. Theoretical Computer Science, 412(33):4247–4259, July 29, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Biely:2011:CWA

[710] Martin Biely and Martin Hutle. Consensus when all processes may be Byzantine for some time. Theoretical Computer Science, 412(33):4260– 4272, July 29, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bonnet:2011:RWF

[711] François Bonnet and Michel Raynal. On the road to the weakest failure detector for k-set agreement in messagepassing systems. Theoretical Computer Science, 412(33):4273–4284, July 29, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cournier:2011:HIS

[712] Alain Cournier, Swan Dubois, and Vincent Villain. How to improve snapstabilizing point-to-point communication space complexity? Theoretical Computer Science, 412(33):4285– 4296, July 29, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dasgupta:2011:FCW

[713] Anurag Dasgupta, Sukumar Ghosh, and Xin Xiao. Fault containment in weakly stabilizing systems. Theoretical Computer Science, 412(33):4297– 4311, July 29, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

GhoshDastidar:2011:SCT

[714] K. GhoshDastidar and T. Herman. Separation of circulating tokens. Theoretical Computer Science, 412(33): 4312–4324, July 29, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gouda:2011:NES

[715] M. G. Gouda and H. B. Acharya. Nash equilibria in stabilizing systems. Theoretical Computer Science, 412(33): 4325–4335, July 29, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kakugawa:2011:ONS

[716] Hirotsugu Kakugawa, Yukiko Yamauchi, Sayaka Kamei, and Toshimitsu Masuzawa. Observations on nonsilent self-stabilizing algorithms in sensor networks with probabilistically intermittent link failures. Theoretical Computer Science, 412(33):4336– 4349, July 29, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Souissi:2011:OBF

[717] Samia Souissi, Taisuke Izumi, and Koichi Wada. Oracle-based flocking of mobile robots in crash-recovery model. Theoretical Computer Science, 412(33): 4350–4360, July 29, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Turau:2011:FCS

[718] Volker Turau and Bernd Hauck. A fault-containing self-stabilizing (3 − 2/(Δ+1))-approximation algorithm for vertex cover in anonymous networks. Theoretical Computer Science, 412(33): 4361–4371, July 29, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBr

[719] Anonymous. Editorial Board. Theoretical Computer Science, 412(33):iii– viii, July 29, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Caillaud:2011:CMC

[720] Benoît Caillaud, Benoît Delahaye, Kim G. Larsen, Axel Legay, Mikkel L. Pedersen, and Andrzej Wasowski. Constraint Markov chains. Theoretical Computer Science, 412(34):4373– 4404, August 5, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gnaedig:2011:PWP

[721] Isabelle Gnaedig and Hélène Kirchner. Proving weak properties of rewriting. Theoretical Computer Science, 412(34): 4405–4438, August 5, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rosa-Velardo:2011:DCP

[722] Fernando Rosa-Velardo and David de Frutos-Escrig. Decidability and complexity of Petri nets with unordered data. Theoretical Computer Science, 412(34):4439–4451, August 5, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Stell:2011:RBA

[723] John Stell and Michael Worboys. Relations between adjacency trees. Theoretical Computer Science, 412(34):4452– 4468, August 5, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBs

[724] Anonymous. Editorial Board. Theoretical Computer Science, 412(34):iii– viii, August 5, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Salvy:2011:PFF

[725] Bruno Salvy, Bob Sedgewick, Michèle Soria, Wojciech Szpankowski, and Brigitte Vallée. Philippe Flajolet, the father of analytic combinatorics. Theoretical Computer Science, 412(35): 4469–4470, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Delaet:2011:DSP

[726] Sylvie Delaët, Partha Sarathi Mandal, Mariusz A. Rokicki, and Sébastien Tixeuil. Deterministic secure positioning in wireless sensor networks. Theoretical Computer Science, 412(35): 4471–4481, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kortsarz:2011:ASN

[727] Guy Kortsarz and Zeev Nutov. Approximating some network design problems with node costs. Theoretical Computer Science, 412(35):4482– 4492, August 12, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cheze:2011:SMC

[728] Guillaume Chèze, André Galligo, Bernard Mourrain, and Jean-Claude Yakoubsohn. A subdivision method for computing nearest gcd with certification. Theoretical Computer Science, 412(35):4493–4503, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wu:2011:AEG

[729] Ro-Yu Wu, Jou-Ming Chang, Kung-Jui Pai, and Yue-Li Wang. Amortized efficiency of generating planar paths in convex position. Theoretical Computer Science, 412(35):4504– 4512, August 12, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shih:2011:OOD

[730] Yuan-Kang Shih and Shin-Shin Kao. One-to-one disjoint path covers on kary n-cubes. Theoretical Computer Science, 412(35):4513–4530, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2011:CDR

[731] Lily Chen, Xueliang Li, and Yongtang Shi. The complexity of determining the rainbow vertex-connection of a graph. Theoretical Computer Science, 412(35):4531–4535, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sevastyanov:2011:TCA

[732] Sergey V. Sevastyanov, Bertrand M. T. Lin, and Hsiao-Lan Huang. Tight complexity analysis of the relocation problem with arbitrary release dates. Theoretical Computer Science, 412(35): 4536–4544, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shoaran:2011:ESS

[733] M. Shoaran and A. Thomo. Evolving schemas for streaming XML. Theoretical Computer Science, 412(35):4545– 4557, August 12, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chiang:2011:DRG

[734] Sheng-Hao Chiang, I-Chen Wu, and Ping-Hung Lin. Drawn k-in-a-row games. Theoretical Computer Science, 412(35):4558–4569, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bodlaender:2011:KBD

[735] Hans L. Bodlaender, Stéphan Thomassé, and Anders Yeo. Kernel bounds for disjoint cycles and disjoint paths. Theoretical Computer Science, 412(35):4570– 4578, August 12, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gebauer:2011:FEH

[736] Heidi Gebauer. Finding and enumerating Hamilton cycles in 4-regular graphs. Theoretical Computer Science, 412(35):4579–4591, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hashemi:2011:AIC

[737] Amir Hashemi and Benyamin M.- Alizadeh. Applying IsRewritten criterion on Buchberger algorithm. Theoretical Computer Science, 412(35):4592– 4603, August 12, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Meduna:2011:WTR

[738] Alexander Meduna and Petr Zemek. Workspace theorems for regularcontrolled grammars. Theoretical Computer Science, 412(35):4604–4612, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Makino:2011:EAB

[739] Kazuhisa Makino, Suguru Tamaki, and Masaki Yamamoto. An exact algorithm for the Boolean connectivity problem for k-CNF. Theoretical Computer Science, 412(35):4613–4618, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Montanaro:2011:UEQ

[740] Ashley Montanaro, Harumichi Nishimura, and Rudy Raymond. Unbounded-error quantum query complexity. Theoretical Computer Science, 412(35):4619– 4628, August 12, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kiran:2011:HSS

[741] S. Kanthi Kiran and N. S. Narayanaswamy. Hardness of subgraph and supergraph problems in c-tournaments. Theoretical Computer Science, 412(35):4629– 4635, August 12, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kim:2011:DPC

[742] Sook-Yeon Kim, Jae-Ha Lee, and Jung-Heum Park. Disjoint path covers in recursive circulants G(2^{m}, 4) with faulty elements. Theoretical Computer Science, 412(35):4636–4649, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Miyata:2011:LBF

[743] Akinobu Miyata, Jun Tarui, and Etsuji Tomita. Learning Boolean functions in AC^{0} on attribute and classification noise — estimating an upper bound on attribute and classification noise. Theoretical Computer Science, 412(35):4650–4660, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Altshuler:2011:SEG

[744] Yaniv Altshuler and Alfred M. Bruckstein. Static and expanding grid coverage with ant robots: Complexity results. Theoretical Computer Science, 412(35):4661–4674, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cohen:2011:TPS

[745] Nathann Cohen, David Coudert, Dorian Mazauric, Napoleão Nepomuceno, and Nicolas Nisse. Tradeoffs in process strategy games with application in the WDM reconfiguration problem. Theoretical Computer Science, 412(35): 4675–4687, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kreutzer:2011:DDM

[746] Stephan Kreutzer and Sebastian Ordyniak. Digraph decompositions and monotonicity in digraph searching. Theoretical Computer Science, 412(35): 4688–4703, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kanj:2011:LAE

[747] Iyad A. Kanj, Andreas Wiese, and Fenghui Zhang. Local algorithms for edge colorings in UDGs. Theoretical Computer Science, 412(35):4704– 4714, August 12, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Perrucci:2011:LSS

[748] Daniel Perrucci. Linear solving for sign determination. Theoretical Computer Science, 412(35):4715–4720, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bogdanowicz:2011:CMC

[749] Damian Bogdanowicz, Krzysztof Giaro, and Robert Janczewski. Consensus models: Computational complexity aspects in modern approaches to the list coloring problem. Theoretical Computer Science, 412(35):4721– 4728, August 12, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lim:2011:NAN

[750] Chae Hoon Lim. A note on the average number of RSA fixed points. Theoretical Computer Science, 412(35):4729– 4737, August 12, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cusick:2011:SDS

[751] Thomas W. Cusick and Lavinia Corina Ciungu. Sum of digits sequences modulo m. Theoretical Computer Science, 412(35):4738–4741, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Currie:2011:LLW

[752] James Currie. Lexicographically least words in the orbit closure of the Rudin–Shapiro word. Theoretical Computer Science, 412(35):4742–4746, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hung:2011:ETE

[753] Ruo-Wei Hung. Embedding two edgedisjoint Hamiltonian cycles into locally twisted cubes. Theoretical Computer Science, 412(35):4747–4753, August 12, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBt

[754] Anonymous. Editorial Board. Theoretical Computer Science, 412(35):iii– viii, August 12, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brlek:2011:ADG

[755] Srečko Brlek and Xavier Provençal. Advances in discrete geometry. Theoretical Computer Science, 412(36): 4755–4756, August 19, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Berthe:2011:ATA

[756] Valérie Berthé. About thin arithmetic discrete planes. Theoretical Computer Science, 412(36):4757–4769, August 19, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kock:2011:ACM

[757] Anders Kock. Affine connections, midpoint formation, and point reflection. Theoretical Computer Science, 412(36): 4770–4777, August 19, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Masse:2011:TIF

[758] A. Blondin Massé, S. Brlek, A. Garon, and S. Labbé. Two infinite families of polyominoes that tile the plane by translation in two distinct ways. Theoretical Computer Science, 412(36): 4778–4786, August 19, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bodini:2011:DRT

[759] Olivier Bodini, Thomas Fernique, Michael Rao, and Éric Rémila. Distances on rhombus tilings. Theoretical Computer Science, 412(36):4787– 4794, August 19, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brocchi:2011:RAS

[760] S. Brocchi, A. Frosini, and S. Rinaldi. A reconstruction algorithm for a subclass of instances of the 2-color problem. Theoretical Computer Science, 412(36):4795–4804, August 19, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Esbelin:2011:CBB

[761] Henri-Alex Esbelin, Rémy Malgouyres, and Colin Cartade. Convergence of binomial-based derivative estimation for C^{2} noisy discretized curves. Theoretical Computer Science, 412(36): 4805–4813, August 19, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Charrier:2011:CEL

[762] Émilie Charrier, Fabien Feschet, and Lilian Buzer. Computing efficiently the lattice width in any dimension. Theoretical Computer Science, 412(36): 4814–4823, August 19, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Tajine:2011:PME

[763] Mohamed Tajine and Alain Daurat. Patterns for multigrid equidistributed functions: Application to general parabolas and length estimation. Theoretical Computer Science, 412(36): 4824–4840, August 19, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Brlek:2011:LTS

[764] Srečko Brlek, Michel Koskas, and Xavier Provençal. A linear time and space algorithm for detecting path intersection in Z^{d}. Theoretical Computer Science, 412(36):4841–4850, August 19, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBu

[765] Anonymous. Editorial Board. Theoretical Computer Science, 412(36):iii– viii, August 19, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ayala-Rincon:2011:P

[766] Mauricio Ayala-Rincón, Elaine Pimentel, and Fairouz Kamareddine. Preface. Theoretical Computer Science, 412(37):4851–4852, August 26, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rocha:2011:FLS

[767] Camilo Rocha, César Muñoz, and Gilles Dowek. A formal library of set relations and its application to synchronous languages. Theoretical Computer Science, 412(37):4853–4866, August 26, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kesner:2011:PFL

[768] Delia Kesner and Fabien Renaud. A prismoid framework for languages with resources. Theoretical Computer Science, 412(37):4867–4892, August 26, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ferreira:2011:RDF

[769] Francicleber Martins Ferreira and Ana Teresa Martins. Recursive definitions and fixed-points on well-founded structures. Theoretical Computer Science, 412(37):4893–4904, August 26, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seldin:2011:SRC

[770] Jonathan P. Seldin. The search for a reduction in combinatory logic equivalent to λ-reduction. Theoretical Computer Science, 412(37):4905– 4918, August 26, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Horozal:2011:RMT

[771] Fulya Horozal and Florian Rabe. Representing model theory in a typetheoretical logical framework. Theoretical Computer Science, 412(37):4919– 4945, August 26, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Benevides:2011:SMD

[772] Mario R. F. Benevides and L. Menasché Schechter. A study on multidimensional products of graphs and hybrid logics. Theoretical Computer Science, 412(37):4946–4966, August 26, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBv

[773] Anonymous. Editorial Board. Theoretical Computer Science, 412(37):iii– viii, August 26, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jacobs:2011:P

[774] Bart Jacobs, Milad Niqui, Jan Rutten, and Alexandra Silva. Preface. Theoretical Computer Science, 412(38):4967– 4968, September 2, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Adamek:2011:SOI

[775] Jirí Adámek, Stefan Milius, and Jirí Velebil. On second-order iterative monads. Theoretical Computer Science, 412(38):4969–4988, September 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Balan:2011:CA

[776] A. Balan and A. Kurz. On coalgebras over algebras. Theoretical Computer Science, 412(38):4989–5005, September 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Capretta:2011:CFP

[777] Venanzio Capretta. Coalgebras in functional programming and type theory. Theoretical Computer Science, 412(38):5006–5024, September 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cirstea:2011:MTP

[778] Corina Cîrstea. Maximal traces and path-based coalgebraic temporal logics. Theoretical Computer Science, 412(38):5025–5042, September 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Klin:2011:BSO

[779] Bartek Klin. Bialgebras for structural operational semantics: an introduction. Theoretical Computer Science, 412(38):5043–5069, September 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kupke:2011:CSM

[780] Clemens Kupke and Dirk Pattinson. Coalgebraic semantics of modal logics: an overview. Theoretical Computer Science, 412(38):5070–5094, September 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Sokolova:2011:PSC

[781] Ana Sokolova. Probabilistic systems coalgebraically: a survey. Theoretical Computer Science, 412(38):5095– 5110, September 2, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBw

[782] Anonymous. Editorial Board. Theoretical Computer Science, 412(38):ii– vii, September 2, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Pralat:2011:EDM

[783] Pawel Pralat and Changping Wang. An edge deletion model for complex networks. Theoretical Computer Science, 412(39):5111–5120, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bienkowski:2011:RCA

[784] Marcin Bienkowski, Marek Chrobak, and Lukasz Jez. Randomized competitive algorithms for online buffer management in the adaptive adversary model. Theoretical Computer Science, 412(39):5121–5131, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Okhotin:2011:EPL

[785] Alexander Okhotin. Expressive power of LL(k) Boolean grammars. Theoretical Computer Science, 412(39):5132– 5155, September 9, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Didimo:2011:DGR

[786] Walter Didimo, Peter Eades, and Giuseppe Liotta. Drawing graphs with right angle crossings. Theoretical Computer Science, 412(39):5156– 5166, September 9, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Murata:2011:MEB

[787] Atsushi Murata and Akira Matsubayashi. Minimum energy broadcast on rectangular grid wireless networks. Theoretical Computer Science, 412(39):5167–5175, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hon:2011:SDS

[788] Wing-Kai Hon, Kunihiko Sadakane, and Wing-Kin Sung. Succinct data structures for Searchable Partial Sums with optimal worst-case performance. Theoretical Computer Science, 412(39): 5176–5186, September 9, 2011. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bui-Xuan:2011:BWG

[789] Binh-Minh Bui-Xuan, Jan Arne Telle, and Martin Vatshelle. Boolean-width of graphs. Theoretical Computer Science, 412(39):5187–5204, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kaminski:2011:SPB

[790] Marcin Kamiński, Paul Medvedev, and Martin Milanic. Shortest paths between shortest paths. Theoretical Computer Science, 412(39):5205–5210, September 9, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lim:2011:IBO

[791] Kyungkuk Lim, Kangbok Lee, and Soo Y. Chang. Improved bounds for online scheduling with eligibility constraints. Theoretical Computer Science, 412(39):5211–5224, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liu:2011:OAO

[792] Ming Liu, Feifeng Zheng, Chengbin Chu, and Yinfeng Xu. Optimal algorithms for online scheduling on parallel machines to minimize the makespan with a periodic availability constraint. Theoretical Computer Science, 412(39):5225–5231, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ghareghani:2011:FFS

[793] N. Ghareghani, M. Mohammad-Noori, and P. Sharifani. On z-factorization and c-factorization of standard episturmian words. Theoretical Computer Science, 412(39):5232–5238, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Amir:2011:EAC

[794] Amihood Amir, Gad M. Landau, Joong Chae Na, Heejin Park, Kunsoo Park, and Jeong Seop Sim. Efficient algorithms for consensus string problems minimizing both distance sum and radius. Theoretical Computer Science, 412(39):5239–5246, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Wang:2011:ORF

[795] Haitao Wang, Amitabh Chaudhary, and Danny Z. Chen. Online rectangle filling. Theoretical Computer Science, 412(39):5247–5275, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Harju:2011:NFB

[796] Tero Harju and Tomi Kärki. On the number of frames in binary words. Theoretical Computer Science, 412(39): 5276–5284, September 9, 2011. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Thomo:2011:RVP

[797] A. Thomo and S. Venkatesh. Rewriting of visibly pushdown languages for XML data integration. Theoretical Computer Science, 412(39):5285–5297, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anshelevich:2011:EAE

[798] Elliot Anshelevich and Bugra Caskurlu. Exact and approximate equilibria for optimal group network formation. Theoretical Computer Science, 412(39): 5298–5314, September 9, 2011. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Monti:2011:RGS

[799] A. Monti and B. Sinaimeri. Rainbow graph splitting. Theoretical Computer Science, 412(39):5315–5324, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bentz:2011:HFN

[800] Cédric Bentz. On the hardness of finding near-optimal multicuts in directed acyclic graphs. Theoretical Computer Science, 412(39):5325–5332, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fang:2011:OBS

[801] Yang Fang, Xiwen Lu, and Peihai Liu. Online batch scheduling on parallel machines with delivery times. Theoretical Computer Science, 412(39):5333– 5339, September 9, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Zhang:2011:ALP

[802] Wen-Qi Zhang and Yong-Jin Liu. Approximating the longest paths in grid graphs. Theoretical Computer Science, 412(39):5340–5350, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hung:2011:ECA

[803] Ruo-Wei Hung and Maw-Shang Chang. An efficient certifying algorithm for the Hamiltonian cycle problem on circulararc graphs. Theoretical Computer Science, 412(39):5351–5373, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Su:2011:AGR

[804] Shenghui Su, Shuwang Lü, and Xiubin Fan. Asymptotic granularity reduction and its application. Theoretical Computer Science, 412(39):5374– 5386, September 9, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shin:2011:PCR

[805] Sujin Shin. Pseudo-cyclic renewal systems. Theoretical Computer Science, 412(39):5387–5399, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Charlier:2011:GFR

[806] Émilie Charlier and Narad Rampersad. The growth function of S-recognizable sets. Theoretical Computer Science, 412(39):5400–5408, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Colussi:2011:CCC

[807] Livio Colussi. The convergence classes of Collatz function. Theoretical Computer Science, 412(39):5409–5419, September 9, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See comment [2458].

Hoefer:2011:CRT

[808] Martin Hoefer, Vahab S. Mirrokni, Heiko Röglin, and Shang-Hua Teng. Competitive routing over time. Theoretical Computer Science, 412(39): 5420–5432, September 9, 2011. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Beckmann:2011:CUB

[809] Arnold Beckmann and Samuel R. Buss. Corrected upper bounds for free-cut elimination. Theoretical Computer Science, 412(39):5433–5445, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Yang:2011:CCK

[810] Guomin Yang and Chik How Tan. Certificateless cryptography with KGC trust level 3. Theoretical Computer Science, 412(39):5446–5457, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cheng:2011:MGB

[811] Qi Cheng and Yu-Hsin Li. On the minimum gap between sums of square roots of small integers. Theoretical Computer Science, 412(39):5458–5465, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heggernes:2011:GLC

[812] Pinar Heggernes, Daniel Meister, and Charis Papadopoulos. Graphs of linear clique-width at most 3. Theoretical Computer Science, 412(39):5466– 5486, September 9, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Steinberg:2011:CCO

[813] Benjamin Steinberg. The Cerný conjecture for one-cluster automata with prime length cycle. Theoretical Computer Science, 412(39):5487– 5491, September 9, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Richomme:2011:FBS

[814] Gwénaël Richomme and Patrice Séébold. On factorially balanced sets of words. Theoretical Computer Science, 412(39):5492–5497, September 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chen:2011:IBT

[815] Xiaofeng Chen, Willy Susilo, Fangguo Zhang, Haibo Tian, and Jin Li. Identity-based trapdoor mercurial commitments and applications. Theoretical Computer Science, 412(39): 5498–5512, September 9, 2011. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBx

[816] Anonymous. Editorial Board. Theoretical Computer Science, 412(39):iii– viii, September 9, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dolev:2011:P

[817] Shlomi Dolev, Sandeep Kulkarni, and André Schiper. Preface. Theoretical Computer Science, 412(40):5513–5514, September 16, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Manne:2011:SSA

[818] Fredrik Manne, Morten Mjelde, Laurence Pilard, and Sébastien Tixeuil. A self-stabilizing 2/3-approximation algorithm for the maximum matching problem. Theoretical Computer Science, 412(40):5515–5526, September 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Turau:2011:NAS

[819] Volker Turau and Bernd Hauck. A new analysis of a self-stabilizing maximum weight matching algorithm with approximation ratio 2. Theoretical Computer Science, 412(40):5527–5540, September 16, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Datta:2011:SSL

[820] Ajoy K. Datta, Lawrence L. Larmore, and Priyanka Vemula. Self-stabilizing leader election in optimal space under an arbitrary scheduler. Theoretical Computer Science, 412(40):5541–5561, September 16, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cobb:2011:SMM

[821] Jorge A. Cobb and Mohamed G. Gouda. Stabilization of max-min fair networks without per-flow state. Theoretical Computer Science, 412(40): 5562–5579, September 16, 2011. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Robinson:2011:ABC

[822] Peter Robinson and Ulrich Schmid. The Asynchronous Bounded-Cycle model. Theoretical Computer Science, 412(40):5580–5601, September 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Biely:2011:SCU

[823] Martin Biely, Ulrich Schmid, and Bettina Weiss. Synchronous consensus under hybrid process and link failures. Theoretical Computer Science, 412 (40):5602–5630, September 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hoepman:2011:SSS

[824] Jaap-Henk Hoepman, Andreas Larsson, Elad M. Schiller, and Philippas Tsigas. Secure and self-stabilizing clock synchronization in sensor networks. Theoretical Computer Science, 412(40):5631–5647, September 16, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBy

[825] Anonymous. Editorial Board. Theoretical Computer Science, 412(40):iii– viii, September 16, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Balkova:2011:BRC

[826] L’. Balková, E. Pelantová, and S. Starosta. On Brlek–Reutenauer conjecture. Theoretical Computer Science, 412(41):5649–5655, September 23, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See note [1270].

Barmpalias:2011:KCI

[827] George Barmpalias and C. S. Vlek. Kolmogorov complexity of initial segments of sequences and arithmetical definability. Theoretical Computer Science, 412(41):5656–5667, September 23, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Calude:2011:FSC

[828] Cristian S. Calude, Kai Salomaa, and Tania K. Roblot. Finite state complexity. Theoretical Computer Science, 412 (41):5668–5677, September 23, 2011.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Cincotti:2011:GPC

[829] Alessandro Cincotti. The game of nplayer Cutcake. Theoretical Computer Science, 412(41):5678–5683, September 23, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

DeSantis:2011:EPS

[830] Alfredo De Santis, Anna Lisa Ferrara, and Barbara Masucci. Efficient provably-secure hierarchical key assignment schemes. Theoretical Computer Science, 412(41):5684–5699, September 23, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Dereniowski:2011:CSW

[831] Dariusz Dereniowski. Connected searching of weighted trees. Theoretical Computer Science, 412(41):5700–5713, September 23, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frougny:2011:PAN

[832] Christiane Frougny, Edita Pelantová, and Milena Svobodová. Parallel addition in non-standard numeration systems. Theoretical Computer Science, 412(41):5714–5727, September 23, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Goldstein:2011:SCU

[833] Ilya Goldstein. Subword complexity of uniform D0L words over finite groups. Theoretical Computer Science, 412(41):5728–5743, September 23, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Gutin:2011:KBU

[834] G. Gutin, M. Jones, and A. Yeo. Kernels for below-upper-bound parameterizations of the hitting set and directed dominating set problems. Theoretical Computer Science, 412(41):5744–5751, September 23, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kosowski:2011:SBH

[835] Adrian Kosowski, Alfredo Navarra, and Cristina M. Pinotti. Synchronous black hole search in directed graphs. Theoretical Computer Science, 412(41):5752– 5759, September 23, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rajsbaum:2011:SPD

[836] Sergio Rajsbaum and Jorge Urrutia. Some problems in distributed computational geometry. Theoretical Computer Science, 412(41):5760–5770, September 23, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shiloni:2011:RAE

[837] Asaf Shiloni, Noa Agmon, and Gal A. Kaminka. Of robot ants and elephants: a computational comparison. Theoretical Computer Science, 412(41):5771– 5788, September 23, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Turaev:2011:NCT

[838] S. Turaev, J. Dassow, and M. Selamat. Nonterminal complexity of tree controlled grammars. Theoretical Computer Science, 412(41):5789–5795, September 23, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Drucker:2011:BSM

[839] Andrew Drucker. Block sensitivity of minterm-transitive functions. Theoretical Computer Science, 412(41):5796– 5801, September 23, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Garcia:2011:RES

[840] Pedro García, Damián López, José Ruiz, and Gloria I. Álvarez. From regular expressions to smaller NFAs. Theoretical Computer Science, 412 (41):5802–5807, September 23, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Heam:2011:CCP

[841] P.-C. Héam. On the complexity of computing the profinite closure of a rational language. Theoretical Computer Science, 412(41):5808–5813, September 23, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2011:EBz

[842] Anonymous. Editorial Board. Theoretical Computer Science, 412(41):iii– viii, September 23, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rozenberg:2011:Pf

[843] G. Rozenberg.  Preface.  Theo-

retical Computer Science, 412(42): 5815, September 30, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751100702X.

Nguyen:2011:RSF

[844] Hung Son Nguyen, Sankar K. Pal, and Andrzej Skowron. Rough sets and fuzzy sets in natural computing. Theoretical Computer Science, 412(42):5816– 5819, September 30, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511004543.

Csajbok:2011:ASB

[845] Zoltán Csajbók. Approximation of sets based on partial covering. Theoretical Computer Science, 412(42):5820– 5833, September 30, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511004555.

Ganivada:2011:FRG

[846] Avatharam Ganivada, Soumitra Dutta, and Sankar K. Pal. Fuzzy rough granular neural networks, fuzzy granules, and classification. Theoretical Computer Science, 412(42):5834– 5853, September 30, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511004567.

Horta:2011:EFC

[847] Danilo Horta, Ivan C. de Andrade, and Ricardo J. G. B. Campello. Evolutionary fuzzy clustering of relational data. Theoretical Computer Science, 412(42):5854–5870, September 30, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511004579.

Jensen:2011:FRN

[848] Richard Jensen and Chris Cornelis. Fuzzy-rough nearest neighbour classification and prediction. Theoretical Computer Science, 412(42):5871– 5884, September 30, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511004580.

Lingras:2011:CAR

[849] P. Lingras and C. J. Butz. Conservative and aggressive rough SVR modeling. Theoretical Computer Science, 412(42):5885–5901, September 30, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511004609.

Lu:2011:CPI

[850] Juan Lu, Sheng-Gang Li, Xiao-Fei Yang, and Wen-Qing Fu. Categorical properties of M-indiscernibility spaces. Theoretical Computer Science, 412(42):5902–5908, September 30, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511004592.

Pizzi:2011:FQE

[851] Nick J. Pizzi. Fuzzy quartile encoding as a preprocessing method for biomedical pattern classification. Theoretical Computer Science, 412(42):5909– 5925, September 30, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511004610.

Ramanna:2011:NIF

[852] Sheela Ramanna, Amir H. Meghdadi, and James F. Peters. Natureinspired framework for measuring visual image resemblance: a near rough set approach. Theoretical Computer Science, 412(42):5926–5938, September 30, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511004622.

Skowron:2011:ISM

[853] Andrzej Skowron and Piotr Wasilewski. Information systems in modeling interactive computations on granules. Theoretical Computer Science, 412(42):5939–5959, September 30, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511004634.

Szczuka:2011:FNN

[854] Marcin Szczuka and Dominik Ślezak. Feedforward neural networks for compound signals. Theoretical Computer Science, 412(42):5960–5973, September 30, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511004646.

Anonymous:2011:EBaa

[855] Anonymous. Editorial Board. Theoretical Computer Science, 412(42):iii– viii, September 30, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511006992.

Rozenberg:2011:Pg

[856] G. Rozenberg.  Preface.  Theo-

retical Computer Science, 412(43): 5975, October 7, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007201.

Barbuti:2011:SCL

[857] Roberto Barbuti, Andrea Maggiolo-Schettini, Paolo Milazzo, and Giovanni Pardini. Spatial Calculus of Looping Sequences. Theoretical Computer Science, 412(43):5976– 6001, October 7, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511000648.

Chicano:2011:ELD

[858] Francisco Chicano, L. Darrell Whitley, Enrique Alba, and Francisco Luna. Elementary landscape decomposition of the frequency assignment problem.

Theoretical Computer Science, 412(43): 6002–6019, October 7, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511001150.

Doerr:2011:EAD

[859] Benjamin Doerr, Anton Eremeev, Frank Neumann, Madeleine Theile, and Christian Thyssen. Evolutionary algorithms and dynamic programming. Theoretical Computer Science, 412(43): 6020–6035, October 7, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751100661X.

Cordero:2011:SCS

[860] Francesca Cordero, András Horváth, Daniele Manini, Lucia Napione, Massimiliano De Pierro, Simona Pavan, Andrea Picco, Andrea Veglio, Matteo Sereno, Federico Bussolino, and Gianfranco Balbo. Simplification of a complex signal transduction model using invariants and flow equivalent servers. Theoretical Computer Science, 412(43): 6036–6057, October 7, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511005068.

Galpin:2011:EBP

[861] Vashti Galpin. Equivalences for a biological process algebra. Theoretical Computer Science, 412(43):6058– 6082, October 7, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511006165.

Kari:2011:PSA

[862] Lila Kari and Benoît Masson. Polyominoes simulating arbitrary-neighborhood zippers and tilings. Theoretical Computer Science, 412(43):6083–6100, October 7, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751100394X.

Lobo:2011:GGS

[863] Daniel Lobo, Francisco J. Vico, and Jürgen Dassow. Graph grammars with string-regulated rewriting. Theoretical Computer Science, 412(43): 6101–6111, October 7, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511005925.

Soloveichik:2011:ECP

[864] David Soloveichik and Erik Winfree. Erratum to “The computational power of Benenson automata” [Theoret. Comput. Sci. 344 (2005) 279–297]. Theoretical Computer Science, 412(43): 6112–6113, October 7, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511003537. See

[3].

Anonymous:2011:EBab

[865] Anonymous. Editorial Board. Theoretical Computer Science, 412(43): iii–viii, October 7, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007250.

Alechina:2011:RAP

[866] Natasha Alechina, Mehdi Dastani, Brian Logan, and John-Jules Ch. Meyer. Reasoning about plan revision in BDI agent programs. Theoretical Computer Science, 412(44): 6115–6134, October 14, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511004920.

Cousot:2011:GSA

[867] Patrick Cousot and Radhia Cousot. Grammar semantics, analysis and parsing by abstract interpretation. Theoretical Computer Science, 412(44): 6135–6192, October 14, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511004919.

Nakazawa:2011:TCT

[868] Koji Nakazawa, Makoto Tatsuta, Yukiyoshi Kameyama, and Hiroshi Nakano. Type checking and typability in domain-free lambda calculi. Theoretical Computer Science, 412(44): 6193–6207, October 14, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511005391.

Sasturkar:2011:PAA

[869] Amit Sasturkar, Ping Yang, Scott D. Stoller, and C. R. Ramakrishnan. Policy analysis for Administrative Role-Based Access Control. Theoretical Computer Science, 412(44):6208– 6234, October 14, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511003914.

Anonymous:2011:EBac

[870] Anonymous. Editorial Board. Theoretical Computer Science, 412(44): iii–viii, October 14, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007468.

Akama:2011:SSO

[871] Yohji Akama. Set systems: Order types, continuous nondeterministic deformations, and quasi-orders. Theoretical Computer Science, 412(45):6235– 6251, October 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511006712.

Balkova:2011:ACI

[872] L’ubomíra Balková, Karel Brinda, and Ondrej Turek. Abelian complexity of infinite words associated with quadratic Parry numbers. Theoretical Computer Science, 412(45):6252– 6260, October 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751100692X.

Bonomo:2011:BCC

[873] Flavia Bonomo, Sara Mattia, and Gianpaolo Oriolo. Bounded coloring of co-comparability graphs and the pickup and delivery tour combination problem. Theoretical Computer Science, 412(45):6261–6268, October 21, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006220.

Chen:2011:OAO

[874] Xin Chen, Yan Lan, Attila Benko, György Dósa, and Xin Han. Optimal algorithms for online scheduling with bounded rearrangement at the end. Theoretical Computer Science, 412(45): 6269–6278, October 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006244.

Cheng:2011:CMP

[875] Eddie Cheng, Marc J. Lipman, László Lipták, and David Sherman. Conditional matching preclusion for the arrangement graphs. Theoretical Computer Science, 412(45):6279–6289, October 21, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006177.

Fernau:2011:EAM

[876] Henning Fernau, Joachim Kneis, Dieter Kratsch, Alexander Langer, Mathieu Liedloff, Daniel Raible, and Peter Rossmanith. An exact algorithm for the Maximum Leaf Spanning Tree problem. Theoretical Computer Science, 412(45):6290–6302, October 21, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511006219.

Grabner:2011:RMW

[877] Peter J. Grabner and Wolfgang Steiner. Redundancy of minimal weight expansions in Pisot bases. Theoretical Computer Science, 412(45):6303– 6315, October 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511006943.

Gui:2011:VFM

[878] Xiangquan Gui, Xiaohong Hao, Yuanping Zhang, and Xuerong Yong. A volume first maxima-finding algorithm. Theoretical Computer Science, 412(45): 6316–6326, October 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751100689X.

Huang:2011:CSW

[879] Yun Bao Huang. The complexity of smooth words on 2-letter alphabets. Theoretical Computer Science, 412(45): 6327–6339, October 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511005901.

Ito:2011:PCS

[880] Takehiro Ito, Marcin Kami

nski, Daniël Paulusma, and Dimitrios M. Thilikos. Parameterizing cut sets in a graph by the number of their components. Theoretical Computer Science, 412(45):6340– 6350, October 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006049.

Jukna:2011:YHK

[881] Stasys Jukna and Georg Schnitger. Yet harder knapsack problems. Theoretical Computer Science, 412(45):6351– 6358, October 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511005913.

Kiwi:2011:LAS

[882] Marcos Kiwi, Gonzalo Navarro, and Claudio Telha. On-line approximate string matching with bounded errors. Theoretical Computer Science, 412(45): 6359–6370, October 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006669.

Kulich:2011:DMR

[883] Tomás Kulich. Dynamic monopolies with randomized starting configuration. Theoretical Computer Science, 412(45):6371–6381, October 21, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006657.

Kushwah:2011:EIB

[884] Prashant Kushwah and Sunder Lal. An efficient identity based generalized signcryption scheme. Theoretical Computer Science, 412(45):6382– 6389, October 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511006700.

Orenstein:2011:TEC

[885] Yaron Orenstein and Dana Ron. Testing Eulerianity and connectivity in directed sparse graphs. Theoretical Computer Science, 412(45):6390– 6408, October 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511006463.

Park:2011:SMP

[886] Jung-Heum Park and Insung Ihm. Strong matching preclusion. Theoretical Computer Science, 412(45):6409– 6419, October 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511006694.

Woryna:2011:CDA

[887] Adam Woryna. The concept of duality for automata over a changing alphabet and generation of a free group by such automata. Theoretical Computer Science, 412(45):6420– 6431, October 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511006931.

Yamakami:2011:IPC

[888] Tomoyuki Yamakami. Immunity and pseudorandomness of contextfree languages. Theoretical Computer Science, 412(45):6432–6450, October 21, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006232.

Ren:2011:NAC

[889] Wei Ren and Qing Zhao. A note on ‘Algorithms for connected set cover problem and fault-tolerant connected set cover problem’. Theoretical Computer Science, 412(45):6451– 6454, October 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006189. See

[6].

Anonymous:2011:EBad

[890] Anonymous. Editorial Board. Theoretical Computer Science, 412(45): iii–viii, October 21, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007596.

Masse:2011:PCC

[891] A. Blondin Massé, S. Brlek, S. Labbé, and L. Vuillon. Palindromic complexity of codings of rotations. Theoretical Computer Science, 412(46):6455– 6463, October 28, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006682.

Bu:2011:BCP

[892] Yuehua Bu and Yulin Li. Backbone coloring of planar graphs without special circles. Theoretical Computer Science, 412(46):6464–6468, October 28, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511005597.

Chatzigiannakis:2011:PMC

[893] Ioannis Chatzigiannakis, Othon Michail, Stavros Nikolaou, Andreas Pavlogiannis, and Paul G. Spirakis. Passively mobile communicating machines that use restricted space. Theoretical Computer Science, 412(46):6469– 6483, October 28, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511005895.

Fomin:2011:GGG

[894] Fedor V. Fomin, Petr A. Golovach, and Daniel Lokshtanov. Guard games on graphs: Keep the intruder out! Theoretical Computer Science, 412(46): 6484–6497, October 28, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007122.

Herranz:2011:RAO

[895] Javier Herranz. Restricted adaptive oblivious transfer. Theoretical Computer Science, 412(46):6498– 6506, October 28, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511005500.

Kolpakov:2011:NDW

[896] Roman Kolpakov and Michaël Rao. On the number of Dejean words over alphabets of 5, 6, 7, 8, 9 and 10 letters. Theoretical Computer Science, 412(46): 6507–6516, October 28, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006670.

Maltais:2011:HRC

[897] Elizabeth Maltais and Lucia Moura. Hardness results for covering arrays avoiding forbidden edges and errorlocating arrays. Theoretical Computer Science, 412(46):6517–6530, October 28, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006207.

Marini:2011:NAL

[898] Claudio Marini, Giulia Simi, Andrea Sorbi, and Marianna Sorrentino. A note on algebras of languages. Theoretical Computer Science, 412(46): 6531–6536, October 28, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007109.

Martinez:2011:ARQ

[899] Conrado Martínez, Alois Panholzer, and Helmut Prodinger. The analysis of Range Quickselect and related problems. Theoretical Computer Science, 412(46):6537–6555, October 28, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511005573.

Moura:2011:WPT

[900] A. Moura. The word problem for ω-terms over DA. Theoretical Computer Science, 412(46):6556– 6569, October 28, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511006645.

Wang:2011:EHP

[901] Shiying Wang and Shurong Zhang. Embedding Hamiltonian paths in k-ary n-cubes with conditional edge faults. Theoretical Computer Science, 412(46): 6570–6584, October 28, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511001502.

Anonymous:2011:EBae

[902] Anonymous. Editorial Board. Theoretical Computer Science, 412(46): iii–viii, October 28, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007821.

Merro:2011:TCW

[903] Massimo Merro, Francesco Ballardin, and Eleonora Sibilio. A timed calculus for wireless systems. Theoretical Computer Science, 412(47):6585– 6611, November 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006268.

Baaz:2011:FOS

[904] Matthias Baaz, Agata Ciabattoni, and Norbert Preining. First-order satisfiability in Gödel logics: an NP-complete fragment. Theoretical Computer Science, 412(47):6612– 6623, November 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006256.

Anonymous:2011:EBaf

[905] Anonymous. Editorial Board. Theoretical Computer Science, 412(47): iii–viii, November 4, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751100805X.

Badkobeh:2011:FRV

[906] Golnaz Badkobeh. Fewest repetitions versus maximal-exponent powers in infinite binary words. Theoretical Computer Science, 412(48):6625– 6633, November 11, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006724.

Bansal:2011:CAP

[907] Mukul S. Bansal, Jianrong Dong, and David Fernández-Baca. Comparing and aggregating partially resolved trees. Theoretical Computer Science, 412(48):6634–6652, November 11, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007158.

Dombek:2011:NRU

[908] D. Dombek, Z. Masáková, and E. Pelantová. Number representation using generalized −βtransformation. Theoretical Computer Science, 412(48):6653–6665, November 11, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751100716X.

Epstein:2011:OSR

[909] Leah Epstein and Hanan Zebedat-Haider. Online scheduling with rejection and withdrawal. Theoretical Computer Science, 412(48):6666– 6674, November 11, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007195.

Gabarro:2011:CGI

[910] Joaquim Gabarró, Alina García, and Maria Serna. The complexity of game isomorphism. Theoretical Computer Science, 412(48):6675– 6695, November 11, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006323.

Gu:2011:EDR

[911] Xiaoyang Gu and Jack H. Lutz. Effective dimensions and relative frequencies. Theoretical Computer Science, 412(48):6696–6711, November 11, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007110.

Jing:2011:MTC

[912] Caixia Jing, Wanzhen Huang, and Guochun Tang. Minimizing total completion time for re-entrant flow shop scheduling problems. Theoretical Computer Science, 412(48):6712– 6719, November 11, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007183.

Kim:2011:QRR

[913] Changwook Kim. Quasi-rocking realtime pushdown automata. Theoretical Computer Science, 412(48):6720– 6735, November 11, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007171.

Nam:2011:CCB

[914] Wonhong Nam, Hyunyoung Kil, and Dongwon Lee. On the computational complexity of behavioral description-based Web service composition. Theoretical Computer Science, 412(48):6736–6749, November 11, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511003185.

Nikoletseas:2011:INH

[915] S. Nikoletseas, C. Raptopoulos, and P. G. Spirakis. On the independence number and Hamiltonicity of uniform random intersection graphs. Theoretical Computer Science, 412(48):6750– 6760, November 11, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007699.

Paulusma:2011:PGT

[916] Daniël Paulusma and Johan M. M. van Rooij. On partitioning a graph into two connected subgraphs. Theoretical Computer Science, 412(48):6761– 6769, November 11, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007675.

Vladimerou:2011:SDH

[917] Vladimeros Vladimerou, Pavithra Prabhakar, Mahesh Viswanathan, and Geir Dullerud. Specifications for decidable hybrid games. Theoretical Computer Science, 412(48):6770– 6785, November 11, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007407.

Yang:2011:PTA

[918] Chaoxia Yang and Guojun Li. A polynomial time approximation scheme for embedding hypergraph in a weighted cycle. Theoretical Computer Science, 412(48):6786–6793, November 11, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006906.

Anonymous:2011:EBag

[919] Anonymous. Editorial Board. Theoretical Computer Science, 412(48):iii– viii, November 11, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008139.

Bozzelli:2011:HPC

[920] Laura Bozzelli, Axel Legay, and Sophie Pinchinat. Hardness of preorder checking for basic formalisms. Theoretical Computer Science, 412(49):6795– 6808, November 18, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007419.

Danicic:2011:UTC

[921] Sebastian Danicic, Richard W. Barraclough, Mark Harman, John D. Howroyd, Ákos Kiss, and Michael R. Laurence. A unifying theory of control dependence and its application to arbitrary program structures. Theoretical Computer Science, 412(49):6809– 6842, November 18, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007377.

Weller:2011:EQF

[922] Daniel Weller. On the elimination of quantifier-free cuts. Theoretical Computer Science, 412(49):6843– 6854, November 18, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007390.

Anonymous:2011:EBah

[923] Anonymous. Editorial Board. Theoretical Computer Science, 412(49):iii– viii, November 18, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511008310.

Alvarez:2011:RSU

[924] Carme Àlvarez, Maria Blesa, and Maria Serna. The robustness of stability under link and node failures. Theoretical Computer Science, 412(50):6855–6878, November 25, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007493.

Cicalese:2011:CST

[925] Ferdinando Cicalese, Tobias Jacobs, Eduardo Laber, and Marco Molinaro. On the complexity of searching in trees and partially ordered structures. Theoretical Computer Science, 412(50):6879–6896, November 25, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007535.

MacLean:2011:MPE

[926] Stuart MacLean and Suprakash Datta. The minimum positional error incurred by any connectivity-based positioning algorithm for mobile wireless systems. Theoretical Computer Science, 412(50):6897–6912, November 25, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007511.

Tan:2011:EAM

[927] Haisheng Tan, Tiancheng Lou, Yuexuan Wang, Qiang-Sheng Hua, and Francis C. M. Lau. Exact algorithms to minimize interference in wireless sensor networks. Theoretical Computer Science, 412(50):6913– 6925, November 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007523.

Czyzowicz:2011:ADR

[928] Jurek Czyzowicz, David Ilcinkas, Arnaud Labourel, and Andrzej Pelc. Asynchronous deterministic rendezvous in bounded terrains. Theoretical Computer Science, 412(50):6926–6937, November 25, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007687.

Huang:2011:EBS

[929] Chao-Wen Huang, Hui-Ling Huang, and Sun-Yuan Hsieh. Edge-bipancyclicity of star graphs with faulty elements. Theoretical Computer Science, 412(50):6938–6947, November 25, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007729.

Massberg:2011:RPA

[930] Jens Maßberg and Jan Schneider. Rectangle packing with additional restrictions. Theoretical Computer Science, 412(50):6948–6958, November 25, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007730.

I:2011:VEP

[931] Tomohiro I., Shunsuke Inenaga, Hideo Bannai, and Masayuki Takeda. Verifying and enumerating parameterized border arrays. Theoretical Computer Science, 412(50):6959–6981, November 25, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007742.

Cygan:2011:DSF

[932] Marek Cygan, Geevarghese Philip, Marcin Pilipczuk, Michal Pilipczuk, and Jakub Onufry Wojtaszczyk. Dominating set is fixed parameter tractable in claw-free graphs. Theoretical Computer Science, 412(50):6982–7000, November 25, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007766.

Golovach:2011:BFG

[933] Petr Golovach, Pinar Heggernes, Dieter Kratsch, Daniel Lokshtanov, Daniel Meister, and Saket Saurabh. Bandwidth on AT-free graphs. Theoretical Computer Science, 412(50):7001– 7008, November 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007778.

Kardos:2011:CMP

[934] Frantisek Kardos, Ján Katrenic, and Ingo Schiermeyer. On computing the minimum 3-path vertex cover and dissociation number of graphs. Theoretical Computer Science, 412(50):7009– 7017, November 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007754.

Adler:2011:FPA

[935] Isolde Adler, Frederic Dorn, Fedor V. Fomin, Ignasi Sau, and Dimitrios M. Thilikos. Faster parameterized algorithms for minor containment. Theoretical Computer Science, 412(50):7018– 7028, November 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007912.

Luo:2011:NCP

[936] Zhongxuan Luo, Wenyu Hu, and Dongwoo Sheen. The nearest complex polynomial with a zero in a given complex domain. Theoretical Computer Science, 412(50):7029– 7043, November 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007924.

Tu:2011:PDA

[937] Jianhua Tu and Wenli Zhou. A primaldual approximation algorithm for the vertex cover P_{3} problem. Theoretical Computer Science, 412(50):7044– 7048, November 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007894.

Cao:2011:SC

[938] Zhigang Cao and Xiaoguang Yang. Selfish bin covering. Theoretical Computer Science, 412(50):7049–7058, November 25, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007936.

Fleiner:2011:ASS

[939] Tamás Fleiner, Robert W. Irving, and David F. Manlove. An algorithm for a super-stable roommates problem. Theoretical Computer Science, 412(50):7059–7065, November 25, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007882.

Anonymous:2011:EBai

[940] Anonymous. Editorial Board. Theoretical Computer Science, 412(50):iii– viii, November 25, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008425.

Rozenberg:2011:Ph

[941] G. Rozenberg.  Preface.  Theo-

retical Computer Science, 412(51): 7067, December 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008528.

Meyer:2011:UQQ

[942] David A. Meyer and James Pommersheim. On the uselessness of quantum queries. Theoretical Computer Science, 412(51):7068–7074, December 2, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511005883.

Manrique:2011:SBA

[943] Daniel Manrique, Alfonso Rodríguez-Patón, and Petr Sosík. On the scalability of biocomputing algorithms: The case of the maximum clique problem. Theoretical Computer Science, 412(51): 7075–7086, December 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007705.

Anonymous:2011:EBaj

[944] Anonymous. Editorial Board. Theoretical Computer Science, 412(51): iii–viii, December 2, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008577.

Kulik:2011:ASD

[945] Ariel Kulik, Hadas Shachnai, Oded Shmueli, and Robert Sayegh. Approximation schemes for deal splitting and covering integer programs with multiplicity constraints. Theoretical Computer Science, 412(52):7087– 7098, December 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007948.

Andreka:2011:ETK

[946] Hajnal Andréka, Szabolcs Mikulás, and István Németi. The equational theory of Kleene lattices. Theoretical Computer Science, 412(52):7099– 7108, December 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511008000.

Flammini:2011:ORC

[947] Michele Flammini, Gianpiero Monaco, Luca Moscardelli, Mordechai Shalom, and Shmuel Zaks. Optimizing regenerator cost in traffic grooming. Theoretical Computer Science, 412(52):7109– 7121, December 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007997.

McKay:2011:CFU

[948] Neil A. McKay. Canonical forms of uptimals. Theoretical Computer Science, 412(52):7122–7132, December 9, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007973.

Barmpalias:2011:NIS

[949] George Barmpalias and T. F. Sterkenburg. On the number of infinite sequences with trivial initial segment complexity. Theoretical Computer Science, 412(52):7133–7146, December 9, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007961.

Bei:2011:BBB

[950] Xiaohui Bei, Wei Chen, Shang-Hua Teng, Jialin Zhang, and Jiajie Zhu. Bounded budget betweenness centrality game for strategic network formations. Theoretical Computer Science, 412(52):7147– 7168, December 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008218.

Goles:2011:ECC

[951] E. Goles, P.-E. Meunier, I. Rapaport, and G. Theyssier. Erratum to: “Communication Complexity and Intrinsic Universality in Cellular Automata” [Theor. Comput. Sci. 412 (1–2) (2011) 2–21]. Theoretical Computer Science, 412(52):7169– 7170, December 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751100795X. See

Anonymous:2011:EBak

[952] Anonymous. Editorial Board. Theoretical Computer Science, 412(52): iii–viii, December 9, 2011. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511008711.

DiPierro:2012:EQA

[953] Alessandra Di Pierro and Gethin Norman. Editorial: Quantitative aspects of programming languages. Theoretical Computer Science, 413 (1):1, January 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751100898X.

Philippou:2012:PAM

[954] Anna Philippou, Insup Lee, and Oleg Sokolsky. PADS: an approach to modeling resource demand and supply for the formal analysis of hierarchical scheduling. Theoretical Computer Science, 413(1): 2–20, January 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007134.

Cerny:2012:SD

[955] Pavol Cerný, Thomas A. Henzinger, and Arjun Radhakrishna. Simulation distances. Theoretical Computer Science, 413(1):21–35, January 6, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511006633.

Katoen:2012:LPA

[956] Joost-Pieter Katoen, Jaco van de Pol, Mariëlle Stoelinga, and Mark Timmer. A linear process-algebraic format with data for probabilistic automata. Theoretical Computer Science, 413(1): 36–57, January 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006311.

Andova:2012:BBC

[957] Suzana Andova, Sonja Georgievska, and Nikola Trcka. Branching bisimulation congruence for probabilistic systems. Theoretical Computer Science, 413(1):58–72, January 6, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751100630X.

Aldini:2012:AMT

[958] Alessandro Aldini. Approximating Markovian testing equivalence. Theoretical Computer Science, 413(1): 73–86, January 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006293.

Heidarian:2012:ACS

[959] Faranak Heidarian, Julien Schmaltz, and Frits Vaandrager. Analysis of a clock synchronization protocol for wireless sensor networks. Theoretical Computer Science, 413(1):87– 105, January 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511006281.

Hayden:2012:FCP

[960] Richard A. Hayden, Anton Stefanek, and Jeremy T. Bradley. Fluid computation of passage-time distributions in large Markov models. Theoretical Computer Science, 413(1):106– 141, January 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751100627X.

Albert:2012:CAO

[961] Elvira Albert, Puri Arenas, Samir Genaim, German Puebla, and Damiano Zanardini. Cost analysis of objectoriented bytecode programs. Theoretical Computer Science, 413(1): 142–159, January 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511006190.

Bianco:2012:QFS

[962] Alessandro Bianco, Marco Faella, Fabio Mogavero, and Aniello Murano. Quantitatively fair scheduling. Theoretical Computer Science, 413(1): 160–175, January 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511005561.

Anonymous:2012:EBa

[963] Anonymous. Editorial Board. Theoretical Computer Science, 413(1): ii–vii, January 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009029.

Angel:2012:RTA

[964] Eric Angel, Evripidis Bampis, and Nicolas Thibault. Randomized truthful algorithms for scheduling selfish tasks on parallel machines. Theoretical Computer Science, 414(1): 1–8, January 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008371.

Broersma:2012:UCS

[965] Hajo Broersma, Petr A. Golovach, Daniël Paulusma, and Jian Song. Updating the complexity status of coloring graphs without a fixed induced linear forest. Theoretical Computer Science, 414(1):9–19, January 13, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751100836X.

Dekking:2012:PMP

[966] Michel Dekking. Paperfolding morphisms, planefilling curves, and fractal tiles. Theoretical Computer Science, 414(1):20–37, January 13, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008085.

Jesse-Jozefczyk:2012:PCG

[967] Katarzyna Jesse-Józefczyk. The possible cardinalities of global secure sets in cographs. Theoretical Computer Science, 414(1):38– 46, January 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008358.

LeGonidec:2012:CFC

[968] Marion Le Gonidec. On the complexity of a family of k-context-free sequences. Theoretical Computer Science, 414(1): 47–54, January 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007985.

Li:2012:CLT

[969] Angsheng Li and Yicheng Pan. Characterizations of locally testable linearand affine-invariant families. Theoretical Computer Science, 414(1): 55–75, January 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008231.

Petitjean:2012:SST

[970] François Petitjean and Pierre Gançarski. Summarizing a set of time series by averaging: From Steiner sequence to compact multiple alignment. Theoretical Computer Science, 414(1): 76–91, January 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751100822X.

Schmied:2012:AED

[971] Richard Schmied and Claus Viehmann. Approximating edge dominating set in dense graphs. Theoretical Computer Science, 414(1):92–99, January 13, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008243.

Anonymous:2012:EBb

[972] Anonymous. Editorial Board. Theoretical Computer Science, 414(1):iii– viii, January 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009224.

Kamide:2012:PTN

[973] Norihiro Kamide and Heinrich Wansing. Proof theory of Nelson’s paraconsistent logic: a uniform perspective. Theoretical Computer Science, 415(1): 1–38, January 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008978.

Petria:2012:GCB

[974] Marius Petria. Generic constructions for behavioral specifications. Theoretical Computer Science, 415(1): 39–59, January 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008954.

Wang:2012:LFB

[975] Yisong Wang, Jia-Huai You, Li Yan Yuan, Yi-Dong Shen, and Mingyi Zhang. The loop formula based semantics of description logic programs. Theoretical Computer Science, 415(1): 60–85, January 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008966.

DAlessandro:2012:QPL

[976] Flavio D’Alessandro, Benedetto Intrigila, and Stefano Varricchio. Quasipolynomials, linear Diophantine equations and semi-linear sets. Theoretical Computer Science, 416(1):1– 16, January 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751100884X.

Esik:2012:MCF

[977] Zoltán Ésik and Szabolcs Iván. On Müller context-free grammars. Theoretical Computer Science, 416(1): 17–32, January 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008826.

Janson:2012:RTA

[978] Svante Janson. Renewal theory in the analysis of tries and strings. Theoretical Computer Science, 416(1): 33–54, January 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008516.

Khandekar:2012:AFT

[979] Rohit Khandekar, Guy Kortsarz, and Zeev Nutov. Approximating faulttolerant group-Steiner problems. Theoretical Computer Science, 416(1): 55–64, January 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007092.

Kong:2012:SRP

[980] Jiangxu Kong, Weifan Wang, and Xuding Zhu. The surviving rate of planar graphs. Theoretical Computer Science, 416(1):65–70, January 27, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008255.

Okhotin:2012:LEC

[981] Alexander Okhotin and Oksana Yakimova. Language equations with complementation: Expressive power. Theoretical Computer Science, 416(1): 71–86, January 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008267.

Poggi:2012:LTA

[982] Marcus Poggi and David Sotelo. A linear time approximation algorithm for permutation flow shop scheduling. Theoretical Computer Science, 416(1): 87–94, January 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008838.

Anonymous:2012:EBc

[983] Anonymous. Editorial Board. Theoretical Computer Science, 416(1):iii– viii, January 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010309.

Hlineny:2012:P

[984] Petr Hlinený and Antonín Kucera. Preface. Theoretical Computer Science, 417(1):1, February 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010498.

Bezakova:2012:CSM

[985] Ivona Bezáková and Adam J. Friedlander. Counting and sampling minimum (s, t)-cuts in weighted planar graphs in polynomial time. Theoretical Computer Science, 417(1): 2–11, February 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511003999.

Bilo:2012:IAN

[986] Davide Bilò, Luciano Gualà, and Guido Proietti. Improved approximability and non-approximability results for graph diameter decreasing problems. Theoretical Computer Science, 417(1): 12–22, February 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511003963.

Bshouty:2012:TDP

[987] Nader H. Bshouty and Hanna Mazzawi. Toward a deterministic polynomial time algorithm with optimal additive query complexity. Theoretical Computer Science, 417(1):23– 35, February 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007717.

Datta:2012:CCF

[988] Samir Datta, Meena Mahajan, B. V. Raghavendra Rao, Michael Thomas, and Heribert Vollmer. Counting classes and the fine structure between NC^{1} and L. Theoretical Computer Science, 417(1): 36–49, February 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511004737.

David:2012:ACM

[989] Julien David. Average complexity of Moore’s and Hopcroft’s algorithms. Theoretical Computer Science, 417(1): 50–65, February 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008814.

Kalyanasundaram:2012:ISN

[990] Subrahmanyam Kalyanasundaram, Richard J. Lipton, Kenneth W. Regan, and Farbod Shokrieh. Improved simulation of nondeterministic Turing machines. Theoretical Computer Science, 417(1):66– 73, February 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511004002.

Manzonetto:2012:SNC

[991] Giulio Manzonetto and Paolo Tranquilli. Strong normalization of ML^{F} via a calculus of coercions. Theoretical Computer Science, 417(1): 74–94, February 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511004749.

Tazari:2012:FAS

[992] Siamak Tazari. Faster approximation schemes and parameterized algorithms on (odd-) H-minor-free graphs. Theoretical Computer Science, 417(1): 95–107, February 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511007900.

Anonymous:2012:EBd

[993] Anonymous. Editorial Board. Theoretical Computer Science, 417(1):iii– viii, February 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751101053X.

Borodin:2012:SCS

[994] Allan Borodin, Ioana Ivan, Yuli Ye, and Bryce Zimny. On sum coloring and sum multi-coloring for restricted families of graphs. Theoretical Computer Science, 418(1):1– 13, February 10, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009170.

Droste:2012:WAM

[995] Manfred Droste and Heiko Vogler. Weighted automata and multi-valued logics over arbitrary bounded lattices. Theoretical Computer Science, 418(1): 14–36, February 10, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009157.

Anta:2012:DRC

[996] Antonio Fernández Anta, Miguel A. Mosteiro, and Christopher Thraves. Deterministic recurrent communication in restricted sensor networks. Theoretical Computer Science, 418(1): 37–47, February 10, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008887.

Fratani:2012:RSE

[997] Séverine Fratani. Regular sets over extended tree structures. Theoretical Computer Science, 418(1):48– 70, February 10, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008905.

Kolpakov:2012:PSR

[998] Roman Kolpakov. On primary and secondary repetitions in words. Theoretical Computer Science, 418(1):71– 81, February 10, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511008942.

Lai:2012:TCR

[999] Cheng-Nan Lai. Two conditions for reducing the maximal length of nodedisjoint paths in hypercubes. Theoretical Computer Science, 418(1):82– 91, February 10, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511009169.

Xu:2012:DBM

[1000] Hongli Xu, Liusheng Huang, Wang Liu, Yindong Zhang, and Yanjing Sun. Degree-bounded minimum spanning tree for unit disk graph. Theoretical Computer Science, 418(1):92– 105, February 10, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511008899.

Zhang:2012:TCS

[1001] Zhifang Zhang, Yeow Meng Chee, San Ling, Mulan Liu, and Huaxiong Wang. Threshold changeable secret sharing schemes revisited. Theoretical Computer Science, 418(1):106– 115, February 10, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511008206.

Seki:2012:ASI

[1002] Shinnosuke Seki. Absoluteness of subword inequality is undecidable. Theoretical Computer Science, 418(1):116– 120, February 10, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008875.

Anonymous:2012:EBe

[1003] Anonymous. Editorial Board. Theoretical Computer Science, 418(1):iii– viii, February 10, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000072.

Rozenberg:2012:Pa

[1004] G. Rozenberg.  Preface.  The-

oretical Computer Science, 419(1): 1, February 17, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000199.

Barbuti:2012:PMC

[1005] Roberto Barbuti, Francesca Levi, Paolo Milazzo, and Guido Scatena. Probabilistic model checking of biological systems with uncertain kinetic rates. Theoretical Computer Science, 419(1):2–16, February 17, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008929.

Ben-Aroya:2012:BSS

[1006] Avraham Ben-Aroya and Amnon Ta-Shma. Better short-seed quantumproof extractors. Theoretical Computer Science, 419(1):17–25, February 17, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009558.

Caravagna:2012:BPN

[1007] Giulio Caravagna and Jane Hillston. Bio-PEPAd: a non-Markovian extension of Bio-PEPA. Theoretical Computer Science, 419(1):26– 49, February 17, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009479.

Finck:2012:PAS

[1008] Steffen Finck and Hans-Georg Beyer. Performance analysis of the simultaneous perturbation stochastic approximation algorithm on the noisy sphere model. Theoretical Computer Science, 419(1):50–72, February 17, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009340.

Li:2012:COW

[1009] Lvzhou Li, Daowen Qiu, Xiangfu Zou, Lvjun Li, Lihua Wu, and Paulo Mateus. Characterizations of one-way general quantum finite automata. Theoretical Computer Science, 419(1):73– 91, February 17, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008917.

Anonymous:2012:EBf

[1010] Anonymous. Editorial Board. Theoretical Computer Science, 419(1):iii– viii, February 17, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000242.

Giraudo:2012:IBB

[1011] Samuele Giraudo. Intervals of balanced binary trees in the Tamari lattice. Theoretical Computer Science, 420(1): 1–27, February 24, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751100939X.

Golovach:2012:IPO

[1012] Petr A. Golovach, Marcin Kamiński, Daniël Paulusma, and Dimitrios M. Thilikos. Induced packing of odd cycles in planar graphs. Theoretical Computer Science, 420(1):28– 35, February 24, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751100911X.

Kamae:2012:BVC

[1013] Teturo Kamae. Behavior of various complexity functions. Theoretical Computer Science, 420(1):36– 47, February 24, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009315.

Kavitha:2012:PGH

[1014] Telikepalli Kavitha. Properties of Gomory–Hu co-cycle bases. Theoretical Computer Science, 420(1):48– 55, February 24, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009108.

Kirsten:2012:DUP

[1015] Daniel Kirsten. Decidability, undecidability, and PSPACE-completeness of the twins property in the tropical semiring. Theoretical Computer Science, 420(1):56–63, February 24, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009133.

Ron:2012:TCW

[1016] Dana Ron and Gilad Tsur. Testing computability by width-two OBDDs. Theoretical Computer Science, 420(1): 64–79, February 24, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009145.

Moothathu:2012:EEN

[1017] T. K. Subrahmonian Moothathu. Eulerian entropy and non-repetitive subword complexity. Theoretical Computer Science, 420(1):80–88, February 24, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009327.

Wu:2012:SMD

[1018] Weiwei Wu, Minming Li, Wanyong Tian, Jason Chun Xue, and Enhong Chen. Single and multiple device DSA problems, complexities and online algorithms. Theoretical Computer Science, 420(1):89– 98, February 24, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009121.

Anonymous:2012:EBg

[1019] Anonymous. Editorial Board. Theoretical Computer Science, 420(1):iii– viii, February 24, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200045X.

Civril:2012:CSS

[1020] A. Çivril and M. Magdon-Ismail.

Column subset selection via sparse approximation of SVD. Theoretical Computer Science, 421(1):1–14, March 2, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009388.

Chitturi:2012:BPT

[1021] Bhadrachalam Chitturi and I. Hal Sudborough. Bounding prefix transposition distance for strings and permutations. Theoretical Computer Science, 421(1):15–24, March 2, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009376.

Connamacher:2012:ETD

[1022] Harold Connamacher. Exact thresholds for DPLL on random XOR-SAT and NP-complete extensions of XOR-SAT. Theoretical Computer Science, 421(1):25–55, March 2, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511009339.

Elbassioni:2012:CAV

[1023] Khaled Elbassioni and Hans Raj Tiwary. Complexity of approximating the vertex centroid of a polyhedron. Theoretical Computer Science, 421(1):56–61, March 2, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511009364.

Wang:2012:FTB

[1024] Shiying Wang and Yuxing Yang. Fault tolerance in bubble-sort graph networks. Theoretical Computer Science, 421(1):62–69, March 2, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511009352.

Guo:2012:WSA

[1025] Alan Guo. Winning strategies for aperiodic subtraction games. Theoretical Computer Science, 421(1): 70–73, March 2, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511009443.

Anonymous:2012:EBh

[1026] Anonymous. Editorial Board. Theoretical Computer Science, 421(1): iii–viii, March 2, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000734.

Amir:2012:QDP

[1027] Amihood Amir, Yonatan Aumann, Avivit Levy, and Yuri Roshko. Quasi-distinct parsing and optimal compression methods. Theoretical Computer Science, 422(1):1–14, March 9, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009418.

Attrapadung:2012:ABE

[1028] Nuttapong Attrapadung, Javier Herranz, Fabien Laguillaumie, Benoît Libert, Elie de Panafieu, and Carla Ràfols. Attribute-based encryption schemes with constant-size ciphertexts. Theoretical Computer Science, 422 (1):15–38, March 9, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009649.

Fu:2012:CSP

[1029] Bin Fu, Yumei Huo, and Hairong Zhao. Coordinated scheduling of production and delivery with production window and delivery capacity constraints. Theoretical Computer Science, 422(1):39–51, March 9, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511009546.

Larsson:2012:OIS

[1030] Urban Larsson. The ⋆-operator and invariant subtraction games. Theoretical Computer Science, 422(1): 52–58, March 9, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511009406.

Lustrek:2012:IVM

[1031] Mitja Lustrek, Ivan Bratko, and Matjaz Gams. Independent-valued minimax: Pathological or beneficial? Theoretical Computer Science, 422(1): 59–77, March 9, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511009522.

Anonymous:2012:EBi

[1032] Anonymous. Editorial Board. Theoretical Computer Science, 422(1): iii–viii, March 9, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512000916.

Broersma:2012:DCN

[1033] Hajo Broersma, Petr A. Golovach, Daniël Paulusma, and Jian Song. Determining the chromatic number of triangle-free 2P_{3}-free graphs in polynomial time. Theoretical Computer Science, 423(1):1–10, March 16, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511010619.

Chan:2012:NSB

[1034] Ho-Leung Chan, Nicole Megow, René Sitters, and Rob van Stee. A note on sorting buffers offline. Theoretical Computer Science, 423(1): 11–18, March 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511010620.

Gorbenko:2012:SPC

[1035] A. A. Gorbenko and V. Yu. Popov.

The set of parameterized k-covers problem. Theoretical Computer Science, 423(1):19–24, March 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511010188.

Lacroix:2012:NCP

[1036] Mathieu Lacroix, A. Ridha Mahjoub, Sébastien Martin, and Christophe Picouleau. On the NP-completeness of the perfect matching free subgraph problem. Theoretical Computer Science, 423(1):25–29, March 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397511010413.

Russo:2012:MPS

[1037] Luís M. S. Russo. Monge properties of sequence alignment. Theoretical Computer Science, 423(1): 30–49, March 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010449.

Tassa:2012:FAM

[1038] Tamir Tassa. Finding all maximallymatchable edges in a bipartite graph. Theoretical Computer Science, 423(1): 50–58, March 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010474. See

addendum [1706].

vanStee:2012:IAO

[1039] Rob van Stee. An improved algorithm for online rectangle filling. Theoretical Computer Science, 423(1): 59–74, March 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010437.

Ben-Amram:2012:CSE

[1040] Amir M. Ben-Amram and Simon Yoffe. Corrigendum to “A simple and efficient Union-Find-Delete algorithm” [Theoret. Comput. Sci. 412 (4–5) 487– 492]. Theoretical Computer Science, 423(1):75, March 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010590. See

[395].

Anonymous:2012:EBj

[1041] Anonymous. Editorial Board. Theoretical Computer Science, 423(1): iii–viii, March 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001119.

Bezhanishvili:2012:STM

[1042] Nick Bezhanishvili and Ian Hodkinson.

Sahlqvist theorem for modal fixed point logic. Theoretical Computer Science, 424(1):1–19, March 23, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009455.

Ehrhard:2012:SML

[1043] Thomas Ehrhard. The Scott model of linear logic is the extensional collapse of its relational model. Theoretical Computer Science, 424(1): 20–45, March 23, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009467.

Golas:2012:AGT

[1044] Ulrike Golas, Leen Lambers, Hartmut Ehrig, and Fernando Orejas. Attributed graph transformation with inheritance: Efficient conflict detection and local confluence analysis using abstract critical pairs. Theoretical Computer Science, 424(1): 46–68, March 23, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000680.

Anonymous:2012:EBk

[1045] Anonymous. Editorial Board. Theoretical Computer Science, 424(1): iii–viii, March 23, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001697.

Rozenberg:2012:Pb

[1046] G. Rozenberg.  Preface.  The-

oretical Computer Science, 425(1): 1, March 30, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001429.

Lehre:2012:ESI

[1047] Per Kristian Lehre, Frank Neumann, Jonathan E. Rowe, and Xin Yao. Editorial to the special issue on “Theoretical Foundations of Evolutionary Computation”. Theoretical Computer Science, 425(1):2–3, March 30, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001430.

Yoon:2012:QGC

[1048] Yourim Yoon, Yong-Hyuk Kim, Alberto Moraglio, and Byung-Ro Moon. Quotient geometric crossovers and redundant encodings. Theoretical Computer Science, 425(1):4–16, March 30, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511006918.

Doerr:2012:CCP

[1049] Benjamin Doerr, Edda Happ, and Christian Klein. Crossover can provably be useful in evolutionary computation. Theoretical Computer Science, 425(1):17–33, March 30, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397510005979.

Vanneschi:2012:SNB

[1050] Leonardo Vanneschi, Yuri Pirola, Giancarlo Mauri, Marco Tomassini, Philippe Collard, and Sébastien Verel. A study of the neutrality of Boolean function landscapes in genetic programming. Theoretical Computer Science, 425(1):34– 57, March 30, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511002416.

Sutton:2012:CMB

[1051] Andrew M. Sutton, L. Darrell Whitley, and Adele E. Howe. Computing the moments of k-bounded pseudo-Boolean functions over Hamming spheres of arbitrary radius in polynomial time. Theoretical Computer Science, 425(1): 58–74, March 30, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511001101.

Auger:2012:HBM

[1052] Anne Auger, Johannes Bader, Dimo Brockhoff, and Eckart Zitzler. Hypervolumebased multiobjective optimization: Theoretical foundations and practical implications. Theoretical Computer Science, 425(1):75–103, March 30, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397511002428.

Bringmann:2012:ALH

[1053] Karl Bringmann and Tobias Friedrich. Approximating the least hypervolume contributor: NP-hard in general, but fast in practice. Theoretical Computer Science, 425(1): 104–116, March 30, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397510005050.

Witt:2012:AIL

[1054] Carsten Witt. Analysis of an iterated local search algorithm for vertex cover in sparse random graphs. Theoretical Computer Science, 425(1): 117–125, March 30, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511000363.

Stalph:2012:RMS

[1055] Patrick O. Stalph, Xavier Llorà, David E. Goldberg, and Martin V. Butz. Resource management and scalability of the XCSF learning classifier system. Theoretical Computer Science, 425(1):126–141, March 30, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397510003932.

Anonymous:2012:EBl

[1056] Anonymous. Editorial Board. Theoretical Computer Science, 425(1): iii–viii, March 30, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200148X.

Chang:2012:SRA

[1057] Ching-Lueh Chang. Some results on approximate 1-median selection in metric spaces. Theoretical Computer Science, 426–427 (1):1–12, April 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009637.

Yen:2012:CCP

[1058] William Chung-Kung Yen. The connected p-center problem on block graphs with forbidden vertices. Theoretical Computer Science, 426–427 (1):13–24, April 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751100973X.

Gagie:2012:NAW

[1059] Travis Gagie, Gonzalo Navarro, and Simon J. Puglisi. New algorithms on wavelet trees and applications to information retrieval. Theoretical Computer Science, 426–427(1): 25–41, April 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009625.

Grunsky:2012:ALR

[1060] Igor Grunsky, Igor Potapov, and Elena Pryanichnikova. On algebra of languages representable by vertex-labeled graphs. Theoretical Computer Science, 426–427(1): 42–48, April 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007389.

Hempel:2012:IHC

[1061] Harald Hempel and Michael Krüger. Inverse Hamiltonian Cycle and inverse 3 Dimensional Matching are coNP-complete. Theoretical Computer Science, 426–427(1):49–65, April 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397511007080.

Kent:2012:DSS

[1062] Carmel Kent, Moshe Lewenstein, and Dafna Sheinwald. On demand string sorting over unbounded alphabets. Theoretical Computer Science, 426–427 (1):66–74, April 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009613.

Lin:2012:AID

[1063] Min Chih Lin, Francisco J. Soulignac, and Jayme L. Szwarcfiter. Arboricity, h-index, and dynamic algorithms. Theoretical Computer Science, 426–427 (1):75–90, April 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009662.

Su:2012:PKC

[1064] Shenghui Su and Shuwang Lü. A public key cryptosystem based on three new provable problems. Theoretical Computer Science, 426–427 (1):91–117, April 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009716.

Anonymous:2012:EBm

[1065] Anonymous. Editorial Board. Theoretical Computer Science, 426–427 (1):iii–viii, April 6, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001740.

Blanchet-Sadri:2012:TSL

[1066] F. Blanchet-Sadri and Robert Mercas. The three-squares lemma for partial words with one hole. Theoretical Computer Science, 428(1):1– 9, April 13, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000400.

Cameron:2012:CVG

[1067] Kathie Cameron, Benjamin Lévêque, and Frédéric Maffray. Coloring vertices of a graph or finding a Meyniel obstruction. Theoretical Computer Science, 428(1):10–17, April 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009844.

Clouser:2012:TSS

[1068] Thomas Clouser, Mikhail Nesterenko, and Christian Scheideler. Tiara: a self-stabilizing deterministic skip list and skip graph. Theoretical Computer Science, 428(1):18– 35, April 13, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000023.

Conidis:2012:EAM

[1069] Chris J. Conidis. Effectively approximating measurable sets by open sets. Theoretical Computer Science, 428(1):36–46, April 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000394.

Dieudonne:2012:SSG

[1070] Yoann Dieudonné and Franck Petit. Self-stabilizing gathering with strong multiplicity detection. Theoretical Computer Science, 428(1): 47–57, April 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009704.

Giancarlo:2012:APS

[1071] R. Giancarlo and F. Utro. Algorithmic paradigms for stability-based cluster validity and model selection statistical methods, with applications to microarray data analysis. Theoretical Computer Science, 428(1): 58–79, April 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000606.

Kamei:2012:SSA

[1072] Sayaka Kamei and Hirotsugu Kakugawa. A self-stabilizing 6-approximation for the minimum connected dominating set with safe convergence in unit disk graphs. Theoretical Computer Science, 428(1):80–90, April 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009674.

Pan:2012:NPM

[1073] Victor Y. Pan. A note on the paper by Murat Cenk and Ferruh Ozbudak “Multiplication of polynomials modulo x^{n}”, Theoret. Comput. Sci. 412 (2011) 3451–3462. Theoretical Computer Science, 428(1): 91, April 13, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009832. See

erratum [641].

Richomme:2012:CCP

[1074] Gwénaël Richomme and Patrice Séébold. Completing a combinatorial proof of the rigidity of Sturmian words generated by morphisms. Theoretical Computer Science, 428 (1):92–97, April 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009686.

Anonymous:2012:EBn

[1075] Anonymous. Editorial Board. Theoretical Computer Science, 428(1): iii–viii, April 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512002009.

Ausiello:2012:P

[1076] Giorgio Ausiello, Hendrik Jan Hoogeboom, Juhani Karhumäki, Ion Petre, and Arto Salomaa. Preface. Theoretical Computer Science, 429 (1):1–20, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511009856.

terBeek:2012:VTA

[1077] Maurice H. ter Beek and Jetty Kleijn. Vector team automata. Theoretical Computer Science, 429(1):21– 29, April 20, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511009868.

Bodlaender:2012:SCN

[1078] Hans L. Bodlaender and Jurriaan Hage. On switching classes, NLCwidth, cliquewidth and treewidth. Theoretical Computer Science, 429(1): 30–35, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751100987X.

Bonizzoni:2012:RPM

[1079] Paola Bonizzoni, Gianluca Della Vedova, and Riccardo Dondi. A randomized PTAS for the minimum Consensus Clustering with a fixed number of clusters. Theoretical Computer Science, 429(1):36–45, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009881.

Brijder:2012:CGT

[1080] Robert Brijder, Joris J. M. Gillis, and Jan Van den Bussche. A comparison of graph-theoretic DNA hybridization models. Theoretical Computer Science, 429(1):46–53, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009893.

Daley:2012:RC

[1081] Mark Daley, Helmut Jürgensen, Lila Kari, and Kalpana Mahalingam. Relativized codes. Theoretical Computer Science, 429(1):54–64, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S030439751100990X.

Diekert:2012:LTP

[1082] Volker Diekert and Steffen Kopecki.

Language theoretical properties of hairpin formations. Theoretical Computer Science, 429(1):65–73, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397511009911.

Frisco:2012:ASA

[1083] Pierluigi Frisco, Gordon Govan, and Alberto Leporati. Asynchronous P systems with active membranes. Theoretical Computer Science, 429(1): 74–86, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009923.

Gabriel:2012:MEC

[1084] Karsten Gabriel and Hartmut Ehrig. Modelling evolution of communication platforms and scenarios based on transformations of high-level nets and processes. Theoretical Computer Science, 429(1):87–97, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009935.

Gao:2012:SCUa

[1085] Yuan Gao, Lila Kari, and Sheng Yu. State complexity of union and intersection of star on k regular languages. Theoretical Computer Science, 429(1): 98–107, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009947.

Genova:2012:FEG

[1086] Daniela Genova and Natasa Jonoska. Forbidding and enforcing on graphs. Theoretical Computer Science, 429(1): 108–117, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009959.

Harel:2012:MMS

[1087] David Harel and Amir Kantor. Multimodal scenarios revisited: a netbased representation. Theoretical Computer Science, 429(1):118–127, April 20, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009960.

Harju:2012:SFW

[1088] Tero Harju. Square-free words obtained from prefixes by permutations. Theoretical Computer Science, 429(1): 128–133, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009972.

Hirvensalo:2012:PQR

[1089] Mika Hirvensalo. On probabilistic and quantum reaction systems. Theoretical Computer Science, 429(1): 134–143, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009984.

Hosoda:2012:AHM

[1090] Jun Hosoda, Juraj Hromkovic, Taisuke Izumi, Hirotaka Ono, Monika Steinová, and Koichi Wada. On the approximability and hardness of minimum topic connected overlay and its special instances. Theoretical Computer Science, 429(1):144– 154, April 20, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511009996. See

corrigendum [2456].

Ibarra:2012:CEP

[1091] Oscar H. Ibarra and Hsu-Chun Yen. On the containment and equivalence problems for two-way transducers. Theoretical Computer Science, 429(1): 155–163, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010000.

Ito:2012:ASA

[1092] Masami Ito. Algebraic structures of automata. Theoretical Computer Science, 429(1):164–168, April 20, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010012.

Jensen:2012:SLS

[1093] Kurt Jensen, Lars M. Kristensen, and Thomas Mailund. The sweepline state space exploration method. Theoretical Computer Science, 429(1): 169–179, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010024.

Kari:2012:UPG

[1094] Jarkko Kari. Universal pattern generation by cellular automata. Theoretical Computer Science, 429(1): 180–184, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511010036.

Kleijn:2012:LSS

[1095] Jetty Kleijn and Maciej Koutny. Localities in systems with a/sync communication. Theoretical Computer Science, 429(1):185–192, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397511010048.

Kreowski:2012:PGT

[1096] Hans-Jörg Kreowski and Sabine Kuske.

Polynomial graph transformability. Theoretical Computer Science, 429(1): 193–201, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751101005X.

vanLeeuwen:2012:CUP

[1097] Jan van Leeuwen and Jirí Wiedermann.

Computation as an unbounded process. Theoretical Computer Science, 429(1): 202–212, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511010061.

Manzoni:2012:DBP

[1098] Luca Manzoni, Leonardo Vanneschi, and Giancarlo Mauri. A distance between populations for one-point crossover in genetic algorithms. Theoretical Computer Science, 429(1): 213–221, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010073.

Maurer:2012:LOM

[1099] H. Maurer. A linear ordering of a multiparameter universe is usually nonsense. Theoretical Computer Science, 429(1): 222–226, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010085.

Mazurkiewicz:2012:PNL

[1100] Antoni Mazurkiewicz. Petri net languages revisited. Theoretical Computer Science, 429(1):227–235, April 20, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010097.

Nielsen:2012:TTE

[1101] Mogens Nielsen and Karl Krukow. Transfer of trust in event-based reputation systems. Theoretical Computer Science, 429(1):236–246, April 20, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010103.

Okubo:2012:RA

[1102] Fumiya Okubo, Satoshi Kobayashi, and Takashi Yokomori. Reaction automata. Theoretical Computer Science, 429(1):247–257, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397511010115.

Paun:2012:TBT

[1103] Gheorghe Păun and Mario J. Pérez-Jiménez. Towards bridging two cellinspired models: P systems and R systems. Theoretical Computer Science, 429(1):258–264, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397511010127.

Perrin:2012:NSW

[1104] Dominique Perrin and Antonio Restivo.

A note on Sturmian words. Theoretical Computer Science, 429(1): 265–272, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511010139.

Piao:2012:SCC

[1105] Xiaoxue Piao and Kai Salomaa. State complexity of the concatenation of regular tree languages. Theoretical Computer Science, 429(1):273– 281, April 20, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511010140.

Restivo:2012:GTA

[1106] Antonio Restivo and Roberto Vaglica.

A graph theoretic approach to automata minimality. Theoretical Computer Science, 429(1):282–291, April 20, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010152.

Thiagarajan:2012:MDT

[1107] P. S. Thiagarajan and Shaofa Yang. Modular discrete time approximations of distributed hybrid automata. Theoretical Computer Science, 429(1): 292–304, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010164.

Anonymous:2012:EBo

[1108] Anonymous. Editorial Board. Theoretical Computer Science, 429(1): iii–viii, April 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512002149.

Abramsky:2012:P

[1109] Samson Abramsky, Michael Mislove, and Catuscia Palamidessi. Preface. Theoretical Computer Science, 430 (1):1–2, April 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512002423.

Hermida:2012:MIC

[1110] C. Hermida and R. D. Tennent. Monoidal indeterminates and categories of possible worlds. Theoretical Computer Science, 430(1):3– 22, April 27, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000163.

Schmidt:2012:ILT

[1111] David A. Schmidt. Inverse-limit and topological aspects of abstract interpretation. Theoretical Computer Science, 430(1):23–42, April 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000175.

Bauer:2012:FFP

[1112] Andrej Bauer. On the failure of fixedpoint theorems for chain-complete lattices in the effective topos. Theoretical Computer Science, 430(1): 43–50, April 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009650.

Anders:2012:ADQ

[1113] J. Anders, E. Andersson, D. E.

Browne, E. Kashefi, and D. K. L. Oi. Ancilla-driven quantum computation with twisted graph states. Theoretical Computer Science, 430 (1):51–72, April 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001260.

Keimel:2012:EAO

[1114] Klaus Keimel and Jimmie D. Lawson. Extending algebraic operations to D-completions. Theoretical Computer Science, 430(1):73– 87, April 27, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511001356.

Kurz:2012:MSA

[1115] Alexander Kurz and Raul Leal. Modalities in the Stone age: a comparison of coalgebraic logics. Theoretical Computer Science, 430(1): 88–116, April 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512002885.

Lanzagorta:2012:TIS

[1116] Marco Lanzagorta and Keye Martin. Teleportation with an imperfect state. Theoretical Computer Science, 430(1): 117–125, April 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512000187.

Anonymous:2012:EBp

[1117] Anonymous. Editorial Board. Theoretical Computer Science, 430(1): iii–viii, April 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512002472.

Rozenberg:2012:Pc

[1118] G. Rozenberg.  Preface.  The-

oretical Computer Science, 431(1): 1, May 4, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512002587.

Ciobanu:2012:MAB

[1119] Gabriel Ciobanu and Maciej Koutny.

Modelling and analysis of biological systems: Based on papers presented at the Workshop on Membrane Computing and Biologically Inspired Process Calculi (MeCBIC) held in 2008 (Iasi), 2009 (Bologna) and 2010 (Jena). Theoretical Computer Science, 431(1):2–3, May 4, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511010401.

Paun:2012:IHL

[1120] Gheorghe Păun and Mario J. Pérez-Jiménez. An infinite hierarchy of languages defined by dP systems. Theoretical Computer Science, 431 (1):4–12, May 4, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751101019X.

Pinna:2012:MDS

[1121] G. Michele Pinna and Andrea Saba.

Modeling dependencies and simultaneity in membrane system computations. Theoretical Computer Science, 431(1):13–39, May 4, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751101036X.

Cardelli:2012:PS

[1122] Luca Cardelli and Philippa Gardner. Processes in space. Theoretical Computer Science, 431(1):40– 55, May 4, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010176.

Delzanno:2012:RPB

[1123] Giorgio Delzanno and Gianluigi Zavattaro. Reachability problems in BioAmbients. Theoretical Computer Science, 431(1):56–74, May 4, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751101022X.

Coppo:2012:STC

[1124] Mario Coppo, Ferruccio Damiani, Maurizio Drocco, Elena Grassi, Eva Sciacca, Salvatore Spinella, and Angelo Troina. Simulation techniques for the calculus of wrapped compartments. Theoretical Computer Science, 431(1):75–95, May 4, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010395.

Barbuti:2012:FAM

[1125] Roberto Barbuti, Giulio Caravagna, Andrea Maggiolo-Schettini, Paolo Milazzo, and Simone Tini. Foundational aspects of multiscale modeling of biological systems with process algebras. Theoretical Computer Science, 431(1):96–116, May 4, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010243.

Bacci:2012:MSB

[1126] Giorgio Bacci and Marino Miculan. Measurable stochastics for Brane Calculus. Theoretical Computer Science, 431(1):117–136, May 4, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511010218.

Feret:2012:LAR

[1127] Jerome Feret, Thomas Henzinger, Heinz Koeppl, and Tatjana Petrov. Lumpability abstractions of rule-based systems. Theoretical Computer Science, 431(1):137–164, May 4, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397511010255.

Bioglio:2012:TSS

[1128] Livio Bioglio, Mariangiola Dezani-Ciancaglini, Paola Giannini, and Angelo Troina. Typed stochastic semantics for the calculus of looping sequences. Theoretical Computer Science, 431(1):165–180, May 4, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511010383.

Lakin:2012:SSM

[1129] Matthew R. Lakin, Loïc Paulevé, and Andrew Phillips. Stochastic simulation of multiple process calculi for biology. Theoretical Computer Science, 431(1):181–206, May 4, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511010231.

Banks:2012:ATQ

[1130] Richard Banks and L. J. Steggles. An abstraction theory for qualitative models of biological systems. Theoretical Computer Science, 431(1): 207–218, May 4, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010371.

Giavitto:2012:IRN

[1131] Jean-Louis Giavitto, Hanna Klaudel, and Franck Pommereau. Integrated regulatory networks (IRNs): Spatially organized biochemical modules. Theoretical Computer Science, 431(1): 219–234, May 4, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010206.

Anonymous:2012:EBq

[1132] Anonymous. Editorial Board. Theoretical Computer Science, 431(1): iii–viii, May 4, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512002630.

Ann:2012:FAC

[1133] Hsing-Yen Ann, Chang-Biau Yang, Chiou-Ting Tseng, and Chiou-Yi Hor. Fast algorithms for computing the constrained LCS of run-length encoded strings. Theoretical Computer Science, 432(1):1–9, May 11, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512000850.

Biro:2012:ASM

[1134] Péter Biró, David F. Manlove, and Eric J. McDermid. “almost stable” matchings in the Roommates problem with bounded preference lists. Theoretical Computer Science, 432 (1):10–20, May 11, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000588.

Blanchet-Sadri:2012:CPW

[1135] F. Blanchet-Sadri, Aleksandar Chakarov, Lucas Manuelli, Jarett Schwartz, and Slater Stich. Constructing partial words with subword complexities not achievable by full words. Theoretical Computer Science, 432 (1):21–27, May 11, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000862.

Chen:2012:ERA

[1136] Kuan-Yu Chen, Ping-Hui Hsu, and Kun-Mao Chao. Efficient retrieval of approximate palindromes in a run-length encoded string. Theoretical Computer Science, 432(1): 28–37, May 11, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200059X.

Elberfeld:2012:ITT

[1137] Michael Elberfeld, Ilka Schnoor, and Till Tantau. Influence of tree topology restrictions on the complexity of haplotyping with missing data.

Theoretical Computer Science, 432 (1):38–51, May 11, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512000515.

Gambini:2012:NLP

[1138] I. Gambini and L. Vuillon. Non-lattice-periodic tilings of R^{3} by single polycubes. Theoretical Computer Science, 432(1):52–57, May 11, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512000503.

Guo:2012:CED

[1139] Longkun Guo and Hong Shen. On the complexity of the edge-disjoint min-min problem in planar digraphs. Theoretical Computer Science, 432 (1):58–63, May 11, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511009698.

Kontogiannis:2012:MCS

[1140] Spyros Kontogiannis and Paul Spirakis. On mutual concavity and strategically-zero-sum bimatrix games. Theoretical Computer Science, 432 (1):64–76, May 11, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512000527.

Matomaki:2012:NGA

[1141] Kaisa Matomäki and Kalle Saari. A new geometric approach to Sturmian words. Theoretical Computer Science, 432(1):77–84, May 11, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000977.

Salomaa:2012:SOW

[1142] Arto Salomaa. Subword occurrences, weighted automata and iterated morphisms, especially the Fibonacci morphism. Theoretical Computer Science, 432(1):85–93, May 11, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000539.

Honkala:2012:MDS

[1143] Juha Honkala. Marked D0L systems and the 2n-conjecture. Theoretical Computer Science, 432(1): 94–97, May 11, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000849.

Anonymous:2012:EBr

[1144] Anonymous. Editorial Board. Theoretical Computer Science, 432(1): iii–viii, May 11, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512002800.

Crole:2012:AEE

[1145] Roy L. Crole. Alpha equivalence equalities. Theoretical Computer Science, 433(1):1–19, May 18, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000667.

Diaconescu:2012:AAS

[1146] Razvan Diaconescu. An axiomatic approach to structuring specifications. Theoretical Computer Science, 433 (1):20–42, May 18, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512002071.

Huang:2012:VPC

[1147] Shuqin Huang, Yongzhi Cao, Hanpin Wang, and Wanling Qu. Valuepassing CCS with noisy channels. Theoretical Computer Science, 433 (1):43–59, May 18, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512002083.

Naumann:2012:RRI

[1148] David A. Naumann, Augusto Sampaio, and Leila Silva. Refactoring and representation independence for class hierarchies. Theoretical Computer Science, 433(1):60–97, May 18, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001284.

Anonymous:2012:EBs

[1149] Anonymous. Editorial Board. Theoretical Computer Science, 433(1): iii–viii, May 18, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200299X.

Bao:2012:AAS

[1150] Xiaoguang Bao and Zhaohui Liu. Approximation algorithms for single vehicle scheduling problems with release and service times on a tree or cycle. Theoretical Computer Science, 434(1):1–10, May 25, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200103X.

Bonato:2012:FCF

[1151] Anthony Bonato, Margaret-Ellen Messinger, and Pawel Pralat. Fighting constrained fires in graphs. Theoretical Computer Science, 434(1): 11–22, May 25, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000989.

Crochemore:2012:IAR

[1152] Maxime Crochemore, Costas S. Iliopoulos, Marcin Kubica, M. Sohel Rahman, German Tischler, and Tomasz Waleń. Improved algorithms for the range next value problem and applications. Theoretical Computer Science, 434(1): 23–34, May 25, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001569.

Karakostas:2012:DAC

[1153] George Karakostas, Jeff Kinne, and Dieter van Melkebeek. On derandomization and average-case complexity of monotone functions. Theoretical Computer Science, 434(1):

35–44, May 25, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001582.

Liao:2012:NSP

[1154] Ching-Jong Liao, Hsin-Hui Tsou, and Kuo-Ling Huang. Neighborhood search procedures for single machine tardiness scheduling with sequence-dependent setups. Theoretical Computer Science, 434(1):45–52, May 25, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001004.

Shang:2012:TCB

[1155] Yun Shang, Xian Lu, and Ruqian Lu. A theory of computation based on unsharp quantum logic: Finite state automata and pushdown automata. Theoretical Computer Science, 434(1):53–86, May 25, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001594.

Ueno:2012:SLB

[1156] Kenya Ueno. A stronger LP bound for formula size lower bounds via clique constraints. Theoretical Computer Science, 434(1):87– 97, May 25, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001247.

Yoshida:2012:TDG

[1157] Yuichi Yoshida and Yusuke Kobayashi. Testing the (s, t)-disconnectivity of graphs and digraphs. Theoretical Computer Science, 434(1):98– 113, May 25, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001028.

Anonymous:2012:EBt

[1158] Anonymous. Editorial Board. Theoretical Computer Science, 434(1): iii–viii, May 25, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003325.

Blume:2012:P

[1159] Matthias Blume and Germán Vidal.

Preface. Theoretical Computer Science, 435(1):1–2, June 1, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001624.

Berardi:2012:IMS

[1160] Stefano Berardi and Makoto Tatsuta. Internal models of system F for decompilation. Theoretical Computer Science, 435(1):3– 20, June 1, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001636.

Danvy:2012:IDS

[1161] Olivier Danvy, Kevin Millikin, Johan Munk, and Ian Zerny. On interderiving small-step and big-step semantics: a case study for storeless call-by-need evaluation. Theoretical Computer Science, 435(1): 21–42, June 1, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001648.

Howe:2012:PSS

[1162] Jacob M. Howe and Andy King. A pearl on SAT and SMT solving in Prolog. Theoretical Computer Science, 435(1):43–55, June 1, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200165X.

Kiselyov:2012:DCO

[1163] Oleg Kiselyov. Delimited control in OCaml, abstractly and concretely. Theoretical Computer Science, 435 (1):56–76, June 1, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001661.

Remy:2012:CSI

[1164] Didier Rémy and Boris Yakobowski. A church-style intermediate language for MLF. Theoretical Computer Science, 435(1):77–105, June 1, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001673.

Saurin:2012:BTB

[1165] Alexis Saurin. Böhm theorem and Böhm trees for the Λμ-calculus. Theoretical Computer Science, 435(1): 106–138, June 1, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001685.

Anonymous:2012:EBu

[1166] Anonymous. Editorial Board. Theoretical Computer Science, 435(1): iii–viii, June 1, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200343X.

Rozenberg:2012:Pd

[1167] G. Rozenberg.  Preface.  The-

oretical Computer Science, 436(1): 1, June 8, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003647.

Berstel:2012:SSC

[1168] Jean Berstel, Luc Boasson, and Isabelle Fagnot. Splicing systems and the Chomsky hierarchy. Theoretical Computer Science, 436(1):2– 22, June 8, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512002289.

Cabessa:2012:EPA

[1169] Jérémie Cabessa and Alessandro E. P.

Villa. The expressive power of analog recurrent neural networks on infinite input streams. Theoretical Computer Science, 436(1):23– 34, June 8, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512000990.

Cattaneo:2012:SPP

[1170] G. Cattaneo, M. Comito, and D. Bianucci. Sand piles: From physics to cellular automata models. Theoretical Computer Science, 436(1):35–53, June 8, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001867.

Chen:2012:LPS

[1171] Tianshi Chen, Ke Tang, Guoliang Chen, and Xin Yao. A large population size can be unhelpful in evolutionary algorithms. Theoretical Computer Science, 436(1):54– 70, June 8, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511001368.

Doerr:2012:NEL

[1172] Benjamin Doerr, Daniel Johannsen, and Carola Winzen. Non-existence of linear universal drift functions. Theoretical Computer Science, 436 (1):71–86, June 8, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001053.

Lettieri:2012:AIF

[1173] Giuseppe Lettieri. an abstract interpretation framework for genotype elimination algorithms. Theoretical Computer Science, 436(1):87– 105, June 8, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512002940.

Zhou:2012:UTI

[1174] Dong Zhou, Dan Luo, Ruqian Lu, and Zhangang Han. The use of tail inequalities on the probable computational time of randomized search heuristics. Theoretical Computer Science, 436(1): 106–117, June 8, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000382.

Anonymous:2012:EBv

[1175] Anonymous. Editorial Board. Theoretical Computer Science, 436(1): iii–viii, June 8, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003696.

Adali:2012:AOL

[1176] Sibel Adali, Tina Liu, and Malik Magdon-Ismail. An analysis of optimal link bombs. Theoretical Computer Science, 437(1):1– 20, June 15, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001600.

Beal:2012:PLP

[1177] Richard Beal and Donald Adjeroh.

Parameterized longest previous factor. Theoretical Computer Science, 437 (1):21–34, June 15, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001235.

Chebolu:2012:CAC

[1178] Prasad Chebolu, Leslie Ann Goldberg, and Russell Martin. The complexity of approximately counting stable matchings. Theoretical Computer Science, 437(1):35–68, June 15, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001818.

Chuan:2012:FWP

[1179] Wai-Fong Chuan, Fang-Yi Liao, Hui-Ling Ho, and Fei Yu. Fibonacci word patterns in two-way infinite Fibonacci words. Theoretical Computer Science, 437(1):69–81, June 15, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001612.

Cui:2012:SCC

[1180] Bo Cui, Yuan Gao, Lila Kari, and Sheng Yu. State complexity of combined operations with two basic operations. Theoretical Computer Science, 437(1):82–102, June 15, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751200182X.

Anonymous:2012:EBw

[1181] Anonymous. Editorial Board. Theoretical Computer Science, 437(1): iii–viii, June 15, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200391X.

Kobayashi:2012:DCS

[1182] Yuji Kobayashi. The derivational complexity of string rewriting systems. Theoretical Computer Science, 438(1):1–12, June 22, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001892.

Christodoulou:2012:CAP

[1183] George Christodoulou, Vahab S. Mirrokni, and Anastasios Sidiropoulos. Convergence and approximation in potential games. Theoretical Computer Science, 438(1):13– 27, June 22, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001855.

Kim:2012:SLA

[1184] Changwook Kim. On the structure of linear apex NLC graph grammars. Theoretical Computer Science, 438 (1):28–33, June 22, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001909.

Mertzios:2012:RTG

[1185] George B. Mertzios. The recognition of triangle graphs. Theoretical Computer Science, 438(1): 34–47, June 22, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001946.

Bang-Jensen:2012:ADS

[1186] Jørgen Bang-Jensen and Anders Yeo. Arc-disjoint spanning sub(di)graphs in digraphs. Theoretical Computer Science, 438(1):48–54, June 22, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397512002204.

Liu:2012:NRS

[1187] Ming Liu, Feifeng Zheng, Chengbin Chu, and Yinfeng Xu. New results on single-machine scheduling with past-sequence-dependent delivery times. Theoretical Computer Science, 438(1):55–61, June 22, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512002290.

Becher:2012:LCM

[1188] Verónica Becher and Pablo Ariel Heiber. A linearly computable measure of string complexity. Theoretical Computer Science, 438(1): 62–73, June 22, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512002277.

Munro:2012:SRP

[1189] J. Ian Munro, Rajeev Raman, Venkatesh Raman, and Srinivasa Rao S. Succinct representations of permutations and functions. Theoretical Computer Science, 438(1):74– 88, June 22, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512002253.

Kaminski:2012:MCC

[1190] Marcin Kamiński. max-cut and containment relations in graphs. Theoretical Computer Science, 438(1): 89–95, June 22, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001880.

Elbassioni:2012:RCS

[1191] Khaled Elbassioni, Slobodan Jelić, and Domagoj Matijević. The relation of Connected Set Cover and Group Steiner Tree. Theoretical Computer Science, 438(1):96– 101, June 22, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001879.

Anonymous:2012:EBx

[1192] Anonymous. Editorial Board. Theoretical Computer Science, 438(1): iii–viii, June 22, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200415X.

Jiang:2012:AAS

[1193] Haitao Jiang, Daming Zhu, and Binhai Zhu. A (1 + ϵ)-approximation algorithm for sorting by short blockmoves. Theoretical Computer Science, 439(1):1–8, June 29, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512002393.

Kaminski:2012:CIS

[1194] Marcin Kamiński, Paul Medvedev, and Martin Milanic. Complexity of independent set reconfigurability problems. Theoretical Computer Science, 439(1):9–15, June 29, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512002241.

Kerivin:2012:CEC

[1195] H. L. M. Kerivin, M. Lacroix, and A. R. Mahjoub. On the complexity of the Eulerian closed walk with precedence path constraints problem. Theoretical Computer Science, 439 (1):16–29, June 29, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512002344.

Miklos:2012:AND

[1196] István Miklós and Eric Tannier. Approximating the number of Double Cut-and-Join scenarios. Theoretical Computer Science, 439(1): 30–40, June 29, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512002265.

Nikolopoulos:2012:FDA

[1197] Stavros D. Nikolopoulos, Leonidas Palios, and Charis Papadopoulos. A fully dynamic algorithm for the recognition of P_{4}-sparse graphs. Theoretical Computer Science, 439(1): 41–57, June 29, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200240X.

Prodinger:2012:ACC

[1198] Helmut Prodinger. Approximate counting with m counters: a detailed analysis. Theoretical Computer Science, 439(1):58–68, June 29, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397512002368.

Xin:2012:AOD

[1199] Qin Xin, Fredrik Manne, Yan Zhang, and Xin Wang. Almost optimal distributed M2M multicasting in wireless mesh networks. Theoretical Computer Science, 439(1): 69–82, June 29, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512002356.

Anonymous:2012:EBy

[1200] Anonymous. Editorial Board. Theoretical Computer Science, 439(1): iii–viii, June 29, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004380.

Balogh:2012:NLB

[1201] János Balogh, József Békési, and Gábor Galambos. New lower bounds for certain classes of bin packing algorithms. Theoretical Computer Science, 440–441:1–13, ???? 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003611.

Bulteau:2012:TAM

[1202] Laurent Bulteau, Guillaume Fertin, Minghui Jiang, and Irena Rusu. Tractability and approximability of maximal strip recovery. Theoretical Computer Science, 440–441:14– 28, ???? 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003994.

Challita:2012:SDA

[1203] Khalil Challita. A semi-dynamical approach for solving qualitative spatial constraint satisfaction problems. Theoretical Computer Science, 440– 441:29–38, ???? 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200326X.

Chuan:2012:LFC

[1204] Wai-Fong Chuan and Hui-Ling Ho. Locating factors of a characteristic word via the generalized Zeckendorf representation of numbers. Theoretical Computer Science, 440– 441:39–51, ???? 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003593.

Czyzowicz:2012:CBA

[1205] Jurek Czyzowicz, Leszek Gasieniec, and Andrzej Pelc. Choosing the best among peers. Theoretical Computer Science, 440–441: 52–59, ???? 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003271.

Pinto:2012:EAA

[1206] Paulo E. D. Pinto, Fábio Protti, and Jayme L. Szwarcfiter. Exact and approximation algorithms for error-detecting even codes. Theoretical Computer Science, 440–441: 60–72, ???? 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003192.

Restivo:2012:EMC

[1207] Antonio Restivo and Roberto Vaglica.

Extremal minimality conditions on automata. Theoretical Computer Science, 440–441:73–84, ???? 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003210.

Anonymous:2012:EBz

[1208] Anonymous. Editorial Board. Theoretical Computer Science, 440–441: iii–viii, ???? 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004501.

Nandy:2012:P

[1209] Subhas C. Nandy and Sandeep Sen. Preface. Theoretical Computer Science, 442(1):1, July 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751200463X.

Alt:2012:SMR

[1210] Helmut Alt and Ludmila Scharf. Shape matching by random sampling. Theoretical Computer Science, 442 (1):2–12, July 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751000160X.

Mahajan:2012:PMP

[1211] Meena Mahajan, Prajakta Nimbhorkar, and Kasturi Varadarajan. The planar k-means problem is NP-hard. Theoretical Computer Science, 442 (1):13–21, July 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397510003269.

Yamanaka:2012:EEO

[1212] Katsuhisa Yamanaka, Yota Otachi, and Shin ichi Nakano. Efficient enumeration of ordered trees with k leaves. Theoretical Computer Science, 442 (1):22–27, July 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751100051X.

Anonymous:2012:EBaa

[1213] Anonymous. Editorial Board. Theoretical Computer Science, 442(1): ii–vii, July 13, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004689.

Albers:2012:SOS

[1214] Susanne Albers and Matthias Hellwig. Semi-online scheduling revisited. Theoretical Computer Science, 443(1):1–9, July 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512002939.

Bang-Jensen:2012:FIS

[1215] Jørgen Bang-Jensen, Frédéric Havet, and Nicolas Trotignon. Finding an induced subdivision of a digraph. Theoretical Computer Science, 443 (1):10–24, July 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200237X.

Bille:2012:SMV

[1216] Philip Bille, Inge Li Gørtz, Hjalte Wedel Vildhøj, and David Kofoed Wind. String matching with variable length gaps. Theoretical Computer Science, 443(1):25–34, July 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512002915.

Blanchet-Sadri:2012:PAC

[1217] F. Blanchet-Sadri, Robert Mercas, Abraham Rashin, and Elara Willett. Periodicity algorithms and a conjecture on overlaps in partial words. Theoretical Computer Science, 443 (1):35–45, July 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003064.

Fedou:2012:ADW

[1218] Jean-Marc Fédou and Gabriele Fici. Automata and differentiable words. Theoretical Computer Science, 443 (1):46–62, July 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003052.

Hsieh:2012:EAC

[1219] Sun-Yuan Hsieh and Ying-Hsuan Chang. Extraconnectivity of kary n-cube networks. Theoretical Computer Science, 443(1):63– 69, July 20, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512002927.

Liu:2012:GDC

[1220] Yun Liu. Groups and decompositions of codes. Theoretical Computer Science, 443(1):70–81, July 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200312X.

Anonymous:2012:EBab

[1221] Anonymous. Editorial Board. Theoretical Computer Science, 443(1): ii–vii, July 20, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004896.

Kutten:2012:P

[1222] Shay Kutten and Janez Zerovnik. Preface. Theoretical Computer Science, 444(1):1–2, July 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000321.

Bermond:2012:GLS

[1223] Jean-Claude Bermond, David Coudert, Joanna Moulierac, Stéphane Pérennes, Ignasi Sau, and Fernando Solano Donado. GMPLS label space minimization through hypergraph layouts. Theoretical Computer Science, 444 (1):3–16, July 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000801.

Nisse:2012:DCE

[1224] Nicolas Nisse, Ivan Rapaport, and Karol Suchan. Distributed computing of efficient routing schemes in generalized chordal graphs. Theoretical Computer Science, 444(1): 17–27, July 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000333.

Blair:2012:ESS

[1225] Jean R. S. Blair and Fredrik Manne. An efficient self-stabilizing distance-2 coloring algorithm. Theoretical Computer Science, 444(1):28– 39, July 27, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000813.

Bar-Noy:2012:OCG

[1226] Amotz Bar-Noy, Panagiotis Cheilaris, Michael Lampis, Valia Mitsou, and Stathis Zachos. Ordered coloring of grids and related graphs. Theoretical Computer Science, 444(1): 40–51, July 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200401X.

Bilo:2012:RVG

[1227] D. Bilò, Y. Disser, M. Mihalák, S. Suri, E. Vicari, and P. Widmayer. Reconstructing visibility graphs with simple robots. Theoretical Computer Science, 444(1):52–59, July 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000357.

Czyzowicz:2012:MEP

[1228] Jurek Czyzowicz, Stefan Dobrev, Leszek Gasieniec, David Ilcinkas, Jesper Jansson, Ralf Klasing, Ioannis Lignos, Russell Martin, Kunihiko Sadakane, and Wing-Kin Sung. More efficient periodic traversal in anonymous undirected graphs. Theoretical Computer Science, 444(1):

60–76, July 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000825.

Yamamoto:2012:OTU

[1229] Kenta Yamamoto, Taisuke Izumi, Yoshiaki Katayama, Nobuhiro Inuzuka, and Koichi Wada. The optimal tolerance of uniform observation error for mobile robot convergence. Theoretical Computer Science, 444(1):77–86, July 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004033.

Even:2012:RRP

[1230] Guy Even and Moti Medina. Revisiting randomized parallel load balancing algorithms. Theoretical Computer Science, 444(1):87– 99, July 27, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000369.

Sudo:2012:LSL

[1231] Yuichi Sudo, Junya Nakamura, Yukiko Yamauchi, Fukuhito Ooshita, Hirotsugu Kakugawa, and Toshimitsu Masuzawa. Loosely-stabilizing leader election in a population protocol model. Theoretical Computer Science, 444(1): 100–112, July 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000345.

Imbs:2012:VWC

[1232] Damien Imbs and Michel Raynal. Virtual world consistency: a condition for STM systems (with a versatile protocol with invisible read operations). Theoretical Computer Science, 444(1): 113–127, July 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004021.

Anonymous:2012:EBac

[1233] Anonymous. Editorial Board. Theoretical Computer Science, 444(1): iii–viii, July 27, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512005014.

Geffert:2012:AHF

[1234] Viliam Geffert. An alternating hierarchy for finite automata. Theoretical Computer Science, 445(1): 1–24, August 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004094.

Hong:2012:BCI

[1235] Yanmei Hong and Liying Kang. Backup 2-center on interval graphs. Theoretical Computer Science, 445(1): 25–35, August 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004240.

Hong:2012:MCS

[1236] Seok-Hee Hong and Hiroshi Nagamochi. Minimum cost star-shaped drawings of plane graphs with a fixed embedding and concave corner constraints. Theoretical Computer Science, 445(1):36–51, August 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004331.

Hu:2012:VMS

[1237] Chunqiang Hu, Xiaofeng Liao, and Xiuzhen Cheng. Verifiable multi-secret sharing based on LFSR sequences. Theoretical Computer Science, 445(1): 52–62, August 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004276.

Klouda:2012:BFC

[1238] Karel Klouda. Bispecial factors in circular non-pushy D0L languages. Theoretical Computer Science, 445(1): 63–74, August 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004288.

Liu:2012:OAO

[1239] Ming Liu, Feifeng Zheng, Shijin Wang, and Jiazhen Huo. Optimal algorithms for online single machine scheduling with deteriorating jobs. Theoretical Computer Science, 445(1): 75–81, August 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004252.

Mezzini:2012:FDA

[1240] Mauro Mezzini. Fully dynamic algorithm for chordal graphs with O(1) query-time and O(n^{2}) update-time. Theoretical Computer Science, 445(1): 82–92, August 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004239.

Anonymous:2012:EBad

[1241] Anonymous. Editorial Board. Theoretical Computer Science, 445(1): iii–viii, August 3, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512005208.

Cao:2012:MBH

[1242] Zining Cao. More on bisimulations for higher order π-calculus. Theoretical Computer Science, 446(1): 1–19, August 10, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512003581.

Khadim:2012:RTD

[1243] U. Khadim and P. J. L. Cuijpers. Repairing time-determinism in the process algebra for hybrid systems ACP_{hs}^{srt}. Theoretical Computer Science, 446(1): 20–47, August 10, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004859.

Motallebi:2012:MVH

[1244] Hassan Motallebi and Mohammad Abdollahi Azgomi. Modeling and verification of hybrid dynamic systems using multisingular hybrid Petri nets. Theoretical Computer Science, 446(1): 48–74, August 10, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004811.

Zhou:2012:AMP

[1245] Chunlai Zhou and Mingsheng Ying.

Approximating Markov processes through filtration. Theoretical Computer Science, 446(1):75–97, August 10, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512002873.

Anonymous:2012:EBae

[1246] Anonymous. Editorial Board. Theoretical Computer Science, 446(1): iii–viii, August 10, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005506.

Wu:2012:P

[1247] Weili Wu and Ovidiu Daescu. Preface. Theoretical Computer Science, 447(1):1, August 17, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005579.

Bollig:2012:SOB

[1248] Beate Bollig. On symbolic OBDDbased algorithms for the minimum spanning tree problem. Theoretical Computer Science, 447(1):2–12, August 17, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009480.

Chen:2012:CFT

[1249] Danny Z. Chen and Ewa Misiolek. Computing feasible toolpaths for 5-axis machines. Theoretical Computer Science, 447(1): 13–25, August 17, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009510.

Chen:2012:PCS

[1250] Xujin Chen, Xiaodong Hu, and Weidong Ma. Pairwise cooperations in selfish ring routing for minimax linear latency. Theoretical Computer Science, 447(1):26– 37, August 17, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009509.

Du:2012:PTA

[1251] Hongwei Du, Qiang Ye, Jiaofei Zhong, Yuexuan Wang, Wonjun Lee, and Haesun Park. Polynomial-time approximation scheme for minimum connected dominating set under routing cost constraint in wireless sensor networks. Theoretical Computer Science, 447(1):

38–43, August 17, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008644.

Eppstein:2012:EDS

[1252] David Eppstein, Michael T. Goodrich, Darren Strash, and Lowell Trott. Extended dynamic subgraph statistics using h-index parameterized data structures. Theoretical Computer Science, 447(1):44–52, August 17, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009534.

Fan:2012:ROG

[1253] Neng Fan, Qipeng P. Zheng, and Panos M. Pardalos. Robust optimization of graph partitioning involving interval uncertainty. Theoretical Computer Science, 447(1): 53–61, August 17, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008851.

He:2012:CVR

[1254] Xin He, Jiun-Jie Wang, and Huaming Zhang. Compact visibility representation of 4-connected plane graphs. Theoretical Computer Science, 447(1): 62–73, August 17, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001296.

Italiano:2012:FSB

[1255] Giuseppe F. Italiano, Luigi Laura, and Federico Santaroni. Finding strong bridges and strong articulation points in linear time. Theoretical Computer Science, 447(1): 74–84, August 17, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511009303.

Langetepe:2012:SAP

[1256] Elmar Langetepe. Searching for an axis-parallel shoreline. Theoretical Computer Science, 447(1):85– 99, August 17, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010450.

Li:2012:CSL

[1257] Deying Li, Zheng Li, Wenkai Ma, Hong Chen, and Wenping Chen. Constrained surface-level gateway placement for underwater acoustic wireless sensor networks. Theoretical Computer Science, 447(1):100– 106, August 17, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511008632.

Shen:2012:NTA

[1258] Yilin Shen, Dung T. Nguyen, Ying Xuan, and My T. Thai. New techniques for approximating optimal substructure problems in power-law graphs. Theoretical Computer Science, 447(1): 107–119, August 17, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008930.

Yamakami:2012:DTA

[1259] Tomoyuki Yamakami. A dichotomy theorem for the approximate counting of complex-weighted boundeddegree Boolean CSPs. Theoretical Computer Science, 447(1):120– 135, August 17, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003088.

Yu:2012:BPN

[1260] Guanglong Yu, Zhengke Miao, and Jinlong Shu. Bases of primitive nonpowerful sign patterns. Theoretical Computer Science, 447(1):136– 143, August 17, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511008863.

Anonymous:2012:EBaf

[1261] Anonymous. Editorial Board. Theoretical Computer Science, 447(1): iii–viii, August 17, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005610.

Chen:2012:NCA

[1262] Lei Chen, Weiming Zeng, and Changhong Lu. NP-completeness and APX-completeness of restrained domination in graphs. Theoretical Computer Science, 448(1):1–8, August 24, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004264.

Fotakis:2012:EMS

[1263] Dimitris Fotakis, Alexis C. Kaporis, and Paul G. Spirakis. Efficient methods for selfish network design. Theoretical Computer Science, 448(1): 9–20, August 24, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512003982.

Freydenberger:2012:WUM

[1264] Dominik D. Freydenberger, Hossein Nevisi, and Daniel Reidenbach. Weakly unambiguous morphisms. Theoretical Computer Science, 448(1):21– 40, August 24, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004100.

Garcia:2012:PCS

[1265] Pedro García, Damián López, and Manuel Vázquez de Parga. Polynomial characteristic sets for DFA identification. Theoretical Computer Science, 448(1):41–46, August 24, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004070.

Ghiyasvand:2012:TAS

[1266] Mehdi Ghiyasvand. An O(m(m + n log n) log(nC))-time algorithm to solve the minimum cost tension problem. Theoretical Computer Science, 448(1):47–55, August 24, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397512003143.

Koiran:2012:ACC

[1267] Pascal Koiran. Arithmetic circuits:

the chasm at depth four gets wider. Theoretical Computer Science, 448(1): 56–65, August 24, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003131.

Nishida:2012:EEA

[1268] Taishin Y. Nishida. An extended Earley’s algorithm for Petri net controlled grammars without λ rules and cyclic rules. Theoretical Computer Science, 448(1):66–79, August 24, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004082.

Strand:2012:DTC

[1269] Robin Strand and Nicolas Normand.

Distance transform computation for digital distance functions. Theoretical Computer Science, 448(1):80– 93, August 24, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200432X.

Basic:2012:NPS

[1270] Bojan Basić. A note on the paper “On Brlek–Reutenauer conjecture”. Theoretical Computer Science, 448(1): 94–96, August 24, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004045. See

[826] and corrigendum [1464].

Anonymous:2012:EBag

[1271] Anonymous. Editorial Board. Theoretical Computer Science, 448(1): iii–viii, August 24, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512005919.

Holzer:2012:P

[1272] Markus Holzer, Martin Kutrib, and Giovanni Pighizzini. Preface. Theoretical Computer Science, 449(1): 1, August 31, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751200607X.

Baldan:2012:EUC

[1273] Paolo Baldan, Alessandro Bruni, Andrea Corradini, Barbara König, César Rodríguez, and Stefan Schwoon. Efficient unfolding of contextual Petri nets. Theoretical Computer Science, 449(1):2–22, August 31, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004318.

Bednarova:2012:SCB

[1274] Zuzana Bednárová, Viliam Geffert, Carlo Mereghetti, and Beatrice Palano. The size-cost of Boolean operations on constant height deterministic pushdown automata. Theoretical Computer Science, 449(1):23– 36, August 31, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004306.

Brzozowski:2012:SCP

[1275] Janusz Brzozowski, Baiyu Li, and Yuli Ye. Syntactic complexity of prefix-, suffix-, bifix-, and factorfree regular languages. Theoretical Computer Science, 449(1):37– 53, August 31, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003556.

Champarnaud:2012:GRL

[1276] Jean-Marc Champarnaud, Jean-Philippe Dubernard, Franck Guingne, and Hadrien Jeanne. Geometrical regular languages and linear Diophantine equations: the strongly connected case. Theoretical Computer Science, 449(1): 54–63, August 31, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003568.

Dassow:2012:ECG

[1277] Jürgen Dassow, Florin Manea, and Bianca Truthe. On external contextual grammars with subregular selection languages. Theoretical Computer Science, 449(1):64– 73, August 31, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512003520.

Ibarra:2012:SMT

[1278] Oscar H. Ibarra and Nicholas Q. Tran. On synchronized multi-tape and multi-head automata. Theoretical Computer Science, 449(1): 74–84, August 31, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512003507.

Jiraskova:2012:RBR

[1279] Galina Jirásková and Juraj Sebej. Reversal of binary regular languages. Theoretical Computer Science, 449(1): 85–92, August 31, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751200429X.

Jiraskova:2012:SPS

[1280] Galina Jirásková and Tomás Masopust. On a structural property in the state complexity of projected regular languages. Theoretical Computer Science, 449(1):93– 105, August 31, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512003532.

Kunc:2012:SCO

[1281] Michal Kunc and Alexander Okhotin.

State complexity of operations on two-way finite automata over a unary alphabet. Theoretical Computer Science, 449(1):106–118, August 31, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003544.

Malcher:2012:DCT

[1282] Andreas Malcher, Carlo Mereghetti, and Beatrice Palano. Descriptional complexity of two-way pushdown automata with restricted head reversals. Theoretical Computer Science, 449(1): 119–133, August 31, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003519.

Turaev:2012:LCG

[1283] Sherzod Turaev, Jürgen Dassow, Florin Manea, and Mohd Hasan Selamat. Language classes generated by tree controlled grammars with bounded nonterminal complexity. Theoretical Computer Science, 449(1):134– 144, August 31, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200357X.

Anonymous:2012:EBah

[1284] Anonymous. Editorial Board. Theoretical Computer Science, 449(1): iii–viii, August 31, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512006123.

Bouchou-Markhoff:2012:OSY

[1285] Béatrice Bouchou-Markhoff, Pascal Caron, Jean-Marc Champarnaud, and Denis Maurel. Obituary for Sheng Yu. Theoretical Computer Science, 450 (1):1, September 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512003970.

Bouchou-Markhoff:2012:P

[1286] B. Bouchou-Markhoff, P. Caron, J.-M. Champarnaud, and D. Maurel. Preface. Theoretical Computer Science, 450 (1):2, September 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512003763.

Martyugin:2012:CPC

[1287] Pavel Martyugin. Complexity of problems concerning reset words for cyclic and Eulerian automata. Theoretical Computer Science, 450(1): 3–9, September 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512003775.

Daciuk:2012:SRF

[1288] Jan Daciuk and Dawid Weiss. Smaller representation of finite state automata. Theoretical Computer Science, 450(1): 10–21, September 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003787.

Bechet:2012:CGI

[1289] Denis Béchet, Alexandre Dikovsky, and Annie Foret. Categorial grammars with iterated types form a strict hierarchy of k-valued languages. Theoretical Computer Science, 450(1):22– 30, September 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003799.

Carnino:2012:SDK

[1290] Vincent Carnino and Sven De Felice.

Sampling different kinds of acyclic automata using Markov chains. Theoretical Computer Science, 450(1):31– 42, September 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003805.

Heam:2012:LOT

[1291] Pierre-Cyrille Héam, Vincent Hugot, and Olga Kouchnarenko. Loops and overloops for Tree-Walking Automata. Theoretical Computer Science, 450(1): 43–53, September 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003817.

Han:2012:ACA

[1292] Yo-Sub Han and Sang-Ki Ko. Analysis of a cellular automaton model for car traffic with a junction. Theoretical Computer Science, 450(1):54– 67, September 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512003829.

Holzer:2012:NSC

[1293] Markus Holzer, Martin Kutrib, and Katja Meckel. Nondeterministic state complexity of star-free languages. Theoretical Computer Science, 450(1):68– 80, September 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512003830.

Castiglione:2012:NMA

[1294] G. Castiglione, A. Restivo, and M. Sciortino. Nondeterministic Moore automata and Brzozowski’s minimization algorithm. Theoretical Computer Science, 450(1):81– 91, September 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512003842.

Castano:2012:FSI

[1295] José M. Castaño and Rodrigo Castaño. A finite state intersection approach to propositional satisfiability. Theoretical Computer Science, 450(1):92– 108, September 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512003854.

Chairungsee:2012:UMA

[1296] Supaporn Chairungsee and Maxime Crochemore. Using minimal absent words to build phylogeny. Theoretical Computer Science, 450(1):109– 116, September 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003866.

Anonymous:2012:EBai

[1297] Anonymous. Editorial Board. Theoretical Computer Science, 450(1):iii– viii, September 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006287.

Meseguer:2012:FCP

[1298] José Meseguer and Peter Csaba Ölveczky. Formalization and correctness of the PALS architectural pattern for distributed real-time systems. Theoretical Computer Science, 451(1): 1–37, September 14, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005348.

Dowek:2012:PHL

[1299] Gilles Dowek and Murdoch J. Gabbay. PNL to HOL: From the logic of nominal sets to the logic of higher-order functions. Theoretical Computer Science, 451(1):38– 69, September 14, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005439.

Hierons:2012:CAM

[1300] Robert M. Hierons. The complexity of asynchronous model based testing. Theoretical Computer Science, 451(1): 70–82, September 14, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512005324.

Anonymous:2012:EBaj

[1301] Anonymous. Editorial Board. Theoretical Computer Science, 451(1):iii– viii, September 14, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512006615.

Alon:2012:BI

[1302] Noga Alon, Yuval Emek, Michal Feldman, and Moshe Tennenholtz. Bayesian ignorance. Theoretical Computer Science, 452(1):1–11, September 21, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004616.

Castel-Taleb:2012:AAB

[1303] Hind Castel-Taleb, Lynda Mokdad, and Nihal Pekergin. An algorithm approach to bounding aggregations of multidimensional Markov chains. Theoretical Computer Science, 452(1):12– 20, September 21, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512005129.

Florencio:2012:FCG

[1304] Christophe Costa Florêncio and Henning Fernau. On families of categorial grammars of bounded value, their learnability and related complexity questions. Theoretical Computer Science, 452(1):21–38, September 21, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004604.

Damaschke:2012:PRA

[1305] Peter Damaschke and Leonid Molokov.

Parameterized reductions and algorithms for a graph editing problem that generalizes vertex cover. Theoretical Computer Science, 452(1):39– 46, September 21, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005130.

Day:2012:PTT

[1306] Adam R. Day. Process and truthtable characterisations of randomness. Theoretical Computer Science, 452(1): 47–55, September 21, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005105.

Emek:2012:COC

[1307] Yuval Emek and Michal Feldman.

Computing optimal contracts in combinatorial agencies. Theoretical Computer Science, 452(1):56– 74, September 21, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004628.

Fleischer:2012:AAH

[1308] Rudolf Fleischer and Gerhard J. Woeginger. An algorithmic analysis of the Honey-Bee game. Theoretical Computer Science, 452(1):75– 87, September 21, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512005142.

Hajiaghayi:2012:CP

[1309] MohammadTaghi Hajiaghayi, Rohit Khandekar, Guy Kortsarz, and Julián Mestre. The checkpoint problem. Theoretical Computer Science, 452(1):88– 99, September 21, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004793.

Ilic:2012:IBW

[1310] Aleksandar Ilić, Sandi Klavzar, and Yoomi Rho. The index of a binary word. Theoretical Computer Science, 452(1):100–106, September 21, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004835.

Kesselman:2012:GTA

[1311] Alex Kesselman and Stefano Leonardi. Game-theoretic analysis of Internet switching with selfish users. Theoretical Computer Science, 452(1):107– 116, September 21, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005117.

Nikolopoulos:2012:ITC

[1312] Stavros D. Nikolopoulos and Leonidas Palios. An O(nm)-time certifying algorithm for recognizing HHD-free graphs. Theoretical Computer Science, 452(1): 117–131, September 21, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005427.

Stewart:2012:GTE

[1313] Iain A. Stewart. A general technique to establish the asymptotic conditional diagnosability of interconnection networks. Theoretical Computer Science, 452(1):132–147, September 21, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004598.

Anonymous:2012:EBak

[1314] Anonymous. Editorial Board. Theoretical Computer Science, 452(1):iii– viii, September 21, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006755.

Dolev:2012:EAA

[1315] Shlomi Dolev and Christian Scheideler. Editorial for algorithmic aspects of wireless sensor networks. Theoretical Computer Science, 453(1): 1, September 28, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006846.

Avin:2012:NUP

[1316] Chen Avin, Zvi Lotker, Francesco Pasquale, and Yvonne-Anne Pignolet. A note on uniform power connectivity in the physical signal to interference plus noise (SINR) model. Theoretical Computer Science, 453(1):2– 13, September 28, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511010425.

Bradonjic:2012:NOR

[1317] Milan Bradonjić, Eddie Kohler, and Rafail Ostrovsky. Near-optimal radio use for wireless network synchronization. Theoretical Computer Science, 453(1):14–28, September 28, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751100819X.

Holzer:2012:MCW

[1318] Stephan Holzer, Yvonne Anne Pignolet, Jasmin Smula, and Roger Wattenhofer. Monitoring churn in wireless networks. Theoretical Computer Science, 453(1):29– 43, September 28, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511006888.

Huc:2012:VRA

[1319] Florian Huc, Aubin Jarry, Pierre Leone, and José Rolim. Virtual raw anchor coordinates: a new localization paradigm. Theoretical Computer Science, 453(1):44– 53, September 28, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751100942X.

Kanj:2012:ILA

[1320] Iyad A. Kanj and Ge Xia. Improved local algorithms for spanner construction. Theoretical Computer Science, 453(1):54–64, September 28, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511004361.

Klonowski:2012:EEA

[1321] M. Klonowski, M. Kutylowski, and J. Zatopiański. Energy efficient alert in single-hop networks of extremely weak devices. Theoretical Computer Science, 453(1):65– 74, September 28, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001016.

Cichon:2012:KPK

[1322] Jacek Cichoń, Zbigniew Golebiewski, and Miroslaw Kutylowski. From key predistribution to key redistribution. Theoretical Computer Science, 453(1): 75–87, September 28, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512005373.

Nutov:2012:IAA

[1323] Zeev Nutov and Michael Segal. Improved approximation algorithms for maximum lifetime problems in wireless networks. Theoretical Computer Science, 453(1):88–97, September 28, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511006621.

Wendeberg:2012:SLB

[1324] Johannes Wendeberg, Thomas Janson, and Christian Schindelhauer. Selflocalization based on Ambient Signals. Theoretical Computer Science, 453(1): 98–109, September 28, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511009431.

Anonymous:2012:EBal

[1325] Anonymous. Editorial Board. Theoretical Computer Science, 453(1):iii– viii, September 28, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512006895.

Rozenberg:2012:Pe

[1326] G. Rozenberg.  Preface.  Theo-

retical Computer Science, 454(1):1, October 5, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006986.

Rozenberg:2012:Pf

[1327] Grzegorz Rozenberg and Arto Salomaa. Preface. Theoretical Computer Science, 454(1):2–4, October 5, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006974.

Angeleska:2012:RRC

[1328] Angela Angeleska, Natasa Jonoska, and Masahico Saito. Rewriting rule chains modeling DNA rearrangement pathways. Theoretical Computer Science, 454(1):5–22, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004069.

Apostolico:2012:PSM

[1329] Alberto Apostolico, Péter L. Erdős, and Alpár Jüttner. Parameterized searching with mismatches for run-length encoded strings. Theoretical Computer Science, 454(1): 23–29, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512002381.

Azimi:2012:SGA

[1330] Sepinoud Azimi, Tero Harju, Miika Langille, and Ion Petre. Simple gene assembly as a rewriting of directed overlap-inclusion graphs. Theoretical Computer Science, 454(1): 30–37, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512003623.

Biegler:2012:ADS

[1331] Franziska Biegler, Mark Daley, and Ian McQuillan. Algorithmic decomposition of shuffle on words. Theoretical Computer Science, 454(1): 38–50, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512003234.

Bonizzoni:2012:BPP

[1332] Paola Bonizzoni, Chiara Braghin, Riccardo Dondi, and Gabriella Trucco. The binary perfect phylogeny with persistent characters. Theoretical Computer Science, 454(1):51–63, October 5, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512005294.

Brijder:2012:PDP

[1333] Robert Brijder, Tero Harju, and Hendrik Jan Hoogeboom. Pivots, determinants, and perfect matchings of graphs. Theoretical Computer Science, 454(1): 64–71, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001831.

Calude:2012:CEI

[1334] Cristian S. Calude, Elena Calude, and Melissa S. Queen. The complexity of Euler’s integer partition theorem.

Theoretical Computer Science, 454(1): 72–80, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512002563.

Chiniforooshan:2012:ORC

[1335] Ehsan Chiniforooshan, Mark Daley, Oscar H. Ibarra, Lila Kari, and Shinnosuke Seki. One-reversal counter machines and multihead automata: Revisited. Theoretical Computer Science, 454(1):81–87, October 5, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003246.

Choffrut:2012:RRH

[1336] Christian Choffrut and Serge Grigorieff. Rational relations having a rational trace on each finite intersection of rational relations. Theoretical Computer Science, 454(1): 88–94, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512002319.

Corolli:2012:ERS

[1337] Luca Corolli, Carlo Maj, Fabrizio Marini, Daniela Besozzi, and Giancarlo Mauri. An excursion in reaction systems: From computer science to biology. Theoretical Computer Science, 454 (1):95–108, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003258.

deLuca:2012:GPM

[1338] Aldo de Luca and Alessandro De Luca. A generalized palindromization map in free monoids. Theoretical Computer Science, 454(1): 109–128, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512000655.

Diekert:2012:SFL

[1339] Volker Diekert, Manfred Kufleitner, and Pascal Weil. Star-free languages are Church–Rosser congruential. Theoretical Computer Science, 454(1): 129–135, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512000643.

Gabow:2012:CID

[1340] Harold N. Gabow. A combinatoric interpretation of dual variables for weighted matching and f-factors. Theoretical Computer Science, 454(1): 136–163, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512006068.

Gao:2012:SCUb

[1341] Yuan Gao, Lila Kari, and Sheng Yu. State complexity of union and intersection of square and reversal on k regular languages. Theoretical Computer Science, 454(1):164– 171, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003180.

Huova:2012:PBW

[1342] Mari Huova, Juhani Karhumäki, and Aleksi Saarela. Problems in between words and abelian words: k-abelian avoidability. Theoretical Computer Science, 454(1):172– 177, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512002307.

Jolivet:2012:CMC

[1343] Timo Jolivet and Jarkko Kari. Consistency of multidimensional combinatorial substitutions. Theoretical Computer Science, 454(1):178– 188, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003222.

Kleijn:2012:RPN

[1344] Jetty Kleijn, Maciej Koutny, and Marta Pietkiewicz-Koutny. Regions of Petri nets with a/sync connections. Theoretical Computer Science, 454(1): 189–198, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200360X.

Mignosi:2012:CSW

[1345] Filippo Mignosi and Antonio Restivo.

Characteristic Sturmian words are extremal for the Critical Factorization Theorem. Theoretical Computer Science, 454(1):199–205, October 5, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512002320.

Okubo:2012:PLC

[1346] Fumiya Okubo, Satoshi Kobayashi, and Takashi Yokomori. On the properties of language classes defined by bounded reaction automata. Theoretical Computer Science, 454(1): 206–221, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512002575.

Paun:2012:AR

[1347] Gheorghe Paun and Mario J. Pérez-Jiménez. P automata revisited. Theoretical Computer Science, 454(1): 222–230, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512000837.

Piao:2012:LBS

[1348] Xiaoxue Piao and Kai Salomaa. Lower bounds for the size of deterministic unranked tree automata. Theoretical Computer Science, 454(1): 231–239, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512003155.

Skowron:2012:IIS

[1349] Andrzej Skowron and Piotr Wasilewski. Interactive information systems: Toward perception based computing.

Theoretical Computer Science, 454(1): 240–260, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003635.

Anonymous:2012:EBam

[1350] Anonymous. Editorial Board. Theoretical Computer Science, 454(1): iii–viii, October 5, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007037.

Cavalcanti:2012:SII

[1351] Ana Cavalcanti and David Déharbe.

Special issue: International Colloquium on Theoretical Aspects of Computing — ICTAC 2010. Theoretical Computer Science, 455(1): 1, October 12, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007219.

Borba:2012:TSP

[1352] Paulo Borba, Leopoldo Teixeira, and Rohit Gheyi. A theory of software product line refinement. Theoretical Computer Science, 455(1): 2–30, October 12, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000679.

Chen:2012:FDS

[1353] Zhenbang Chen, Zhiming Liu, and Ji Wang. Failure-divergence semantics and refinement of long running transactions. Theoretical Computer Science, 455(1):31–65, October 12, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004057.

Aiguier:2012:FAF

[1354] Marc Aiguier, Frédéric Boulanger, and Bilal Kanso. A formal abstract framework for modelling and testing complex software systems. Theoretical Computer Science, 455(1): 66–97, October 12, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010486.

Castro:2012:EDB

[1355] Pablo F. Castro and Thomas S. E. Maibaum. Encapsulating deontic and branching time specifications. Theoretical Computer Science, 455(1):98– 122, October 12, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511009820.

Anonymous:2012:EBan

[1356] Anonymous. Editorial Board. Theoretical Computer Science, 455(1):iii– viii, October 12, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007268.

Rozenberg:2012:Pg

[1357] G. Rozenberg.  Preface.  Theo-

retical Computer Science, 456(1):1, October 19, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007517.

Berghammer:2012:CSB

[1358] Rudolf Berghammer, Tobias Friedrich, and Frank Neumann. Convergence of set-based multi-objective optimization, indicators and deteriorative cycles. Theoretical Computer Science, 456(1): 2–17, October 19, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005300.

Campos:2012:ANG

[1359] Marcelino Campos and José M. Sempere. Accepting Networks of Genetic Processors are computationally complete. Theoretical Computer Science, 456(1):18–29, October 19, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006238.

Donaldson:2012:MMS

[1360] Robin Donaldson and Muffy Calder.

Modular modelling of signalling pathways and their cross-talk. Theoretical Computer Science, 456(1):30– 50, October 19, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006482.

Iwama:2012:QCC

[1361] Kazuo Iwama, Harumichi Nishimura, Rudy Raymond, and Junichi Teruyama. Quantum counterfeit coin problems. Theoretical Computer Science, 456(1): 51–64, October 19, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512005336.

Manea:2012:CRD

[1362] Florin Manea. Complexity results for deciding Networks of Evolutionary Processors. Theoretical Computer Science, 456(1):65–79, October 19, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512006391.

Petre:2012:MID

[1363] Ion Petre and Sergey Verlan. Matrix insertion-deletion systems. Theoretical Computer Science, 456(1): 80–88, October 19, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512006470.

Sun:2012:GSR

[1364] Zhengya Sun, Wei Jin, and Jue Wang. Generic subset ranking using binary classifiers. Theoretical Computer Science, 456(1):89– 99, October 19, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004847.

Xing:2012:ACF

[1365] Hongyan Xing, Qiansheng Zhang, and Kaisheng Huang. Analysis and control of fuzzy discrete event systems using bisimulation equivalence. Theoretical Computer Science, 456(1):100– 111, October 19, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005312.

Anonymous:2012:EBao

[1366] Anonymous. Editorial Board. Theoretical Computer Science, 456(1):iii– viii, October 19, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007566.

Acuna:2012:TSE

[1367] Vicente Acuña, Etienne Birmelé, Ludovic Cottret, Pierluigi Crescenzi, Fabien Jourdan, Vincent Lacroix, Alberto Marchetti-Spaccamela, Andrea Marino, Paulo Vieira Milreu, Marie-France Sagot, and Leen Stougie. Telling stories: Enumerating maximal directed acyclic graphs with a constrained set of sources and targets. Theoretical Computer Science, 457(1):1–9, October 26, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007207.

Bermond:2012:OGR

[1368] Jean-Claude Bermond and Joseph G.

Peters. Optimal gathering in radio grids with interference. Theoretical Computer Science, 457(1):

10–26, October 26, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007165.

Breslauer:2012:SES

[1369] Dany Breslauer and Giuseppe F. Italiano. On suffix extensions in suffix trees. Theoretical Computer Science, 457(1):27–34, October 26, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751200713X.

Cao:2012:SSO

[1370] Qian Cao, T. C. E. Cheng, Guohua Wan, and Yi Li. Several semi-online scheduling problems on two identical machines with combined information. Theoretical Computer Science, 457(1): 35–44, October 26, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512006500.

Comi:2012:CPC

[1371] Marco Comi, Bhaskar DasGupta, Michael Schapira, and Venkatakumar Srinivasan. On communication protocols that compute almost privately. Theoretical Computer Science, 457(1): 45–58, October 26, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512006676.

Elor:2012:SAM

[1372] Yotam Elor and Alfred M. Bruckstein. A ‘thermodynamic’ approach to multi-robot cooperative localization. Theoretical Computer Science, 457(1): 59–75, October 26, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200655X.

Elor:2012:TRS

[1373] Yotam Elor and Alfred M. Bruckstein. Two-robot source seeking with point measurements. Theoretical Computer Science, 457(1):76– 85, October 26, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006536.

Golovach:2012:CVS

[1374] Petr A. Golovach, Daniël Paulusma, and Jian Song. Computing vertexsurjective homomorphisms to partially reflexive trees. Theoretical Computer Science, 457(1):86– 100, October 26, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006561.

Graca:2012:CBC

[1375] Daniel S. Graça, Ning Zhong, and H. S. Dumas. The connection between computability of a nonlinear problem and its linearization: the Hartman–Grobman theorem revisited. Theoretical Computer Science, 457(1): 101–110, October 26, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512006962.

Heinz:2012:LLL

[1376] Jeffrey Heinz, Anna Kasprzik, and Timo Kötzing. Learning in the limit with lattice-structured hypothesis spaces. Theoretical Computer Science, 457(1):111–127, October 26, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007128.

Higgins:2012:BWT

[1377] Peter M. Higgins. Burrows–Wheeler transformations and de Bruijn words. Theoretical Computer Science, 457(1): 128–136, October 26, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007141.

Jacob:2012:THD

[1378] Riko Jacob, Stephan Ritscher, Christian Scheideler, and Stefan Schmid. Towards higher-dimensional topological self-stabilization: a distributed algorithm for Delaunay graphs. Theoretical Computer Science, 457(1):137– 148, October 26, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007372.

Okhotin:2012:PBG

[1379] Alexander Okhotin and Christian Reitwießner. Parsing Boolean grammars over a one-letter alphabet using online convolution. Theoretical Computer Science, 457(1):149– 157, October 26, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006421.

Wang:2012:SRP

[1380] Weifan Wang, Jiangxu Kong, and Lianzhu Zhang. The 2-surviving rate of planar graphs without 4-cycles. Theoretical Computer Science, 457(1):158– 165, October 26, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006706.

Xue:2012:MPC

[1381] Yu-Mei Xue and Teturo Kamae. Maximal pattern complexity, dual system and pattern recognition. Theoretical Computer Science, 457(1):166– 173, October 26, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006433.

Yuan:2012:BPO

[1382] Junling Yuan, Wenhua Li, and Jinjiang Yuan. A best possible online algorithm for scheduling equallength jobs on two machines with chain precedence constraints. Theoretical Computer Science, 457(1):174– 180, October 26, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006688.

Anonymous:2012:EBap

[1383] Anonymous. Editorial Board. Theoretical Computer Science, 457(1):iii– viii, October 26, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007761.

Aceto:2012:RFD

[1384] Luca Aceto, Matteo Cimini, Anna Ingolfsdottir, MohammadReza Mousavi, and Michel A. Reniers. Rule formats for distributivity. Theoretical Computer Science, 458(1):1–28, November 2, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007451.

Bezem:2012:TSC

[1385] Marc Bezem, Dag Hovland, and Hoang Truong. A type system for counting instances of software components. Theoretical Computer Science, 458(1): 29–48, November 2, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007414.

Chatterjee:2012:EPG

[1386] Krishnendu Chatterjee and Laurent Doyen. Energy parity games. Theoretical Computer Science, 458(1):49– 60, November 2, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007475.

Choi:2012:SBM

[1387] Young ri Choi and Mohamed G.

Gouda. A state-based model of sensor protocols. Theoretical Computer Science, 458(1):61– 75, November 2, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007864.

Liu:2012:CSB

[1388] Jia Liu and Huimin Lin. A complete symbolic bisimulation for full applied pi calculus. Theoretical Computer Science, 458(1):76– 112, November 2, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007438.

Millo:2012:PSM

[1389] Jean-Vivien Millo and Robert de Simone. Periodic scheduling of marked graphs using balanced binary words. Theoretical Computer Science, 458(1): 113–130, November 2, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007852.

Hesselink:2012:FII

[1390] Wim H. Hesselink and Gerard R. Renardel de Lavalette. Finite and infinite implementation of transition systems. Theoretical Computer Science, 458(1): 131–135, November 2, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007876.

Anonymous:2012:EBaq

[1391] Anonymous. Editorial Board. Theoretical Computer Science, 458(1):iii– viii, November 2, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512008122.

Ananth:2012:CGB

[1392] Prabhanjan V. Ananth and Ambedkar Dukkipati. Complexity of Gröbner basis detection and border basis detection. Theoretical Computer Science, 459(1):1–15, November 9, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007505.

Aubrun:2012:TSF

[1393] Nathalie Aubrun and Marie-Pierre Béal. Tree-shifts of finite type. Theoretical Computer Science, 459(1):16– 25, November 9, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007153.

Barbay:2012:LTC

[1394] Jérémy Barbay, Johannes Fischer, and Gonzalo Navarro. LRM-Trees: Compressed indices, adaptive sorting, and compressed permutations. Theoretical Computer Science, 459(1):26– 41, November 9, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007712.

Bienkowski:2012:RPU

[1395] Marcin Bienkowski and Jaroslaw Kutylowski. The k-resource problem in uniform metric spaces. Theoretical Computer Science, 459(1):42– 54, November 9, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006214.

Bienvenu:2012:SRE

[1396] Laurent Bienvenu and Christopher Porter. Strong reductions in effective randomness. Theoretical Computer Science, 459(1):55–68, November 9, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S030439751200641X.

Bouda:2012:IHE

[1397] Jan Bouda, Matej Pivoluska, and Martin Plesch. Improving the Hadamard extractor. Theoretical Computer Science, 459(1):69– 76, November 9, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007384.

Bourgeois:2012:ADC

[1398] N. Bourgeois, F. Della Croce, B. Escoffier, and V. Th. Paschos. Algorithms for dominating clique problems. Theoretical Computer Science, 459(1): 77–88, November 9, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007116.

Chang:2012:CDP

[1399] Gerard J. Chang, B. S. Panda, and D. Pradhan. Complexity of distance paired-domination problem in graphs. Theoretical Computer Science, 459(1): 89–99, November 9, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007980.

Cicalese:2012:BIP

[1400] Ferdinando Cicalese, Tobias Jacobs, Eduardo Laber, and Caio Valentim. The binary identification problem for weighted trees. Theoretical Computer Science, 459(1):100– 112, November 9, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512006056.

Cincotti:2012:LSP

[1401] Alessandro Cincotti. The lattice structure of n-player games. Theoretical Computer Science, 459(1):113– 119, November 9, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512005865.

Anonymous:2012:EBar

[1402] Anonymous. Editorial Board. Theoretical Computer Science, 459(1):iii– viii, November 9, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512008341.

Downey:2012:LBR

[1403] Rod Downey and Keng Meng Ng. Lowness for bounded randomness. Theoretical Computer Science, 460(1):1– 9, November 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005403.

Wang:2012:CPM

[1404] Zhenbo Wang and Zhenhua Cui. Combination of parallel machine scheduling and vertex cover. Theoretical Computer Science, 460(1):10– 15, November 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005397.

Dourado:2012:RIG

[1405] Mitre C. Dourado, Lucia Draque Penso, Dieter Rautenbach, and Jayme L. Szwarcfiter. Reversible iterative graph processes. Theoretical Computer Science, 460(1):16– 25, November 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005798.

Lozin:2012:MRI

[1406] Vadim V. Lozin and Raffaele Mosca.

Maximum regular induced subgraphs in 2P_{3}-free graphs. Theoretical Computer Science, 460(1):26–33, November 16, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200583X.

Wang:2012:FTA

[1407] Shiying Wang, Guozhen Zhang, and Kai Feng. Fault tolerance in kary n-cube networks. Theoretical Computer Science, 460(1):34– 41, November 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512005828.

Shachnai:2012:MCR

[1408] Hadas Shachnai, Gal Tamir, and Tami Tamir. Minimal cost reconfiguration of data placement in a storage area network. Theoretical Computer Science, 460(1):42– 53, November 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512006007.

Wu:2012:CRE

[1409] Yongcheng Wu. Computability on random events and variables in a computable probability space. Theoretical Computer Science, 460(1):54– 69, November 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512006226.

Elbassioni:2012:CHP

[1410] Khaled Elbassioni, Rajiv Raman, Saurabh Ray, and René Sitters. On the complexity of the highway problem. Theoretical Computer Science, 460(1): 70–77, November 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007360.

Lee:2012:DTS

[1411] Jia Lee, Rui-Long Yang, and Kenichi Morita. Design of 1-tape 2-symbol reversible Turing machines based on reversible logic elements. Theoretical Computer Science, 460(1):78– 88, November 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007359.

Nedev:2012:LBB

[1412] Zhivko Nedev. Lower bound for balanced sets. Theoretical Computer Science, 460(1):89–93, November 16, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005853.

Rombo:2012:ESM

[1413] Simona E. Rombo. Extracting string motif bases for quorum higher than two. Theoretical Computer Science, 460(1):94–103, November 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397512006032.

Anonymous:2012:EBas

[1414] Anonymous. Editorial Board. Theoretical Computer Science, 460(1):iii– viii, November 16, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008481.

Du:2012:PC

[1415] Ding-Zhu Du and Bin Fu. Preface — COCOON’2011. Theoretical Computer Science, 461(1): 1, November 23, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008778.

Cai:2012:SSR

[1416] Jin-Yi Cai and Michael Kowalczyk. Spin systems on k-regular graphs with complex edge functions. Theoretical Computer Science, 461(1):2– 16, November 23, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000576.

DeMarco:2012:CMT

[1417] Gianluca De Marco, Evangelos Kranakis, and Gábor Wiener. Computing majority with triple queries. Theoretical Computer Science, 461(1):17– 26, November 23, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003118.

Jiang:2012:PCM

[1418] Minghui Jiang and Yong Zhang. Parameterized complexity in multipleinterval graphs: Domination, partition, separation, irredundancy. Theoretical Computer Science, 461(1):27– 44, November 23, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000618.

Liu:2012:CPA

[1419] Yunlong Liu, Jianxin Wang, Jiong Guo, and Jianer Chen. Complexity and parameterized algorithms for Cograph Editing. Theoretical Computer Science, 461(1):45– 54, November 23, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009595.

Messner:2012:KCP

[1420] Jochen Messner and Thomas Thierauf.

A Kolmogorov complexity proof of the Lovász Local Lemma for satisfiability. Theoretical Computer Science, 461(1): 55–64, November 23, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005415.

Misra:2012:PIF

[1421] Neeldhara Misra, Geevarghese Philip, Venkatesh Raman, and Saket Saurabh. On parameterized independent feedback vertex set. Theoretical Computer Science, 461(1):65–75, November 23, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001417.

Tian:2012:EAA

[1422] Cong Tian, Zhenhua Duan, and Nan Zhang. An efficient approach for abstraction-refinement in model checking. Theoretical Computer Science, 461(1):76–85, November 23, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511009741.

Yamakami:2012:ACC

[1423] Tomoyuki Yamakami. Approximation complexity of complex-weighted degree-two counting constraint satisfaction problems. Theoretical Computer Science, 461(1):86–105, November 23, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751100956X.

Zeng:2012:ECA

[1424] Shengke Zeng, Shaoquan Jiang, and Zhiguang Qin. An efficient conditionally anonymous ring signature in the random oracle model. Theoretical Computer Science, 461(1):106– 114, November 23, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512000631.

Anonymous:2012:EBat

[1425] Anonymous. Editorial Board. Theoretical Computer Science, 461(1):iii– viii, November 23, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512008870.

Didier:2012:VLL

[1426] Gilles Didier, Eduardo Corel, Ivan Laprevotte, Alex Grossmann, and Claudine Landès-Devauchelle. Variable length local decoding and alignmentfree sequence comparison. Theoretical Computer Science, 462(1):1– 11, November 30, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007669.

Edmonds:2012:CAT

[1427] Jeff Edmonds. On the competitiveness of AIMD–TCP within a general network. Theoretical Computer Science, 462(1):12–22, November 30, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007426.

Epstein:2012:MCP

[1428] Leah Epstein and Asaf Levin. On the max coloring problem. Theoretical Computer Science, 462(1):23– 38, November 30, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007463.

Fang:2012:CCS

[1429] Liming Fang, Willy Susilo, Chunpeng Ge, and Jiandong Wang. Chosenciphertext secure anonymous conditional proxy re-encryption with keyword search. Theoretical Computer Science, 462(1):39–58, November 30, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007906.

Jerabek:2012:RFT

[1430] Emil Jerábek. Root finding with threshold circuits. Theoretical Computer Science, 462(1):59–69, November 30, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512008006.

Lemoine:2012:NI

[1431] Julien Lemoine and Simon Viennot. Nimbers are inevitable. Theoretical Computer Science, 462(1):70– 79, November 30, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512008018.

Miao:2012:SDJ

[1432] Cuixia Miao, Yuzhong Zhang, and Cuilian Wu. Scheduling of deteriorating jobs with release dates to minimize the maximum lateness. Theoretical Computer Science, 462(1):80– 87, November 30, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007967.

Michael:2012:EFT

[1433] Loizos Michael. Evolvability via the Fourier transform. Theoretical Computer Science, 462(1):88– 98, November 30, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007700.

Anonymous:2012:EBau

[1434] Anonymous. Editorial Board. Theoretical Computer Science, 462(1):iii– viii, November 30, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009012.

Fomin:2012:FSI

[1435] Fedor V. Fomin, Pierre Fraigniaud, Stephan Kreutzer, and Dimitrios M. Thilikos. Foreword: Special Issue on Theory and Applications of Graph Searching Problems. Theoretical Computer Science, 463(1): 1, December 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009139.

Berwanger:2012:ECD

[1436] Dietmar Berwanger, Erich Grädel, Lukasz Kaiser, and Roman Rabinovich. Entanglement and the complexity of directed graphs. Theoretical Computer Science, 463(1):2– 25, December 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200669X.

Safari:2012:HWH

[1437] MohammadAli Safari. Hyper-T-width and hyper-D-width: Stable connectivity measures for hypergraphs. Theoretical Computer Science, 463(1):26– 34, December 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512008602.

Butterfield:2012:RSS

[1438] Jane V. Butterfield, Daniel W. Cranston, Gregory J. Puleo, Douglas B. West, and Reza Zamani. Revolutionaries and spies: Spygood and spy-bad graphs. Theoretical Computer Science, 463(1):35– 53, December 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751200535X.

Carraher:2012:LRG

[1439] James Carraher, Ilkyoo Choi, Michelle Delcourt, Lawrence H. Erickson, and Douglas B. West. Locating a robber on a graph via distance queries. Theoretical Computer Science, 463(1): 54–61, December 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512006469.

Megow:2012:OGE

[1440] Nicole Megow, Kurt Mehlhorn, and Pascal Schweitzer. Online graph exploration: New results on old and new algorithms. Theoretical Computer Science, 463(1):62–72, December 7, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512006445.

Meister:2012:CD

[1441] Daniel Meister and Jan Arne Telle. Chordal digraphs. Theoretical Computer Science, 463(1):73–83, December 7, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006019.

Adler:2012:HDM

[1442] Isolde Adler, Tomás Gavenciak, and Tereza Klimosová. Hypertree-depth and minors in hypergraphs. Theoretical Computer Science, 463(1):84– 95, December 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008274.

Dereniowski:2012:ASS

[1443] Dariusz Dereniowski. Approximate search strategies for weighted trees. Theoretical Computer Science, 463(1): 96–113, December 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006512.

Alvarez:2012:CMD

[1444] C. Àlvarez, J. Díaz, D. Mitsche, and M. Serna. Continuous monitoring in the dynamic sensor field model. Theoretical Computer Science, 463(1):114– 122, December 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004446.

Gordinowicz:2012:PLU

[1445] Przemyslaw Gordinowicz, Richard J.

Nowakowski, and Pawel Pralat. POL-ISH — let us play the cleaning game.

Theoretical Computer Science, 463(1): 123–132, December 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004586.

Kehagias:2012:SRC

[1446] Athanasios Kehagias and Pawel Pralat. Some remarks on cops and drunk robbers. Theoretical Computer Science, 463(1):133–147, December 7, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751200789X.

Anonymous:2012:EBav

[1447] Anonymous. Editorial Board. Theoretical Computer Science, 463(1):iii– viii, December 7, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009188.

Aoto:2012:P

[1448] Takahito Aoto and Aart Middeldorp. Preface. Theoretical Computer Science, 464(1):1–2, December 14, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512008298.

Gramlich:2012:MTR

[1449] Bernhard Gramlich. Modularity in term rewriting revisited. Theoretical Computer Science, 464(1):3– 19, December 14, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008286.

Jouannaud:2012:DCM

[1450] Jean-Pierre Jouannaud and Jiaxiang Liu. From diagrammatic confluence to modularity. Theoretical Computer Science, 464(1):20– 34, December 14, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008067.

Dershowitz:2012:JEM

[1451] Nachum Dershowitz. Jumping and escaping: Modular termination and the abstract path ordering. Theoretical Computer Science, 464(1):35– 47, December 14, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008638.

Endrullis:2012:HIR

[1452] Jörg Endrullis, Dimitri Hendriks, and Jan Willem Klop. Highlights in infinitary rewriting and lambda calculus. Theoretical Computer Science, 464(1): 48–71, December 14, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200792X.

Nagashima:2012:DCT

[1453] Masanori Nagashima, Masahiko Sakai, and Toshiki Sakabe. Determinization of conditional term rewriting systems. Theoretical Computer Science, 464(1): 72–89, December 14, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512008250.

Futatsugi:2012:PPS

[1454] Kokichi Futatsugi, Daniel Gaina, and Kazuhiro Ogata. Principles of proof scores in CafeOBJ. Theoretical Computer Science, 464(1):90– 112, December 14, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007918.

Dezani-Ciancaglini:2012:TWW

[1455] Mariangiola Dezani-Ciancaglini, Ross Horne, and Vladimiro Sassone. Tracing where and who provenance in Linked Data: a calculus. Theoretical Computer Science, 464(1):113– 129, December 14, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512006020.

Anonymous:2012:EBaw

[1456] Anonymous. Editorial Board. Theoretical Computer Science, 464(1):iii– viii, December 14, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009620.

Jacobs:2012:AAT

[1457] Tobias Jacobs. Analytical aspects of tie breaking. Theoretical Computer Science, 465(??):1– 9, December 21, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008808.

Lu:2012:MPB

[1458] Huazhong Lü, Xianyue Li, and Heping Zhang. Matching preclusion for balanced hypercubes. Theoretical Computer Science, 465(??):10– 20, December 21, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008705.

Nakprasit:2012:ECP

[1459] Keaitsuda Nakprasit and Kittikorn Nakprasit. Equitable colorings of planar graphs without short cycles. Theoretical Computer Science, 465(??): 21–27, December 21, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200864X.

Takabatake:2012:NDP

[1460] Toshinori Takabatake and Tomoki Nakamigawa. Node-disjoint paths in a level block of generalized hierarchical completely connected networks. Theoretical Computer Science, 465(??): 28–34, December 21, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008663.

Tholey:2012:LTA

[1461] Torsten Tholey. Linear time algorithms for two disjoint paths problems on directed acyclic graphs. Theoretical Computer Science, 465(??):

35–48, December 21, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008791.

Vakhania:2012:BLC

[1462] Nodari Vakhania. Branch less, cut more and minimize the number of late equal-length jobs on identical machines. Theoretical Computer Science, 465(??):49–60, December 21, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008079.

Wang:2012:ACI

[1463] Yan Wang, Jianxi Fan, Xiaohua Jia, and He Huang. An algorithm to construct independent spanning trees on parity cubes. Theoretical Computer Science, 465(??):61– 72, December 21, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007943.

Balkova:2012:CSB

[1464] L’. Balková, E. Pelantová, and S. Starosta. Corrigendum: “On Brlek–Reutenauer conjecture”. Theoretical Computer Science, 465(??): 73–74, December 21, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008237. See

[1270].

Anonymous:2012:EBax

[1465] Anonymous. Editorial Board. Theoretical Computer Science, 465(??):iii– viii, December 21, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009942.

Rozenberg:2012:Ph

[1466] G. Rozenberg.  Preface.  Theo-

retical Computer Science, 466(??): 1, December 28, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010158.

Chollet:2012:FAM

[1467] Agathe Chollet, Guy Wallet, Laurent Fuchs, Eric Andres, and Gaëlle Largeteau-Skapin. Foundational aspects of multiscale digitization. Theoretical Computer Science, 466(??): 2–19, December 28, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007347.

Ganguly:2012:DDE

[1468] Niloy Ganguly, Saptarshi Ghosh, Tyll Krueger, and Ajitesh Srivastava. Degree distributions of evolving alphabetic bipartite networks and their projections. Theoretical Computer Science, 466(??):20–36, December 28, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007682.

Ganivada:2012:FRG

[1469] Avatharam Ganivada, Shubhra Sankar Ray, and Sankar K. Pal. Fuzzy rough granular self-organizing map and fuzzy rough entropy. Theoretical Computer Science, 466(??):37– 63, December 28, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007955.

Gravner:2012:RPS

[1470] Janko Gravner and David Griffeath. Robust periodic solutions and evolution from seeds in one-dimensional edge cellular automata. Theoretical Computer Science, 466(??):64– 86, December 28, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512008043.

Salomaa:2012:FSG

[1471] Arto Salomaa. Functions and sequences generated by reaction systems. Theoretical Computer Science, 466(??): 87–96, December 28, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007177.

Anonymous:2012:EBay

[1472] Anonymous. Editorial Board. Theoretical Computer Science, 466(??):iii– viii, December 28, 2012. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512010225.

Bova:2013:PSH

[1473] Simone Bova and Franco Montagna.

Polynomial space hardness without disjunction property. Theoretical Computer Science, 467(??): 1–11, January 7, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007992.

Delzanno:2013:CRL

[1474] Giorgio Delzanno and Fernando Rosa-Velardo. On the coverability and reachability languages of monotonic extensions of Petri nets. Theoretical Computer Science, 467(??): 12–29, January 7, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008717.

Pechoux:2013:SSI

[1475] Romain Péchoux. Synthesis of supinterpretations: a survey. Theoretical Computer Science, 467(??): 30–52, January 7, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010043.

Rabe:2013:OTA

[1476] Markus N. Rabe and Sven Schewe.

Optimal time-abstract schedulers for CTMDPs and continuous-time Markov games. Theoretical Computer Science, 467(??):53–67, January 7, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751200881X.

vanderMeyden:2013:IFSa

[1477] Ron van der Meyden and Chenyi Zhang. Information flow in systems with schedulers, Part I: Definitions. Theoretical Computer Science, 467(??): 68–88, January 7, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009796.

Wang:2013:EMC

[1478] Farn Wang. Efficient model-checking of dense-time systems with timeconvexity analysis. Theoretical Computer Science, 467(??):89–108, January 7, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512008699.

Anonymous:2013:EBa

[1479] Anonymous. Editorial Board. Theoretical Computer Science, 467(??): iii–viii, January 7, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751201064X.

Briceno:2013:LAB

[1480] Raimundo Briceño and Ivan Rapaport. Letting Alice and Bob choose which problem to solve: Implications to the study of cellular automata. Theoretical Computer Science, 468(??): 1–11, January 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010419.

Bucci:2013:EST

[1481] Michelangelo Bucci, Alessandro De Luca, and Gabriele Fici. Enumeration and structure of trapezoidal words. Theoretical Computer Science, 468(??): 12–22, January 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010377.

Calamoneri:2013:EPC

[1482] T. Calamoneri, E. Montefusco, R. Petreschi, and B. Sinaimeri. Exploring pairwise compatibility graphs. Theoretical Computer Science, 468(??): 23–36, January 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010456.

Chang:2013:BSD

[1483] Ching-Lueh Chang and Yuh-Dauh Lyuu. Bounding the sizes of dynamic monopolies and convergent sets for threshold-based cascades. Theoretical Computer Science, 468(??): 37–49, January 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010468.

Chen:2013:OFT

[1484] Yebin Chen, Ying Li, and Tao Chen.

Optimal fault-tolerant routing algorithm and fault-tolerant diameter in directed double-loop networks. Theoretical Computer Science, 468(??): 50–58, January 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010389.

David:2013:PEA

[1485] J. David and L. Nourine. A parameterizable enumeration algorithm for sequence mining. Theoretical Computer Science, 468(??):59– 68, January 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010353.

Guillaume:2013:AGU

[1486] Delalleau Guillaume. Ancestors graph and an upper bound for the subword complexity function. Theoretical Computer Science, 468(??):69– 82, January 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010444.

Gao:2013:SCB

[1487] Yihan Gao, Jieming Mao, Xiaoming Sun, and Song Zuo. On the sensitivity complexity of bipartite graph properties. Theoretical Computer Science, 468 (??):83–91, January 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010365.

Lee:2013:FCO

[1488] Taehyung Lee, Joong Chae Na, Heejin Park, Kunsoo Park, and Jeong Seop Sim. Finding consensus and optimal alignment of circular strings. Theoretical Computer Science, 468(??): 92–101, January 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010481.

Mou:2013:DPS

[1489] Chenqi Mou, Dongming Wang, and Xiaoliang Li. Decomposing polynomial sets into simple sets over finite fields: the positive-dimensional case. Theoretical Computer Science, 468(??): 102–113, January 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010390.

Zhao:2013:AST

[1490] Kejun Zhao and Xiwen Lu. Approximation schemes for two-agent scheduling on parallel machines. Theoretical Computer Science, 468(??):114– 121, January 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010031.

Anonymous:2013:EBb

[1491] Anonymous. Editorial Board. Theoretical Computer Science, 468(??): iii–viii, January 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010961.

Au:2013:SIB

[1492] Man Ho Au, Joseph K. Liu, Willy Susilo, and Tsz Hon Yuen. Secure ID-based linkable and revocable-iff-linked ring signature with constantsize construction. Theoretical Computer Science, 469(??):1–14, January 21, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009528.

Brandstadt:2013:CTP

[1493] Andreas Brandstädt, Synara Brito, Sulamita Klein, Loana Tito Nogueira, and Fábio Protti. Cycle transversals in perfect graphs and cographs. Theoretical Computer Science, 469(??): 15–23, January 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009516. See

corrigendum [1678].

Dumrauf:2013:CMC

[1494] Dominic Dumrauf and Burkhard Monien. On the PLS-complexity of maximum constraint assignment. Theoretical Computer Science, 469(??): 24–52, January 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009760.

Flocchini:2013:ETV

[1495] Paola Flocchini, Bernard Mans, and Nicola Santoro. On the exploration of time-varying networks. Theoretical Computer Science, 469(??):53– 68, January 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009504.

Fujioka:2013:MCI

[1496] Kaoru Fujioka. Morphic characterizations with insertion systems controlled by a context of length one. Theoretical Computer Science, 469(??): 69–76, January 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009103.

Turek:2013:ACA

[1497] Ondrej Turek. Abelian complexity and Abelian co-decomposition. Theoretical Computer Science, 469(??): 77–91, January 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009553.

Xiao:2013:CSA

[1498] Mingyu Xiao and Hiroshi Nagamochi.

Confining sets and avoiding bottleneck cases: a simple maximum independent set algorithm in degree-3 graphs. Theoretical Computer Science, 469(??): 92–104, January 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008729.

Xiao:2013:FTW

[1499] Mingyu Xiao, Takuro Fukunaga, and Hiroshi Nagamochi. FPTASs for trimming weighted trees. Theoretical Computer Science, 469(??):105– 118, January 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751200878X.

Yu:2013:CDD

[1500] Xiaojiang Yu. Certain discrete dynamical systems, number systems and related integral self-affine sets. Theoretical Computer Science, 469(??): 119–126, January 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512008675.

Yu:2013:LSC

[1501] Fuxiang Yu and Ker-I Ko. On logarithmic-space computable real numbers. Theoretical Computer Science, 469(??):127–133, January 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397512009115.

Anonymous:2013:EBc

[1502] Anonymous. Editorial Board. Theoretical Computer Science, 469(??): iii–viii, January 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512011176.

Agnarsson:2013:SBA

[1503] Geir Agnarsson, Magnús M. Halldórsson, and Elena Losievskaja. SDP-based algorithms for maximum independent set problems on hypergraphs. Theoretical Computer Science, 470(??): 1–9, January 28, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010559.

Akutsu:2013:APA

[1504] Tatsuya Akutsu, Daiji Fukagawa, Magnús M. Halldórsson, Atsuhiro Takasu, and Keisuke Tanaka. Approximation and parameterized algorithms for common subtrees and edit distance between unordered trees. Theoretical Computer Science, 470(??): 10–22, January 28, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751201047X.

Bar-Noy:2013:PMU

[1505] Amotz Bar-Noy, Panagiotis Cheilaris, Yi Feng, and Mordecai J. Golin. Paging mobile users in cellular networks: Optimality versus complexity and simplicity. Theoretical Computer Science, 470 (??):23–35, January 28, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010729.

Brzozowski:2013:QCI

[1506] Janusz Brzozowski, Galina Jirásková, and Baiyu Li. Quotient complexity of ideal languages. Theoretical Computer Science, 470(??):36– 52, January 28, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009875.

Cao:2013:CCM

[1507] Zhigang Cao and Xiaoguang Yang. Complementary cooperation, minimal winning coalitions, and power indices. Theoretical Computer Science, 470(??): 53–92, January 28, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512010766.

Chan:2013:SWF

[1508] Sze-Hang Chan, Tak-Wah Lam, and Lap-Kei Lee. Scheduling for weighted flow time and energy with rejection penalty. Theoretical Computer Science, 470(??):93–104, January 28, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512010511.

Chen:2013:PTI

[1509] Jianer Chen, Iyad A. Kanj, Jie Meng, Ge Xia, and Fenghui Zhang. Parameterized top-K algorithms. Theoretical Computer Science, 470(??): 105–119, January 28, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751200984X.

Anonymous:2013:EBd

[1510] Anonymous. Editorial Board. Theoretical Computer Science, 470(??): iii–viii, January 28, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512011309.

Tang:2013:MAR

[1511] Jianguo Tang, Kun She, Fan Min, and William Zhu. A matroidal approach to rough set theory. Theoretical Computer Science, 471(??): 1–11, February 3, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010067.

Doerr:2013:MEC

[1512] Benjamin Doerr, Daniel Johannsen, Timo Kötzing, Frank Neumann, and Madeleine Theile. More effective crossover operators for the all-pairs shortest path problem. Theoretical Computer Science, 471(??):12– 26, February 3, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010055.

Gori:2013:APP

[1513] Roberta Gori and Francesca Levi. An analysis for proving probabilistic termination of biological systems. Theoretical Computer Science, 471(??): 27–73, February 3, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751201002X.

Su:2013:NGT

[1514] Guifu Su, Liming Xiong, Yi Sun, and Daobin Li. Nordhaus–Gaddumtype inequality for the hyper-Wiener index of graphs when decomposing into three parts. Theoretical Computer Science, 471(??):74– 83, February 3, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009814.

Doerr:2013:BBC

[1515] Benjamin Doerr, Timo Kötzing, Johannes Lengler, and Carola Winzen. Black-box complexities of combinatorial problems. Theoretical Computer Science, 471(??):84–106, February 3, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009681.

Anonymous:2013:EBe

[1516] Anonymous. Editorial Board. Theoretical Computer Science, 471(??): iii–viii, February 3, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513000212.

Angjeli:2013:LMF

[1517] Ariana Angjeli, Eddie Cheng, and László Lipták. Linearly many faults in dual-cube-like networks. Theoretical Computer Science, 472(??): 1–8, February 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512011024.

Demange:2013:SCP

[1518] Marc Demange and Dominique De Werra. On some coloring problems in grids. Theoretical Computer Science, 472(??):9–27, February 11, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009784.

Dobrev:2013:EUD

[1519] Stefan Dobrev, Paola Flocchini, Rastislav Královic, and Nicola Santoro. Exploring an unknown dangerous graph using tokens. Theoretical Computer Science, 472(??): 28–45, February 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010523.

Dong:2013:HCR

[1520] Qiang Dong, Junlin Zhou, Yan Fu, and Hui Gao. Hamiltonian connectivity of restricted hypercube-like networks under the conditional fault model. Theoretical Computer Science, 472(??): 46–59, February 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010778.

Farach-Colton:2013:OMA

[1521] Martín Farach-Colton, Antonio Fernández Anta, and Miguel A. Mosteiro. Optimal memory-aware Sensor Network Gossiping (or how to break the Broadcast lower bound). Theoretical Computer Science, 472(??):60– 80, February 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010730.

Fellows:2013:CSP

[1522] Michael Fellows, Tobias Friedrich, Danny Hermelin, Nina Narodytska, and Frances Rosamond. Constraint satisfaction problems: Convexity makes AllDifferent constraints tractable. Theoretical Computer Science, 472(??):81–89, February 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397512010870.

Fotakis:2013:WIS

[1523] Dimitris Fotakis and Christos Tzamos. Winner-imposing strategy-proof mechanisms for multiple Facility Location games. Theoretical Computer Science, 472(??):90–103, February 11, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512010791.

Du:2013:CDD

[1524] Ye Du. On the complexity of deciding degeneracy in a bimatrix game with sparse payoff matrix. Theoretical Computer Science, 472(??):104– 109, February 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009851.

Anonymous:2013:EBf

[1525] Anonymous. Editorial Board. Theoretical Computer Science, 472(??): iii–viii, February 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000352.

Hutter:2013:GEF

[1526] Marcus Hutter, Frank Stephan, Vladimir Vovk, and Thomas Zeugmann. Guest Editors’ foreword. Theoretical Computer Science, 473(??): 1–3, February 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009280.

Lever:2013:TPB

[1527] Guy Lever, François Laviolette, and John Shawe-Taylor. Tighter PAC-Bayes bounds through distributiondependent priors. Theoretical Computer Science, 473(??):4–28, February 18, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009346.

Pestov:2013:PLU

[1528] Vladimir Pestov. PAC learnability under non-atomic measures: a problem by Vidyasagar. Theoretical Computer Science, 473(??):29– 45, February 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200936X.

Balle:2013:LPA

[1529] Borja Balle, Jorge Castro, and Ricard Gavaldà. Learning probabilistic automata: a study in state distinguishability. Theoretical Computer Science, 473(??):46– 60, February 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009309.

Koolen:2013:SI

[1530] Wouter M. Koolen and Steven de Rooij. Switching investments. Theoretical Computer Science, 473(??): 61–76, February 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009334.

Antos:2013:TCF

[1531] András Antos, Gábor Bartók, Dávid Pál, and Csaba Szepesvári. Toward a classification of finite partialmonitoring games. Theoretical Computer Science, 473(??):77–99, February 18, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009292.

Case:2013:MLN

[1532] John Case and Timo Kötzing. Memory-limited non-U-shaped learning with solved open problems. Theoretical Computer Science, 473(??): 100–123, February 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009310.

Jain:2013:LCa

[1533] Sanjay Jain, Samuel E. Moelius III, and Sandra Zilles. Learning without coding. Theoretical Computer Science, 473(??):

124–148, February 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009322.

Orseau:2013:ANL

[1534] Laurent Orseau. Asymptotic nonlearnability of universal agents with computable horizon functions. Theoretical Computer Science, 473(??): 149–156, February 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009358.

Zhdanov:2013:IKR

[1535] Fedor Zhdanov and Yuri Kalnishkan.

An identity for kernel ridge regression. Theoretical Computer Science, 473(??): 157–178, February 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009371.

Anonymous:2013:EBg

[1536] Anonymous. Editorial Board. Theoretical Computer Science, 473(??): iii–viii, February 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000509.

Berard:2013:EPT

[1537] B. Bérard, F. Cassez, S. Haddad, D. Lime, and O. H. Roux. The expressive power of time Petri nets. Theoretical Computer Science, 474(??): 1–20, February 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512010845.

Barcelo:2013:PRE

[1538] Pablo Barceló, Juan Reutter, and Leonid Libkin. Parameterized regular expressions and their languages. Theoretical Computer Science, 474(??): 21–45, February 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513000029.

Gaina:2013:ILC

[1539] Daniel Gaina. Interpolation in logics with constructors. Theoretical Computer Science, 474(??):46– 59, February 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751201081X.

Genaim:2013:RBA

[1540] Samir Genaim and Damiano Zanardini. Reachability-based acyclicity analysis by Abstract Interpretation. Theoretical Computer Science, 474(??): 60–79, February 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512011115.

Lescanne:2013:CUL

[1541] Pierre Lescanne. On counting untyped lambda terms. Theoretical Computer Science, 474(??):80– 97, February 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010493.

Miller:2013:FFS

[1542] Dale Miller and Elaine Pimentel. A formal framework for specifying sequent calculus proof systems. Theoretical Computer Science, 474(??): 98–116, February 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010894.

Anonymous:2013:EBh

[1543] Anonymous. Editorial Board. Theoretical Computer Science, 474(??): iii–viii, February 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000741.

Araujo:2013:HNS

[1544] J. Araujo, V. Campos, F. Giroire, N. Nisse, L. Sampaio, and R. Soares. On the hull number of some graph classes. Theoretical Computer Science, 475(??):1–12, March 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011504.

Bang-Jensen:2013:PAD

[1545] Jørgen Bang-Jensen, Daniel Gonçalves, and Anders Yeo. Partitioning the arcs of a digraph into a star forest of the underlying graph with prescribed orientation properties. Theoretical Computer Science, 475(??): 13–20, March 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512010869.

Bednarska-Bzdega:2013:WFM

[1546] Małgorzata Bednarska-Bzdega. On weight function methods in Chooser– Picker games. Theoretical Computer Science, 475(??):21–33, March 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397513000030.

Blanchet-Sadri:2013:DRS

[1547] F. Blanchet-Sadri and Sean Simmons. Deciding representability of sets of words of equal length. Theoretical Computer Science, 475(??): 34–46, March 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751300008X.

Bretto:2013:FPH

[1548] Alain Bretto, Yannick Silvestre, and Thierry Vallée. Factorization of products of hypergraphs: Structure and algorithms. Theoretical Computer Science, 475(??):47– 58, March 4, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513000297.

Dolev:2013:UPH

[1549] Shlomi Dolev, Limor Lahiani, and Yinnon Haviv. Unique permutation hashing. Theoretical Computer Science, 475 (??):59–65, March 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000133.

Ewert:2013:NCS

[1550] Sigrid Ewert and Andries van der Walt. Necessary conditions for subclasses of random context languages. Theoretical Computer Science, 475 (??):66–72, March 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000042.

Brocchi:2013:TSC

[1551] S. Brocchi, A. Frosini, R. Pinzani, and S. Rinaldi. A tiling system for the class of L-convex polyominoes. Theoretical Computer Science, 475 (??):73–81, March 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011486.

Kuo:2013:FTP

[1552] Che-Nan Kuo, Hsin-Hung Chou, Nai-Wen Chang, and Sun-Yuan Hsieh. Fault-tolerant path embedding in folded hypercubes with both node and edge faults. Theoretical Computer Science, 475(??): 82–91, March 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010584.

Bienkowski:2013:ICA

[1553] Marcin Bienkowski, Marek Chrobak, Christoph Dürr, Mathilde Hurand, Artur Jez, Lukasz Jez, and Grzegorz Stachowiak. A ϕ-competitive algorithm for collecting items with increasing weights from a dynamic queue. Theoretical Computer Science, 475(??): 92–102, March 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000121.

Karim:2013:GBF

[1554] S. Karim, J. Sawada, Z. Alamgir, and S. M. Husnine. Generating bracelets with fixed content. Theoretical Computer Science, 475(??): 103–112, March 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010547.

Hon:2013:FCD

[1555] Wing-Kai Hon, Tsung-Han Ku, Rahul Shah, Sharma V. Thankachan, and Jeffrey Scott Vitter. Faster compressed dictionary matching. Theoretical Computer Science, 475(??): 113–119, March 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009826.

Balkova:2013:PBR

[1556] L’. Balková, E. Pelantová, and Š. Starosta. Proof of the Brlek– Reutenauer conjecture. Theoretical Computer Science, 475(??):120– 125, March 4, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011395.

Blumensath:2013:ESS

[1557] Achim Blumensath. Erratum to “On the structure of graphs in the Caucal hierarchy” [Theoret. Comput. Sci. 400 (2008) 19–45]. Theoretical Computer Science, 475(??): 126–127, March 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000108. See

[5].

Anonymous:2013:EBi

[1558] Anonymous. Editorial Board. Theoretical Computer Science, 475(??): iii–viii, March 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000996.

AhmadiPourAnari:2013:EPP

[1559] Nima AhmadiPourAnari, Shayan Ehsani, Mohammad Ghodsi, Nima Haghpanah, Nicole Immorlica, Hamid Mahini, and Vahab Mirrokni. Equilibrium pricing with positive externalities. Theoretical Computer Science, 476 (??):1–15, March 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000649.

Chang:2013:TCP

[1560] Jian Chang, Hui-Juan Wang, Jian-Liang Wu, and Yong-Ga A. Total colorings of planar graphs with maximum degree 8 and without 5cycles with two chords. Theoretical Computer Science, 476(??):16– 23, March 11, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000650.

Goldberg:2013:RGC

[1561] Leslie Ann Goldberg, Paul W. Goldberg, Piotr Krysta, and Carmine Ventre. Ranking games that have competitiveness-based strategies. Theoretical Computer Science, 476(??): 24–37, March 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000637.

Golovach:2013:OPC

[1562] Petr A. Golovach, Pim van ’t Hof, and Daniël Paulusma. Obtaining planarity by contracting few edges. Theoretical Computer Science, 476(??): 38–46, March 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000078.

Habib:2013:UPP

[1563] Michel Habib and Juraj Stacho. Unique perfect phylogeny is intractable. Theoretical Computer Science, 476(??): 47–66, March 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000145.

Huang:2013:SEC

[1564] Qiong Huang and Duncan S. Wong.

Short and efficient convertible undeniable signature schemes without random oracles. Theoretical Computer Science, 476(??):67– 83, March 11, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000601.

Lan:2013:MDP

[1565] James K. Lan and Gerard Jennhwa Chang. On the mixed domination problem in graphs. Theoretical Computer Science, 476(??): 84–93, March 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751201078X.

Li:2013:DDL

[1566] Si Li, Kwok Pui Choi, and Taoyang Wu. Degree distribution of large networks generated by the partial duplication model. Theoretical Computer Science, 476(??):94– 108, March 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300011X.

Li:2013:CAA

[1567] Yu Li, Donglei Du, Naihua Xiu, and Dachuan Xu. A combinatorial 2.375-approximation algorithm for the facility location problem with submodular penalties. Theoretical Computer Science, 476(??):109– 117, March 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512010808.

Anonymous:2013:EBj

[1568] Anonymous. Editorial Board. Theoretical Computer Science, 476(??): iii–viii, March 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751300131X.

Liang:2013:AML

[1569] Yu-Chang Liang and Xuding Zhu. Anti-magic labelling of Cartesian product of graphs. Theoretical Computer Science, 477(??):1–5, March 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397512011383.

Loridant:2013:TIC

[1570] B. Loridant, A. Messaoudi, P. Surer, and J. M. Thuswaldner. Tilings induced by a class of cubic Rauzy fractals. Theoretical Computer Science, 477(??):6–31, March 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512011127.

MacGillivray:2013:WNU

[1571] Gary MacGillivray and Jacobus Swarts. Weak near-unanimity functions and digraph homomorphism problems. Theoretical Computer Science, 477(??):32–47, March 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397512011401.

Mamino:2013:CCG

[1572] Marcello Mamino. On the computational complexity of a game of cops and robbers. Theoretical Computer Science, 477(??):48– 56, March 18, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010912.

Mastrolilli:2013:SMS

[1573] Monaldo Mastrolilli, Nikolaus Mutsanas, and Ola Svensson. Single machine scheduling with scenarios. Theoretical Computer Science, 477(??): 57–66, March 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010857.

Meduna:2013:GST

[1574] Alexander Meduna and Petr Zemek. On the generation of sentences with their parses by propagating regular-controlled grammars. Theoretical Computer Science, 477(??): 67–75, March 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000066.

Mertzios:2013:NME

[1575] George B. Mertzios, Sotiris Nikoletseas, Christoforos Raptopoulos, and Paul G. Spirakis. Natural models for evolution on networks. Theoretical Computer Science, 477(??):

76–95, March 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512010754.

Moser:2013:PDV

[1576] Philippe Moser. On the polynomial depth of various sets of random strings. Theoretical Computer Science, 477(??): 96–108, March 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009772.

Becher:2013:NNF

[1577] Verónica Becher and Pablo Ariel Heiber. Normal numbers and finite automata. Theoretical Computer Science, 477(??):109–116, March 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397513000698.

Anonymous:2013:EBk

[1578] Anonymous. Editorial Board. Theoretical Computer Science, 477(??): iii–viii, March 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513001412.

Blumensath:2013:APR

[1579] Achim Blumensath. An algebraic proof of Rabin’s Tree Theorem. Theoretical Computer Science, 478(??): 1–21, March 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000881.

Dutting:2013:BOA

[1580] Paul Dütting, Monika Henzinger, and Ingmar Weber. Bidder optimal assignments for general utilities. Theoretical Computer Science, 478(??): 22–32, March 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000923.

Dabrowski:2013:NRM

[1581] Konrad K. Dabrowski, Marc Demange, and Vadim V. Lozin. New results on maximum induced matchings in bipartite graphs and beyond. Theoretical Computer Science, 478(??): 33–40, March 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000893.

Bazgan:2013:SAB

[1582] Cristina Bazgan, Laurent Gourvès, Jérôme Monnot, and Fanny Pascual. Single approximation for the biobjective Max TSP. Theoretical Computer Science, 478(??):41– 50, March 25, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000832.

Kanj:2013:ING

[1583] Iyad Kanj and Fenghui Zhang. On the independence number of graphs with maximum degree 3. Theoretical Computer Science, 478(??):

51–75, March 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513000935.

Libert:2013:ASN

[1584] Benoît Libert and Moti Yung. Adaptively secure non-interactive threshold cryptosystems. Theoretical Computer Science, 478(??):76–100, March 25, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513000157.

Nevisi:2013:UUM

[1585] Hossein Nevisi and Daniel Reidenbach. Unambiguous 1-uniform morphisms. Theoretical Computer Science, 478(??): 101–117, March 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512010535.

Baba:2013:LTS

[1586] Daisuke Baba, Tomoko Izumi, Fukuhito Ooshita, Hirotsugu Kakugawa, and Toshimitsu Masuzawa. Linear time and space gathering of anonymous mobile agents in asynchronous trees. Theoretical Computer Science, 478(??): 118–126, March 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513000844.

Chen:2013:IRM

[1587] Wenbin Chen, Lingxi Peng, Jianxiong Wang, Fufang Li, Maobin Tang, Wei Xiong, and Songtao Wang. Inapproximability results for the minimum integral solution problem with preprocessing over ℓ_{∞} norm. Theoretical Computer Science, 478(??): 127–131, March 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300090X.

Anonymous:2013:EBl

[1588] Anonymous. Editorial Board. Theoretical Computer Science, 478(??): iii–viii, March 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001679.

Kotsireas:2013:P

[1589] Ilias S. Kotsireas, Bernard Mourrain, Victor Y. Pan, and Lihong Zhi. Preface. Theoretical Computer Science, 479(??):1–3, April 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009383.

Borwein:2013:MMS

[1590] Jonathan M. Borwein and Armin Straub. Mahler measures, short walks and log-sine integrals. Theoretical Computer Science, 479(??): 4–21, April 1, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009462.

Nuel:2013:SAE

[1591] Gregory Nuel and Jean-Guillaume Dumas. Sparse approaches for the exact distribution of patterns in long state sequences generated by a Markov source. Theoretical Computer Science, 479(??):22–42, April 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009401.

Iwane:2013:EIS

[1592] Hidenao Iwane, Hitoshi Yanami, Hirokazu Anai, and Kazuhiro Yokoyama. An effective implementation of symbolicnumeric cylindrical algebraic decomposition for quantifier elimination. Theoretical Computer Science, 479 (??):43–69, April 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009413.

Corless:2013:PEM

[1593] Robert M. Corless. Pseudospectra of exponential matrix polynomials. Theoretical Computer Science, 479 (??):70–80, April 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009425.

Emiris:2013:ICH

[1594] Ioannis Z. Emiris, Tatjana Kalinka, Christos Konaxis, and Thang Luu Ba. Implicitization of curves and (hyper)surfaces using predicted support. Theoretical Computer Science, 479(??):81–98, April 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009395.

Hodorog:2013:RAE

[1595] Madalina Hodorog and Josef Schicho.

A regularization approach for estimating the type of a plane curve singularity. Theoretical Computer Science, 479 (??):99–119, April 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009474.

Tsigaridas:2013:IBC

[1596] Elias P. Tsigaridas. Improved bounds for the CF algorithm. Theoretical Computer Science, 479(??):120– 126, April 1, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009437.

Terui:2013:GIM

[1597] Akira Terui. GPGCD: an iterative method for calculating approximate GCD of univariate polynomials. Theoretical Computer Science, 479(??): 127–149, April 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009449.

Li:2013:CNS

[1598] Zijia Li and Lihong Zhi. Computing the nearest singular univariate polynomials with given root multiplicities. Theoretical Computer Science, 479 (??):150–162, April 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009486.

Li:2013:VEB

[1599] Nan Li and Lihong Zhi. Verified error bounds for isolated singular solutions of polynomial systems: Case of breadth one. Theoretical Computer Science, 479 (??):163–173, April 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009498.

Fassino:2013:SVL

[1600] Claudia Fassino and Maria-Laura Torrente. Simple varieties for limited precision points. Theoretical Computer Science, 479(??):174– 186, April 1, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009450.

Anonymous:2013:EBm

[1601] Anonymous. Editorial Board. Theoretical Computer Science, 479(??): iii–viii, April 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513001874.

Badouel:2013:AGT

[1602] E. Badouel, R. Tchougong, C. Nkuimi-Jugnia, and B. Fotsing. Attribute grammars as tree transducers over cyclic representations of infinite trees and their descriptional composition. Theoretical Computer Science, 480 (??):1–25, April 8, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001163.

Ciabattoni:2013:PTL

[1603] Agata Ciabattoni and Franco Montagna. Proof theory for locally finite many-valued logics: Semi-projective logics. Theoretical Computer Science, 480(??):26–42, April 8, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001126.

DAgostino:2013:SPT

[1604] Marcello D’Agostino, Marcelo Finger, and Dov Gabbay. Semantics and prooftheory of depth bounded Boolean logics. Theoretical Computer Science, 480 (??):43–68, April 8, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001230.

Gurov:2013:RBS

[1605] Dilian Gurov and Marieke Huisman.

Reducing behavioural to structural properties of programs with procedures. Theoretical Computer Science, 480(??):69–103, April 8, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001151.

Aminof:2013:RAD

[1606] Benjamin Aminof, Orna Kupferman, and Robby Lampert. Rigorous approximated determinization of weighted automata. Theoretical Computer Science, 480(??):104– 117, April 8, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751300114X.

Anonymous:2013:EBn

[1607] Anonymous. Editorial Board. Theoretical Computer Science, 480(??): iii–viii, April 8, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513002053.

Fici:2013:LNP

[1608] Gabriele Fici and Luca Q. Zamboni. On the least number of palindromes contained in an infinite word. Theoretical Computer Science, 481 (??):1–8, April 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513001229.

Lu:2013:NFM

[1609] Chia Wei Lu, Chin Lung Lu, and R. C. T. Lee. A new filtration method and a hybrid strategy for approximate string matching. Theoretical Computer Science, 481(??): 9–17, April 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513001540.

Chen:2013:FHR

[1610] Lily Chen, Xueliang Li, and Huishu Lian. Further hardness results on the rainbow vertex-connection number of graphs. Theoretical Computer Science, 481(??):18–23, April 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001217.

Li:2013:BNC

[1611] Yuan Li, John O. Adeyeye, David Murrugarra, Boris Aguilar, and Reinhard Laubenbacher. Boolean nested canalizing functions: a comprehensive analysis. Theoretical Computer Science, 481 (??):24–36, April 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001527.

Ganchev:2013:TCS

[1612] Antoniy Ganchev, Lata Narayanan, and Sunil Shende. A tight characterization of strategic games with a unique equilibrium. Theoretical Computer Science, 481(??):37– 50, April 15, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001175.

Owens:2013:RDA

[1613] Nick D. L. Owens, Andrew Greensted, Jon Timmis, and Andy Tyrrell. The Receptor Density Algorithm. Theoretical Computer Science, 481(??): 51–73, April 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009899.

Golovach:2013:IMD

[1614] Petr A. Golovach, Marcin Kamiński, Daniël Paulusma, and Dimitrios M. Thilikos. Increasing the minimum degree of a graph by contractions. Theoretical Computer Science, 481 (??):74–84, April 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751300162X.

Ordyniak:2013:SAA

[1615] Sebastian Ordyniak, Daniel Paulusma, and Stefan Szeider. Satisfiability of acyclic and almost acyclic CNF formulas. Theoretical Computer Science, 481 (??):85–99, April 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513000054.

Kehagias:2013:CIR

[1616] Athanasios Kehagias, Dieter Mitsche, and Pawel Pralat. Cops and invisible robbers: the cost of drunkenness. Theoretical Computer Science, 481(??): 100–120, April 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513001096.

Anonymous:2013:EBo

[1617] Anonymous. Editorial Board. Theoretical Computer Science, 481(??): iii–viii, April 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513002193.

Currie:2013:ITS

[1618] James Currie. Infinite ternary squarefree words concatenated from permutations of a single word. Theoretical Computer Science, 482(??): 1–8, April 22, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001801.

Chen:2013:MML

[1619] Xujin Chen, Leah Epstein, Elena Kleiman, and Rob van Stee. Maximizing the minimum load: the cost of selfishness. Theoretical Computer Science, 482(??):9–19, April 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001795.

Golovach:2013:DIM

[1620] Petr A. Golovach, Dieter Kratsch, and Daniël Paulusma. Detecting induced minors in AT-free graphs. Theoretical Computer Science, 482(??): 20–32, April 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001618.

Kawase:2013:NEM

[1621] Yasushi Kawase and Kazuhisa Makino.

Nash equilibria with minimum potential in undirected broadcast games. Theoretical Computer Science, 482 (??):33–47, April 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001771.

Roura:2013:FBN

[1622] Salvador Roura. Fibonacci BSTs: a new balancing method for binary search trees. Theoretical Computer Science, 482(??):48–59, April 22, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512010572.

Navarro:2013:SED

[1623] Gonzalo Navarro, Yakov Nekrich, and Luís M. S. Russo. Space-efficient data-analysis queries on grids. Theoretical Computer Science, 482(??): 60–72, April 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512010742.

Jain:2013:LC

[1624] Sanjay Jain, Eric Martin, and Frank Stephan. Learning and classifying. Theoretical Computer Science, 482 (??):73–85, April 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513001576.

Thang:2013:IHP

[1625] Nguyen Kim Thang. N P-hardness of pure Nash equilibrium in scheduling and network design games. Theoretical Computer Science, 482(??): 86–95, April 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009838.

Woryna:2013:CSS

[1626] Adam Woryna. The concept of selfsimilar automata over a changing alphabet and lamplighter groups generated by such automata. Theoretical Computer Science, 482(??): 96–110, April 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300159X.

Wang:2013:TMI

[1627] Xiumei Wang and Yixun Lin. Threematching intersection conjecture for perfect matching polytopes of small dimensions. Theoretical Computer Science, 482(??):111–114, April 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397513001552.

Anonymous:2013:EBp

[1628] Anonymous. Editorial Board. Theoretical Computer Science, 482(??): iii–viii, April 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002417.

Giancarlo:2013:F

[1629] Raffaele Giancarlo and Giovanni Manzini. Foreword. Theoretical Computer Science, 483(??):1, April 29, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011255.

Breslauer:2013:SRT

[1630] Dany Breslauer, Roberto Grossi, and Filippo Mignosi. Simple real-time constant-space string matching. Theoretical Computer Science, 483(??): 2–9, April 29, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010900.

Dondi:2013:FAC

[1631] Riccardo Dondi, Guillaume Fertin, and Stéphane Vialette. Finding approximate and constrained motifs in graphs. Theoretical Computer Science, 483(??):10–21, April 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007979.

Thachuk:2013:CIT

[1632] Chris Thachuk. Compressed indexes for text with wildcards. Theoretical Computer Science, 483(??): 22–35, April 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007840.

Gagie:2013:CRQ

[1633] Travis Gagie, Juha Kärkkäinen, Gonzalo Navarro, and Simon J. Puglisi. Colored range queries and document retrieval. Theoretical Computer Science, 483(??):36– 50, April 29, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007657.

Schirmer:2013:FAA

[1634] Stefanie Schirmer and Robert Giegerich.

Forest alignment with affine gaps and anchors, applied in RNA structure comparison. Theoretical Computer Science, 483(??):51–67, April 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397512007499.

Clifford:2013:SLB

[1635] Raphaël Clifford, Markus Jalsenius, Ely Porat, and Benjamin Sach. Space lower bounds for online pattern matching. Theoretical Computer Science, 483 (??):68–74, April 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005816.

Ferragina:2013:WPS

[1636] Paolo Ferragina. On the weak prefix-search problem. Theoretical Computer Science, 483(??):75– 84, April 29, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005804.

Cantone:2013:ESM

[1637] Domenico Cantone, Salvatore Cristofaro, and Simone Faro. Efficient string-matching allowing for nonoverlapping inversions. Theoretical Computer Science, 483(??):85– 95, April 29, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005452.

Elberfeld:2013:AAO

[1638] Michael Elberfeld, Danny Segev, Colin R. Davidson, Dana Silverbush, and Roded Sharan. Approximation algorithms for orienting mixed graphs. Theoretical Computer Science, 483(??): 96–103, April 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512003167.

Brown:2013:FET

[1639] Daniel G. Brown and Jakub Truszkowski. Fast error-tolerant quartet phylogeny algorithms. Theoretical Computer Science, 483(??):104–114, April 29, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001570.

Kreft:2013:CIR

[1640] Sebastian Kreft and Gonzalo Navarro. On compressing and indexing repetitive sequences. Theoretical Computer Science, 483(??):115–133, April 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397512001259.

Bauer:2013:LAC

[1641] Markus J. Bauer, Anthony J. Cox, and Giovanna Rosone. Lightweight algorithms for constructing and inverting the BWT of string collections. Theoretical Computer Science, 483(??): 134–148, April 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001211.

Daykin:2013:LPA

[1642] David E. Daykin, Jacqueline W.

Daykin, and W. F. Smyth. A linear partitioning algorithm for Hybrid Lyndons using v-order. Theoretical Computer Science, 483(??): 149–161, April 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200120X.

I:2013:PPM

[1643] Tomohiro I, Shunsuke Inenaga, and Masayuki Takeda. Palindrome pattern matching. Theoretical Computer Science, 483(??):162–170, April 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397512001041.

Christou:2013:ESC

[1644] M. Christou, M. Crochemore, C. S.

Iliopoulos, M. Kubica, S. P. Pissis, J. Radoszewski, W. Rytter, B. Szreder, and T. Waleń. Efficient seed computation revisited. Theoretical Computer Science, 483(??): 171–181, April 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511010632.

Anonymous:2013:EBq

[1645] Anonymous. Editorial Board. Theoretical Computer Science, 483(??): iii–viii, April 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513002569.

Carmona:2013:CMC

[1646] Josep Carmona and Jetty Kleijn. Compatibility in a multi-component environment. Theoretical Computer Science, 484(??):1–15, May 6, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397513001977.

Nikolic:2013:ICI

[1647] Durica Nikolić and Fausto Spoto. Inferring complete initialization of arrays. Theoretical Computer Science, 484(??):16–40, May 6, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513000455.

Dosen:2013:GPC

[1648] Kosta Dosen and Zoran Petrić. Graphs of plural cuts. Theoretical Computer Science, 484(??):41– 55, May 6, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513000947.

Worboys:2013:UBM

[1649] M. F. Worboys. Using bigraphs to model topological graphs embedded in orientable surfaces. Theoretical Computer Science, 484(??): 56–69, May 6, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001503.

vanderMeyden:2013:IFSb

[1650] Ron van der Meyden and Chenyi Zhang. Information flow in systems with schedulers, Part II: Refinement. Theoretical Computer Science, 484 (??):70–92, May 6, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000285.

Anonymous:2013:EBr

[1651] Anonymous. Editorial Board. Theoretical Computer Science, 484(??): iii–viii, May 6, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002764.

Bjorklund:2013:SRP

[1652] Johanna Björklund and Lars-Daniel Öhman. Simulation relations for pattern matching in directed graphs. Theoretical Computer Science, 485 (??):1–15, May 13, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300234X.

Blanchet-Sadri:2013:APP

[1653] F. Blanchet-Sadri and Nathan Fox.

Abelian-primitive partial words. Theoretical Computer Science, 485(??): 16–37, May 13, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001989.

Droste:2013:WFA

[1654] Manfred Droste and Werner Kuich. Weighted finite automata over hemirings. Theoretical Computer Science, 485(??):38–48, May 13, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513001606.

Beaulieu:2013:ICG

[1655] G. Beaulieu, K. Burke, and E. Duchêne. Impartial coloring games. Theoretical Computer Science, 485(??): 49–60, May 13, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513001783.

Feldmann:2013:FBP

[1656] Andreas Emil Feldmann. Fast balanced partitioning is hard even on grids and trees. Theoretical Computer Science, 485(??):61–68, May 13, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513002259.

Broersma:2013:TCB

[1657] Hajo Broersma, Petr A. Golovach, and Viresh Patel. Tight complexity bounds for FPT subgraph problems parameterized by the cliquewidth. Theoretical Computer Science, 485(??):69–84, May 13, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513001990.

Kasprzik:2013:FOS

[1658] Anna Kasprzik. Four one-shot learners for regular tree languages and their polynomial characterizability. Theoretical Computer Science, 485(??): 85–106, May 13, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000303.

Anonymous:2013:EBs

[1659] Anonymous. Editorial Board. Theoretical Computer Science, 485(??): iii–viii, May 13, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002922.

Mor:2013:P

[1660] Tal Mor and Alain Tapp. Preface. Theoretical Computer Science, 486(??):1–3, May 20, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011036.

Brassard:2013:SHA

[1661] Gilles Brassard and André Allan Méthot. Strict hierarchy among Bell Theorems. Theoretical Computer Science, 486(??):4–10, May 20, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397512011048.

Cleve:2013:QEC

[1662] Richard Cleve, Wim van Dam, Michael Nielsen, and Alain Tapp. Quantum entanglement and the communication complexity of the inner product function. Theoretical Computer Science, 486(??):11–19, May 20, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751201105X.

Galliard:2013:DQN

[1663] Viktor Galliard, Alain Tapp, and Stefan Wolf. Deterministic quantum nonlocality and graph colorings. Theoretical Computer Science, 486(??): 20–26, May 20, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011061.

Hanggi:2013:INS

[1664] Esther Hänggi, Renato Renner, and Stefan Wolf. The impossibility of non-signaling privacy amplification. Theoretical Computer Science, 486 (??):27–42, May 20, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011073.

LeGall:2013:QWN

[1665] François Le Gall. Quantum weakly nondeterministic communication complexity. Theoretical Computer Science, 486(??):43–49, May 20, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011085.

Renner:2013:TCN

[1666] Renato Renner and Stefan Wolf. Towards characterizing the non-locality of entangled quantum states. Theoretical Computer Science, 486(??): 50–60, May 20, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011097.

Brassard:2013:CQN

[1667] Gilles Brassard, Anne Broadbent, Esther Hänggi, André Allan Méthot, and Stefan Wolf. Classical, quantum and nonsignalling resources in bipartite games. Theoretical Computer Science, 486(??):61–72, May 20, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011103.

Anonymous:2013:EBt

[1668] Anonymous. Editorial Board. Theoretical Computer Science, 486(??): iii–viii, May 20, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003071.

Fried:2013:CCT

[1669] Dror Fried, Solomon Eyal Shimony, Amit Benbassat, and Cenny Wenner. Complexity of Canadian traveler problem variants. Theoretical Computer Science, 487(??):1– 16, May 27, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002296.

deParga:2013:PDR

[1670] Manuel Vázquez de Parga, Pedro García, and Damián López. A polynomial double reversal minimization algorithm for deterministic finite automata. Theoretical Computer Science, 487(??):17–22, May 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513001965.

Carrasco:2013:EEC

[1671] Nicolás Carrasco, Jean-Marie Le Bars, and Alfredo Viola. Enumerative encoding of correlationimmune Boolean functions. Theoretical Computer Science, 487(??): 23–36, May 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513001588.

Ndione:2013:AMR

[1672] Antoine Ndione, Aurélien Lemay, and Joachim Niehren. Approximate membership for regular languages modulo the edit distance. Theoretical Computer Science, 487(??):37– 49, May 27, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513001953.

Campos:2013:BCT

[1673] Victor Campos, Frédéric Havet, Rudini Sampaio, and Ana Silva. Backbone colouring: Tree backbones with small diameter in planar graphs. Theoretical Computer Science, 487(??):

50–64, May 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001941.

Early:2013:RTT

[1674] D. Early and M. Schellekens. Running time of the Treapsort algorithm. Theoretical Computer Science, 487 (??):65–73, May 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002132.

Shao:2013:LMP

[1675] Zhendong Shao and Roberto Solis-Oba. L(2, 1)-labelings on the modular product of two graphs. Theoretical Computer Science, 487(??): 74–81, May 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001114.

Couturier:2013:MDS

[1676] Jean-François Couturier, Pinar Heggernes, Pim van ’t Hof, and Dieter Kratsch. Minimal dominating sets in graph classes: Combinatorial bounds and enumeration. Theoretical Computer Science, 487(??): 82–94, May 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002508.

Yu:2013:OSS

[1677] Sheng Yu and Prudence W. H. Wong.

Online scheduling of simple linear deteriorating jobs to minimize the total general completion time. Theoretical Computer Science, 487(??): 95–102, May 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001564.

Brandstadt:2013:CST

[1678] Andreas Brandstädt, Synara Brito, Sulamita Klein, Loana Tito Nogueira, and Fábio Protti. Corrigendum to “Cycle transversals in perfect graphs and cographs” [Theoret. Comput. Sci. 469 (2013) 15–23]. Theoretical Computer Science, 487(??): 103–105, May 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002351. See

[1493].

Anonymous:2013:EBu

[1679] Anonymous. Editorial Board. Theoretical Computer Science, 487(??): iii–viii, May 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003228.

Rozenberg:2013:Pa

[1680] G. Rozenberg.  Preface.  Theo-

retical Computer Science, 488(??): 1, June 3, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003526.

Pal:2013:CGS

[1681] Sankar K. Pal and Romi Banerjee.

Context granulation and subjectiveinformation quantification. Theoretical Computer Science, 488(??):2– 14, June 3, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011498.

Degasperi:2013:PAF

[1682] A. Degasperi and M. Calder. A process algebra framework for multiscale modelling of biological systems. Theoretical Computer Science, 488 (??):15–45, June 3, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002314.

Diker:2013:CRS

[1683] Murat Diker. Categories of rough sets and textures. Theoretical Computer Science, 488(??):46–65, June 3, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011231.

Jung:2013:CPF

[1684] Achim Jung and Jonathan E. Rowe.

Convergence of preference functions. Theoretical Computer Science, 488 (??):66–77, June 3, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002363.

Sosik:2013:SPM

[1685] Petr Sosík, Andrei Paun, and Alfonso Rodríguez-Patón. P systems with proteins on membranes characterize PSPACE. Theoretical Computer Science, 488(??):78– 95, June 3, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513002004.

Anonymous:2013:EBv

[1686] Anonymous. Editorial Board. Theoretical Computer Science, 488(??): iii–viii, June 3, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513003460.

Berglund:2013:SLR

[1687] Martin Berglund, Henrik Björklund, and Johanna Björklund. Shuffled languages — representation and recognition. Theoretical Computer Science, 489–490(??):1–20, June 10, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397513003307.

Chan:2013:OCK

[1688] Aldar C.-F. Chan. On optimal cryptographic key derivation. Theoretical Computer Science, 489–490(??): 21–36, June 10, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513002843.

Jain:2013:MCS

[1689] Sanjay Jain and Efim Kinber. Mind change speed-up for learning languages from positive data. Theoretical Computer Science, 489–490 (??):37–47, June 10, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002715.

Yu:2013:PCA

[1690] Fuxiang Yu and Ker-I Ko. On parallel complexity of analytic functions. Theoretical Computer Science, 489–490 (??):48–57, June 10, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002703.

Lu:2013:SOS

[1691] Xinrong Lu and Zhaohui Liu. Semionline scheduling problems on two uniform machines under a grade of service provision. Theoretical Computer Science, 489–490(??):58– 66, June 10, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002879.

Cohen:2013:IAA

[1692] Nachshon Cohen and Zeev Nutov. A 1 + ln 2-approximation algorithm for minimum-cost 2-edge-connectivity augmentation of trees with constant radius. Theoretical Computer Science, 489–490(??): 67–74, June 10, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513002661.

Kesselman:2013:CBM

[1693] Alex Kesselman, Boaz Patt-Shamir, and Gabriel Scalosub. Competitive buffer management with packet dependencies. Theoretical Computer Science, 489–490(??):75–87, June 10, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513002697.

Christodoulou:2013:TCA

[1694] George Christodoulou, Annamária Kovács, and Rob van Stee. A truthful constant approximation for maximizing the minimum load on related machines. Theoretical Computer Science, 489–490 (??):88–98, June 10, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751300265X.

Anonymous:2013:EBw

[1695] Anonymous. Editorial Board. Theoretical Computer Science, 489–490 (??):iii–viii, June 10, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513003782.

Berberich:2013:ESN

[1696] Eric Berberich, Pavel Emeliyanenko, Alexander Kobel, and Michael Sagraloff. Exact symbolic-numeric computation of planar algebraic curves. Theoretical Computer Science, 491(??):

1–32, June 17, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002983.

Cerny:2013:PSE

[1697] Anton Cerný. On Prouhet’s solution to the equal powers problem. Theoretical Computer Science, 491 (??):33–46, June 17, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002636.

Krenn:2013:AWI

[1698] Daniel Krenn. Analysis of the widthw non-adjacent form in conjunction with hyperelliptic curve cryptography and with lattices. Theoretical Computer Science, 491(??): 47–70, June 17, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002685.

Cheng:2013:SMP

[1699] Eddie Cheng, Shalin Shah, Vyom Shah, and Daniel E. Steffy. Strong matching preclusion for augmented cubes. Theoretical Computer Science, 491 (??):71–77, June 17, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003393.

Yang:2013:FFH

[1700] Yuxing Yang and Shiying Wang.

Fault-free Hamiltonian cycles passing through a linear forest in ternary n-cubes with faulty edges. Theoretical Computer Science, 491(??): 78–82, June 17, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009802.

Seo:2013:PIC

[1701] Jae Woo Seo, Dae Hyun Yum, and Pil Joong Lee. Proxy-invisible CCAsecure type-based proxy re-encryption without random oracles. Theoretical Computer Science, 491(??):83– 93, June 17, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010560.

Dong:2013:AAT

[1702] Jianming Dong, An Zhang, Yong Chen, and Qifan Yang. Approximation algorithms for two-machine open shop scheduling with batch and delivery coordination. Theoretical Computer Science, 491(??): 94–102, June 17, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003332.

Zhang:2013:MMD

[1703] Shurong Zhang and Shiying Wang. Many-to-many disjoint path covers in k-ary n-cubes. Theoretical Computer Science, 491(??):103– 118, June 17, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002648.

Jin:2013:CIC

[1704] Jing Jin, Baogang Xu, and Xiaoyan Zhang. On the complexity of injective colorings and its generalizations. Theoretical Computer Science, 491(??): 119–126, June 17, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003344.

Yin:2013:SGG

[1705] Xiao Yin and Daming Zhu. Sorting genomes by generalized translocations. Theoretical Computer Science, 491(??): 127–135, June 17, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000388.

Tassa:2013:ASA

[1706] Tamir Tassa. Addendum to “Finding all maximally-matchable edges in a bipartite graph” [Theoret. Comput. Sci. 423 (2012) 50–58]. Theoretical Computer Science, 491(??): 136, June 17, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002284. See

[1038].

Anonymous:2013:EBx

[1707] Anonymous. Editorial Board. Theoretical Computer Science, 491(??): iii–viii, June 17, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003915.

Cheval:2013:DEB

[1708] Vincent Cheval, Véronique Cortier, and Stéphanie Delaune. Deciding equivalence-based properties using constraint solving. Theoretical Computer Science, 492(??):1– 39, June 24, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513003009.

Kamide:2013:TBP

[1709] Norihiro Kamide. Temporal BI: Proof system, semantics and translations. Theoretical Computer Science, 492 (??):40–69, June 24, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751300337X.

Kullmann:2013:DPR

[1710] Oliver Kullmann and Xishun Zhao. On Davis–Putnam reductions for minimally unsatisfiable clause-sets. Theoretical Computer Science, 492(??): 70–87, June 24, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513003174.

Colazzo:2013:EAI

[1711] D. Colazzo, G. Ghelli, L. Pardini, and C. Sartiani. Efficient asymmetric inclusion of regular expressions with interleaving and counting for XML type-checking. Theoretical Computer Science, 492(??): 88–116, June 24, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003319.

vanBreugel:2013:ASD

[1712] Franck van Breugel, Claudio Hermida, Michael Makkai, and James Worrell. Addendum to “Recursively defined metric spaces without contraction” [Theoret. Comput. Sci. 380 (1/2) (2007) 143–163]. Theoretical Computer Science, 492(??): 117–122, June 24, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300340X. See

[4].

Anonymous:2013:EBy

[1713] Anonymous. Editorial Board. Theoretical Computer Science, 492(??): iii–viii, June 24, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004076.

DAgostino:2013:GAL

[1714] Giovanna D’Agostino and Salvatore La Torre. Games, Automata, Logic, and Formal Verification (GandALF 2011). Theoretical Computer Science, 493(??):1, July 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004209.

Gradel:2013:MCG

[1715] Erich Grädel. Model-checking games for logics of imperfect information. Theoretical Computer Science, 493 (??):2–14, July 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009541.

Bresolin:2013:GTA

[1716] Davide Bresolin and Marta Capiluppi. A game-theoretic approach to fault diagnosis and identification of hybrid systems. Theoretical Computer Science, 493(??):15–29, July 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007396.

Zimmermann:2013:OBP

[1717] Martin Zimmermann. Optimal bounds in parametric LTL games. Theoretical Computer Science, 493(??): 30–45, July 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007487.

Gruner:2013:CGD

[1718] Sten Grüner, Frank G. Radmacher, and Wolfgang Thomas. Connectivity games over dynamic networks. Theoretical Computer Science, 493 (??):46–65, July 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512010596.

Haar:2013:CRR

[1719] Stefan Haar, Christian Kern, and Stefan Schwoon. Computing the reveals relation in occurrence nets. Theoretical Computer Science, 493 (??):66–79, July 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003368.

Brozek:2013:DOS

[1720] Václav Brozek. Determinacy and optimal strategies in infinite-state stochastic reachability games. Theoretical Computer Science, 493(??): 80–97, July 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200967X.

Bresolin:2013:ODP

[1721] Davide Bresolin, Angelo Montanari, Pietro Sala, and Guido Sciavicco. Optimal decision procedures for MPNL over finite structures, the natural numbers, and the integers. Theoretical Computer Science, 493(??): 98–115, July 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009759.

Benerecetti:2013:ASS

[1722] Massimo Benerecetti, Marco Faella, and Stefano Minopoli. Automatic synthesis of switching controllers for linear hybrid systems: Safety control. Theoretical Computer Science, 493(??):116–138, July 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009747.

Anonymous:2013:EBz

[1723] Anonymous. Editorial Board. Theoretical Computer Science, 493(??): iii–viii, July 1, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513004180.

Fiala:2013:P

[1724] Jiří Fiala, Jan Kratochvíl, and Angsheng Li. Preface. Theoretical Computer Science, 494(??): 1, July 8, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513004337.

Ambos-Spies:2013:NET

[1725] Klaus Ambos-Spies and Timur Bakibayev. Nontriviality for exponential time w.r.t. weak reducibilities. Theoretical Computer Science, 494 (??):2–12, July 8, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751200480X.

Babu:2013:SAL

[1726] Ajesh Babu, Nutan Limaye, Jaikumar Radhakrishnan, and Girish Varma. Streaming algorithms for language recognition problems. Theoretical Computer Science, 494(??):13– 23, July 8, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512011437.

Feng:2013:TJI

[1727] Rongquan Feng, Menglong Nie, and Hongfeng Wu. Twisted Jacobi intersections curves. Theoretical Computer Science, 494(??):24– 35, July 8, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011425.

Ambainis:2013:SNG

[1728] Andris Ambainis, Dmitry Kravchenko, Nikolay Nahimov, Alexander Rivosh, and Madars Virza. On symmetric nonlocal games. Theoretical Computer Science, 494(??):36– 48, July 8, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002120.

Cibulka:2013:GSG

[1729] Josef Cibulka, Jan Kyncl, Viola Mészáros, Rudolf Stolar, and Pavel Valtr. Graph sharing games: Complexity and connectivity. Theoretical Computer Science, 494(??):49– 62, July 8, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011449.

Cai:2013:PFI

[1730] Jin-Yi Cai and Michael Kowalczyk. Partition functions on kregular graphs with {0, 1}-vertex assignments and real edge functions. Theoretical Computer Science, 494 (??):63–74, July 8, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513000091.

Crespelle:2013:TAM

[1731] Christophe Crespelle and Ioan Todinca. An O(n^{2})-time algorithm for the minimal interval completion problem. Theoretical Computer Science, 494(??):75–85, July 8, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512011462.

Hartung:2013:ILC

[1732] Sepp Hartung and Rolf Niedermeier. Incremental list coloring of graphs, parameterized by conservation. Theoretical Computer Science, 494(??): 86–98, July 8, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513000169.

Uhlmann:2013:TLP

[1733] Johannes Uhlmann and Mathias Weller. Two-Layer Planarization parameterized by feedback edge set. Theoretical Computer Science, 494 (??):99–111, July 8, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513000911.

He:2013:SRA

[1734] Xin He and Huaming Zhang. A simple routing algorithm based on Schnyder coordinates. Theoretical Computer Science, 494(??):112– 121, July 8, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000674.

Bulteau:2013:RMB

[1735] Laurent Bulteau, Guillaume Fertin, and Irena Rusu. Revisiting the Minimum Breakpoint Linearization Problem. Theoretical Computer Science, 494(??):122–133, July 8, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011413.

Guillemot:2013:KFA

[1736] Sylvain Guillemot and Matthias Mnich. Kernel and fast algorithm for dense triplet inconsistency. Theoretical Computer Science, 494(??): 134–143, July 8, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011474.

Kao:2013:DPT

[1737] Ming-Yang Kao, Henry C. M. Leung, He Sun, and Yong Zhang. Deterministic polynomial-time algorithms for designing short DNA words. Theoretical Computer Science, 494(??): 144–160, July 8, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011450.

Anonymous:2013:EBaa

[1738] Anonymous. Editorial Board. Theoretical Computer Science, 494(??): iii–viii, July 8, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513004301.

Cantone:2013:FAR

[1739] Domenico Cantone and Micha Hofri. Further analysis of the remedian algorithm. Theoretical Computer Science, 495(??):1–16, July 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513004519.

Bucci:2013:RCF

[1740] Michelangelo Bucci, Alessandro De Luca, and Luca Q. Zamboni. Reversible Christoffel factorizations. Theoretical Computer Science, 495(??): 17–24, July 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513004544.

Dehghan:2013:ACP

[1741] Ali Dehghan, Mohammad-Reza Sadeghi, and Arash Ahadi. Algorithmic complexity of proper labeling problems. Theoretical Computer Science, 495 (??):25–36, July 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513003988.

Dereniowski:2013:MCE

[1742] Dariusz Dereniowski and Danny Dyer. On minimum cost edge searching. Theoretical Computer Science, 495 (??):37–49, July 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004635.

Elzinga:2013:VSK

[1743] Cees H. Elzinga and Hui Wang.

Versatile string kernels. Theoretical Computer Science, 495(??):50– 65, July 15, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300460X.

Adar:2013:SPC

[1744] Ron Adar and Leah Epstein. Selfish bin packing with cardinality constraints. Theoretical Computer Science, 495(??):66–80, July 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004532.

Cordon-Franco:2013:CPG

[1745] Andrés Cordón-Franco, Hans van Ditmarsch, David Fernández-Duque, and Fernando Soler-Toscano. A colouring protocol for the generalized Russian cards problem. Theoretical Computer Science, 495(??): 81–95, July 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003617.

Gu:2013:EAR

[1746] Manzhan Gu and Xiwen Lu. The expected asymptotical ratio for preemptive stochastic online problem. Theoretical Computer Science, 495(??):

96–112, July 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513004477.

Anonymous:2013:EBab

[1747] Anonymous. Editorial Board. Theoretical Computer Science, 495(??): iii–viii, July 15, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513004738.

Datta:2013:P

[1748] Ajoy K. Datta and Stéphane Devismes. Preface. Theoretical Computer Science, 496(??):1, July 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513004842.

Mehta:2013:IBI

[1749] Abhinav Mehta, Shashank Agrawal, and Kannan Srinathan. Interplay between (im)perfectness, synchrony and connectivity: the case of reliable message transmission. Theoretical Computer Science, 496(??):2– 16, July 22, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512010833.

Chen:2013:MFA

[1750] Jingshu Chen and Sandeep Kulkarni. MR4UM: a framework for adding fault tolerance to UML state diagrams. Theoretical Computer Science, 496(??):17–33, July 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009127.

Delporte-Gallet:2013:BAH

[1751] Carole Delporte-Gallet, Hugues Fauconnier, and Hung Tran-The. Byzantine agreement with homonyms in synchronous systems. Theoretical Computer Science, 496(??):34– 49, July 22, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010420.

Ebnenasir:2013:FDF

[1752] Ali Ebnenasir, Reza Hajisheykhi, and Sandeep S. Kulkarni. Facilitating the design of fault tolerance in transaction level SystemC programs. Theoretical Computer Science, 496(??): 50–68, July 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010407.

Kentros:2013:SMO

[1753] Sotirios Kentros and Aggelos Kiayias.

Solving the at-most-once problem with nearly optimal effectiveness. Theoretical Computer Science, 496(??): 69–88, July 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003010.

Cournier:2013:SSM

[1754] Alain Cournier, Swan Dubois, Anissa Lamani, Franck Petit, and Vincent Villain. The snap-stabilizing message forwarding algorithm on tree topologies. Theoretical Computer Science, 496(??):89–112, July 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002867.

Pal:2013:EMC

[1755] Ranjan Pal and Pan Hui. Economic models for cloud service markets: Pricing and capacity planning. Theoretical Computer Science, 496(??): 113–124, July 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010018.

Peri:2013:CCE

[1756] Sathya Peri and K. Vidyasankar. Correctness of concurrent executions of closed nested transactions in transactional memory systems. Theoretical Computer Science, 496(??): 125–153, July 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000868.

Crain:2013:TUC

[1757] Tyler Crain, Damien Imbs, and Michel Raynal. Towards a universal construction for transactionbased multiprocess programs. Theoretical Computer Science, 496(??): 154–169, July 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008614.

Santos:2013:OPB

[1758] Nuno Santos and André Schiper. Optimizing Paxos with batching and pipelining. Theoretical Computer Science, 496(??):170–183, July 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397512009097.

Even:2013:CDE

[1759] Guy Even, Moti Medina, Gregor Schaffrath, and Stefan Schmid. Competitive and deterministic embeddings of virtual networks. Theoretical Computer Science, 496(??): 184–194, July 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009577.

Anonymous:2013:EBac

[1760] Anonymous. Editorial Board. Theoretical Computer Science, 496(??): iii–viii, July 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004891.

Chen:2013:AAB

[1761] Lin Chen and Guochuan Zhang. Approximation algorithms for a bilevel knapsack problem. Theoretical Computer Science, 497(??):1– 12, July 29, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007694.

Umesato:2013:CSF

[1762] Takuya Umesato, Toshiki Saitoh, Ryuhei Uehara, Hiro Ito, and Yoshio Okamoto. The complexity of the stamp folding problem. Theoretical Computer Science, 497(??):13– 19, July 29, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007670.

Zhu:2013:IAF

[1763] Yongding Zhu and Jinhui Xu. Improved algorithms for the farthest colored Voronoi diagram of segments. Theoretical Computer Science, 497 (??):20–30, July 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005440.

Li:2013:CSO

[1764] Fei Li. A comprehensive study of an online packet scheduling algorithm. Theoretical Computer Science, 497 (??):31–38, July 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005385.

Chen:2013:TMM

[1765] Zhixiang Chen, Bin Fu, Yang Liu, and Robert Schweller. On testing monomials in multivariate polynomials. Theoretical Computer Science, 497 (??):39–54, July 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003106.

Yang:2013:ECT

[1766] Boting Yang. Euclidean chains and their shortcuts. Theoretical Computer Science, 497(??):55– 67, July 29, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512002411.

Zhang:2013:CCM

[1767] Nan Zhang, Zhenhua Duan, and Cong Tian. A cylinder computation model for many-core parallel computing. Theoretical Computer Science, 497(??): 68–83, July 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001405.

Duan:2013:CPS

[1768] Zhenhua Duan, Nan Zhang, and Maciej Koutny. A complete proof system for propositional projection temporal logic. Theoretical Computer Science, 497(??):84–107, July 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200062X.

Liu:2013:AAP

[1769] Ming Liu, Feifeng Zheng, Shijin Wang, and Yinfeng Xu. Approximation algorithms for parallel machine scheduling with linear deterioration. Theoretical Computer Science, 497(??): 108–111, July 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000564.

ElShawi:2013:FQS

[1770] Radwa El Shawi and Joachim Gudmundsson. Fast query structures in anisotropic media. Theoretical Computer Science, 497(??):112– 122, July 29, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512000552.

Gao:2013:ADR

[1771] Xiaofeng Gao, Zaixin Lu, Weili Wu, and Bin Fu. Algebraic data retrieval algorithms for multi-channel wireless data broadcast. Theoretical Computer Science, 497(??):123– 130, July 29, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511010462.

Liu:2013:ORE

[1772] Qinghai Liu, Xiaohui Huang, and Zhao Zhang. Optimally restricted edge connected elementary Harary graphs. Theoretical Computer Science, 497(??): 131–138, July 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511009753.

Liu:2013:AAM

[1773] Hui Liu, Zewen Liu, Deying Li, Xianling Lu, and Hongwei Du. Approximation algorithms for minimum latency data aggregation in wireless sensor networks with directional antenna. Theoretical Computer Science, 497(??): 139–153, July 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009728.

Cheng:2013:SPA

[1774] Yukun Cheng, Wei Yu, and Guochuan Zhang. Strategy-proof approximation mechanisms for an obnoxious facility game on networks. Theoretical Computer Science, 497(??): 154–163, July 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009601.

Li:2013:NOM

[1775] Fei Li. A near-optimal memoryless online algorithm for FIFO buffering two packet classes. Theoretical Computer Science, 497(??):164– 172, July 29, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009583.

Li:2013:MEM

[1776] Deying Li, Zewen Liu, Yi Hong, Wenping Chen, and Huan Ma. Minimum energy multicast/broadcast routing with reception cost in wireless sensor networks. Theoretical Computer Science, 497(??):173– 180, July 29, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511009492.

Anonymous:2013:EBad

[1777] Anonymous. Editorial Board. Theoretical Computer Science, 497(??): iii–viii, July 29, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751300501X.

Araujo:2013:GST

[1778] João Araújo, Wolfram Bentz, and Peter J. Cameron. Groups synchronizing a transformation of non-uniform kernel. Theoretical Computer Science, 498(??):1–9, August 5, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513004829.

Devismes:2013:OPR

[1779] Stéphane Devismes, Franck Petit, and Sébastien Tixeuil. Optimal probabilistic ring exploration by semisynchronous oblivious robots. Theoretical Computer Science, 498(??): 10–27, August 5, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751300443X.

Hung:2013:DNC

[1780] Ruo-Wei Hung. DVcube: a novel compound architecture of disc-ring graph and hypercube-like graph. Theoretical Computer Science, 498(??): 28–45, August 5, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513004556.

An:2013:BDM

[1781] Min Kyung An, Nhat X. Lam, Dung T.

Huynh, and Trac N. Nguyen. Boundeddegree minimum-radius spanning trees in wireless sensor networks. Theoretical Computer Science, 498(??): 46–57, August 5, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004453.

Cui:2013:SLP

[1782] Cewei Cui, Zhe Dang, Thomas R. Fischer, and Oscar H. Ibarra. Similarity in languages and programs. Theoretical Computer Science, 498(??): 58–75, August 5, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004520.

delaHiguera:2013:PAO

[1783] Colin de la Higuera, Jean-Christophe Janodet, Émilie Samuel, Guillaume Damiand, and Christine Solnon. Polynomial algorithms for open plane graph and subgraph isomorphisms. Theoretical Computer Science, 498(??): 76–99, August 5, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003976.

Leiss:2013:WHG

[1784] Ernst L. Leiss. The worst Hanoi graphs. Theoretical Computer Science, 498(??):100–106, August 5, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S030439751300371X.

Boucher:2013:ASS

[1785] Christina Boucher, Gad M. Landau, Avivit Levy, David Pritchard, and Oren Weimann. On approximating string selection problems with outliers. Theoretical Computer Science, 498(??): 107–114, August 5, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513004490.

Liu:2013:PCC

[1786] Hong Liu and Daming Zhu. Parameterized complexity of control by voter selection in Maximin, Copeland, Borda, Bucklin, and Approval election systems. Theoretical Computer Science, 498(??):115– 123, August 5, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513003708.

Anonymous:2013:EBae

[1787] Anonymous. Editorial Board. Theoretical Computer Science, 498(??): iii–viii, August 5, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513005264.

Rozenberg:2013:Pb

[1788] G. Rozenberg.  Preface.  Theo-

retical Computer Science, 499(??): 1, August 12, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005550.

Costa:2013:AMR

[1789] José Félix Costa and Raimundo Leong. The ARNN model relativises P = NP and P ≠ NP. Theoretical Computer Science, 499(??): 2–22, August 12, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003733.

Czeizler:2013:SMT

[1790] Eugen Czeizler and Alexandru Popa.

Synthesizing minimal tile sets for complex patterns in the framework of patterned DNA self-assembly. Theoretical Computer Science, 499(??): 23–37, August 12, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003605.

Gao:2013:SCS

[1791] Yuan Gao and Lila Kari. State complexity of star of union and square of union on k regular languages. Theoretical Computer Science, 499(??): 38–50, August 12, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300457X.

Kobayashi:2013:EAC

[1792] Satoshi Kobayashi. Enumeration approach to computing chemical equilibria. Theoretical Computer Science, 499 (??):51–87, August 12, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751300368X.

Jiang:2013:USS

[1793] Keqin Jiang, Tao Song, and Linqiang Pan. Universality of sequential spiking neural P systems based on minimum spike number. Theoretical Computer Science, 499(??): 88–97, August 12, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513005094.

Zheng:2013:SST

[1794] Shenggen Zheng, Daowen Qiu, Jozef Gruska, Lvzhou Li, and Paulo Mateus. State succinctness of twoway finite automata with quantum and classical states. Theoretical Computer Science, 499(??):98– 112, August 12, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513004593.

Vaz:2013:NIS

[1795] Daniel Vaz, Luís Paquete, and Aníbal Ponte. A note on the ϵindicator subset selection. Theoretical Computer Science, 499(??):113– 116, August 12, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513003642.

Anonymous:2013:EBaf

[1796] Anonymous. Editorial Board. Theoretical Computer Science, 499(??): iii–viii, August 12, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005483.

Huang:2013:WCF

[1797] Wenqi Huang and Kun He. On the weak computability of a four dimensional orthogonal packing and time scheduling problem. Theoretical Computer Science, 501(??): 1–10, August 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005392.

Ibarra:2013:OPG

[1798] Oscar H. Ibarra and Shinnosuke Seki. On the open problem of Ginsburg concerning semilinear sets and related problems. Theoretical Computer Science, 501(??):11– 19, August 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005410.

Behle:2013:LCT

[1799] Christoph Behle, Andreas Krebs, and Mark Mercer. Linear circuits, two-variable logic and weakly blocked monoids. Theoretical Computer Science, 501(??):20–33, August 27, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513005082.

Rajasingh:2013:EIC

[1800] Indra Rajasingh, R. Bharati, Albert Muthumalai, and A. S. Shanthi. On excessive index of certain networks. Theoretical Computer Science, 501(??): 34–40, August 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513003678.

Glen:2013:TRL

[1801] Amy Glen and Jamie Simpson. The total run length of a word. Theoretical Computer Science, 501(??): 41–48, August 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513004581.

Klingler:2013:UPT

[1802] Lee Klingler, Rainer Steinwandt, and Dominique Unruh. On using probabilistic Turing machines to model participants in cryptographic protocols. Theoretical Computer Science, 501(??): 49–51, August 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513003666.

Tomasevic:2013:ACC

[1803] Violeta Tomasević and Milo Tomasević. An analysis of chain characteristics in the cryptanalytic TMTO method. Theoretical Computer Science, 501(??):52–61, August 27, 2013.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397513003691.

Tsai:2013:FII

[1804] Chang-Hsiung Tsai and Jheng-Cheng Chen. Fault isolation and identification in general biswapped networks under the PMC diagnostic model. Theoretical Computer Science, 501(??): 62–71, August 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004568.

Vakhania:2013:MML

[1805] Nodari Vakhania and Frank Werner.

Minimizing maximum lateness of jobs with naturally bounded job data on a single machine in polynomial time. Theoretical Computer Science, 501(??): 72–81, August 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004830.

vanAlten:2013:PAC

[1806] C. J. van Alten. Partial algebras and complexity of satisfiability and universal theory for distributive lattices, boolean algebras and Heyting algebras. Theoretical Computer Science, 501(??): 82–92, August 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003630.

Vitanyi:2013:CKC

[1807] Paul M. B. Vitányi. Conditional Kolmogorov complexity and universal probability. Theoretical Computer Science, 501(??):93–100, August 27, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513005124.

Apon:2013:ORM

[1808] Daniel Apon, Jonathan Katz, and Alex J. Malozemoff. One-round multi-party communication complexity of distinguishing sums. Theoretical Computer Science, 501(??): 101–108, August 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513005422.

Anonymous:2013:EBag

[1809] Anonymous. Editorial Board. Theoretical Computer Science, 501(??): iii–viii, August 27, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751300594X.

Brlek:2013:GCS

[1810] Srecko Brlek, Christophe Reutenauer, and Jean-Guy Penaud. Generation of combinatorial structures. Theoretical Computer Science, 502(??): 1–3, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751300618X.

Walsh:2013:CMD

[1811] Timothy R. Walsh. Counting maps on doughnuts. Theoretical Computer Science, 502(??):4–15, September 2, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511007146.

Brinkmann:2013:GVC

[1812] Gunnar Brinkmann, Nico Van Cleemput, and Tomaz Pisanski. Generation of various classes of trivalent graphs. Theoretical Computer Science, 502(??): 16–29, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000540.

Abdo:2013:EGI

[1813] Mohamed Abdo. Efficient generation of the ideals of a poset in Gray code order, part II. Theoretical Computer Science, 502(??):30– 45, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001272.

Sawada:2013:GCF

[1814] J. Sawada and A. Williams. A Gray code for fixed-density necklaces and Lyndon words in constant amortized time. Theoretical Computer Science, 502(??):46– 54, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512000497.

Bodini:2013:BSI

[1815] Olivier Bodini and Alice Jacquot. Boltzmann samplers for ν-balanced cycles. Theoretical Computer Science, 502(??):55–63, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397512003209.

Metivier:2013:TBC

[1816] Y. Métivier, J. M. Robson, N. Saheb-Djahromi, and A. Zemmari. On the time and the bit complexity of distributed randomised anonymous ring colouring. Theoretical Computer Science, 502(??):64– 75, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512002903.

Fortier:2013:EGG

[1817] Jérôme Fortier, Alain Goupil, Jonathan Lortie, and Jérôme Tremblay. Exhaustive generation of gominoes. Theoretical Computer Science, 502(??): 76–87, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001843.

Fedou:2013:ESP

[1818] J. M. Fedou, A. Frosini, and S. Rinaldi. Enumeration of 4-stack polyominoes. Theoretical Computer Science, 502(??): 88–97, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397511008346.

Masse:2013:CPD

[1819] A. Blondin Massé, A. Garon, and S. Labbé. Combinatorial properties of double square tiles. Theoretical Computer Science, 502(??):98– 117, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009723.

Berthe:2013:SJP

[1820] Valérie Berthé, Annie Lacasse, Geneviève Paquin, and Xavier Provençal. A study of Jacobi–Perron boundary words for the generation of discrete planes. Theoretical Computer Science, 502(??):118– 142, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512003179.

Aval:2013:HTS

[1821] Jean-Christophe Aval and Philippe Duchon. Half-turn symmetric FPLs with rare couplings and tilings of hexagons. Theoretical Computer Science, 502(??):143–152, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397512001958.

Mantaci:2013:EGP

[1822] R. Mantaci and P. Massazza. On the exhaustive generation of plane partitions. Theoretical Computer Science, 502(??):153–164, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397512010432.

Oudinet:2013:NDA

[1823] Johan Oudinet, Alain Denise, and Marie-Claude Gaudel. A new dichotomic algorithm for the uniform random generation of words in regular languages. Theoretical Computer Science, 502(??):165– 176, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007335.

Lorenz:2013:NRR

[1824] William Andrew Lorenz and Yann Ponty. Non-redundant random generation algorithms for weighted context-free grammars. Theoretical Computer Science, 502(??):177– 194, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513000443.

Aval:2013:DT

[1825] Jean-Christophe Aval, Adrien Boussicault, and Sandrine Dasse-Hartaut. Dyck tableaux. Theoretical Computer Science, 502(??):195–209, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397511009571.

Barcucci:2013:GSP

[1826] Elena Barcucci and Vincent Vajnovszki. Generalized Schröder permutations. Theoretical Computer Science, 502(??):210–216, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397512001910.

DAdderio:2013:EFS

[1827] M. D’Adderio and A. M. Garsia.

Explicit formulas for a sequence of binary tree polynomials. Theoretical Computer Science, 502(??):217– 226, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005154.

Bodini:2013:ARS

[1828] O. Bodini, D. Gardy, and A. Jacquot.

Asymptotics and random sampling for BCI and BCK lambda terms. Theoretical Computer Science, 502(??):227– 238, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000583.

Bilotta:2013:CSC

[1829] S. Bilotta, F. Disanto, R. Pinzani, and S. Rinaldi. Catalan structures and Catalan pairs. Theoretical Computer Science, 502(??):239– 248, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006524.

Anonymous:2013:EBah

[1830] Anonymous. Editorial Board. Theoretical Computer Science, 502(??): iii–viii, September 2, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006154.

Fiadeiro:2013:ITS

[1831] José Luiz Fiadeiro and Antónia Lopes. An interface theory for service-oriented design. Theoretical Computer Science, 503(??):1– 30, September 9, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003629.

Howells:2013:STC

[1832] Paul Howells and Mark d’Inverno. Specifying termination in CSP. Theoretical Computer Science, 503(??): 31–60, September 9, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003599.

Konur:2013:CMC

[1833] Savas Konur, Michael Fisher, and Sven Schewe. Combined model checking for temporal, probabilistic, and real-time logics. Theoretical Computer Science, 503(??):61– 88, September 9, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300515X.

Chen:2013:PBO

[1834] Wei Chen, Guangda Hu, and Jialin Zhang. On the power of breakable objects. Theoretical Computer Science, 503(??):89–108, September 9, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397513004489.

Krishna:2013:QCZ

[1835] Siddharth Krishna and Anca Muscholl.

A quadratic construction for Zielonka automata with acyclic communication structure. Theoretical Computer Science, 503(??):109–114, September 9, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397513005185.

Genaim:2013:CSB

[1836] Samir Genaim and Damiano Zanardini. Corrigendum to “Reachabilitybased acyclicity analysis by abstract interpretation” [Theoretical Computer Science 474 (2013) 60–79]. Theoretical Computer Science, 503(??): 115, September 9, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005628.

Anonymous:2013:EBai

[1837] Anonymous. Editorial Board. Theoretical Computer Science, 503(??): iii–viii, September 9, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006294.

Ehmsen:2013:BBO

[1838] Martin R. Ehmsen and Kim S. Larsen.

Better bounds on online unit clustering. Theoretical Computer Science, 500 (??):1–24, August 19, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005112.

Meeks:2013:CFF

[1839] Kitty Meeks and Alexander Scott. The complexity of Free-Flood-It on 2 × n boards. Theoretical Computer Science, 500(??):25–43, August 19, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004647.

Bartal:2013:BLD

[1840] Yair Bartal, Douglas E. Carroll, Adam Meyerson, and Ofer Neiman. Bandwidth and low dimensional embedding. Theoretical Computer Science, 500(??): 44–56, August 19, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004507.

Peltomaki:2013:IPW

[1841] Jarkko Peltomäki. Introducing privileged words: Privileged complexity of Sturmian words. Theoretical Computer Science, 500(??):57– 67, August 19, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004027.

Csaba:2013:ORM

[1842] Béla Csaba, Thomas A. Plick, and Ali Shokoufandeh. Optimal random matchings, tours, and spanning trees in hierarchically separated trees. Theoretical Computer Science, 500(??):

68–89, August 19, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003721.

Blin:2013:SSI

[1843] Lélia Blin, Maria Potop-Butucaru, and Stephane Rovedakis. A superstabilizing log(n)-approximation algorithm for dynamic Steiner trees. Theoretical Computer Science, 500(??): 90–112, August 19, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004969.

Anonymous:2013:EBaj

[1844] Anonymous. Editorial Board. Theoretical Computer Science, 500(??): iii–viii, August 19, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005689.

Cosnard:2013:CBE

[1845] Michel Cosnard. Celebrating the 60th birthday of Eric Goles. Theoretical Computer Science, 504(??): 1–2, September 16, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006592.

Calude:2013:DMS

[1846] Cristian Calude, Bruno Durand, Anahí Gajardo, Dominique Perrin, Ivan Rapaport, and Sergio Rica. Discrete mathematical structures: From dynamics to complexity. Theoretical Computer Science, 504(??):3– 4, September 16, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006580.

Bermond:2013:DAG

[1847] Jean-Claude Bermond, Michel Cosnard, and Stéphane Pérennes. Directed acyclic graphs with the unique dipath property. Theoretical Computer Science, 504(??):5–11, September 16, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005841.

Noual:2013:ANM

[1848] Mathilde Noual, Damien Regnault, and Sylvain Sené. About non-monotony in Boolean automata networks. Theoretical Computer Science, 504(??):12– 25, September 16, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512005282.

Macauley:2013:ALS

[1849] M. Macauley and H. S. Mortveit. An atlas of limit set dynamics for asynchronous elementary cellular automata. Theoretical Computer Science, 504(??): 26–37, September 16, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008651.

Dennunzio:2013:LRD

[1850] Alberto Dennunzio, Enrico Formenti, and Julien Provillard. Local rule distributions, language complexity and non-uniform cellular automata. Theoretical Computer Science, 504(??):38– 51, September 16, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004458.

Perrot:2013:KSP

[1851] Kévin Perrot and Éric Rémila.

Kadanoff sand pile model. Avalanche structure and wave shape. Theoretical Computer Science, 504(??):52– 72, September 16, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001102.

Goles:2013:CBP

[1852] Eric Goles, Pedro Montealegre-Barba, and Ioan Todinca. The complexity of the bootstraping percolation and other problems. Theoretical Computer Science, 504(??):73– 82, September 16, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007402.

Martin:2013:GUM

[1853] Bruno Martin. On Goles’ universal machines: a computational point of view. Theoretical Computer Science, 504(??): 83–88, September 16, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513003320.

Anonymous:2013:EBak

[1854] Anonymous. Editorial Board. Theoretical Computer Science, 504(??):iii– viii, September 16, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513006518.

Ogihara:2013:TAM

[1855] Mitusnori Ogihara and Jun Tarui. Theory and applications of models of computation 2011. Theoretical Computer Science, 505(??): 1, September 23, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513006725.

Althaus:2013:LDS

[1856] Ernst Althaus, Joschka Kupilas, and Rouven Naujoks. On the lowdimensional Steiner minimum tree problem in Hamming metric. Theoretical Computer Science, 505(??):2– 10, September 23, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513001205.

Chin:2013:NAC

[1857] Francis Y. L. Chin, Henry C. M. Leung, and S. M. Yiu. Nonadaptive complex group testing with multiple positive sets. Theoretical Computer Science, 505(??):11– 18, September 23, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002855.

Brunsch:2013:BII

[1858] Tobias Brunsch and Heiko Röglin. A bad instance for k-means++. Theoretical Computer Science, 505(??):19– 26, September 23, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001806.

Liskiewicz:2013:GBS

[1859] Maciej Liśkiewicz, Rüdiger Reischuk, and Ulrich Wölfel. Greybox steganography. Theoretical Computer Science, 505(??):27–41, September 23, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200309X.

Junosza-Szaniawski:2013:FEA

[1860] Konstanty Junosza-Szaniawski, Jan Kratochvíl, Mathieu Liedloff, Peter Rossmanith, and Pawel Rzazewski. Fast exact algorithm for L(2, 1)- labeling of graphs. Theoretical Computer Science, 505(??):42–54, September 23, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006548.

Meier:2013:GSD

[1861] Arne Meier and Thomas Schneider. Generalized satisfiability for the description logic ALC. Theoretical Computer Science, 505(??):55– 73, September 23, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513001187.

Suzuki:2013:EFL

[1862] Akira Suzuki, Kei Uchizawa, and Xiao Zhou. Energy and fan-in of logic circuits computing symmetric Boolean functions. Theoretical Computer Science, 505(??):74– 80, September 23, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512010882.

Jeandel:2013:TDM

[1863] Emmanuel Jeandel and Pascal Vanier. Turing degrees of multidimensional SFTs. Theoretical Computer Science, 505(??):81–92, September 23, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397512008031.

Anonymous:2013:EBal

[1864] Anonymous. Editorial Board. Theoretical Computer Science, 505(??):iii– viii, September 23, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513006658.

Bera:2013:CDD

[1865] Sahadev Bera, Partha Bhowmick, Peer Stelldinger, and Bhargab B. Bhattacharya. On covering a digital disc with concentric circles in Z^{2}. Theoretical Computer Science, 506(??):1– 16, September 30, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005793.

Blanchet-Sadri:2013:SBP

[1866] F. Blanchet-Sadri and Brent Woodhouse. Strict bounds for pattern avoidance. Theoretical Computer Science, 506(??):17–28, September 30, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006075.

Alcon:2013:SCG

[1867] Liliana Alcón, Luerbio Faria, Celina M. H. de Figueiredo, and Marisa Gutierrez. Split clique graph complexity. Theoretical Computer Science, 506(??):29–42, September 30, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397513005367.

Dieudonne:2013:DGE

[1868] Yoann Dieudonné, Florence Levé, Franck Petit, and Vincent Villain. Deterministic geoleader election in disoriented anonymous systems. Theoretical Computer Science, 506(??):43– 54, September 30, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300563X.

Duh:2013:TSD

[1869] Dyi-Rong Duh, Yao-Chung Lin, Cheng-Nan Lai, and Yue-Li Wang.

Two spanning disjoint paths with required length in generalized hypercubes. Theoretical Computer Science, 506(??):55–78, September 30, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397513005823.

Ceccherini-Silberstein:2013:CAB

[1870] Tullio Ceccherini-Silberstein, Michel Coornaert, Francesca Fiorenzi, and Zoran Sunić. Cellular automata between sofic tree shifts. Theoretical Computer Science, 506(??):79– 101, September 30, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005100.

Flouri:2013:ESC

[1871] Tomás Flouri, Costas S. Iliopoulos, Tomasz Kociumaka, Solon P. Pissis, Simon J. Puglisi, W. F. Smyth, and Wojciech Tyczyński. Enhanced string covering. Theoretical Computer Science, 506(??):102– 114, September 30, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006105.

Misra:2013:SMO

[1872] Neeldhara Misra, N. S. Narayanaswamy, Venkatesh Raman, and Bal Sri Shankar. Solving min ones 2-sat as fast as vertex cover. Theoretical Computer Science, 506(??):115– 121, September 30, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005355.

Anonymous:2013:EBam

[1873] Anonymous. Editorial Board. Theoretical Computer Science, 506(??):iii– viii, September 30, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006841.

Lin:2013:P

[1874] Guohui Lin. Preface. Theoretical Computer Science, 507(??):1, October 7, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007032.

He:2013:AAL

[1875] Jing (Selena) He, Shouling Ji, Yi Pan, and Zhipeng Cai. Approximation algorithms for load-balanced virtual backbone construction in wireless sensor networks. Theoretical Computer Science, 507(??):2–16, October 7, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751201050X.

Chen:2013:RPA

[1876] Xujin Chen, Xiaodong Hu, Weidong Ma, and Changjun Wang. Reducing price of anarchy of selfish task allocation with more selfishness. Theoretical Computer Science, 507(??): 17–33, October 7, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001515.

Zhang:2013:MTW

[1877] Qiang Zhang, Weiwei Wu, and Minming Li. Minimizing the total weighted completion time of fully parallel jobs with integer parallel units. Theoretical Computer Science, 507(??): 34–40, October 7, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001266.

Jiang:2013:FVS

[1878] Wei Jiang, Tian Liu, Chaoyi Wang, and Ke Xu. Feedback vertex sets on restricted bipartite graphs. Theoretical Computer Science, 507(??): 41–51, October 7, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512011243.

Kosowski:2013:MMM

[1879] Adrian Kosowski, Alfredo Navarra, Dominik Pajak, and Cristina M. Pinotti. Maximum matching in multi-interface networks. Theoretical Computer Science, 507(??):52– 60, October 7, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000686.

Paquete:2013:BSP

[1880] Luís Paquete, Mathias Jaschob, Kathrin Klamroth, and Jochen Gorski. On a biobjective search problem in a line: Formulations and algorithms. Theoretical Computer Science, 507(??): 61–71, October 7, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001254.

Poon:2013:SMS

[1881] Chung Keung Poon and Binhai Zhu. Streaming with minimum space: an algorithm for covering by two congruent balls. Theoretical Computer Science, 507(??):72– 82, October 7, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001138.

Zhu:2013:CPP

[1882] Yongding Zhu and Jinhui Xu. On the central path problem. Theoretical Computer Science, 507(??): 83–99, October 7, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005136.

Yang:2013:FMS

[1883] Boting Yang. Fast-mixed searching and related problems on graphs. Theoretical Computer Science, 507(??): 100–113, October 7, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002995.

Anonymous:2013:EBan

[1884] Anonymous. Editorial Board. Theoretical Computer Science, 507(??): iii–viii, October 7, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513007019.

Zhu:2013:P

[1885] Binhai Zhu and Jun Luo. Preface. Theoretical Computer Science, 508 (??):1, October 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513007317.

Xiao:2013:PED

[1886] Mingyu Xiao and Hiroshi Nagamochi. Parameterized edge dominating set in graphs with degree bounded by 3. Theoretical Computer Science, 508(??): 2–15, October 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512007888.

Wang:2013:PCM

[1887] Jianxin Wang, Weizhong Luo, Qilong Feng, and Jiong Guo. Parameterized complexity of Min-power multicast problems in wireless ad hoc networks. Theoretical Computer Science, 508(??): 16–25, October 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001922.

Zhang:2013:GRE

[1888] Huaming Zhang and Swetha Govindaiah. Greedy routing via embedding graphs onto semi-metric spaces. Theoretical Computer Science, 508(??):

26–34, October 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001065.

Lan:2013:KPS

[1889] Yan Lan, György Dósa, Xin Han, Chenyang Zhou, and Attila Benko. 2D knapsack: Packing squares. Theoretical Computer Science, 508(??): 35–40, October 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200744X.

Chen:2013:AIS

[1890] Danny Z. Chen and Ewa Misiolek. Algorithms for interval structures with applications. Theoretical Computer Science, 508(??):41– 53, October 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397511010607.

Birks:2013:TAO

[1891] Martin Birks and Stanley P. Y. Fung.

Temperature aware online algorithms for scheduling equal length jobs. Theoretical Computer Science, 508(??): 54–65, October 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512001223.

Tan:2013:AAC

[1892] Xuehou Tan and Gangshan Wu. Approximation algorithms for cutting a convex polyhedron out of a sphere.

Theoretical Computer Science, 508(??): 66–73, October 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512003076.

Mukherjee:2013:MWR

[1893] Joydeep Mukherjee, Priya Ranjan Sinha Mahapatra, Arindam Karmakar, and Sandip Das. Minimumwidth rectangular annulus. Theoretical Computer Science, 508(??): 74–80, October 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512001934.

Anonymous:2013:EBao

[1894] Anonymous. Editorial Board. Theoretical Computer Science, 508(??): iii–viii, October 14, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513007299.

Kosowski:2013:P

[1895] Adrian Kosowski and Masafumi Yamashita. Preface. Theoretical Computer Science, 509(??):1–2, October 21, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513007482.

Herlihy:2013:PLD

[1896] Maurice Herlihy, Sergio Rajsbaum, and Michel Raynal. Power and limits of distributed computing shared memory models. Theoretical Computer Science, 509(??):3–24, October 21, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001813.

Nowak:2013:PRB

[1897] Thomas Nowak, Matthias Függer, and Alexander Kößler. On the performance of a retransmission-based synchronizer. Theoretical Computer Science, 509(??): 25–39, October 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004008.

Schneider:2013:SBD

[1898] Johannes Schneider, Michael Elkin, and Roger Wattenhofer. Symmetry breaking depending on the chromatic number or the neighborhood growth. Theoretical Computer Science, 509(??): 40–50, October 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008249.

Cooper:2013:CTR

[1899] Colin Cooper, Alan Frieze, and Tomasz Radzik. The cover times of random walks on random uniform hypergraphs. Theoretical Computer Science, 509(??): 51–69, October 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000820.

Chalopin:2013:TBB

[1900] Jérémie Chalopin, Shantanu Das, Arnaud Labourel, and Euripides Markou. Tight bounds for black hole search with scattered agents in synchronous rings. Theoretical Computer Science, 509(??): 70–85, October 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513001199.

Guilbault:2013:GAO

[1901] Samuel Guilbault and Andrzej Pelc. Gathering asynchronous oblivious agents with local vision in regular bipartite graphs. Theoretical Computer Science, 509(??):86– 96, October 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512006494.

Brandes:2013:EES

[1902] Philipp Brandes, Bastian Degener, Barbara Kempkes, and Friedhelm Meyer auf der Heide. Energy-efficient strategies for building short chains of mobile robots locally. Theoretical Computer Science, 509(??):97– 112, October 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512009887.

Brejova:2013:ERC

[1903] Brona Brejová, Stefan Dobrev, Rastislav Královic, and Tomás Vinar. Efficient routing in carrier-based mobile networks. Theoretical Computer Science, 509(??):113–121, October 21, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300087X.

Bermond:2013:OTD

[1904] Jean-Claude Bermond, Luisa Gargano, Stephane Perénnes, Adele A. Rescigno, and Ugo Vaccaro. Optimal time data gathering in wireless networks with multidirectional antennas. Theoretical Computer Science, 509(??):122– 139, October 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002302.

Anonymous:2013:EBap

[1905] Anonymous. Editorial Board. Theoretical Computer Science, 509(??): iii–viii, October 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007445.

Bonsma:2013:CRS

[1906] Paul Bonsma. The complexity of rerouting shortest paths. Theoretical Computer Science, 510(??):1– 12, October 28, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006920.

Dosa:2013:TAB

[1907] György Dósa, Rongheng Li, Xin Han, and Zsolt Tuza. Tight absolute bound for First Fit Decreasing bin-packing:

FFD(L) ≤ (11/9)OPT(L) + 6/9. Theoretical Computer Science, 510(??): 13–61, October 28, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513006774.

DePrisco:2013:CVC

[1908] Roberto De Prisco and Alfredo De Santis. Color visual cryptography schemes for black and white secret images. Theoretical Computer Science, 510(??): 62–86, October 28, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513006750.

Eom:2013:SCC

[1909] Hae-Sung Eom and Yo-Sub Han. State complexity of combined operations for suffix-free regular languages. Theoretical Computer Science, 510(??): 87–93, October 28, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513006944.

Cheng:2013:CEF

[1910] Dongqin Cheng, Rong-Xia Hao, and Yan-Quan Feng. Conditional edgefault pancyclicity of augmented cubes. Theoretical Computer Science, 510(??): 94–101, October 28, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513006907.

Dosa:2013:GSM

[1911] György Dósa and Csanád Imreh.

The generalization of scheduling with machine cost. Theoretical Computer Science, 510(??):102–110, October 28, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006798.

Boros:2013:DMC

[1912] Endre Boros, Ondrej Cepek, and Petr Kucera. A decomposition method for CNF minimality proofs. Theoretical Computer Science, 510(??): 111–126, October 28, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006968.

Dourado:2013:CNI

[1913] Mitre C. Dourado, Dieter Rautenbach, Vinícius Fernandes dos Santos, Philipp M. Schäfer, and Jayme L. Szwarcfiter. On the Carathéodory number of interval and graph convexities. Theoretical Computer Science, 510(??):127– 135, October 28, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006749.

Anonymous:2013:EBaq

[1914] Anonymous. Editorial Board. Theoretical Computer Science, 510(??): iii–viii, October 28, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513007743.

Paschos:2013:P

[1915] Vangelis Th. Paschos. Preface. Theoretical Computer Science, 511(??): 1, November 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513007822.

Luo:2013:ILP

[1916] Weizhong Luo, Jianxin Wang, Qilong Feng, Jiong Guo, and Jianer Chen. Improved linear problem kernel for planar connected dominating set. Theoretical Computer Science, 511(??): 2–12, November 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513004659.

Lente:2013:ESS

[1917] Ch. Lenté, M. Liedloff, A. Soukhal, and V. T’Kindt. On an extension of the Sort & Search method with application to scheduling theory. Theoretical Computer Science, 511(??): 13–22, November 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513003848.

Furer:2013:ETA

[1918] Martin Fürer, Serge Gaspers, and Shiva Prasad Kasiviswanathan. An exponential time 2-approximation algorithm for bandwidth. Theoretical Computer Science, 511(??):23– 31, November 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300248X.

Bougeret:2013:MEA

[1919] Marin Bougeret, Pierre-François Dutot, and Denis Trystram. Moderately exponential approximation for makespan minimization on related machines. Theoretical Computer Science, 511(??):32–41, November 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397513002338.

Lampis:2013:PMP

[1920] Michael Lampis. Parameterized maximum path coloring. Theoretical Computer Science, 511(??):42– 53, November 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000625.

Belmonte:2013:GCS

[1921] Rémy Belmonte and Martin Vatshelle.

Graph classes with structured neighborhoods and algorithmic applications. Theoretical Computer Science, 511(??): 54–65, November 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000613.

Bui-Xuan:2013:FDP

[1922] Binh-Minh Bui-Xuan, Jan Arne Telle, and Martin Vatshelle. Fast dynamic programming for locally checkable vertex subset and vertex partitioning problems. Theoretical Computer Science, 511(??):66– 76, November 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000595.

Crowston:2013:PCM

[1923] Robert Crowston, Gregory Gutin, Mark Jones, Venkatesh Raman, and Saket Saurabh. Parameterized complexity of MaxSat Above Average. Theoretical Computer Science, 511(??): 77–84, November 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000431.

Brankovic:2013:NPA

[1924] Ljiljana Brankovic and Henning Fernau. A novel parameterised approximation algorithm for minimum vertex cover. Theoretical Computer Science, 511(??):85–108, November 4, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512010821.

Kim:2013:EAP

[1925] Eun Jung Kim and Daniel Gonçalves. On exact algorithms for the permutation CSP. Theoretical Computer Science, 511(??):109–116, November 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397512009565.

Bodlaender:2013:KBP

[1926] Hans L. Bodlaender, Bart M. P.

Jansen, and Stefan Kratsch. Kernel bounds for path and cycle problems. Theoretical Computer Science, 511(??): 117–136, November 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008262.

Damaschke:2013:SSS

[1927] Peter Damaschke. Sparse solutions of sparse linear systems: Fixed-parameter tractability and an application of complex group testing. Theoretical Computer Science, 511(??):137– 146, November 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006457.

Xiao:2013:NPA

[1928] Mingyu Xiao, Ton Kloks, and Sheung-Hung Poon. New parameterized algorithms for the edge dominating set problem. Theoretical Computer Science, 511(??):147–158, November 4, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006044.

Dondi:2013:IDP

[1929] Riccardo Dondi, Giancarlo Mauri, and Italo Zoppis. The l-Diversity problem: Tractability and approximability. Theoretical Computer Science, 511(??): 159–171, November 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512004823.

Heggernes:2013:PCV

[1930] Pinar Heggernes, Pim van ’t Hof, Bart M. P. Jansen, Stefan Kratsch, and Yngve Villanger. Parameterized complexity of vertex deletion into perfect graph classes. Theoretical Computer Science, 511(??):172– 180, November 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397512002332.

Anonymous:2013:EBar

[1931] Anonymous. Editorial Board. Theoretical Computer Science, 511(??): iii–viii, November 4, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513007871.

Defago:2013:P

[1932] Xavier Défago, Franck Petit, and Vincent Villain. Preface. Theoretical Computer Science, 512(??): 1, November 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513008074.

Berns:2013:BSS

[1933] Andrew Berns, Sukumar Ghosh, and Sriram V. Pemmaraju. Building selfstabilizing overlay networks with the transitive closure framework. Theoretical Computer Science, 512(??): 2–14, November 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513001539.

Beauquier:2013:TCA

[1934] J. Beauquier, P. Blanchard, J. Burman, and S. Delaët. Tight complexity analysis of population protocols with cover times — The ZebraNet example. Theoretical Computer Science, 512(??): 15–27, November 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200953X.

Baldoni:2013:PIB

[1935] Roberto Baldoni, Silvia Bonomi, and Amir Soltani Nezhad. A protocol for implementing Byzantine storage in churn-prone distributed systems. Theoretical Computer Science, 512(??): 28–40, November 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002673.

Attiya:2013:NTP

[1936] Hagit Attiya and Armando Castañeda.

A non-topological proof for the impossibility of k-set agreement. Theoretical Computer Science, 512(??): 41–48, November 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008626.

Datta:2013:SSL

[1937] Ajoy K. Datta, Stéphane Devismes, Lawrence L. Larmore, and Yvan Rivierre. Self-stabilizing labeling and ranking in ordered trees. Theoretical Computer Science, 512(??):49– 66, November 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300604X.

Frey:2013:TAP

[1938] Davide Frey, Arnaud Jégou, Anne-Marie Kermarrec, Michel Raynal, and Julien Stainer. Trust-aware peer sampling: Performance and privacy tradeoffs. Theoretical Computer Science, 512(??):67–83, November 11, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000856.

Klappenecker:2013:DRR

[1939] Andreas Klappenecker, Hyunyoung Lee, and Jennifer L. Welch. Dynamic regular registers in systems with churn. Theoretical Computer Science, 512(??): 84–97, November 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512007931.

Chatzigiannakis:2013:CPS

[1940] Ioannis Chatzigiannakis, Othon Michail, Stavros Nikolaou, and Paul G. Spirakis. The computational power of simple protocols for self-awareness on graphs. Theoretical Computer Science, 512(??): 98–118, November 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751200802X.

Nor:2013:CSD

[1941] Rizal Mohd Nor, Mikhail Nesterenko, and Christian Scheideler. Corona: a stabilizing deterministic messagepassing skip list. Theoretical Computer Science, 512(??):119–129, November 11, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008055.

Imbs:2013:WFD

[1942] Damien Imbs and Michel Raynal. The weakest failure detector to implement a register in asynchronous systems with hybrid communication. Theoretical Computer Science, 512(??):130– 142, November 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512006408.

Anonymous:2013:EBas

[1943] Anonymous. Editorial Board. Theoretical Computer Science, 512(??):iii– viii, November 11, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008128.

Jo:2013:PMM

[1944] Shinhaeng Jo, Jung-Heum Park, and Kyung-Yong Chwa. Paired many-to-many disjoint path covers in faulty hypercubes. Theoretical Computer Science, 513(??):1– 24, November 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513007603.

Dash:2013:SAC

[1945] Santanu Kumar Dash, Sven-Bodo Scholz, Stephan Herhut, and Bruce Christianson. A scalable approach to computing representative lowest common ancestor in directed acyclic graphs. Theoretical Computer Science, 513(??):25–37, November 18, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513007226.

Farzan:2013:SEA

[1946] Arash Farzan and J. Ian Munro. Succinct encoding of arbitrary graphs. Theoretical Computer Science, 513(??): 38–52, November 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513007238.

Crowston:2013:MBS

[1947] R. Crowston, G. Gutin, M. Jones, and G. Muciaccia. Maximum balanced subgraph problem parameterized above lower bound. Theoretical Computer Science, 513(??):53– 64, November 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513008062.

Eades:2013:LTA

[1948] Peter Eades, Seok-Hee Hong, Naoki Katoh, Giuseppe Liotta, Pascal Schweitzer, and Yusuke Suzuki. A linear time algorithm for testing maximal 1-planarity of graphs with a rotation system. Theoretical Computer Science, 513(??):65– 76, November 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007214.

Krysztowiak:2013:IAR

[1949] Przemyslaw Krysztowiak. An improved approximation ratio for the jump number problem on interval orders. Theoretical Computer Science, 513(??): 77–84, November 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007639.

Liu:2013:RBP

[1950] Jian Liu and Lusheng Chen. On the relationships between perfect nonlinear functions and universal hash families. Theoretical Computer Science, 513(??): 85–95, November 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300769X.

Monti:2013:DWR

[1951] Angelo Monti. Deciding the winner in k rounds for DISJOINT ARROWS, a new combinatorial partizan game. Theoretical Computer Science, 513(??): 96–108, November 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513007561.

Barbay:2013:CPA

[1952] Jérémy Barbay and Gonzalo Navarro. On compressing permutations and adaptive sorting. Theoretical Computer Science, 513(??):109–123, November 18, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513007962.

Gutin:2013:PCC

[1953] Gregory Gutin, Gabriele Muciaccia, and Anders Yeo. Parameterized complexity of k-Chinese Postman Problem. Theoretical Computer Science, 513(??): 124–128, November 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513007640.

Linz:2013:OTC

[1954] Simone Linz, Katherine St. John, and Charles Semple. Optimizing tree and character compatibility across several phylogenetic trees. Theoretical Computer Science, 513(??):129– 136, November 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513007676.

Anonymous:2013:EBat

[1955] Anonymous. Editorial Board. Theoretical Computer Science, 513(??):iii– viii, November 18, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008323.

Demetrescu:2013:P

[1956] Camil Demetrescu, Stefano Leonardi, and Alberto Marchetti-Spaccamela. Preface. Theoretical Computer Science, 514(??):1, November 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300858X.

Angelini:2013:TMP

[1957] Patrizio Angelini, Pier Francesco Cortese, Giuseppe Di Battista, and Maurizio Patrignani. Topological morphing of planar graphs. Theoretical Computer Science, 514(??):2– 20, November 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006427.

Bilo:2013:SCC

[1958] Vittorio Bilò, Alessandro Celi, Michele Flammini, and Vasco Gallotti. Social context congestion games. Theoretical Computer Science, 514(??): 21–35, November 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009735.

Bilardi:2013:OEP

[1959] Gianfranco Bilardi and Francesco Versaci. Optimal eviction policies for stochastic address traces. Theoretical Computer Science, 514(??):36– 60, November 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513000662.

Boria:2013:RMW

[1960] Nicolas Boria, Jérôme Monnot, and Vangelis Th. Paschos. Reoptimization of maximum weight induced hereditary subgraph problems. Theoretical Computer Science, 514(??):61– 74, November 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009668.

Chrobak:2013:BBI

[1961] Marek Chrobak, Lukasz Jez, and Jirí Sgall. Better bounds for incremental frequency allocation in bipartite graphs. Theoretical Computer Science, 514(??):75–83, November 25, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512004781.

Crescenzi:2013:CDR

[1962] Pilu Crescenzi, Roberto Grossi, Michel Habib, Leonardo Lanzi, and Andrea Marino. On computing the diameter of real-world undirected graphs. Theoretical Computer Science, 514(??): 84–95, November 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512008687.

Eppstein:2013:CBR

[1963] David Eppstein, Michael T. Goodrich, Maarten Löffler, Darren Strash, and Lowell Trott. Category-based routing in social networks: Membership dimension and the small-world phenomenon. Theoretical Computer Science, 514(??): 96–104, November 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003356.

Nutov:2013:SNA

[1964] Zeev Nutov. Survivable network activation problems. Theoretical Computer Science, 514(??):105–115, November 25, 2013. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397512009863.

Anonymous:2013:EBau

[1965] Anonymous. Editorial Board. Theoretical Computer Science, 514(??):iii– viii, November 25, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008517.

Tkachev:2014:CCI

[1966] Ilya Tkachev and Alessandro Abate.

Characterization and computation of infinite-horizon specifications over Markov processes. Theoretical Computer Science, 515(??):1–18, January 2, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300724X.

Burrell:2014:SRR

[1967] Mike Burrell, Robin Cockett, and Brian Redmond. Safe recursion revisited I: Categorical semantics for lower complexity. Theoretical Computer Science, 515(??):19– 45, January 2, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513007536.

Faella:2014:ATD

[1968] Marco Faella, Salvatore La Torre, and Aniello Murano. Automatatheoretic decision of timed games. Theoretical Computer Science, 515(??): 46–63, January 2, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513006452.

Ma:2014:EID

[1969] Shuai Ma, Wenfei Fan, and Loreto Bravo. Extending inclusion dependencies with conditions. Theoretical Computer Science, 515(??):64– 95, January 2, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513008207.

Larsen:2014:RSR

[1970] Kim G. Larsen, Axel Legay, Louis-Marie Traonouez, and Andrzej Wasowski. Robust synthesis for realtime systems. Theoretical Computer Science, 515(??):96–122, January 2, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006397.

Anonymous:2014:EBa

[1971] Anonymous. Editorial Board. Theoretical Computer Science, 515(??): iii–viii, January 2, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008852.

DAngelo:2014:LFS

[1972] Gianlorenzo D’Angelo, Mattia D’Emidio, and Daniele Frigioni. A loopfree shortest-path routing algorithm for dynamic networks. Theoretical Computer Science, 516(??):1–19, January 9, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008190.

Duchene:2014:VNP

[1973] Eric Duchêne and Gabriel Renault.

Vertex Nim played on graphs. Theoretical Computer Science, 516(??): 20–27, January 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008694.

Duh:2014:IEF

[1974] Dyi-Rong Duh, Tzu-Lung Chen, and Yue-Li Wang. n − 3-edge-fault-tolerant weak-pancyclicity of (n, k)-star graphs. Theoretical Computer Science, 516(??): 28–39, January 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300844X.

Dennunzio:2014:MCA

[1975] Alberto Dennunzio, Enrico Formenti, and Michael Weiss. Multidimensional cellular automata: closing property, quasi-expansivity, and (un)decidability issues. Theoretical Computer Science, 516(??):40– 59, January 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008232.

Gast:2014:AVC

[1976] Mikael Gast and Mathias Hauptmann. Approximability of the vertex cover problem in power-law graphs. Theoretical Computer Science, 516(??): 60–70, January 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008220.

Abbes:2014:ETI

[1977] F. Abbes, M. Hbaib, and M. Merkhi. β-expansion and transcendence in F_{q}((x^{−1})). Theoretical Computer Science, 516(??):71–77, January 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S030439751300861X.

Karthick:2014:ASI

[1978] T. Karthick. On atomic structure of P_{5}-free subclasses and Maximum Weight Independent Set problem. Theoretical Computer Science, 516(??):

78–85, January 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008633.

Kowalik:2014:KNI

[1979] Lukasz Kowalik and Marcin Mucha.

A 9k kernel for nonseparating independent set in planar graphs. Theoretical Computer Science, 516(??): 86–95, January 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008657.

Maffray:2014:FRD

[1980] Frédéric Maffray. Fast recognition of doubled graphs. Theoretical Computer Science, 516(??):96– 100, January 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008645.

Okhotin:2014:PMM

[1981] Alexander Okhotin. Parsing by matrix multiplication generalized to Boolean grammars. Theoretical Computer Science, 516(??):101–120, January 9, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006919.

Levin:2014:ASB

[1982] Asaf Levin and Aleksander Vainer.

Adaptivity in the stochastic blackjack knapsack problem. Theoretical Computer Science, 516(??):121– 126, January 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513008736.

Meduna:2014:OSR

[1983] Alexander Meduna and Petr Zemek. One-sided random context grammars with a limited number of right random context rules. Theoretical Computer Science, 516(??): 127–132, January 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751300827X.

Ni:2014:ECP

[1984] Wei-Ping Ni and Jian-Liang Wu. Edge coloring of planar graphs which any two short cycles are adjacent at most once. Theoretical Computer Science, 516(??): 133–138, January 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513008426.

Anonymous:2014:EBb

[1985] Anonymous. Editorial Board. Theoretical Computer Science, 516(??): iii–viii, January 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513008992.

Tschaikowski:2014:TCS

[1986] Max Tschaikowski and Mirco Tribastone. Tackling continuous state-space explosion in a Markovian process algebra. Theoretical Computer Science, 517 (??):1–33, January 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006403.

Rowe:2014:STA

[1987] Reuben N. S. Rowe and S. J. van Bakel. Semantic types and approximation for Featherweight Java. Theoretical Computer Science, 517(??): 34–74, January 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006415.

Honda:2014:OCP

[1988] Kohei Honda, Nobuko Yoshida, and Martin Berger. An observationally complete program logic for imperative higher-order functions. Theoretical Computer Science, 517(??): 75–101, January 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008219.

Tutu:2014:PAS

[1989] Ionut Tutu. Parameterisation for abstract structured specifications. Theoretical Computer Science, 517(??): 102–142, January 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008268.

Anonymous:2014:EBc

[1990] Anonymous. Editorial Board. Theoretical Computer Science, 517(??): iii–viii, January 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513009146.

Wang:2014:TCE

[1991] Huijuan Wang, Bin Liu, Jianliang Wu, and Guizhen Liu. Total coloring of embedded graphs with maximum degree at least seven. Theoretical Computer Science, 518(??): 1–9, January 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513003563.

Zhang:2014:CCT

[1992] Yong Zhang, Francis Y. L. Chin, and Hing-Fung Ting. Constantcompetitive tree node assignment. Theoretical Computer Science, 518(??): 10–21, January 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513003654.

Wang:2014:SRP

[1993] Weifan Wang, Stephen Finbow, and Jiangxu Kong. The 2-surviving rate of planar graphs without 6-cycles. Theoretical Computer Science, 518(??): 22–31, January 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513003861.

Luo:2014:PCM

[1994] Weizhong Luo, Jianxin Wang, Jiong Guo, and Jianer Chen. Parameterized complexity of Max-lifetime Target Coverage in wireless sensor networks. Theoretical Computer Science, 518(??): 32–41, January 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004623.

Pelantova:2014:PRL

[1995] Edita Pelantová and Stepán Starosta.

Palindromic richness for languages invariant under more symmetries. Theoretical Computer Science, 518(??): 42–63, January 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005379.

Ventre:2014:TOU

[1996] Carmine Ventre. Truthful optimization using mechanisms with verification. Theoretical Computer Science, 518(??): 64–79, January 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005768.

Reidenbach:2014:RCF

[1997] Daniel Reidenbach and Markus L.

Schmid. Regular and context-free pattern languages over small alphabets. Theoretical Computer Science, 518(??): 80–95, January 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300577X.

Olsen:2014:ALB

[1998] Martin Olsen and Anastasios Viglas. On the approximability of the link building problem. Theoretical Computer Science, 518(??):96– 116, January 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005811.

Wan:2014:PSB

[1999] Long Wan, Ran Ma, and Jinjiang Yuan. Primary-secondary bicriteria scheduling on identical machines to minimize the total completion time of all jobs and the maximum T - time of all machines. Theoretical Computer Science, 518(??):117– 123, January 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004465.

Wu:2014:ECP

[2000] Jian-Liang Wu and Ling Xue. Edge colorings of planar graphs without 5cycles with two chords. Theoretical Computer Science, 518(??):124– 127, January 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005434.

Ruggieri:2014:CQL

[2001] Salvatore Ruggieri, Pavlos Eirinakis, K. Subramani, and Piotr Wojciechowski. On the complexity of quantified linear systems. Theoretical Computer Science, 518(??):128– 134, January 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005781.

Anonymous:2014:EBd

[2002] Anonymous. Editorial Board. Theoretical Computer Science, 518(??): iii–viii, January 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513009377.

Kivinen:2014:GEI

[2003] Jyrki Kivinen, Csaba Szepesvári, and Thomas Zeugmann. Guest Editors’ introduction. Theoretical Computer Science, 519(??):1–3, January 30, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007068.

Gerchinovitz:2014:AOO

[2004] Sébastien Gerchinovitz and Jia Yuan Yu. Adaptive and optimal online linear regression on ℓ^{1}-balls. Theoretical Computer Science, 519(??): 4–28, January 30, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007159.

Warmuth:2014:CIS

[2005] Manfred K. Warmuth, Wouter M.

Koolen, and David P. Helmbold. Combining initial segments of lists. Theoretical Computer Science, 519(??): 29–45, January 30, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007123.

Salomon:2014:RSB

[2006] Antoine Salomon and Jean-Yves Audibert. Robustness of stochastic bandit policies. Theoretical Computer Science, 519(??):46–67, January 30, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300710X.

Darnstadt:2014:SLC

[2007] Malte Darnstädt, Hans Ulrich Simon, and Balázs Szörényi. Supervised learning and co-training. Theoretical Computer Science, 519(??): 68–87, January 30, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007111.

Zhang:2014:ATM

[2008] Xinhua Zhang, Ankan Saha, and S. V. N. Vishwanathan. Accelerated training of max-margin Markov networks with kernels. Theoretical Computer Science, 519(??):88– 102, January 30, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007172.

Cortes:2014:DAS

[2009] Corinna Cortes and Mehryar Mohri. Domain adaptation and sample bias correction theory and algorithm for regression. Theoretical Computer Science, 519(??):103–126, January 30, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007184.

Orseau:2014:UKS

[2010] Laurent Orseau. Universal knowledgeseeking agents. Theoretical Computer Science, 519(??):127–139, January 30, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007160.

Lattimore:2014:GTC

[2011] Tor Lattimore and Marcus Hutter.

General time consistent discounting. Theoretical Computer Science, 519(??): 140–154, January 30, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007135.

Kotzing:2014:ILP

[2012] Timo Kötzing. Iterative learning from positive data and counters. Theoretical Computer Science, 519(??): 155–169, January 30, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007147.

Anonymous:2014:EBe

[2013] Anonymous. Editorial Board. Theoretical Computer Science, 519(??): iii–viii, January 30, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513009481.

Xu:2014:MMT

[2014] Hui Xu, Jing Tian, and Xianzhong Zhao. Monoid-matrix type automata. Theoretical Computer Science, 520(??): 1–10, February 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513006038.

Vaya:2014:IDU

[2015] Shailesh Vaya. Information dissemination in unknown radio networks with large labels. Theoretical Computer Science, 520(??):11– 26, February 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513006063.

Berenbrink:2014:RBM

[2016] Petra Berenbrink, Robert Elsässer, and Thomas Sauerwald. Randomised broadcasting: Memory vs. randomness. Theoretical Computer Science, 520(??): 27–42, February 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513006087.

Takaoka:2014:SIA

[2017] Tadao Takaoka. Sharing information for the all pairs shortest path problem. Theoretical Computer Science, 520(??): 43–50, February 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006762.

Yang:2014:IBI

[2018] Guomin Yang, Chik How Tan, Yi Mu, Willy Susilo, and Duncan S. Wong. Identity based identification from algebraic coding theory. Theoretical Computer Science, 520(??):51– 61, February 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006786.

Wagner:2014:ICP

[2019] Urs Wagner and Gérard Maze. Improvements in closest point search based on dual HKZ-bases. Theoretical Computer Science, 520(??): 62–74, February 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007573.

Hong:2014:VSE

[2020] Zhen-Mu Hong and Jun-Ming Xu.

Vulnerability of super edge-connected networks. Theoretical Computer Science, 520(??):75–86, February 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397513007986.

Bang-Jensen:2014:CMM

[2021] Jørgen Bang-Jensen and Anders Yeo.

The complexity of multicut and mixed multicut problems in (di)graphs. Theoretical Computer Science, 520(??): 87–96, February 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513007998.

Wang:2014:SMP

[2022] Shiying Wang and Kai Feng. Strong matching preclusion for torus networks. Theoretical Computer Science, 520(??): 97–110, February 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513008001.

Ozbudak:2014:IPD

[2023] Ferruh Özbudak and Oguz Yayla. Improved probabilistic decoding of interleaved Reed–Solomon codes and folded Hermitian codes. Theoretical Computer Science, 520(??):111– 123, February 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513008050.

Xu:2014:TCP

[2024] Renyu Xu and Jian-Liang Wu. Total coloring of planar graphs with 7cycles containing at most two chords. Theoretical Computer Science, 520(??): 124–129, February 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513006439.

Vrba:2014:CAF

[2025] Zeljko Vrba, Pål Halvorsen, and Carsten Griwodz. A correction to Andersson’s fusion tree construction. Theoretical Computer Science, 520(??): 130–132, February 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007202.

Baudon:2014:PPT

[2026] Olivier Baudon, Julien Bensmail, Jakub Przybylo, and Mariusz Woźniak. Partitioning powers of traceable or Hamiltonian graphs. Theoretical Computer Science, 520(??):133–137, February 6, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007688.

Anonymous:2014:EBf

[2027] Anonymous. Editorial Board. Theoretical Computer Science, 520(??): iii–viii, February 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000073.

Cappelle:2014:RSC

[2028] Márcia R. Cappelle, Lucia Penso, and Dieter Rautenbach. Recognizing some complementary products. Theoretical Computer Science, 521(??): 1–7, February 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008244.

Vagvolgyi:2014:TS

[2029] Sándor Vágvölgyi. Tree shuffle. Theoretical Computer Science, 521(??): 8–28, February 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008256.

Crochemore:2014:EPP

[2030] M. Crochemore, C. S. Iliopoulos, M. Kubica, J. Radoszewski, W. Rytter, and T. Waleń. Extracting powers and periods in a word from its runs structure. Theoretical Computer Science, 521(??):29– 41, February 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008621.

Wang:2014:ODA

[2031] Hung-Lung Wang and Kuan-Yu Chen. One-dimensional approximate point set pattern matching with L_{p}-norm. Theoretical Computer Science, 521(??): 42–50, February 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008669.

Demaine:2014:HES

[2032] Erik D. Demaine, Martin L. Demaine, Nicholas J. A. Harvey, Ryuhei Uehara, Takeaki Uno, and Yushi Uno. UNO is hard, even for a single player. Theoretical Computer Science, 521(??): 51–61, February 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008670.

Luo:2014:SOS

[2033] Taibo Luo, Yinfeng Xu, Li Luo, and Changzheng He. Semi-online scheduling with two GoS levels and unit processing time. Theoretical Computer Science, 521(??):62– 72, February 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008700.

Grinshpun:2014:ATM

[2034] Andrey Grinshpun, Pakawat Phalitnonkiat, Sasha Rubin, and Andrei Tarfulea. Alternating traps in Muller and parity games. Theoretical Computer Science, 521(??):73– 91, February 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008761.

Samal:2014:GGC

[2035] R. Sámal and T. Valla. The guarding game is E-complete. Theoretical Computer Science, 521(??):92– 106, February 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008785.

Fotakis:2014:HND

[2036] Dimitris Fotakis, Alexis C. Kaporis, Thanasis Lianeas, and Paul G. Spirakis. On the hardness of network design for bottleneck routing games. Theoretical Computer Science, 521(??): 107–122, February 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008797.

Lin:2014:INE

[2037] Ling Lin and Zhiyi Tan. Inefficiency of Nash Equilibrium for scheduling games with constrained jobs: a parametric analysis. Theoretical Computer Science, 521(??):123– 134, February 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513008438.

Anonymous:2014:EBg

[2038] Anonymous. Editorial Board. Theoretical Computer Science, 521(??): iii–viii, February 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514000401.

Crespelle:2014:NTB

[2039] Christophe Crespelle and Philippe Gambette. (nearly-)tight bounds on the contiguity and linearity of cographs. Theoretical Computer Science, 522(??):1–12, February 20, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513008803.

Bopardikar:2014:CMP

[2040] Shaunak D. Bopardikar and Subhash Suri. k-Capture in multiagent pursuit evasion, or the lion and the hyenas. Theoretical Computer Science, 522(??): 13–23, February 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513008918.

Zhang:2014:MBC

[2041] Yi-Hao Zhang, Yue-Hong Shen, Min Dong, and Yi-Min Jiang. A method of batching conflict routings in shuffle-exchange networks. Theoretical Computer Science, 522(??): 24–33, February 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008943.

Dabrowski:2014:CGR

[2042] Konrad K. Dabrowski, Petr A. Golovach, and Daniel Paulusma. Colouring of graphs with Ramsey-type forbidden subgraphs. Theoretical Computer Science, 522(??):34– 43, February 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513009067.

Plociennik:2014:PPS

[2043] Kai Plociennik. A probabilistic PTAS for shortest common superstring. Theoretical Computer Science, 522(??): 44–53, February 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513009079.

Wang:2014:TCP

[2044] Huijuan Wang, Lidong Wu, and Jianliang Wu. Total coloring of planar graphs with maximum degree 8. Theoretical Computer Science, 522(??): 54–61, February 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513009080.

Egidi:2014:DAP

[2045] Lavinia Egidi and Giovanni Manzini. Design and analysis of periodic multiple seeds. Theoretical Computer Science, 522(??):62–76, February 20, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513009092.

Duval:2014:LCU

[2046] J.-P. Duval, T. Lecroq, and A. Lefebvre. Linear computation of unbordered conjugate on unordered alphabet. Theoretical Computer Science, 522 (??):77–84, February 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513009109.

Feng:2014:MWI

[2047] Qilong Feng, Jianxin Wang, and Jianer Chen. Matching and weighted P_{2}- packing: Algorithms and kernels. Theoretical Computer Science, 522(??): 85–94, February 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513009237.

Greco:2014:TPS

[2048] Gianluigi Greco and Francesco Scarcello. Tree projections and structural decomposition methods: Minimality and game-theoretic characterization. Theoretical Computer Science, 522(??): 95–114, February 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513009249.

Anonymous:2014:EBh

[2049] Anonymous. Editorial Board. Theoretical Computer Science, 522(??): iii–viii, February 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000553.

Bozzelli:2014:VGO

[2050] Laura Bozzelli and Sophie Pinchinat.

Verification of gap-order constraint abstractions of counter systems. Theoretical Computer Science, 523(??): 1–36, February 27, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300892X.

Cockett:2014:RCE

[2051] Robin Cockett and Richard Garner. Restriction categories as enriched categories. Theoretical Computer Science, 523(??):37–55, February 27, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513009304.

Adamek:2014:BMP

[2052] Jirí Adámek, Stefan Milius, and Jirí Velebil. Base modules for parametrized iterativity. Theoretical Computer Science, 523(??):56– 85, February 27, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513009316.

Balabanov:2014:HQB

[2053] Valeriy Balabanov, Hui-Ju Katherine Chiang, and Jie-Hong R. Jiang. Henkin quantifiers and Boolean formulae: a certification perspective of DQBF. Theoretical Computer Science, 523(??): 86–100, February 27, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513009328.

Anonymous:2014:EBi

[2054] Anonymous. Editorial Board. Theoretical Computer Science, 523(??): iii–viii, February 27, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514000693.

Ballerstein:2014:ARS

[2055] Kathrin Ballerstein and Dirk Oliver Theis. An algorithm for random signed 3-SAT with intervals. Theoretical Computer Science, 524(??): 1–26, March 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513007974.

Le:2014:CRR

[2056] Van Bang Le and Florian Pfender. Complexity results for rainbow matchings. Theoretical Computer Science, 524(??):27–33, March 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513009250.

Calamoneri:2014:PCGa

[2057] T. Calamoneri and R. Petreschi. On pairwise compatibility graphs having Dilworth number two. Theoretical Computer Science, 524(??): 34–40, March 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513009274. See

corrigendum [2457].

Blanchet-Sadri:2014:RIP

[2058] F. Blanchet-Sadri, Bob Chen, and Sinziana Munteanu. Recurrence in infinite partial words. Theoretical Computer Science, 524(??):41– 47, March 6, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513009286.

Kobert:2014:PMA

[2059] Kassian Kobert, Jörg Hauser, and Alexandros Stamatakis. Is the Protein Model Assignment problem under linked branch lengths NP-hard? Theoretical Computer Science, 524 (??):48–58, March 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000140.

Dieudonne:2014:PAM

[2060] Yoann Dieudonné and Andrzej Pelc.

Price of asynchrony in mobile agents computing. Theoretical Computer Science, 524(??):59–67, March 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397514000152.

Yu:2014:OWA

[2061] Cui Yu, Xiaofan Yang, Li He, and Jing Zhang. Optimal wavelength assignment in the implementation of parallel algorithms with ternary n-cube communication pattern on mesh optical network. Theoretical Computer Science, 524(??): 68–77, March 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000206.

Wei:2014:FTS

[2062] Qi Wei, Erfang Shan, and Liying Kang. A FPTAS for a two-stage hybrid flow shop problem and optimal algorithms for identical jobs. Theoretical Computer Science, 524(??): 78–89, March 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000218.

Mahajan:2014:MMI

[2063] Meena Mahajan, B. V. Raghavendra Rao, and Karteek Sreenivasaiah. Monomials, multilinearity and identity testing in simple read-restricted circuits. Theoretical Computer Science, 524(??): 90–102, March 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400022X.

Rybarczyk:2014:CIS

[2064] Katarzyna Rybarczyk. Constructions of independent sets in random intersection graphs. Theoretical Computer Science, 524(??): 103–125, March 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000231.

Eirinakis:2014:BAS

[2065] Pavlos Eirinakis, Dimitrios Magos, and Ioannis Mourtos. Blockers and antiblockers of stable matchings. Theoretical Computer Science, 524(??): 126–133, March 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513009298.

Anonymous:2014:EBj

[2066] Anonymous. Editorial Board. Theoretical Computer Science, 524(??): iii–viii, March 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000905.

Anonymous:2014:EBk

[2067] Anonymous. Editorial Board. Theoretical Computer Science, 525(??): iii–viii, March 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001182.

Christodoulakis:2014:ANR

[2068] Manolis Christodoulakis, Michalis Christou, Maxime Crochemore, and Costas S. Iliopoulos. On the average number of regularities in a word. Theoretical Computer Science, 525 (??):3–9, March 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007597.

Edwards:2014:PAB

[2069] James A. Edwards and Uzi Vishkin. Parallel algorithms for Burrows– Wheeler compression and decompression. Theoretical Computer Science, 525(??):10–22, March 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397513007615.

Apostolico:2014:MSS

[2070] Alberto Apostolico, Péter L. Erdős, István Miklós, and Johannes Siemons. Modulated string searching. Theoretical Computer Science, 525(??): 23–29, March 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007652.

Guo:2014:LSS

[2071] Jiong Guo, Danny Hermelin, and Christian Komusiewicz. Local search for string problems: Brute-force is essentially optimal. Theoretical Computer Science, 525(??):30– 41, March 13, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003575.

Keller:2014:GSC

[2072] Orgad Keller, Tsvi Kopelowitz, Shir Landau Feibish, and Moshe Lewenstein. Generalized substring compression. Theoretical Computer Science, 525(??):42–54, March 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397513007627.

Crochemore:2014:NGP

[2073] Maxime Crochemore, Alessio Langiu, and Filippo Mignosi. Note on the greedy parsing optimality for dictionary-based text compression. Theoretical Computer Science, 525(??): 55–59, March 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000309.

Amir:2014:DAP

[2074] Amihood Amir, Alberto Apostolico, Estrella Eisenberg, Gad M. Landau, Avivit Levy, and Noa Lewenstein. Detecting approximate periodic patterns. Theoretical Computer Science, 525(??): 60–67, March 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003381.

Kim:2014:OPM

[2075] Jinil Kim, Peter Eades, Rudolf Fleischer, Seok-Hee Hong, Costas S. Iliopoulos, Kunsoo Park, Simon J.

Puglisi, and Takeshi Tokuyama. Order-preserving matching. Theoretical Computer Science, 525(??): 68–79, March 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007585.

Barton:2014:EAM

[2076] Carl Barton, Costas S. Iliopoulos, Inbok Lee, Laurent Mouchard, Kunsoo Park, and Solon P. Pissis. Extending alignments with kmismatches and l-gaps. Theoretical Computer Science, 525(??):80– 88, March 13, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004660.

Parida:2014:ITM

[2077] Laxmi Parida, Cinzia Pizzi, and Simona E. Rombo. Irredundant tandem motifs. Theoretical Computer Science, 525(??):89–102, March 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397513006099.

Sokol:2014:SDT

[2078] Dina Sokol and Justin Tojeira. Speeding up the detection of tandem repeats over the edit distance. Theoretical Computer Science, 525(??): 103–110, March 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003290.

Ben-Kiki:2014:TOP

[2079] Oren Ben-Kiki, Philip Bille, Dany Breslauer, Leszek Gasieniec, Roberto Grossi, and Oren Weimann. Towards optimal packed string matching. Theoretical Computer Science, 525(??): 111–129, March 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004672.

Cicalese:2014:AMC

[2080] Ferdinando Cicalese, Eduardo Laber, Oren Weimann, and Raphael Yuster. Approximating the maximum consecutive subsums of a sequence. Theoretical Computer Science, 525(??): 130–137, March 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004441.

Matarazzo:2014:EAP

[2081] Ury Matarazzo, Dekel Tsur, and Michal Ziv-Ukelson. Efficient all path score computations on grid graphs. Theoretical Computer Science, 525(??): 138–149, March 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005215.

Anonymous:2014:EBl

[2082] Anonymous. Editorial Board. Theoretical Computer Science, 526(??): iii–viii, March 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001388.

Fomin:2014:SIW

[2083] Fedor V. Fomin, Frédéric Giroire, Alain Jean-Marie, Dorian Mazauric, and Nicolas Nisse. To satisfy impatient Web surfers is hard. Theoretical Computer Science, 526(??): 1–17, March 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514000267.

Golovnev:2014:NEA

[2084] Alexander Golovnev and Konstantin Kutzkov. New exact algorithms for the 2-constraint satisfaction problem. Theoretical Computer Science, 526(??): 18–27, March 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514000279.

Bang-Jensen:2014:ADF

[2085] Jørgen Bang-Jensen and Stéphane Bessy. (arc-)disjoint flows in networks. Theoretical Computer Science, 526(??): 28–40, March 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514000280.

Nikolopoulos:2014:CST

[2086] Stavros D. Nikolopoulos, Leonidas Palios, and Charis Papadopoulos. Counting spanning trees using modular decomposition. Theoretical Computer Science, 526(??):41–57, March 20, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000292.

Brodal:2014:DSP

[2087] Gerth Stølting Brodal, Alexis C. Kaporis, Apostolos N. Papadopoulos, Spyros Sioutas, Konstantinos Tsakalidis, and Kostas Tsichlas. Dynamic 3-sided planar range queries with expected doubly-logarithmic time. Theoretical Computer Science, 526(??): 58–74, March 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000310.

Binucci:2014:UQU

[2088] Carla Binucci, Walter Didimo, and Maurizio Patrignani. Upward and quasi-upward planarity testing of embedded mixed graphs. Theoretical Computer Science, 526(??):75– 89, March 20, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000322.

Kumar:2014:AAN

[2089] Mrinal Kumar, Sounaka Mishra, N. Safina Devi, and Saket Saurabh. Approximation algorithms for node deletion problems on bipartite graphs with finite forbidden subgraph characterization. Theoretical Computer Science, 526(??):90–96, March 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397514000474.

Genkin:2014:NEP

[2090] Daniel Genkin, Michael Kaminski, and Liat Peterfreund. A note on the emptiness problem for alternating finite-memory automata. Theoretical Computer Science, 526(??): 97–107, March 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514000486.

Ota:2014:DCA

[2091] Takahiro Ota, Hirotada Fukae, and Hiroyoshi Morita. Dynamic construction of an antidictionary with linear complexity. Theoretical Computer Science, 526(??):108–119, March 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397514000498.

Chang:2014:TCP

[2092] Jian Chang, Jian-Liang Wu, and Yong-Ga A. Total colorings of planar graphs with sparse triangles. Theoretical Computer Science, 526(??): 120–129, March 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514000504.

Anonymous:2014:EBm

[2093] Anonymous. Editorial Board. Theoretical Computer Science, 527(??): iii–viii, March 27, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514001510.

Fuchs:2014:AAA

[2094] Michael Fuchs, Hsien-Kuei Hwang, and Vytas Zacharovas. An analytic approach to the asymptotic variance of trie statistics and related structures. Theoretical Computer Science, 527(??):1–36, March 27, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000619.

Carvalho:2014:LCG

[2095] Alda Carvalho, Carlos Pereira dos Santos, Cátia Dias, Francisco Coelho, João Pedro Neto, Richard Nowakowski, and Sandra Vinagre. On lattices from combinatorial game theory modularity and a representation theorem: Finite case. Theoretical Computer Science, 527(??): 37–49, March 27, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000620.

Ghasemi:2014:PCC

[2096] Taha Ghasemi and Mohammadreza Razzazi. A PTAS for the cardinality constrained covering with unit balls. Theoretical Computer Science, 527(??): 50–60, March 27, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000632.

Bockenhauer:2014:OKP

[2097] Hans-Joachim Böckenhauer, Dennis Komm, Richard Královic, and Peter Rossmanith. The online knapsack problem: Advice and randomization. Theoretical Computer Science, 527(??): 61–72, March 27, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514000644.

Herranz:2014:ABS

[2098] Javier Herranz. Attribute-based signatures from RSA. Theoretical Computer Science, 527(??):73– 82, March 27, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514000772.

Wang:2014:ABO

[2099] Yang Wang, Man Ho Au, and Willy Susilo. Attribute-based optimistic fair exchange: How to restrict brokers with policies. Theoretical Computer Science, 527(??): 83–96, March 27, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514000784.

Lin:2014:POT

[2100] Fenggen Lin, Lianzhu Zhang, and Fuliang Lu. Pfaffian orientations for a type of bipartite graph. Theoretical Computer Science, 527(??): 97–101, March 27, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514000796.

Lu:2014:HOM

[2101] Jiqiang Lu, Yongzhuang Wei, Jongsung Kim, and Enes Pasalic. The higher-order meet-in-the-middle attack and its application to the Camellia block cipher. Theoretical Computer Science, 527(??):102– 122, March 27, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000802.

Anonymous:2014:EBn

[2102] Anonymous. Editorial Board. Theoretical Computer Science, 528(??): iii–viii, April 3, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001947.

Le:2014:CAR

[2103] Van Bang Le and Ragnar Nevries.

Complexity and algorithms for recognizing polar and monopolar graphs. Theoretical Computer Science, 528 (??):1–11, April 3, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000814.

Ly:2014:ECG

[2104] Olivier Ly and Zhilin Wu. On effective construction of the greatest solution of language inequality XA ⊆ BX. Theoretical Computer Science, 528(??):12–31, April 3, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000838.

Prunescu:2014:TVR

[2105] Mihai Prunescu. A two-valued recurrent double sequence that is not automatic. Theoretical Computer Science, 528(??):32–39, April 3, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751400084X.

Ramirez:2014:GFW

[2106] José L. Ramírez, Gustavo N. Rubiano, and Rodrigo De Castro. A generalization of the Fibonacci word fractal and the Fibonacci snowflake. Theoretical Computer Science, 528 (??):40–56, April 3, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514000851.

Ei:2014:SSR

[2107] Hiromi Ei. Stepped surfaces and Rauzy fractals induced from automorphisms on the free group of rank 2. Theoretical Computer Science, 528 (??):57–78, April 3, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514000978.

Xin:2014:LOC

[2108] Qin Xin, Fredrik Manne, and Xiaolan Yao. Latency-optimal communication in wireless mesh networks. Theoretical Computer Science, 528 (??):79–84, April 3, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514001017.

Broda:2014:HGD

[2109] Sabine Broda, António Machiavelo, Nelma Moreira, and Rogério Reis. A Hitchhiker’s Guide to descriptional complexity through analytic combinatorics. Theoretical Computer Science, 528(??):85–100, April 3, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001066.

Rampersad:2014:NAR

[2110] Narad Rampersad, Michel Rigo, and Pavel Salimov. A note on abelian returns in rotation words. Theoretical Computer Science, 528(??): 101–107, April 3, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000826.

Anonymous:2014:EBo

[2111] Anonymous. Editorial Board. Theoretical Computer Science, 529(??): iii–ix, April 10, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002114.

Rozenberg:2014:P

[2112] G. Rozenberg.  Preface.  Theo-

retical Computer Science, 529(??): 1, April 10, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002175.

Anthony:2014:LBS

[2113] Martin Anthony and Joel Ratsaby. Learning bounds via sample width for classifiers on finite metric spaces. Theoretical Computer Science, 529 (??):2–10, April 10, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005070.

Barbuti:2014:SSP

[2114] Roberto Barbuti, Andrea Maggiolo-Schettini, Paolo Milazzo, and Giovanni Pardini. Simulation of Spatial P system models. Theoretical Computer Science, 529(??):11–45, April 10, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300580X.

Caraiman:2014:HBS

[2115] Simona Caraiman and Vasile I. Manta. Histogram-based segmentation of quantum images. Theoretical Computer Science, 529(??): 46–60, April 10, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005835.

Song:2014:TFS

[2116] Tao Song, Luis F. Macías-Ramos, Linqiang Pan, and Mario J. Pérez-Jiménez. Time-free solution to SAT problem using P systems with active membranes. Theoretical Computer Science, 529(??):61– 68, April 10, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008451.

Alhazov:2014:SCE

[2117] Artiom Alhazov, Alberto Leporati, Giancarlo Mauri, Antonio E. Porreca, and Claudio Zandron. Space complexity equivalence of P systems with active membranes and Turing machines. Theoretical Computer Science, 529(??):69–81, April 10, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008463.

Song:2014:SNP

[2118] Tao Song, Linqiang Pan, and Gheorghe Paun. Spiking neural P systems with rules on synapses. Theoretical Computer Science, 529(??): 82–95, April 10, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000188.

Shah:2014:CNA

[2119] Tariq Shah and Nasreen Kausar. Characterizations of non-associative ordered semigroups by their fuzzy bi-ideals. Theoretical Computer Science, 529(??): 96–110, April 10, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001005.

Anonymous:2014:EBp

[2120] Anonymous. Editorial Board. Theoretical Computer Science, 530(??): iii–ix, April 17, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002266.

Chen:2014:SMD

[2121] Wenbin Chen, Zhengzhang Chen, Nagiza F. Samatova, Lingxi Peng, Jianxiong Wang, and Maobin Tang. Solving the maximum duo-preservation string mapping problem with linear programming. Theoretical Computer Science, 530(??):1–11, April 17, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514001108.

Mansour:2014:CRS

[2122] Yishay Mansour, Boaz Patt-Shamir, and Dror Rawitz. Competitive router scheduling with structured data. Theoretical Computer Science, 530(??): 12–22, April 17, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514001121.

Hanada:2014:ACL

[2123] Hiroyuki Hanada, Mineichi Kudo, and Atsuyoshi Nakamura. Average-case linear-time similar substring searching by the q-gram distance. Theoretical Computer Science, 530(??): 23–41, April 17, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514001285.

Blanchet-Sadri:2014:SPW

[2124] F. Blanchet-Sadri, Yang Jiao, John M.

Machacek, J. D. Quigley, and Xufan Zhang. Squares in partial words. Theoretical Computer Science, 530 (??):42–57, April 17, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001297.

Lin:2014:DTF

[2125] Cheng-Kuan Lin and Yuan-Hsiang Teng. The diagnosability of trianglefree graphs. Theoretical Computer Science, 530(??):58–65, April 17, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397514001303.

Blin:2014:CIM

[2126] Guillaume Blin, Paola Bonizzoni, Riccardo Dondi, Romeo Rizzi, and Florian Sikora. Complexity insights of the Minimum Duplication problem. Theoretical Computer Science, 530 (??):66–79, April 17, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001315.

Liao:2014:CCT

[2127] Chung-Shou Liao and Yamming Huang. The Covering Canadian Traveller Problem. Theoretical Computer Science, 530(??):80–88, April 17, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001327.

Chierichetti:2014:CLF

[2128] Flavio Chierichetti, Ravi Kumar, and Mohammad Mahdian. The complexity of LSH feasibility. Theoretical Computer Science, 530(??): 89–101, April 17, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514001467.

Anonymous:2014:EBq

[2129] Anonymous. Editorial Board. Theoretical Computer Science, 531(??): iii–ix, April 24, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002394.

Bensch:2014:GTI

[2130] Suna Bensch, Frank Drewes, Helmut Jürgensen, and Brink van der Merwe. Graph transformation for incremental natural language analysis. Theoretical Computer Science, 531 (??):1–25, April 24, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514000991.

Kim:2014:MMT

[2131] Sook-Yeon Kim and Jung-Heum Park. Many-to-many two-disjoint path covers in restricted hypercube-like graphs. Theoretical Computer Science, 531 (??):26–36, April 24, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514001078.

Wu:2014:OAS

[2132] Yong Wu, T. C. E. Cheng, and Min Ji.

Optimal algorithms for semi-online machine covering on two hierarchical machines. Theoretical Computer Science, 531(??):37–46, April 24, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400108X.

Li:2014:CDO

[2133] Xianyong Li, Xiaofan Yang, Li He, Jing Zhang, and Cui Yu. Conditional diagnosability of optical multimesh hypercube networks under the comparison diagnosis model. Theoretical Computer Science, 531(??): 47–53, April 24, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001091.

Angelelli:2014:CAT

[2134] Enrico Angelelli, Cristina Bazgan, M. Grazia Speranza, and Zsolt Tuza. Complexity and approximation for Traveling Salesman Problems with profits. Theoretical Computer Science, 531(??):54–65, April 24, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001728.

Furuse:2014:RSC

[2135] Masanobu Furuse and Koichi Yamazaki. A revisit of the scheme for computing treewidth and minimum fillin. Theoretical Computer Science, 531 (??):66–76, April 24, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002047.

Daykin:2014:BVB

[2136] Jacqueline W. Daykin and W. F. Smyth. A bijective variant of the Burrows–Wheeler Transform using V - order. Theoretical Computer Science, 531(??):77–89, April 24, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002059.

Anonymous:2014:EBr

[2137] Anonymous. Editorial Board. Theoretical Computer Science, 532(??): iii–ix, May 1, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002618.

Lu:2014:FA

[2138] Pinyan Lu and Jack Snoeyink. Frontiers in algorithmics. Theoretical Computer Science, 532(??): 1, May 1, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002655.

Boyar:2014:CPM

[2139] Joan Boyar, Kim S. Larsen, and Abyayananda Maiti. A comparison of performance measures via online search. Theoretical Computer Science, 532(??):2–13, May 1, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005380.

Do:2014:FRL

[2140] Huy Hoang Do, Jesper Jansson, Kunihiko Sadakane, and Wing-Kin Sung. Fast relative Lempel–Ziv selfindex for similar sequences. Theoretical Computer Science, 532(??): 14–30, May 1, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005409.

Fan:2014:VDV

[2141] Chenglin Fan, Jun Luo, Wencheng Wang, and Binhai Zhu. Voronoi diagram with visual restriction. Theoretical Computer Science, 532(??): 31–39, May 1, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006051.

He:2014:SOB

[2142] Bryan Dawei He. A simple optimal binary representation of mosaic floorplans and Baxter permutations. Theoretical Computer Science, 532 (??):40–50, May 1, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003587.

Kloks:2014:CBW

[2143] Ton Kloks, Sheung-Hung Poon, Feng-Ren Tsai, and Yue-Li Wang. On the complexity of the black-and-white coloring problem on some classes of perfect graphs. Theoretical Computer Science, 532(??):51– 63, May 1, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513008013.

Nishikawa:2014:BCM

[2144] Kazuhide Nishikawa, Takao Nishizeki, and Xiao Zhou. Bandwidth consecutive multicolorings of graphs. Theoretical Computer Science, 532(??): 64–72, May 1, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513001242.

Tan:2014:MMD

[2145] Xuehou Tan and Bo Jiang. Minimization of the maximum distance between the two guards patrolling a polygonal region. Theoretical Computer Science, 532(??):73–79, May 1, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513002326.

Wang:2014:SSC

[2146] Jiun-Jie Wang and Xin He. Succinct strictly convex greedy drawing of 3-connected plane graphs. Theoretical Computer Science, 532(??): 80–90, May 1, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751300385X.

Anonymous:2014:EBs

[2147] Anonymous. Editorial Board. Theoretical Computer Science, 533(??): iii–ix, May 8, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002801.

Formenti:2014:FPF

[2148] Enrico Formenti, Trung Van Pham, Thi Ha Duong Phan, and Thi Thu Huong Tran. Fixed-point forms of the parallel symmetric sandpile model. Theoretical Computer Science, 533(??):1–14, May 8, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001893.

Campelo:2014:HIC

[2149] Manoel Campêlo, Cristiana Huiban, Rudini M. Sampaio, and Yoshiko Wakabayashi. Hardness and inapproximability of convex recoloring problems. Theoretical Computer Science, 533(??):15–25, May 8, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002199.

Amir:2014:CPV

[2150] Amihood Amir, Estrella Eisenberg, Avivit Levy, and Noa Lewenstein. Closest periodic vectors in L_{p} spaces. Theoretical Computer Science, 533 (??):26–36, May 8, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002217.

Jajcayova:2014:PCU

[2151] Tatiana Jajcayová, Edita Pelantová, and Stepán Starosta. Palindromic closures using multiple antimorphisms. Theoretical Computer Science, 533 (??):37–45, May 8, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002321.

Kuske:2014:ISA

[2152] Dietrich Kuske. Isomorphisms of scattered automatic linear orders. Theoretical Computer Science, 533(??): 46–63, May 8, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002345.

Wang:2014:FTA

[2153] Shiying Wang and Kai Feng. Fault tolerance in the arrangement graphs. Theoretical Computer Science, 533 (??):64–71, May 8, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002473.

Anonymous:2014:EBt

[2154] Anonymous. Editorial Board. Theoretical Computer Science, 534(??): iii–ix, May 15, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514003016.

Droste:2014:P

[2155] Manfred Droste and Heiko Vogler. Preface. Theoretical Computer Science, 534(??):1, May 15, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001571.

Esik:2014:AWS

[2156] Z. Ésik. Axiomatizing weighted synchronization trees and weighted bisimilarity. Theoretical Computer Science, 534(??):2–23, May 15, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001595.

Gastin:2014:APW

[2157] Paul Gastin and Benjamin Monmege.

Adding pebbles to weighted automata: Easy specification & efficient evaluation. Theoretical Computer Science, 534(??):24–44, May 15, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001601.

Kirsten:2014:ACS

[2158] Daniel Kirsten. An algebraic characterization of semirings for which the support of every recognizable series is recognizable. Theoretical Computer Science, 534(??):45– 52, May 15, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001613.

Lehmann:2014:CCB

[2159] Karsten Lehmann and Rafael Peñaloza.

The complexity of computing the behaviour of lattice automata on infinite trees. Theoretical Computer Science, 534(??):53–68, May 15, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514001625.

Meinecke:2014:PMC

[2160] Ingmar Meinecke and Karin Quaas. Parameterized model checking of weighted networks. Theoretical Computer Science, 534(??):69–85, May 15, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514001637.

Damljanovic:2014:BWA

[2161] Nada Damljanović, Miroslav ’Cirić, and Jelena Ignjatović. Bisimulations for weighted automata over an additively idempotent semiring. Theoretical Computer Science, 534(??): 86–100, May 15, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514001583.

Anonymous:2014:EBu

[2162] Anonymous. Editorial Board. Theoretical Computer Science, 535(??): iii–ix, May 22, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751400320X.

Cicalese:2014:LBT

[2163] Ferdinando Cicalese, Gennaro Cordasco, Luisa Gargano, Martin Milanic, and Ugo Vaccaro. Latency-bounded target set selection in social networks. Theoretical Computer Science, 535 (??):1–15, May 22, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001339.

Oum:2014:FAV

[2164] Sang il Oum, Sigve Hortemo Sæther, and Martin Vatshelle. Faster algorithms for vertex partitioning problems parameterized by clique-width. Theoretical Computer Science, 535 (??):16–24, May 22, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002461.

Song:2014:IPA

[2165] Yinglei Song. An improved parameterized algorithm for the independent feedback vertex set problem. Theoretical Computer Science, 535 (??):25–30, May 22, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002539.

Abrams:2014:FNL

[2166] Lowell Abrams and Dena S. Cowen-Morton. A family of Nim-like arrays: the Locator Theorem. Theoretical Computer Science, 535(??): 31–37, May 22, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002540.

Deng:2014:TNI

[2167] Lunzhi Deng and Jiwen Zeng. Two new identity-based threshold ring signature schemes. Theoretical Computer Science, 535(??):38–45, May 22, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002692.

Hu:2014:ACB

[2168] Fu-Tao Hu and Moo Young Sohn. The algorithmic complexity of bondage and reinforcement problems in bipartite graphs. Theoretical Computer Science, 535(??):46–53, May 22, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002722.

Zhang:2014:AAM

[2169] Yaping Zhang, Yishuo Shi, and Zhao Zhang. Approximation algorithm for the minimum weight connected k-subgraph cover problem. Theoretical Computer Science, 535(??): 54–58, May 22, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002485.

Allahbakhshi:2014:CDC

[2170] Mahsa Allahbakhshi. Computing degree and class degree. Theoretical Computer Science, 535(??):59– 64, May 22, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002758.

Anonymous:2014:EBv

[2171] Anonymous. Editorial Board. Theoretical Computer Science, 536(??): iii–ix, May 29, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003429.

Ballester-Bolinches:2014:DBL

[2172] A. Ballester-Bolinches, E. Cosme-Llópez, and R. Esteban-Romero. A description based on languages of the final non-deterministic automaton. Theoretical Computer Science, 536(??):1–20, May 29, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000358.

Nigam:2014:FLA

[2173] Vivek Nigam. A framework for linear authorization logics. Theoretical Computer Science, 536(??): 21–41, May 29, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400111X.

Kahl:2014:WMH

[2174] Thomas Kahl. Weak morphisms of higher dimensional automata. Theoretical Computer Science, 536(??): 42–61, May 29, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001443.

Tremblay:2014:EGA

[2175] Hugo Tremblay, Gilbert Labelle, Srecko Brlek, and Alexandre Blondin Massé. Exhaustive generation of atomic combinatorial differential operators. Theoretical Computer Science, 536(??): 62–69, May 29, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001455.

Anonymous:2014:EBw

[2176] Anonymous. Editorial board. Theoretical Computer Science, 537(??): iii–ix, June 5, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003648.

Cerone:2014:TAC

[2177] Antonio Cerone and Pekka Pihlajasaari. Theoretical aspects of computing (ICTAC 2011). Theoretical Computer Science, 537(??):1– 2, June 5, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003673.

Byg:2014:TPT

[2178] Joakim Byg, Morten Jacobsen, Lasse Jacobsen, Kenneth Yrke Jørgensen, Mikael Harkjær Møller, and Jirí Srba. TCTL-preserving translations from timed-arc Petri nets to networks of timed automata. Theoretical Computer Science, 537(??):3– 28, June 5, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005148.

Diekert:2014:TMP

[2179] Volker Diekert and Martin Leucker.

Topology, monitorable properties and runtime verification. Theoretical Computer Science, 537(??):29– 41, June 5, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002035.

Aceto:2014:AWS

[2180] Luca Aceto, David de Frutos Escrig, Carlos Gregorio-Rodríguez, and Anna Ingolfsdottir. Axiomatizing weak simulation semantics over BCCSP. Theoretical Computer Science, 537 (??):42–71, June 5, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002144.

Areces:2014:CDS

[2181] Carlos Areces, Facundo Carreiro, and Santiago Figueira. Characterization, definability and separation via saturated models. Theoretical Computer Science, 537(??): 72–86, June 5, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400173X.

Cojocaru:2014:SDM

[2182] Liliana Cojocaru and Erkki Mäkinen.

On some derivation mechanisms and the complexity of their Szilard languages. Theoretical Computer Science, 537(??):87–96, June 5, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001741.

vanderMerwe:2014:ASA

[2183] A. B. van der Merwe, L. van Zijl, and J. Geldenhuys. Ambiguity and structural ambiguity of symmetric difference NFAs. Theoretical Computer Science, 537(??):97–104, June 5, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513003162.

Tarau:2014:TGV

[2184] Paul Tarau. Towards a generic view of primality through multiset decompositions of natural numbers. Theoretical Computer Science, 537(??): 105–124, June 5, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003302.

Anonymous:2014:EBx

[2185] Anonymous. Editorial board. Theoretical Computer Science, 538(??): iii–ix, June 12, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003788.

Massink:2014:QAP

[2186] Mieke Massink, Gethin Norman, and Herbert Wiklicky. Quantitative aspects of programming languages and systems (2011–12). Theoretical Computer Science, 538(??): 1, June 12, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003818.

Adao:2014:PIF

[2187] Pedro Adão, Paulo Mateus, and Luca Viganò. Protocol insecurity with a finite number of sessions and a cost-sensitive guessing intruder is NP-complete. Theoretical Computer Science, 538(??):2– 15, June 12, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006956.

Deng:2014:RRT

[2188] Yuxin Deng, Rob van Glabbeek, Matthew Hennessy, and Carroll Morgan. Real-reward testing for probabilistic processes. Theoretical Computer Science, 538(??):16– 36, June 12, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005197.

Drager:2014:LAR

[2189] Klaus Dräger, Marta Kwiatkowska, David Parker, and Hongyang Qu. Local abstraction refinement for probabilistic timed programs. Theoretical Computer Science, 538(??): 37–53, June 12, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513005161.

Fahrenberg:2014:QLT

[2190] Uli Fahrenberg and Axel Legay. The quantitative linear-time-branching-time spectrum. Theoretical Computer Science, 538(??):54–69, June 12, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397513005604.

Giro:2014:OSV

[2191] Sergio Giro. Optimal schedulers vs optimal bases: An approach for efficient exact solving of Markov decision processes. Theoretical Computer Science, 538(??):70–83, June 12, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513006440.

Giro:2014:DPI

[2192] Sergio Giro, Pedro R. D’Argenio, and Luis María Ferrer Fioriti. Distributed probabilistic input/output automata: Expressiveness, (un)decidability and algorithms. Theoretical Computer Science, 538(??):84–102, June 12, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397513005203.

Hansen:2014:CCA

[2193] Henri Hansen and Mark Timmer. A comparison of confluence and ample sets in probabilistic and nonprobabilistic branching time. Theoretical Computer Science, 538(??): 103–123, June 12, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005173.

Klebanov:2014:PQI

[2194] Vladimir Klebanov. Precise quantitative information flow analysis — a symbolic approach. Theoretical Computer Science, 538(??):124– 139, June 12, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003144.

Tschaikowski:2014:EFL

[2195] Max Tschaikowski and Mirco Tribastone. Exact fluid lumpability in Markovian process algebra. Theoretical Computer Science, 538(??): 140–166, June 12, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005598.

Yasuoka:2014:QIF

[2196] Hirotoshi Yasuoka and Tachio Terauchi. Quantitative information flow as safety and liveness hyperproperties. Theoretical Computer Science, 538 (??):167–182, June 12, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513005616.

Anonymous:2014:EBy

[2197] Anonymous. Editorial board. Theoretical Computer Science, 539(??): iii–ix, June 19, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514003922.

Deng:2014:RMB

[2198] Xiaotie Deng, Paul Goldberg, Bo Tang, and Jinshan Zhang. Revenue maximization in a Bayesian double auction market. Theoretical Computer Science, 539(??):1–12, June 19, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002928.

Brzozowski:2014:TA

[2199] Janusz Brzozowski and Hellis Tamm. Theory of átomata. Theoretical Computer Science, 539(??):13– 27, June 19, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002953.

Seager:2014:LBR

[2200] Suzanne Seager. Locating a backtracking robber on a tree. Theoretical Computer Science, 539(??): 28–37, June 19, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514003119.

Benedikt:2014:PCC

[2201] Michael Benedikt, Gabriele Puppis, and Cristian Riveros. The percharacter cost of repairing word languages. Theoretical Computer Science, 539(??):38–67, June 19, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003132.

Fotakis:2014:EIE

[2202] Dimitris Fotakis and Paris Siminelakis.

On the efficiency of influence-and-exploit strategies for revenue maximization under positive externalities. Theoretical Computer Science, 539 (??):68–86, June 19, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003314.

Liang:2014:CCS

[2203] Kaitai Liang, Cheng-Kang Chu, Xiao Tan, Duncan S. Wong, Chunming Tang, and Jianying Zhou. Chosenciphertext secure multi-hop identitybased conditional proxy re-encryption with constant-size ciphertexts. Theoretical Computer Science, 539(??): 87–105, June 19, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003326.

Borzyszkowski:2014:CMM

[2204] Andrzej M. Borzyszkowski. Computing majority via multiple queries. Theoretical Computer Science, 539(??): 106–111, June 19, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003120.

Anonymous:2014:EBz

[2205] Anonymous. Editorial board. Theoretical Computer Science, 540–541 (??):iii–ix, June 26, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751400406X.

Markakis:2014:SIC

[2206] Vangelis Markakis, Ioannis Milis, and Vangelis Th. Paschos. Special issue: “Combinatorial Optimization: Theory of Algorithms and Complexity”. Theoretical Computer Science, 540– 541(??):1, June 26, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514004095.

Angelopoulos:2014:MTR

[2207] Spyros Angelopoulos, Alejandro López-Ortiz, and Konstantinos Panagiotou. Multi-target ray searching problems. Theoretical Computer Science, 540–541 (??):2–12, June 26, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002503.

Boria:2014:RMS

[2208] Nicolas Boria and Federico Della Croce. Reoptimization in machine scheduling. Theoretical Computer Science, 540–541(??):13–26, June 26, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002710.

Fotakis:2014:OSR

[2209] Dimitris Fotakis and Paraschos Koutris. Online sum-radii clustering. Theoretical Computer Science, 540–541 (??):27–39, June 26, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513002119.

Gozupek:2014:CCM

[2210] Didem Gözüpek, Mordechai Shalom, Ariella Voloshin, and Shmuel Zaks. On the complexity of constructing minimum changeover cost arborescences. Theoretical Computer Science, 540–541 (??):40–52, June 26, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400245X.

Biro:2014:SSR

[2211] Péter Biró, Matthijs Bomhoff, Petr A.

Golovach, Walter Kern, and Daniël Paulusma. Solutions for the stable roommates problem with payments. Theoretical Computer Science, 540–541 (??):53–61, June 26, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300251X.

Han:2014:ORK

[2212] Xin Han, Yasushi Kawase, Kazuhisa Makino, and He Guo. Online removable knapsack problem under convex function. Theoretical Computer Science, 540–541(??):62– 69, June 26, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006932.

daFonseca:2014:ESA

[2213] Guilherme D. da Fonseca, Celina M. H. de Figueiredo, Vinícius G. Pereira de Sá, and Raphael C. S. Machado. Efficient sub-5 approximations for minimum dominating sets in unit disk graphs. Theoretical Computer Science, 540–541(??):70–81, June 26, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514000607.

Angel:2014:ODP

[2214] Eric Angel, Evripidis Bampis, Gerasimos G. Pollatos, and Vassilis Zissimopoulos. Optimal data placement on networks with a constant number of clients. Theoretical Computer Science, 540–541(??): 82–88, June 26, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513002491.

Demange:2014:CSG

[2215] Marc Demange, Jérôme Monnot, Petrica Pop, and Bernard Ries. On the complexity of the selective graph coloring problem in some special classes of graphs. Theoretical Computer Science, 540–541(??): 89–102, June 26, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513003022.

Fernandes:2014:SPA

[2216] Cristina G. Fernandes and Rafael C. S. Schouery. Second-price ad auctions with binary bids and markets with good competition. Theoretical Computer Science, 540–541(??): 103–114, June 26, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513006464.

Mastrolilli:2014:BCA

[2217] Monaldo Mastrolilli and Georgios Stamoulis. Bi-criteria and approximation algorithms for restricted matchings. Theoretical Computer Science, 540–541 (??):115–132, June 26, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008712.

Tran:2014:PSS

[2218] Van Du T. Tran, Philippe Chassignet, and Jean-Marc Steyaert. On permuted super-secondary structures of transmembrane β-barrel proteins. Theoretical Computer Science, 540–541(??): 133–142, June 26, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007391.

Watrigant:2014:SMG

[2219] Rémi Watrigant, Marin Bougeret, Rodolphe Giroudeau, and Jean-Claude König. On the sum-max graph partitioning problem. Theoretical Computer Science, 540–541(??):143– 155, June 26, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008682.

Weyland:2014:CCP

[2220] D. Weyland. On the computational complexity of the probabilistic traveling salesman problem with deadlines. Theoretical Computer Science, 540–541 (??):156–168, June 26, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513008724.

Cao:2014:FGN

[2221] Zhigang Cao and Xiaoguang Yang. The fashion game: Network extension of matching pennies. Theoretical Computer Science, 540–541(??): 169–181, June 26, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514003582.

Anonymous:2014:EBaa

[2222] Anonymous. Editorial board. Theoretical Computer Science, 542(??): iii–ix, July 3, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514004484.

Zhang:2014:PSS

[2223] Jiang Zhang, Zhenfeng Zhang, and Yu Chen. PRE: Stronger security notions and efficient construction with non-interactive opening. Theoretical Computer Science, 542 (??):1–16, July 3, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514003338.

Bouchy:2014:DCC

[2224] Florent Bouchy, Alain Finkel, and Pierluigi San Pietro. Dense-choice Counter Machines revisited. Theoretical Computer Science, 542(??): 17–31, July 3, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400334X.

Kurka:2014:ERA

[2225] Petr Kůrka. Exact real arithmetic for interval number systems. Theoretical Computer Science, 542(??): 32–43, July 3, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003351.

Seo:2014:RHI

[2226] Jae Hong Seo and Keita Emura. Revocable hierarchical identity-based encryption. Theoretical Computer Science, 542(??):44–62, July 3, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397514003363.

Ju:2014:DVH

[2227] Ming-Yi Ju, Jeng-Jung Wang, and Shu-Hao Chang. Diameter variability of hypercubes. Theoretical Computer Science, 542(??):63– 70, July 3, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003478.

Dornfelder:2014:PCC

[2228] Martin Dörnfelder, Jiong Guo, Christian Komusiewicz, and Mathias Weller. On the parameterized complexity of consensus clustering. Theoretical Computer Science, 542(??):71– 82, July 3, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400351X.

Navarro:2014:NST

[2229] Gonzalo Navarro and Sharma V. Thankachan. New space/time tradeoffs for top-k document retrieval on sequences. Theoretical Computer Science, 542(??):83–97, July 3, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397514003545.

Goldstein:2014:QGC

[2230] Isaac Goldstein and Moshe Lewenstein. Quick greedy computation for minimum common string partition. Theoretical Computer Science, 542 (??):98–107, July 3, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003557.

R:2014:LTA

[2231] Sundara Rajan R., Indra Rajasingh, Parthiban N., and Rajalaxmi T. M. A linear time algorithm for embedding hypercube into cylinder and torus. Theoretical Computer Science, 542 (??):108–115, July 3, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003569.

Anonymous:2014:EBab

[2232] Anonymous. Editorial Board. Theoretical Computer Science, 543(??): iii–ix, July 10, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004708.

Popa:2014:BLU

[2233] Alexandru Popa. Better lower and upper bounds for the minimum rainbow subgraph problem. Theoretical Computer Science, 543(??):1– 8, July 10, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003570.

Chechik:2014:RFT

[2234] Shiri Chechik and David Peleg. Robust fault tolerant uncapacitated facility location. Theoretical Computer Science, 543(??):9–23, July 10, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003879.

Liu:2014:OSE

[2235] Hailing Liu and Jinjiang Yuan. Online scheduling of equal length jobs on a bounded parallel batch machine with restart or limited restart. Theoretical Computer Science, 543 (??):24–36, July 10, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004204.

Choi:2014:TAS

[2236] Byung-Cheon Choi and Jibok Chung. Two-agent single-machine scheduling problem with just-in-time jobs. Theoretical Computer Science, 543(??): 37–45, July 10, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004216.

Li:2014:AAB

[2237] Kenli Li, Chubo Liu, and Keqin Li. An approximation algorithm based on game theory for scheduling simple linear deteriorating jobs. Theoretical Computer Science, 543(??): 46–51, July 10, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004228.

Frougny:2014:BPA

[2238] Christiane Frougny, Pavel Heller, Edita Pelantová, and Milena Svobodová. k-Block parallel addition versus 1-block parallel addition in nonstandard numeration systems. Theoretical Computer Science, 543(??): 52–67, July 10, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004253.

Arnoux:2014:RPS

[2239] Pierre Arnoux, Masahiro Mizutani, and Tarek Sellami. Random product of substitutions with the same incidence matrix. Theoretical Computer Science, 543(??):68–78, July 10, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004265.

Chen:2014:DSE

[2240] Jie Chen and Hoeteck Wee. Doubly spatial encryption from DBDH. Theoretical Computer Science, 543 (??):79–89, July 10, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004277.

Rusu:2014:MPU

[2241] Irena Rusu. MinMax-profiles: a unifying view of common intervals, nested common intervals and conserved intervals of K permutations. Theoretical Computer Science, 543(??): 90–111, July 10, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004289.

Howard:2014:DMS

[2242] David Howard. Determining membership with 2 simultaneous queries. Theoretical Computer Science, 543(??): 112–119, July 10, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004198.

Li:2014:OSI

[2243] Wenhua Li, Jinjiang Yuan, and Sufang Yang. Online scheduling of incompatible unit-length job families with lookahead. Theoretical Computer Science, 543(??):120– 125, July 10, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751400423X.

Anonymous:2014:EBac

[2244] Anonymous. Editorial Board. Theoretical Computer Science, 544(??): iii–ix, August 7, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751400485X.

Chao:2014:ACI

[2245] Kun-Mao Chao, Tsan sheng Hsu, and D. T. Lee. Algorithms and computation (ISAAC 2012). Theoretical Computer Science, 544(??): 1–2, August 7, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514004927.

Paluch:2014:PCP

[2246] Katarzyna Paluch. Popular and clan-popular b-matchings. Theoretical Computer Science, 544(??): 3–13, August 7, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002965.

Ito:2014:AGU

[2247] Takehiro Ito, Shin ichi Nakano, Yoshio Okamoto, Yota Otachi, Ryuhei Uehara, Takeaki Uno, and Yushi Uno. A 4.31-approximation for the geometric unique coverage problem on unit disks. Theoretical Computer Science, 544(??): 14–31, August 7, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400293X.

Frati:2014:NUP

[2248] Fabrizio Frati, Joachim Gudmundsson, and Emo Welzl. On the number of upward planar orientations of maximal planar graphs. Theoretical Computer Science, 544(??): 32–59, August 7, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003296.

Kociumaka:2014:ECS

[2249] Tomasz Kociumaka, Jakub Pachocki, Jakub Radoszewski, Wojciech Rytter, and Tomasz Waleń. Efficient counting of square substrings in a tree. Theoretical Computer Science, 544(??): 60–73, August 7, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002941.

Kappmeier:2014:AFT

[2250] Jan-Philipp W. Kappmeier, Jannik Matuschke, and Britta Peis. Abstract flows over time: a first step towards solving dynamic packing problems. Theoretical Computer Science, 544(??): 74–83, August 7, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002916.

Ito:2014:RLI

[2251] Takehiro Ito, Kazuto Kawamura, Hirotaka Ono, and Xiao Zhou. Reconfiguration of list L(2, 1)-labelings in a graph. Theoretical Computer Science, 544(??): 84–97, August 7, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002904.

Anonymous:2014:EBad

[2252] Anonymous. Editorial Board. Theoretical Computer Science, 545(??): iii–ix, August 14, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005064.

Friedrich:2014:GEC

[2253] Tobias Friedrich and Jonathan Rowe. Genetic and evolutionary computation. Theoretical Computer Science, 545(??):1, August 14, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005192.

Oliveto:2014:RAS

[2254] Pietro S. Oliveto and Carsten Witt. On the runtime analysis of the Simple Genetic Algorithm. Theoretical Computer Science, 545(??):2– 19, August 14, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513004817.

Rowe:2014:COP

[2255] Jonathan E. Rowe and Dirk Sudholt. The choice of the offspring population size in the (1, λ) evolutionary algorithm. Theoretical Computer Science, 545(??):20–38, August 14, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751300755X.

Jansen:2014:PAR

[2256] Thomas Jansen and Christine Zarges.

Performance analysis of randomised search heuristics operating with a fixed budget. Theoretical Computer Science, 545(??):39–58, August 14, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513004611.

Whitley:2014:CME

[2257] Darrell Whitley, Andrew M. Sutton, Gabriela Ochoa, and Francisco Chicano. The component model for elementary landscapes and partial neighborhoods. Theoretical Computer Science, 545(??):59– 75, August 14, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003594.

Chicano:2014:ECE

[2258] Francisco Chicano, Darrell Whitley, and Enrique Alba. Exact computation of the expectation surfaces for uniform crossover along with bit-flip mutation.

Theoretical Computer Science, 545(??): 76–93, August 14, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751400019X.

Kotzing:2014:MPR

[2259] Timo Kötzing, Andrew M. Sutton, Frank Neumann, and Una-May O’Reilly. The Max problem revisited: the importance of mutation in genetic programming. Theoretical Computer Science, 545(??):94– 107, August 14, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513004684.

Doerr:2014:RAU

[2260] Benjamin Doerr and Carola Winzen. Reducing the arity in unbiased blackbox complexity. Theoretical Computer Science, 545(??):108–121, August 14, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513003411.

Anonymous:2014:EBae

[2261] Anonymous. Editorial Board. Theoretical Computer Science, 546(??): iii–ix, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005295.

Cattani:2014:P

[2262] Gian Luca Cattani, Thomas T. Hildebrandt, and Daniele Varacca. Preface. Theoretical Computer Science, 546 (??):1–2, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001649.

Nielsen:2014:TES

[2263] Mogens Nielsen. Trust in event structures. Theoretical Computer Science, 546(??):3–6, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001650.

Abramsky:2014:EC

[2264] Samson Abramsky. Events in context. Theoretical Computer Science, 546(??):7–9, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001662.

Panangaden:2014:CPC

[2265] Prakash Panangaden. Causality in physics and computation. Theoretical Computer Science, 546(??): 10–16, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001674.

Bauer:2014:CCC

[2266] Andrej Bauer, Gordon D. Plotkin, and Dana S. Scott. Cartesian closed categories of separable Scott domains. Theoretical Computer Science, 546(??): 17–29, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001686.

Benveniste:2014:FHG

[2267] Albert Benveniste, Claude Jard, and Samy Abbes. Foreword in honor of Glynn Winskel. Theoretical Computer Science, 546(??): 30–31, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001698.

Abbes:2014:BCA

[2268] Samy Abbes. Branching cells for asymmetric event structures. Theoretical Computer Science, 546(??): 32–51, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001704.

Benveniste:2014:ABC

[2269] Albert Benveniste, Claude Jard, and Samy Abbes. Application of branching cells to QoS aware service orchestrations. Theoretical Computer Science, 546(??):52–62, August 21, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001753.

Bernardo:2014:RSB

[2270] Marco Bernardo, Rocco De Nicola, and Michele Loreti. Relating strong behavioral equivalences for processes with nondeterminism and probabilities. Theoretical Computer Science, 546 (??):63–92, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001765.

Bucalo:2014:SES

[2271] Anna Bucalo and Giuseppe Rosolini.

Sobriety for equilogical spaces. Theoretical Computer Science, 546(??): 93–98, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001777.

Curien:2014:RCI

[2272] Pierre-Louis Curien, Richard Garner, and Martin Hofmann. Revisiting the categorical interpretation of dependent type theory. Theoretical Computer Science, 546(??):99– 119, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001789.

Fiore:2014:AFB

[2273] Marcelo Fiore. Analytic functors between presheaf categories over groupoids. Theoretical Computer Science, 546(??):120–131, August 21, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001790.

Hyland:2014:ETA

[2274] J. M. E. Hyland. Elements of a theory of algebraic theories. Theoretical Computer Science, 546(??):

132–144, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514001807.

Kinoshita:2014:CTS

[2275] Yoshiki Kinoshita and John Power. Category theoretic structure of setoids. Theoretical Computer Science, 546(??): 145–163, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514001819.

Larsen:2014:CPS

[2276] Kim G. Larsen and Radu Mardare. Complete proof systems for weighted modal logic. Theoretical Computer Science, 546(??):164– 175, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514001820.

Mislove:2014:ADC

[2277] Michael Mislove. Anatomy of a domain of continuous random variables I. Theoretical Computer Science, 546(??): 176–187, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514001832.

Montanari:2014:NCC

[2278] Ugo Montanari and Matteo Sammartino. A network-conscious πcalculus and its coalgebraic semantics. Theoretical Computer Science, 546(??): 188–224, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001844.

Thiagarajan:2014:RTC

[2279] P. S. Thiagarajan and Shaofa Yang.

Rabin’s theorem in the concurrency setting: a conjecture. Theoretical Computer Science, 546(??):225– 236, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001856.

vanOosten:2014:RLO

[2280] Jaap van Oosten. Realizability with a local operator of A. M. Pitts. Theoretical Computer Science, 546(??): 237–243, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001868.

Clausen:2014:CMS

[2281] Christian Clausen, Olivier Danvy, and Moe Masuko. A characterization of Moessner’s sieve. Theoretical Computer Science, 546(??):244– 256, August 21, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400187X.

Anonymous:2014:EBaf

[2282] Anonymous. Editorial Board. Theoretical Computer Science, 547(??): iii–ix, August 28, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005490.

Dragan:2014:CAT

[2283] Feodor F. Dragan and Muad Abu-Ata. Collective additive tree spanners of bounded tree-breadth graphs with generalizations and consequences. Theoretical Computer Science, 547(??): 1–17, August 28, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004319.

Heuberger:2014:SDS

[2284] Clemens Heuberger and Michela Mazzoli. Symmetric digit sets for elliptic curve scalar multiplication without precomputation. Theoretical Computer Science, 547(??):18– 33, August 28, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004344.

Chen:2014:NPW

[2285] Hung-Hsun Chen, Wen-Guei Hu, De-Jan Lai, and Song-Sun Lin. Nonemptiness problems of Wang tiles with three colors. Theoretical Computer Science, 547(??):34– 45, August 28, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004393.

Kobayashi:2014:MFT

[2286] Kojiro Kobayashi. The minimum firing time of the generalized firing squad synchronization problem for squares. Theoretical Computer Science, 547(??): 46–69, August 28, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400440X.

Milanic:2014:SGI

[2287] Martin Milanic, Romeo Rizzi, and Alexandru I. Tomescu. Set graphs. II. Complexity of set graph recognition and similar problems. Theoretical Computer Science, 547(??): 70–81, August 28, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004411.

Calamoneri:2014:PCGb

[2288] Tiziana Calamoneri and Rossella Petreschi. On pairwise compatibility graphs having Dilworth number k. Theoretical Computer Science, 547(??): 82–89, August 28, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004605.

Elor:2014:RCG

[2289] Yotam Elor and Alfred M. Bruckstein. “Robot Cloud” gradient climbing with point measurements. Theoretical Computer Science, 547(??): 90–103, August 28, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004617.

Chatterjee:2014:AMC

[2290] Krishnendu Chatterjee, Monika Henzinger, Sebastian Krinninger, Veronika Loitzenbauer, and Michael A. Raskin. Approximating the minimum cycle mean. Theoretical Computer Science, 547(??):104–116, August 28, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397514004782.

Shparlinski:2014:EPS

[2291] Igor E. Shparlinski. Evasive properties of sparse graphs and some linear equations in primes. Theoretical Computer Science, 547(??):117– 121, August 28, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514004290.

Fraenkel:2014:CEN

[2292] Aviezri S. Fraenkel and Jamie Simpson. Corrigendum to “The exact number of squares in Fibonacci words” [Theoret. Comput. Sci. 218 (1) (1999) 95– 106]. Theoretical Computer Science, 547(??):122, August 28, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514004356. See

[2].

Anonymous:2014:EBag

[2293] Anonymous. Editorial Board. Theoretical Computer Science, 548(??): iii–ix, September 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005726.

Giaquinta:2014:MMU

[2294] Emanuele Giaquinta, Kimmo Fredriksson, Szymon Grabowski, Alexandru I. Tomescu, and Esko Ukkonen. Motif matching using gapped patterns. Theoretical Computer Science, 548(??): 1–13, September 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004794.

Vesti:2014:ERW

[2295] Jetro Vesti. Extensions of rich words.

Theoretical Computer Science, 548(??): 14–24, September 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004800.

Dumas:2014:AEL

[2296] Philippe Dumas. Asymptotic expansions for linear homogeneous divideand-conquer recurrences: Algebraic and analytic approaches collated. Theoretical Computer Science, 548(??): 25–53, September 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004964.

Tan:2014:ECS

[2297] Xiao Tan, Qiong Huang, and Duncan S. Wong. Extending concurrent signature to multiple parties. Theoretical Computer Science, 548(??): 54–67, September 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514004988.

Lelarge:2014:STA

[2298] Marc Lelarge and Hang Zhou. Sublinear-time algorithms for monomerdimer systems on bounded degree graphs. Theoretical Computer Science, 548(??):68–78, September 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397514005003.

Lin:2014:CDA

[2299] Limei Lin, Shuming Zhou, Li Xu, and Dajin Wang. Conditional diagnosability of arrangement graphs under the PMC model. Theoretical Computer Science, 548(??):79– 97, September 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005015.

Kim:2014:ACD

[2300] Hee-Chul Kim, Hyeong-Seok Lim, and Jung-Heum Park. An approach to conditional diagnosability analysis under the PMC model and its application to torus networks. Theoretical Computer Science, 548(??):98– 116, September 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005234.

Richomme:2014:MCE

[2301] Gwenaël Richomme. Minimal critical exponent of quasiperiodic words.

Theoretical Computer Science, 548(??): 117–122, September 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400499X.

Anonymous:2014:EBah

[2302] Anonymous. Editorial Board. Theoretical Computer Science, 549(??): iii–ix, September 11, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005866.

Hetzl:2014:AIQ

[2303] Stefan Hetzl, Alexander Leitsch, Giselle Reis, and Daniel Weller. Algorithmic introduction of quantified cuts. Theoretical Computer Science, 549(??): 1–16, September 11, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004174.

Fujita:2014:ETS

[2304] Ken etsu Fujita and Aleksy Schubert. Existential type systems between Church and Curry style (type-free style). Theoretical Computer Science, 549(??):17–35, September 11, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004186.

Bilkova:2014:PSM

[2305] Marta Bílková, Alessandra Palmigiano, and Yde Venema. Proof systems for Moss’ coalgebraic logic. Theoretical Computer Science, 549(??):36– 60, September 11, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004423.

Wimmer:2014:MCL

[2306] Ralf Wimmer, Nils Jansen, Erika Ábrahám, Joost-Pieter Katoen, and Bernd Becker. Minimal counterexamples for linear-time probabilistic verification. Theoretical Computer Science, 549(??):61–100, September 11, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400454X.

Schewe:2014:TFK

[2307] Klaus-Dieter Schewe and Qing Wang. A theoretical framework for knowledgebased entity resolution. Theoretical Computer Science, 549(??):101– 126, September 11, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004770.

Song:2014:ECM

[2308] Fu Song and Tayssir Touili. Efficient CTL model-checking for pushdown systems. Theoretical Computer Science, 549(??):127–145, September 11, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005155.

Chilton:2014:ATI

[2309] Chris Chilton, Bengt Jonsson, and Marta Kwiatkowska. An algebraic theory of interface automata. Theoretical Computer Science, 549(??):146– 174, September 11, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005611.

Anonymous:2014:EBai

[2310] Anonymous. Editorial Board. Theoretical Computer Science, 550(??): iii–ix, September 18, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006033.

Sun:2014:EEM

[2311] Hongyang Sun, Yuxiong He, Wen-Jing Hsu, and Rui Fan. Energyefficient multiprocessor scheduling for flow time and makespan. Theoretical Computer Science, 550(??):1– 20, September 18, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005246.

Asahiro:2014:CFM

[2312] Yuichi Asahiro, Hiroshi Eto, Takehiro Ito, and Eiji Miyano. Complexity of finding maximum regular induced subgraphs with prescribed degree. Theoretical Computer Science, 550(??):21– 35, September 18, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005386.

Fazli:2014:NPS

[2313] MohammadAmin Fazli, Mohammad Ghodsi, Jafar Habibi, Pooya Jalaly, Vahab Mirrokni, and Sina Sadeghian. On non-progressive spread of influence through social networks. Theoretical Computer Science, 550(??):36– 50, September 18, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005398.

Bille:2014:CQG

[2314] Philip Bille, Patrick Hagge Cording, and Inge Li Gørtz. Compact q-gram profiling of compressed strings. Theoretical Computer Science, 550(??):51– 58, September 18, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005404.

Chen:2014:AEA

[2315] Zhixiang Chen, Bin Fu, Randy Goebel, Guohui Lin, Weitian Tong, Jinhui Xu, Boting Yang, Zhiyu Zhao, and Binhai Zhu. On the approximability of the exemplar adjacency number problem for genomes with gene repetitions. Theoretical Computer Science, 550(??): 59–65, September 18, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005416.

Simpson:2014:PCW

[2316] Jamie Simpson. Palindromes in circular words. Theoretical Computer Science, 550(??):66–78, September 18, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005428.

Kari:2014:MLC

[2317] Lila Kari, Stavros Konstantinidis, and Steffen Kopecki. On the maximality of languages with combined types of code properties. Theoretical Computer Science, 550(??):79– 89, September 18, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005581.

Cho:2014:DIH

[2318] Da-Jung Cho, Yo-Sub Han, and Sang-Ki Ko. Decidability of involution hypercodes. Theoretical Computer Science, 550(??):90–99, September 18, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005593.

Haslegrave:2014:BHA

[2319] John Haslegrave. Bounds on Herman’s algorithm. Theoretical Computer Science, 550(??):100–106, September 18, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005660.

Dekking:2014:STM

[2320] Michel Dekking. On the structure of Thue–Morse subwords, with an application to dynamical systems. Theoretical Computer Science, 550(??):107– 112, September 18, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005672.

Anonymous:2014:EBaj

[2321] Anonymous. Editorial Board. Theoretical Computer Science, 551(??): iii–ix, September 25, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514006252.

Pardini:2014:CSB

[2322] Giovanni Pardini, Roberto Barbuti, Andrea Maggiolo-Schettini, Paolo Milazzo, and Simone Tini. Compositional semantics and behavioural equivalences for reaction systems with restriction. Theoretical Computer Science, 551(??): 1–21, September 25, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002898.

Chattopadhyay:2014:FTG

[2323] Swarup Chattopadhyay, C. A. Murthy, and Sankar K. Pal. Fitting truncated geometric distributions in large scale real world networks. Theoretical Computer Science, 551(??):22– 38, September 25, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514003521.

Maroosi:2014:AEP

[2324] Ali Maroosi and Ravie Chandren Muniyandi. Accelerated execution of P systems with active membranes to solve the N-queens problem. Theoretical Computer Science, 551(??):39– 54, September 25, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003533.

Si:2014:ECT

[2325] Chengyong Si, Jing An, Tian Lan, Thomas Ußmüller, Lei Wang, and Qidi Wu. On the equality constraints tolerance of Constrained Optimization Problems. Theoretical Computer Science, 551(??):55– 65, September 25, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003867.

Lassig:2014:ASP

[2326] Jörg Lässig and Dirk Sudholt. Analysis of speedups in parallel evolutionary algorithms and (1 + λ) EAs for combinatorial optimization. Theoretical Computer Science, 551(??):66– 83, September 25, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004976.

Bonet:2014:RAU

[2327] Blai Bonet, Patrik Haslum, Victor Khomenko, Sylvie Thiébaux, and Walter Vogler. Recent advances in unfolding technique. Theoretical Computer Science, 551(??):84– 101, September 25, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005209.

Bianchi:2014:SLB

[2328] Maria Paola Bianchi, Carlo Mereghetti, and Beatrice Palano. Size lower bounds for quantum automata. Theoretical Computer Science, 551(??):102– 115, September 25, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005210.

Mondal:2014:UCA

[2329] Nabarun Mondal and Partha Pratim Ghosh. Universal computation is ‘almost surely’ chaotic. Theoretical Computer Science, 551(??):116– 131, September 25, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005222.

Anonymous:2014:EBak

[2330] Anonymous. Editorial Board. Theoretical Computer Science, 552(??): iii–ix, October 2, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514006446.

Guo:2014:AHD

[2331] Ying-Jun Guo and Zhi-Xiong Wen. Automaticity of the Hankel determinants of difference sequences of the Thue–Morse sequence. Theoretical Computer Science, 552(??): 1–12, October 2, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005799.

Chen:2014:PDG

[2332] Fei Chen, Kwok wo Wong, Xiaofeng Liao, and Tao Xiang. Period distribution of generalized discrete Arnold cat map. Theoretical Computer Science, 552(??):13– 25, October 2, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005805.

Abdulla:2014:MRM

[2333] Parosh Aziz Abdulla, Yu-Fang Chen, Lukás Holík, and Tomás Vojnar. Mediating for reduction (on minimizing alternating Büchi automata). Theoretical Computer Science, 552(??): 26–43, October 2, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005817.

Cormode:2014:SLC

[2334] Graham Cormode and Hossein Jowhari.

A second look at counting triangles in graph streams. Theoretical Computer Science, 552(??):44– 51, October 2, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005945. See

corrigendum [2461].

Miller:2014:EMG

[2335] Zevi Miller, Dan Pritikin, and I. H. Sudborough. Embedding multidimensional grids into optimal hypercubes. Theoretical Computer Science, 552(??): 52–82, October 2, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005957.

Miklos:2014:CSS

[2336] István Miklós, Sándor Z. Kiss, and Eric Tannier. Counting and sampling SCJ small parsimony solutions. Theoretical Computer Science, 552(??): 83–98, October 2, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005969.

Izumi:2014:SES

[2337] Tomoko Izumi, Keigo Kinpara, Taisuke Izumi, and Koichi Wada. Spaceefficient self-stabilizing counting population protocols on mobile sensor networks. Theoretical Computer Science, 552(??):99–108, October 2, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005970.

Yuan:2014:NSS

[2338] Chen Yuan and Haibin Kan. A note on sparse solutions of sparse linear systems. Theoretical Computer Science, 552(??):109–111, October 2, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005982.

Anonymous:2014:EBal

[2339] Anonymous. Editorial Board. Theoretical Computer Science, 553(??): iii–ix, October 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006628.

Bar-Noy:2014:EAS

[2340] Amotz Bar-Noy, Thomas Erlebach, Magnús M. Halldórsson, Sotiris Nikoletseas, and Pekka Orponen. Editorial for Algorithms for Sensor Systems, Wireless Ad Hoc Networks and Autonomous Mobile Entities. Theoretical Computer Science, 553(??): 1, October 9, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006690.

Cohen:2014:SND

[2341] Asaf Cohen, Shlomi Dolev, and Guy Leshem. Sensor networks: From dependence analysis via matroid bases to online synthesis. Theoretical Computer Science, 553(??): 2–17, October 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005441.

Li:2014:EAN

[2342] Meng Li, Yota Otachi, and Takeshi Tokuyama. Efficient algorithms for network localization using cores of underlying graphs. Theoretical Computer Science, 553(??):18– 26, October 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001133.

Wendeberg:2014:PTA

[2343] Johannes Wendeberg and Christian Schindelhauer. Polynomial-time approximation algorithms for anchorfree TDoA localization. Theoretical Computer Science, 553(??):27– 36, October 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002746.

Yu:2014:DIC

[2344] Dongxiao Yu, Yuexuan Wang, Qiang-Sheng Hua, and Francis C. M. Lau. Distributed (Δ + 1)-coloring in the physical model. Theoretical Computer Science, 553(??):37– 56, October 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004149.

Halldorsson:2014:WCA

[2345] Magnús M. Halldórsson and Pradipta Mitra. Wireless capacity with arbitrary gain matrix. Theoretical Computer Science, 553(??):57– 63, October 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513007548.

Belke:2014:CSA

[2346] Lukas Belke, Thomas Kesselheim, Arie M. C. A. Koster, and Berthold Vöcking. Comparative study of approximation algorithms and heuristics for SINR scheduling with power control. Theoretical Computer Science, 553(??): 64–73, October 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004010.

Chan:2014:MPA

[2347] David Yu Cheng Chan and David Kirkpatrick. Multi-path algorithms for minimum-colour path problems with applications to approximating barrier resilience. Theoretical Computer Science, 553(??):74– 90, October 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002886.

Albagli-Kim:2014:PRI

[2348] Sivan Albagli-Kim, Leah Epstein, Hadas Shachnai, and Tami Tamir. Packing resizable items with application to video delivery over wireless networks. Theoretical Computer Science, 553(??):91–105, October 9, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513009213.

Disser:2014:MPH

[2349] Yann Disser, Subir Kumar Ghosh, Matús Mihalák, and Peter Widmayer. Mapping a polygon with holes using a compass. Theoretical Computer Science, 553(??):106– 113, October 9, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513009225.

Anonymous:2014:EBam

[2350] Anonymous. Editorial Board. Theoretical Computer Science, 554(??): iii–ix, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514006823.

Du:2014:CC

[2351] Ding-Zhu Du and Guochuan Zhang. Computing and combinatorics. Theoretical Computer Science, 554(??): 1, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514006896.

Gao:2014:SMA

[2352] Hongbiao Gao, Yuichi Goto, and Jingde Cheng. A systematic methodology for automated theorem finding. Theoretical Computer Science, 554(??): 2–21, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514004642.

Bianchi:2014:ACO

[2353] Maria Paola Bianchi, Hans-Joachim Böckenhauer, Juraj Hromkovic, Sacha Krug, and Björn Steffen. On the advice complexity of the online L(2, 1)- coloring problem on paths and cycles. Theoretical Computer Science, 554(??): 22–39, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004630.

Hell:2014:CDB

[2354] Pavol Hell and Aurosish Mishra. Hcoloring degree-bounded (acyclic) digraphs. Theoretical Computer Science, 554(??):40–49, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397514004381.

Baron:2014:LSP

[2355] Joshua Baron, Yuval Ishai, and Rafail Ostrovsky. On linear-size pseudorandom generators and hardcore functions. Theoretical Computer Science, 554(??): 50–63, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400437X.

Elmasry:2014:SRO

[2356] Amr Elmasry, Daniel Dahl Juhl, Jyrki Katajainen, and Srinivasa Rao Satti. Selection from read-only memory with limited workspace. Theoretical Computer Science, 554(??): 64–73, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004368.

Kobayashi:2014:LEK

[2357] Yasuaki Kobayashi, Hirokazu Maruta, Yusuke Nakae, and Hisao Tamaki. A linear edge kernel for two-layer crossing minimization. Theoretical Computer Science, 554(??):74– 81, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514004332.

Braverman:2014:HCI

[2358] Vladimir Braverman, Ran Gelles, and Rafail Ostrovsky. How to catch L_{2}- heavy-hitters on sliding windows. Theoretical Computer Science, 554(??): 82–94, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514004320.

Bockenhauer:2014:SGP

[2359] Hans-Joachim Böckenhauer, Juraj Hromkovic, Dennis Komm, Sacha Krug, Jasmin Smula, and Andreas Sprock. The string guessing problem as a method to prove lower bounds on the advice complexity. Theoretical Computer Science, 554(??):95– 108, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514004307.

Chlebik:2014:CBC

[2360] Miroslav Chlebík and Janka Chlebíková. Connection between conjunctive capacity and structural properties of graphs. Theoretical Computer Science, 554(??):

109–118, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003491.

Babu:2014:COG

[2361] Jasine Babu, Manu Basavaraju, L. Sunil Chandran, and Deepak Rajendraprasad. 2-connecting outerplanar graphs without blowing up the pathwidth. Theoretical Computer Science, 554(??):119–134, October 16, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003375.

Zhang:2014:OAS

[2362] Yong Zhang, Francis Y. L. Chin, Hing-Fung Ting, Xin Han, Chung Keung Poon, Yung H. Tsin, and Deshi Ye. Online algorithms for 1-space bounded 2-dimensional bin packing and square packing. Theoretical Computer Science, 554(??):135– 149, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001716.

Lin:2014:NPS

[2363] Yu-An Lin and Sheung-Hung Poon.

Non-planar square-orthogonal drawing with few-bend edges. Theoretical Computer Science, 554(??):150– 168, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001054.

Duan:2014:PDP

[2364] Zhenhua Duan and Cong Tian. A practical decision procedure for Propositional Projection Temporal Logic with infinite models. Theoretical Computer Science, 554(??):169–190, October 16, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514001042.

Cao:2014:SMP

[2365] Zhigang Cao, Xujin Chen, and Changjun Wang. Schedules for marketing products with negative externalities. Theoretical Computer Science, 554(??):191–206, October 16, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514000255.

Shi:2014:APM

[2366] Feng Shi, Jianxin Wang, Jianer Chen, Qilong Feng, and Jiong Guo. Algorithms for parameterized maximum agreement forest problem on multiple trees. Theoretical Computer Science, 554(??):207–216, October 16, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514000176.

Mans:2014:TDG

[2367] Bernard Mans and Luke Mathieson. On the treewidth of dynamic graphs. Theoretical Computer Science, 554(??): 217–228, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000164.

Zhang:2014:FPD

[2368] Nan Zhang, Zhenhua Duan, Cong Tian, and Dingzhu Du. A formal proof of the deadline driven scheduler in PPTL axiomatic system. Theoretical Computer Science, 554(??): 229–253, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513009262.

Cheong:2014:FAD

[2369] O. Cheong, R. El Shawi, and J. Gudmundsson. A fast algorithm for data collection along a fixed track. Theoretical Computer Science, 554(??): 254–262, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008931.

Georgiou:2014:SEN

[2370] Konstantinos Georgiou, George Karakostas, Jochen Könemann, and Zuzanna Stamirowska. Social exchange networks with distant bargaining. Theoretical Computer Science, 554(??): 263–274, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513008773.

Tong:2014:AMI

[2371] Weitian Tong, Randy Goebel, and Guohui Lin. Approximating the minimum independent dominating set in perturbed graphs. Theoretical Computer Science, 554(??):275– 282, October 16, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513008414.

Anonymous:2014:EBan

[2372] Anonymous. Editorial Board. Theoretical Computer Science, 555(??): iii–ix, October 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751400718X.

Ghosh:2014:GEF

[2373] Subir Kumar Ghosh and Takeshi Tokuyama. Guest Editors’ foreword. Theoretical Computer Science, 555 (??):1, October 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007257.

Bera:2014:AAP

[2374] Suman K. Bera, Shalmoli Gupta, Amit Kumar, and Sambuddha Roy. Approximation algorithms for the partition vertex cover problem. Theoretical Computer Science, 555(??): 2–8, October 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002734.

Bhattacharya:2014:CSP

[2375] Binay Bhattacharya, Soudipta Chakraborty, Ehsan Iranmanesh, and Ramesh Krishnamurti. The cyclical scheduling problem. Theoretical Computer Science, 555(??):9–22, October 23, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002564.

ElOuali:2014:RAA

[2376] Mourad El Ouali, Helena Fohlin, and Anand Srivastav. A randomised approximation algorithm for the hitting set problem. Theoretical Computer Science, 555(??):23– 34, October 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002515.

Uchizawa:2014:GRC

[2377] Kei Uchizawa, Takanori Aoki, Takehiro Ito, and Xiao Zhou. Generalized rainbow connectivity of graphs. Theoretical Computer Science, 555(??): 35–42, October 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000243.

Fekete:2014:CSD

[2378] Sándor P. Fekete, Jan-Marc Reinhardt, and Nils Schweer. A competitive strategy for distance-aware online shape allocation. Theoretical Computer Science, 555(??):43– 54, October 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001881.

Babu:2014:FOE

[2379] Jasine Babu, Ahmad Biniaz, Anil Maheshwari, and Michiel Smid. Fixedorientation equilateral triangle matching of point sets. Theoretical Computer Science, 555(??):55–70, October 23, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751300875X.

Chun:2014:BOL

[2380] Jinhee Chun, Takashi Horiyama, Takehiro Ito, Natsuda Kaothanthong, Hirotaka Ono, Yota Otachi, Takeshi Tokuyama, Ryuhei Uehara, and Takeaki Uno. Base-object location problems for base-monotone regions. Theoretical Computer Science, 555(??): 71–84, October 23, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397513008748.

Anonymous:2014:EBao

[2381] Anonymous. Editorial Board. Theoretical Computer Science, 556(??): iii–ix, October 30, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007415.

Widmayer:2014:COA

[2382] Peter Widmayer, Yinfeng Xu, and Binhai Zhu. Combinatorial optimization and applications. Theoretical Computer Science, 556(??): 1, October 30, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007488.

Kobayashi:2014:ICM

[2383] Yuki Kobayashi, Yuya Higashikawa, Naoki Katoh, and Naoyuki Kamiyama. An inductive construction of minimally rigid body-hinge simple graphs. Theoretical Computer Science, 556(??): 2–12, October 30, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006148.

Ma:2014:MHL

[2384] Huan Ma, Zaixin Lu, Deying Li, Yuqing Zhu, Lidan Fan, and Weili Wu. Mining hidden links in social networks to achieve equilibrium. Theoretical Computer Science, 556(??): 13–24, October 30, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006136.

Wu:2014:LAG

[2385] Ro-Yu Wu, Jou-Ming Chang, Hung-Chang Chan, and Kung-Jui Pai. A loopless algorithm for generating multiple binary tree sequences simultaneously. Theoretical Computer Science, 556(??):25–33, October 30, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006124.

Wylie:2014:FCD

[2386] Tim Wylie and Binhai Zhu. Following a curve with the discrete Fréchet distance. Theoretical Computer Science, 556(??):34–44, October 30, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514004629.

Ahadi:2014:TSD

[2387] Arash Ahadi, Amirhossein Mozafari, and Alireza Zarei. Touring a sequence of disjoint polygons: Complexity and extension. Theoretical Computer Science, 556(??):45– 54, October 30, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514004435.

Liu:2014:CCB

[2388] Tian Liu, Min Lu, Zhao Lu, and Ke Xu. Circular convex bipartite graphs: Feedback vertex sets. Theoretical Computer Science, 556(??): 55–62, October 30, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514003508.

Tong:2014:AMM

[2389] Weitian Tong, Randy Goebel, Tian Liu, and Guohui Lin. Approximating the maximum multiple RNA interaction problem. Theoretical Computer Science, 556(??):63–70, October 30, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003107.

Christ:2014:OCE

[2390] Marie G. Christ, Lene M. Favrholdt, and Kim S. Larsen. Online bin covering: Expectations vs. guarantees. Theoretical Computer Science, 556(??): 71–84, October 30, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004654.

Chen:2014:PAA

[2391] Zhi-Zhong Chen, Ying Fan, and Lusheng Wang. Parameterized and approximation algorithms for finding two disjoint matchings. Theoretical Computer Science, 556(??): 85–93, October 30, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002527.

Anonymous:2014:EBap

[2392] Anonymous. Editorial Board. Theoretical Computer Science, 557(??): iii–ix, November 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007543.

Honorat:2014:GFM

[2393] Anthony Honorat, Maria Potop-Butucaru, and Sébastien Tixeuil. Gathering fat mobile robots with slim omnidirectional cameras. Theoretical Computer Science, 557(??): 1–27, November 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005921.

Zhao:2014:ACE

[2394] Yancai Zhao, Zuhua Liao, and Lianying Miao. On the algorithmic complexity of edge total domination. Theoretical Computer Science, 557(??): 28–33, November 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005933.

Ahmed:2014:SMP

[2395] Pritom Ahmed, Costas S. Iliopoulos, A. S. M. Sohidull Islam, and M. Sohel Rahman. The swap matching problem revisited. Theoretical Computer Science, 557(??):34– 49, November 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514006173.

Borowiecki:2014:FVS

[2396] Mieczyslaw Borowiecki, Ewa Drgas-Burchardt, and Elzbieta Sidorowicz. A feedback vertex set of 2-degenerate graphs. Theoretical Computer Science, 557(??):50–58, November 6, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514006379.

Pai:2014:ICH

[2397] Kung-Jui Pai, Jou-Ming Chang, Jinn-Shyong Yang, and Ro-Yu Wu. Incidence coloring on hypercubes. Theoretical Computer Science, 557(??): 59–65, November 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006380. See

corrigendum [2460].

Wang:2014:SCR

[2398] Shiying Wang and Lei Zhang. Sufficient conditions for k-restricted edge connected graphs. Theoretical Computer Science, 557(??):66– 75, November 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006392.

Borowiecki:2014:BAC

[2399] Piotr Borowiecki, Dariusz Dereniowski, and Pawel Pralat. Brushing with additional cleaning restrictions. Theoretical Computer Science, 557(??): 76–86, November 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006707.

Bjorklund:2014:CFS

[2400] Henrik Björklund, Johanna Björklund, and Niklas Zechner. Compression of finite-state automata through failure transitions. Theoretical Computer Science, 557(??):87–100, November 6, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006720.

Cheng:2014:CDM

[2401] Eddie Cheng, Ke Qiu, and Zhizhang Shen. On the conditional diagnosability of matching composition networks. Theoretical Computer Science, 557(??): 101–114, November 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514006732.

Honkala:2014:RCF

[2402] Juha Honkala. Remarks concerning the freeness problem over morphism and matrix semigroups. Theoretical Computer Science, 557(??):115– 119, November 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514006343.

Lewenstein:2014:LSI

[2403] Moshe Lewenstein, J. Ian Munro, Venkatesh Raman, and Sharma V. Thankachan. Less space: Indexing for queries with wildcards. Theoretical Computer Science, 557(??):120– 127, November 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514006562.

Anonymous:2014:EBaq

[2404] Anonymous. Editorial Board. Theoretical Computer Science, 558(??): iii–ix, November 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007956.

Bshouty:2014:GEF

[2405] Nader H. Bshouty, Gilles Stoltz, Nicolas Vayatis, and Thomas Zeugmann. Guest Editors’ foreword. Theoretical Computer Science, 558(??):1– 4, November 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006999.

Gao:2014:CCP

[2406] Ziyuan Gao and Frank Stephan. Confident and consistent partial learning of recursive functions. Theoretical Computer Science, 558(??):5– 17, November 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007002.

Florencio:2014:RIV

[2407] Christophe Costa Florêncio and Sicco Verwer. Regular inference as vertex coloring. Theoretical Computer Science, 558(??):18–34, November 13, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007014.

Samei:2014:AMP

[2408] Rahim Samei, Pavel Semukhin, Boting Yang, and Sandra Zilles. Algebraic methods proving Sauer’s bound for teaching complexity. Theoretical Computer Science, 558(??):35– 50, November 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007026.

Ben-David:2014:DSC

[2409] Shalev Ben-David and Lev Reyzin. Data stability in clustering: a closer look. Theoretical Computer Science, 558(??):51–61, November 13, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007038.

Ortner:2014:RBR

[2410] Ronald Ortner, Daniil Ryabko, Peter Auer, and Rémi Munos. Regret bounds for restless Markov bandits. Theoretical Computer Science, 558(??): 62–76, November 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751400704X.

Carpentier:2014:MNS

[2411] Alexandra Carpentier and Rémi Munos. Minimax number of strata for online stratified sampling: the case of noisy samples. Theoretical Computer Science, 558(??):77– 106, November 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007051.

Moroshko:2014:WLS

[2412] Edward Moroshko and Koby Crammer. Weighted last-step min-max algorithm with improved sub-logarithmic regret. Theoretical Computer Science, 558(??): 107–124, November 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007063.

Lattimore:2014:NOP

[2413] Tor Lattimore and Marcus Hutter. Near-optimal PAC bounds for discounted MDPs. Theoretical Computer Science, 558(??):125– 143, November 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007075.

Koolen:2014:BLS

[2414] Wouter M. Koolen and Vladimir Vovk. Buy low, sell high. Theoretical Computer Science, 558(??):144– 158, November 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007087.

Warmuth:2014:KMU

[2415] Manfred K. Warmuth, Wojciech Kotlowski, and Shuisheng Zhou. Kernelization of matrix updates, when and how? Theoretical Computer Science, 558(??): 159–178, November 13, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007099.

Anonymous:2014:EBar

[2416] Anonymous. Editorial Board. Theoretical Computer Science, 559(??): iii–ix, November 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400807X.

Das:2014:NUC

[2417] Sukanta Das, Enrico Formenti, and Jarkko Kari. Non-uniform Cellular Automata. Theoretical Computer Science, 559(??):1–2, November 20, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008147.

Goles:2014:CCT

[2418] Eric Goles and Pedro Montealegre. Computational complexity of threshold automata networks under different updating schemes. Theoretical Computer Science, 559(??):3– 19, November 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006756.

Kuhlman:2014:ASN

[2419] Chris J. Kuhlman and Henning S. Mortveit. Attractor stability in nonuniform Boolean networks. Theoretical Computer Science, 559(??):20– 33, November 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006185.

Wu:2014:LCS

[2420] Sichao Wu, Abhijin Adiga, and Henning S. Mortveit. Limit cycle structure for dynamic bi-threshold systems. Theoretical Computer Science, 559(??):

34–41, November 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005088.

Mairesse:2014:APC

[2421] Jean Mairesse and Irène Marcovici.

Around probabilistic cellular automata. Theoretical Computer Science, 559(??):42–72, November 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397514006744.

Dennunzio:2014:TRD

[2422] Alberto Dennunzio, Enrico Formenti, and Julien Provillard. Three research directions in non-uniform cellular automata. Theoretical Computer Science, 559(??):73–90, November 20, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005143.

Salo:2014:RPN

[2423] Ville Salo. Realization problems for nonuniform cellular automata. Theoretical Computer Science, 559(??):91– 107, November 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006161.

Manzoni:2014:FSS

[2424] Luca Manzoni and Hiroshi Umeo. The firing squad synchronization problem on CA with multiple updating cycles. Theoretical Computer Science, 559(??):

108–117, November 20, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514006203.

Anonymous:2014:EBas

[2425] Anonymous. Editorial Board. Theoretical Computer Science, 560 (part 1) (??):iii–ix, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008676.

Anonymous:2014:EBat

[2426] Anonymous. Editorial Board. Theoretical Computer Science, 560 (part 2) (??):iii–ix, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008780.

Anonymous:2014:EBau

[2427] Anonymous. Editorial Board. Theoretical Computer Science, 560 (part 3) (??):iii–ix, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008901.

Mor:2014:P

[2428] Tal Mor and Renato Renner. Preface. Theoretical Computer Science, 560 (part 1)(??):1–6, December 4, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007828.

Bennett:2014:QCP

[2429] Charles H. Bennett and Gilles Brassard. Quantum cryptography: Public key distribution and coin tossing. Theoretical Computer Science, 560 (part 1) (??):7–11, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004241.

Damgaard:2014:SIQ

[2430] Ivan Damgård, Serge Fehr, Louis Salvail, and Christian Schaffner. Secure identification and QKD in the bounded-quantum-storage model. Theoretical Computer Science, 560 (part 1) (??):12–26, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006926.

Scarani:2014:BPQ

[2431] Valerio Scarani and Christian Kurtsiefer. The black paper of quantum cryptography: Real implementation problems. Theoretical Computer Science, 560 (part 1)(??):27– 32, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006938.

Ioannou:2014:PKC

[2432] Lawrence M. Ioannou and Michele Mosca. Public-key cryptography based on bounded quantum reference frames. Theoretical Computer Science, 560 (part 1)(??):33– 45, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400694X.

Lucamarini:2014:QKD

[2433] Marco Lucamarini and Stefano Mancini. Quantum key distribution using a twoway quantum channel. Theoretical Computer Science, 560 (part 1)(??): 46–61, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006951.

Alleaume:2014:UQK

[2434] R. Alléaume, C. Branciard, J. Bouda, T. Debuisschert, M. Dianati, N. Gisin, M. Godfrey, P. Grangier, T. Länger, N. Lütkenhaus, C. Monyk, P. Painchault, M. Peev, A. Poppe, T. Pornin, J. Rarity, R. Renner, G. Ribordy, M. Riguidel, L. Salvail, A. Shields, H. Weinfurter, and A. Zeilinger. Using quantum key distribution for cryptographic purposes: a survey. Theoretical Computer Science, 560 (part 1)(??): 62–81, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006963.

Svozil:2014:NCC

[2435] Karl Svozil. Non-contextual chocolate balls versus value indefinite quantum cryptography. Theoretical Computer Science, 560 (part 1)(??):82– 90, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006975.

Watanabe:2014:OAC

[2436] Shun Watanabe, Ryutaroh Matsumoto, and Tomohiko Uyematsu. Optimal axis compensation in quantum key distribution protocols over unital channels. Theoretical Computer Science, 560 (part 1)(??):91– 106, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006987.

Barmpalias:2014:TAM

[2437] George Barmpalias, Manindra Agrawal, and S. Barry Cooper. Theory and applications of models of computation at the Turing Centenary in China. Theoretical Computer Science, 560 (part 2) (??):107, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008858.

Baumann:2014:WCB

[2438] Hervé Baumann, Pierre Fraigniaud, Hovhannes A. Harutyunyan, and Rémi de Verclos. The worst case behavior of randomized gossip protocols. Theoretical Computer Science, 560 (part 2)(??): 108–120, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400615X.

Bi:2014:LBS

[2439] Jingguo Bi and Qi Cheng. Lower bounds of shortest vector lengths in random NTRU lattices. Theoretical Computer Science, 560 (part 2)(??): 121–130, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007737.

Bollig:2014:ICM

[2440] Beate Bollig, Marc Gillé, and Tobias Pröger. Implicit computation of maximum bipartite matchings by sublinear functional operations. Theoretical Computer Science, 560 (part 2)(??): 131–146, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005635.

Escoffier:2014:AMS

[2441] Bruno Escoffier, Vangelis Th. Paschos, and Emeric Tourniaire. Approximating MAX SAT by moderately exponential and parameterized algorithms. Theoretical Computer Science, 560 (part 2)(??):147–157, December 4, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008317.

Feng:2014:IPA

[2442] Qilong Feng, Jianxin Wang, Chao Xu, Jinyi Yao, and Jianer Chen. Improved parameterized algorithms for minimum link-length rectilinear spanning path problem. Theoretical Computer Science, 560 (part 2)(??):158– 171, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005647.

Jiang:2014:TEA

[2443] Shaoquan Jiang. Timed encryption with application to deniable key exchange. Theoretical Computer Science, 560 (part 2)(??):172– 189, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400098X.

Shalom:2014:OOB

[2444] Mordechai Shalom, Ariella Voloshin, Prudence W. H. Wong, Fencol C. C. Yung, and Shmuel Zaks. Online optimization of busy time on parallel machines. Theoretical Computer Science, 560 (part 2)(??):190– 206, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400560X.

Xiao:2014:REA

[2445] Mingyu Xiao and Hiroshi Nagamochi.

A refined exact algorithm for Edge Dominating Set. Theoretical Computer Science, 560 (part 2)(??):207– 216, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005623.

Faella:2014:PSI

[2446] Marco Faella and Aniello Murano.

Preface to the special issue on GandALF 2012. Theoretical Computer Science, 560 (part 3)(??):217–218, December 4, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008974.

Neider:2014:BHS

[2447] Daniel Neider, Roman Rabinovich, and Martin Zimmermann. Down the Borel hierarchy: Solving Muller games via safety games. Theoretical Computer Science, 560 (part 3)(??):219– 234, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000346.

Canavoi:2014:DSI

[2448] Felix Canavoi, Erich Grädel, and Roman Rabinovich. The discrete strategy improvement algorithm for parity games and complexity measures for directed graphs. Theoretical Computer Science, 560 (part 3)(??):235– 250, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000024.

Solimando:2014:VXD

[2449] Alessandro Solimando, Giorgio Delzanno, and Giovanna Guerrini. Validating XML document adaptations via Hedge Automata transformations. Theoretical Computer Science, 560 (part 3)(??): 251–268, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514003156.

Bresolin:2014:ITL

[2450] Davide Bresolin, Dario Della Monica, Angelo Montanari, Pietro Sala, and Guido Sciavicco. Interval temporal logics over strongly discrete linear orders: Expressiveness and complexity. Theoretical Computer Science, 560 (part 3)(??):269–291, December 4, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002552.

Gutierrez:2014:CAH

[2451] Julian Gutierrez, Felix Klaedtke, and Martin Lange. The μ-calculus alternation hierarchy collapses over structures with restricted connectivity. Theoretical Computer Science, 560 (part 3)(??): 292–306, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002497.

Cantone:2014:DTS

[2452] Domenico Cantone and Cristiano Longo. A decidable two-sorted quantified fragment of set theory with ordered pairs and some undecidable extensions. Theoretical Computer Science, 560 (part 3)(??):307– 325, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002333.

Lange:2014:MCP

[2453] Martin Lange, Etienne Lozes, and Manuel Vargas Guzmán. Modelchecking process equivalences. Theoretical Computer Science, 560 (part 3)(??):326–347, December 4, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397514006574.

Cerny:2014:ISD

[2454] Pavol Cerný, Martin Chmelík, Thomas A. Henzinger, and Arjun Radhakrishna. Interface simulation distances. Theoretical Computer Science, 560 (part 3)(??):348–363, December 4, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006550.

Dimovski:2014:PVU

[2455] Aleksandar S. Dimovski. Program verification using symbolic game semantics. Theoretical Computer Science, 560 (part 3)(??):364–379, December 4, 2014. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514000334.

Hosoda:2015:CAH

[2456] Jun Hosoda, Juraj Hromkovic, Taisuke Izumi, Hirotaka Ono, Monika Steinová, and Koichi Wada. Corrigendum to “On the approximability and hardness of minimum topic connected overlay and its special instances” [Theoret. Comput. Sci. 429 (2012) 144–154]. Theoretical Computer Science, 562(??): 660–661, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008445. See

[1090].

Calamoneri:2015:CPC

[2457] T. Calamoneri and R. Petreschi. Corrigendum to “On pairwise compatibility graphs having Dilworth number two” [Theoret. Comput. Sci. 524 (2014) 34–40]. Theoretical Computer Science, 602(??):158–159, October 18, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007677. See

[2057].

Hew:2016:WBP

[2458] Patrick Chisan Hew. Working in binary protects the repetends of 1/3^{h}: Comment on Colussi’s ‘The convergence classes of Collatz function’. Theoretical Computer Science, 618(??): 135–141, March 7, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515012037. See

[807].

Feigenblat:2016:CFI

[2459] Guy Feigenblat, Ofra Itzhaki, and Ely Porat. Corrigendum to “The frequent items problem, under polynomial decay, in the streaming model” [Theoret. Comput. Sci. 411(34-36) (2010) 3048– 3054]. Theoretical Computer Science, 648(??):116, October 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/

                                        article/pii/S030439751630456X. See

[237].

Pai:2016:CIC

[2460] Kung-Jui Pai, Jou-Ming Chang, Jinn-Shyong Yang, and Ro-Yu Wu. Corrigendum to “Incidence coloring on hypercubes” [Theoret. Comput. Sci. 557 (2014) 59–65]. Theoretical Computer Science, 651(??):87–88, October 25, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303887. See

[2397].

Cormode:2017:CSL

[2461] Graham Cormode and Hossein Jowhari. Corrigendum to “A second look at counting triangles in graph streams” [Theoret. Comput. Sci. 552 (2014) 44– 51]. Theoretical Computer Science, 683 (??):31–32, June 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302924.